A Complete Bibliography of Publications in Journal of Fish Diseases (2000–2009)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

22 August 2024 Version 1.04

Title word cross-reference

+ [DW05a]. _{3} [GWF^{+}02]. _{4} [XNS04]. _{50} [KM01]. α [BWD^{+}07]. β [BCMN05, CHWK^{+}04, DW05b, EKG^{+}02, GMS06, GMS07, KS06, LSDB06, PZS02]. κ [HGH^{+}08]. × [BWCM04, MOG09, SBG08]. x_{L} [CYL^{+}06].

-1 [GMS06, GMS07, KS06]. -3 [HHW05]. -actin [DW05b]. -affected [VMN06]. -ATPase [DW05a]. -glucan [BCMN05]. -haemolytic [EKG^{+}02]. -K [DW05a]. -oestradiol [CHWK^{+}04].

. [YCR^{+}09].

001 [HFW01].

1 [BOB^{+}06, HHCBR^{+}07, LSDB06, NPF05, PZS02]. 16S [RLL^{+}07]. 18S [SKR^{+}08].

2 [GSMM09, HLC00, TFN01]. 200 [KM01]. 20E [ASEK03]. 28S [WCT^{+}01].

3 [SMG08]. 3/1 [GMS06, GMS07]. 33 [ASEK03].

51 [IRA03]. 53 [BSTW08].

6-glucan [GMS06].

= [DW05a, DW05b].

A. [FLGS00, IRA03]. A199 [LG03]. A3 [IRA03]. A3-51 [IRA03]. abalone [CCTL06]. abdominal [AGM^{+}08]. aberrant [SSM00]. ability [CN02]. abraded [BNS03]. Abramis [MCS08]. absence [GER^{+}06, NDH05a]. abundance [SKCD07, TWS09]. AC [SRYC02, STV06]. acanthias [BF02a, BF02b, CKK02]. Acanthocephala [DGS^{+}02]. accumulation [KHH^{+}05]. accuracy [GEB^{+}08]. achromogenes [BGBG05]. Achtheres [KPWK06]. acid [ATT^{+}09, AL09, MST^{+}04, RAT^{+}02, RDT^{+}03, SBE03]. acids [ASEK03]. Acipenser [DLS^{+}07]. acquired [SSM01]. acquisition [MEBD05]. across [JOG^{+}06, JSG^{+}08]. actin [DW05b]. Actinobdella [WCB08]. actinospore [AHB07, KEEM^{+}05]. activated [HHW05]. activates [HGH^{+}08]. activating [CTAR01]. active [PZS02]. activities [DRCA^{+}06]. Activity [CUHN08, BCMN05, CFA^{+}00, DRCA^{+}08, LG03, NHA07, PZS02, SBE03, SRYC02, STV06, TFN01, WKES08]. aculeatus [JKD06]. acus [KBPD04, LGF04]. Acute [KM01, TJG04a, DM09, JGS^{+}04]. adenoviral [OLR03]. Adherence [SKI^{+}06]. adherent [MGE04]. adhesin [FLGS00]. Adhesion [CRBM05, VSNS08]. Adjuvant [KMJM06, BCM^{+}02, SMS02]. adjuvanted [SSL^{+}08b]. administered [ZM01a]. administration [BCMN05, CN04, DRCA^{+}08, FHY^{+}08, GMS06, GMS07, IRA03, PWS^{+}08, RLNB03, SSL^{+}08b, VNSB00]. Adriatic [RA08]. adult [BB01, CR02, CRM^{+}03, GKOF09, GPRCH09, MRM^{+}00, MALC00, ØSKH06, RWSK09]. aeglefinus [TBLC07]. aeromonads [Koz07, OBAA09]. Aeromonas [BBLC00, BGBG05, BF07, CFA^{+}00, CL05, FLGS00, GGCFB09, GCWFB08, GHBW03, GPS^{+}08, HKV^{+}06, IRA03, JOG^{+}06, JSG^{+}08, KMJM06, KFM^{+}00, LG03, MPP^{+}03, MBG^{+}02, NLO08, NM02, NA09a, NA09b, OR05, OL06, TBLC07, VSNS08, WBP^{+}05, YCR^{+}09]. aetiological [MBD^{+}02]. aetiology [RST^{+}07, SGRM07, SRT^{+}07]. affect [BBBW01, BBC^{+}05, RAG^{+}07]. affected [AN01, LHHP07, MN08, VMN06]. affecting [KaVT06, KaVKT09, LAZ05]. affects [GMS07]. after [AN04b, CIK^{+}04, CN04, EN03, GPS^{+}08, HGP^{+}07, JGS^{+}04, KKD^{+}09, OL06, PWS^{+}08, RTNA02, XKS04]. against [AW01, AW02, ARC^{+}05, ABC^{+}08, BVG^{+}03, BBN03, BNN03, CUHN08, CPRL09, DM09, DMS09, DRCMB03, FLGS00, GLMG03, GMS06, GMS07, GER^{+}06, HBB09, ILAW03, JAM^{+}01, JTM^{+}01, LLJC03, LCM^{+}01, LG03, LBGR08, LMX^{+}05, MLB^{+}06, MEC^{+}06, MBB^{+}06, PSP^{+}09, RLNB03, SBE03, SPC^{+}08, SWTC01, SB01, SSS^{+}00, TFN01, TMA^{+}01, VSST09, WM04, XK02, XKS04, XKP06, YMKN09, YMN09, YMNH04, ZWW^{+}06, ZWZS02]. agalactiae [DAMQ^{+}04, EKG^{+}02, OFKEA08, PEP^{+}05]. agar [TO02]. Agassiz [EKS09]. AGD [AEN04, SKR^{+}08, VMN06]. agent [BBLC00, CWJL07, GEC09, HMO^{+}08, HWZ^{+}00, KSKEM07, LLJC03, MMA02a, VTA06, WBP^{+}05, YDNM08, ZM01b]. agents [MMTPDR^{+}08, PBW06]. air [FF07, LCH^{+}02]. Alaska [AAH^{+}07]. Alburnus [MEM^{+}09]. alga [CCTL06]. algae [ZRT04]. alginolyticus [CWJL07, JPPS06, YML^{+}08, ZWZS02]. alimentary [BSG^{+}05]. alkaline [CWJL07, WMA^{+}02]. alleles [RLL^{+}07, XCD09]. Allium [NA09b]. allogynogenetic [WYWL01]. almost [MGAE06]. alone [SSL^{+}08b]. Alopias [BT09, BKSB09]. Alphavirus [MG07, GJR^{+}03, GJR^{+}05, GRF^{+}07, GCWR07, JTR^{+}04, PSS09]. alphaviruses [FRW^{+}08, GSW^{+}07, GWJR08]. alpinus [MPP^{+}03, Sta01]. alterations [CIK^{+}04]. Altered [SSM01]. alternate [ESMB00, LYW03]. alternative [GTBA09, PB04]. alters [MDB^{+}09]. altivelis [BTSH03, KIIK04, YKR^{+}02]. aluminium [SI06]. alutus [WWK03]. alvei [PRG^{+}05]. amago [ILAW03]. amamiensis [BBA08]. Amazonian [PD01]. America [CDI^{+}01]. American [BWS^{+}08, EFT^{+}06, MFLC01]. americanum [BF02b]. americanus [BCB02, BWS^{+}08, DSBB05, Dov05]. amines [LMG^{+}03]. amoeba [SKR^{+}08]. Amoebic [AN03, VMN06, AN01, AN04a, AEN04, AN04b, AVN08, BM04, BM06, BN04a, BBN03, BCMN05, FZM00, FBP09, LHHP07, MCN04, MN08, PNA02, RP03, SKR^{+}08, YDNM08, ZM00, ZM01b, ZGM01]. among [CLN05, MOO^{+}09, WSKC03]. amongst [GDG00, NBB08]. amphizoic [CAAN07]. amplification [AL09, CGMS03, EMS05, FNE09, GEC09, GKVS04, KJC^{+}02, MBLK06, PBW06, SKIS05, SSK^{+}08, TYH^{+}01, WCSC07, WCK^{+}09, XMM^{+}09]. anaemia [BLL^{+}07, GMS^{+}09, GER^{+}07, GEB^{+}08, HD05a, HD05b, HO03, MEC^{+}06, MHD^{+}04, MBLK06, NDH^{+}05b, NDH05a, RMG^{+}09, RW03, SBM^{+}00, SFPF01, SLC^{+}06, SRM^{+}03, SMM09, WMA^{+}02]. anal [CN04]. analyses [ASEK03, FRW^{+}08]. Analysis [BF07, GHBW03, GBB^{+}06, BGR08, GDG00, HSK^{+}08, IFA05, KJC^{+}02, LNZL08, MG09, NHA05, OLP01, RBBC^{+}09, SGEM03, TSL^{+}01, WCSC07, WCK^{+}09, WCT^{+}01]. Anarhichas [áKBH06]. anarhichatis [áKBH06]. angelfish [PDOZ08]. anglerfish [FYO08]. Anguilla [CL02, GKB05, JNS^{+}09, LG03]. anguillae [JNS^{+}09]. anguillarum [FLGS00, GOS04, KBH^{+}08, RHS^{+}06, RMHA08, SBE03, SBM^{+}06, SMLZ05, SRAHJ^{+}08, SHN^{+}00, XMM^{+}09, ZWZS02]. Anguillicola [GKB05]. anguilliculture [FLA06]. anguilliseptica [BFR07, FCM^{+}04]. animal [BSM^{+}06]. Animals [Rob05a]. anko [FYO08]. Announcement [Ano01a]. Antagonistic [LG03]. anterior [MN08]. anti [GCN04, SKSE02, ZM01b]. anti- [GCN04, SKSE02, ZM01b]. antibacterial [DBR04, MMTPDR^{+}08, SBE03]. antibiotics [BDSY07]. antibodies [AW01, GCN04, JAM^{+}01, JTM^{+}01, LCM^{+}01, LMX^{+}05, LLWW02, LLZ08, SWTC01, XKS02, XKS05, ZM01b]. Antibody [SHBJ^{+}09, XK02, CPRL09, GLC04, HFW01, LLJ^{+}02, SPC^{+}08, SB01, XK03, YMKN09]. antibody-titres [SB01]. antigen [BPV^{+}03, FDG^{+}06, MMAP00, SLC08]. Antigenic [LLWW02, SSO^{+}09, JTR^{+}04]. Antigenicity [PEP^{+}05]. antigens [ZM01a]. antimicrobial [CUHN08, SXSN07]. antioxidants [WHB^{+}03]. antisera [MEC^{+}06]. antiserum [ILAW03, YMNH04]. antonactinomyxon [AHB07]. Aphanomyces [JZKS04, LHP^{+}03]. API [ASEK03]. Apicomplexa [GA05]. Aplodinotus [WCB08]. Apoptosis [WM04, XKS06, CCL09, HHW05, HHL^{+}05, HCFM04, HLC00, LCC^{+}08, VNSB00, XKS05]. apparently [IGH00]. appearance [WCB08]. applicable [SMM09]. Application [CFA^{+}09, KBH^{+}08, MG09, KJL07, MEC^{+}06, SWTC01, YMNH04]. Applied [Rob00]. approach [DGLS09, OB07]. Approaches [Pen04]. aquabirnavirus [YMN09]. Aquabirnaviruses [ZS04, RBBC^{+}09]. Aquaculture [Pen04, CLN05, HOA09, IA02a, ØSKH06, RGTR02, TJG04a, VLK^{+}04, WGM^{+}08, WMB02]. aquacultured [JDBN08]. AQUAEIA [GOS04]. AQUAEIA-Va [GOS04]. AQUARAPID [GOS04]. AQUARAPID-Va [GOS04]. aquarium [BMD^{+}06]. Aquatic [HHW05, HGH^{+}08, Rob05a, PTT^{+}08, PTAR09]. Arcachon [DGMD^{+}09]. Archipelago [SKCD07]. Arctic [MPP^{+}03, Sta01]. Ardéche [EWBB02]. area [LBO^{+}01, OMLF^{+}06]. areas [BEL09, JO00]. argenteus [DAMQ^{+}04]. Argulus [TSW06, TWS09, VNSB00]. ariakensis [WP00]. aroA [TRY^{+}05, YTLN01]. Artemia [BDSY07, CR02, CRM^{+}03, MRM^{+}00, SBSH08, SAH^{+}07, SAC^{+}09]. arteries [SSL^{+}08a]. Arthromitus [MBD^{+}02]. artificial [IKKS06]. ascidian [HOA09]. ascites [KIIK04]. Asian [GKCF04, KS06, MMB02, PSP^{+}09]. aspects [CMD02, GMS^{+}09, MBD^{+}02]. assay [AOA^{+}09, AL09, CWS02, CL05, CAAN07, DZP^{+}00, EMS05, FNE09, GJR^{+}03, GJR^{+}05, PTT^{+}08, RLL^{+}07, RB03, YWBH05]. assays [GOS04, PGH^{+}07, SB01, VCLS04]. assess [CAAN07]. Assessment [MHD^{+}04, BCB02, PFL^{+}08, PD09, RRG^{+}05, SPL05]. associated [BM04, BM06, BSQW05, BD00, BKNS02, BF02b, BC06, BB07, CKW^{+}01, DGMD^{+}09, DPS^{+}07, EWBB02, GCG^{+}04, GER^{+}07, HD05b, HSD07, IOK^{+}07, JO00, Kha04a, KHH^{+}05, LBO^{+}01, LBGR08, LHW08, MFLC01, MBD^{+}02, MST^{+}04, MO06, OMR^{+}06, PSS09, QSZ^{+}03, RJB00, RMHA08, VFT^{+}03, WCSC07, WP00, XNS04, XKS06, XCD09, ZB07]. Association [AGM^{+}08, APR03, LMG^{+}03]. Asterropteryx [TL09]. asymptomatic [GUE06]. Atlantic [AN01, AN03, AN04a, AEN04, AN04b, AVN08, ATT^{+}09, AW02, AGM^{+}08, BLGR09, BMPB^{+}00, BMKG^{+}07, BMFL^{+}07, BM06, BBLC00, BBFB07, BBBW01, BF02a, BF02b, BT09, BKSB09, BSE02, Bow03, BMS^{+}00, BCM^{+}00, BBC^{+}05, BBN03, BCMN05, BBK04, BLT01, BN04b, CBDC^{+}03, CGW04, CAAN07, DCCF^{+}03, FZM00, FM00, FBP09, FTE^{+}03, GME04, GUE06, GA05, GKOF09, GWF^{+}02, GCR^{+}02, GJR^{+}05, GMW^{+}07, GMS^{+}09, GMBN04, GCN04, GPB^{+}05, GLMG03, GER^{+}06, GER^{+}07, GBB^{+}06, HRG03, JTR^{+}04, JGS^{+}04, JHKB08, Kha04a, KKT^{+}04, KHTF06, KT09, KFM^{+}00, KHH^{+}05, KKD^{+}09, LHHP07, LGR08, LJM08, LSDB06, MASS07, MHD^{+}04, MNM^{+}02, MOR^{+}09, MN08, MGE04, MGAE06, ME08, NFR08, NDB^{+}04, ØSKH06, OLB03, OMR^{+}06, ØGW04, ONI^{+}08, ONS07, PMZD08, PSS09, PTB07, PNA02, RM08, RGTR02, RRG^{+}05, RMG^{+}09, RHS01, RP03]. Atlantic [RP05, Rod07, RM07, RDT^{+}03, RAG^{+}07, SSH^{+}07, SKCD07, SNNB01, SBK^{+}05, SPL05, SPK^{+}05, SSL^{+}08a, SBM^{+}00, SFPF01, SSL^{+}08b, SBB^{+}06, SM08, SWS07, SRM^{+}03, SKGR05, SKR^{+}08, SSS^{+}00, SBSR02, SHS05, SGRM07, SRT^{+}07, SHRM07, SOF^{+}09, TOB^{+}07, TSW00, USR^{+}09, UMG^{+}08, UBB^{+}09, VJR^{+}09, VHM^{+}00, VCN09, VMN06, WHB^{+}03, WGM^{+}08, ZM00, ZM01a]. Atlas [Hum07]. ATP [HKV^{+}06]. ATP-dependent [HKV^{+}06]. ATPase [DW05a, KL07]. Atractoscion [CDI^{+}01]. attachment [BD00, BC06, CAAN07, VLK^{+}04]. Attempts [RF01]. attenuated [SMLZ05, TRY^{+}05]. attributed [HD05b]. Atypical [HKPKS06, MBG^{+}02, TBLC07, BF07, GHBW03, GLMG03, IRA03, MPP^{+}03]. aurantiactinomyxon [HAEM02]. aurata [AVP^{+}09, ATNF02a, CIK^{+}04, CFA^{+}09, CRA^{+}05, CVGG06, MMTPDR^{+}08, MST^{+}04, PRG^{+}05, RKG^{+}00, RAT^{+}02, SBDPAP07]. auratum [AHB07]. auratus [GRL04, WYWL01, ZWW^{+}06, AHB07, EKG^{+}02, GRL04, IRA03, JBB^{+}07, MGB^{+}07, PB04, SMG08, ZWW^{+}06]. aureus [HLC00]. auriga [LVA^{+}06]. Australia [BTSH03, BBA08, JDBN08, NBB08, WCJJ08]. Australian [PGH^{+}07, SCJ^{+}01, SCJ^{+}02]. australis [LG03]. Austropotamobius [EWBB02]. Author [Ano01b, Ano02a, Ano03a, Ano04a, Ano05a]. autochthonous [MDB^{+}09]. autogenous [GLMG03]. auxotrophic [TRY^{+}05]. avian [HLWH07]. avidus [SSO^{+}09]. avirulent [FW01]. avoid [MD08]. awoara [LMJ^{+}00, LMC^{+}01, LCM^{+}01, LJL^{+}03, MWG^{+}02]. ayu [KIIK04].

B [CAB^{+}07, EKG^{+}02, HGH^{+}08, XCD09]. B-cell [CAB^{+}07]. B. [RLNB03]. Bacillary [FTS^{+}01]. bacilliform [EWBB02]. Bacillus [BDSY07, CL02, HBB09, RLNB03]. bacteraemia [BWS^{+}08, NM02, RST^{+}07]. Bacteria [Rob05a, BDSY07, BN04a, GCWFB08, VLK^{+}04]. Bacterial [DM00, KHK^{+}04, BHP^{+}04, BWCM04, BNFA07, CL02, CHWK^{+}04, FH03, GEC09, JLSA08, JSG^{+}08, KWD^{+}01, KIIK04, LLJC03, LB00, PSTV06, SAC^{+}09, TH08, VHM^{+}00]. bactericidal [PZS02]. bacterin [CN04]. bacteriophage [APR03]. bacterium [AJA03, CDGF07, MHKA08, MBD^{+}02, OMR^{+}06]. baculovirus [SBSH08]. balloon [PNS^{+}05]. Baltic [ARKK08, LSDB06, LBN02]. bar [Cvi04a]. Barbatula [KBPD04]. barramundi [BTSH03, BB07, BCM^{+}02, CCL09, CN04, FDG^{+}06, HTC^{+}01, LCC^{+}08, PC04]. barrier [SOF^{+}09]. Baryancistrus [PD01]. based [DZP^{+}00, FRW^{+}08, GJR^{+}03, GJR^{+}05, HSK^{+}08, HTC^{+}01, MCN04, WDC^{+}03, WMA^{+}02, CLVO^{+}05, FBS03, GSM^{+}04, IFA05]. Basilewsky [HWZ^{+}00, PWS^{+}08, SXSN07]. bass [BPV^{+}03, BVG^{+}03, BWCM04, BPBT02, CLL^{+}00, CZP01, GMB^{+}07, GPN05, HO03, MOG09, PSP^{+}09, RTNA02, RNAA04, RP04, SBDPAP07, TMM01, WBNG02]. bath [AN04b, SBSR02]. bathing [RP03]. batrachus [KS06]. Bay [DGMD^{+}09, WCB08]. beads [TL09]. beakperch [JO00]. before [CVGG06]. behaviour [LB09]. belly [GKCF04]. belong [SXSN07]. belonging [CDT^{+}07, DGA^{+}03, OMR^{+}06]. Benedenia [EWCT05, WCT^{+}01]. benthic [IFA05]. benzoate [GER^{+}06, LBGR08, SSS^{+}00]. Berthe [HDP^{+}02]. best [ME08]. beta [BCM^{+}02]. betanodavirus [BOB^{+}06, BCB^{+}09, CAB^{+}07, FDG^{+}06, PSP^{+}09, YMKN09, YMN09, YKR^{+}02, MKM02]. betanodaviruses [CDT^{+}07, GSM^{+}04, SMN^{+}09]. Betta [PTKA02]. betulinol [CHWK^{+}04]. between [BBLC00, BN04b, CGP08, DCCF^{+}03, GPRCH09, KJB07, KH06, KaVKT09, LGR08, ME08, PFL^{+}08, TF04, XSK07, ZASK06]. between-lake [KaVKT09]. binding [BAS^{+}09, BHO05]. Bioavailability [RNAA04]. Bioencapsulation [CR02, MRM^{+}00]. biogenic [LMG^{+}03]. Biography [Eve04, RH04, Win01]. BiographyBiography [Hof00]. Biolog [ASEK03]. biological [MES^{+}02]. biologically [PZS02]. Biology [Rob05b]. biomarkers [SSH^{+}07]. Biophysical [GSW^{+}07]. BioPlus2B [RLNB03]. Biotechnology [Pen04]. birnavirus [HHW05, HGH^{+}08, INKS04, XLZS08]. bithionol [FBP09]. bivalve [LB00, NPF05, RA01]. bivalves [ALAC09]. black [KHA04b, KL07, ZB07]. blackspot [ACGR^{+}05]. bladder [PDOZ08]. bleak [MEM^{+}09]. blister [PNS^{+}05]. bloat [And06]. Bloch [BB07, BCM^{+}02, CWJL07, CCL09, CN04, FDG^{+}06, GKCF04, HTC^{+}01, IGH01, KBPD04, LCC^{+}08, MOO^{+}09, MKW^{+}09, PSP^{+}09, PC04, WYWL01]. blood [CMD01, GAMR^{+}04, HO03, MGE04]. blot [DCTCB^{+}06, DHCHN01, GOS04, PTKA02, SGEM03, WDC^{+}03]. Blue [BWD^{+}07, ACP01, AVN08, BD00, BKNS02, BHB03, BHRB04, BSTW08, BT09, BKSB09, CLQAGCRP08, FLGS00, NSK00, PDOZ08, SBG08, Shi02]. Blue-Cell [Shi02]. bluefin [CMD01, DBJN05, HBN09, JDBN08, MMSM06, RA08, SMN^{+}09, WMB02]. bluespot [GCG^{+}04]. bluespot-like [GCG^{+}04]. body [PSB06, Rod07]. bogaraveo [ACGR^{+}05]. Bonamia [HDP^{+}02]. bone [CM04, DGLS09, GWF^{+}02]. Bonnaterre [BT09, BKSB09]. Book [Ano02b, Rob02a, Rob02b, Bru00, Tac00]. Boone [MBB^{+}06, RBA^{+}03]. borne [SRM^{+}03]. Borneo [BC06]. boundaries [SSM00]. bovine [KMJM06]. brachymystacis [YYY^{+}06, YWS^{+}08]. Brachymystax [YWS^{+}08]. brain [LMC^{+}01, SKI^{+}06]. brama [AVN08, MCS08]. bramae [ESMB00]. branchialis [SWS07]. Branchiomycosis [PD01]. branchiostegal [PNS^{+}05]. branchiphila [DNC^{+}05]. Branchiura [TSW06]. Brazil [FKA^{+}05, LSM^{+}02]. Brazilian [EKS09]. bream [AVP^{+}09, ACGR^{+}05, ATNF02a, CIK^{+}04, CRA^{+}05, CVGG06, DW05a, DW05b, KJL07, KJL08, MCS08, MST^{+}04, RKG^{+}00, RAT^{+}02, SBDPAP07, WSKC03, YIT05, ZWZS02]. brine [MRM^{+}00]. British [SKCD07, GPRCH09, JD02, SGS01]. brooders [SAH^{+}07]. broodfish [BPBT02, ME08]. broodstock [MMD05]. broodstocks [JH01]. Broughton [SKCD07]. brown [DGMD^{+}09, DGA^{+}03, GMB^{+}07, GDG00, GDG01, MALC00, PFL^{+}08]. Brünnich [ACGR^{+}05]. Brunswick [HD05a, HD05b, MHD^{+}04]. bryosalmonae [DGF02, LLHF02, MMBA05, MAF^{+}00, MMA02a, MMA02b, MA04, OW02]. bryosalmonae-infected [DGF02]. Bryozoa [MMBA05]. Bryozoans [OW02, LLHF02]. Bt [SSH^{+}07]. bullhead [Wan02]. bullheads [FLHM04]. burden [BLGR09]. burst [BCMN05, DRCA^{+}08]. butyrium [PWS^{+}08].

C [NMJ^{+}04, NUS^{+}06, SBE03, SMS02]. cage [AVN08, CKW^{+}01, GER^{+}07]. cage-level [GER^{+}07]. calcarifer [BB07, BCM^{+}02, CCL09, CN04, FDG^{+}06, GKCF04, HTC^{+}01, LCC^{+}08, PSP^{+}09, PC04]. calcinosis [Dov05]. calcium [GWF^{+}02]. California [CLQAGCRP08, CDI^{+}01]. californiensis [JH01, RJC04]. Caligidae [LB09, PSS09]. Caligus [BEL09, HBN09, ØSKH06, ØH07, RGTR02]. Callinectes [NSK00]. Calretinin [MGB^{+}07]. Cambarus [LMX^{+}05]. campanula [VSST09]. campbellii [SAC^{+}09]. Canada [Bro09, GMB^{+}07, GDR^{+}07, GPRCH09, HD05a, HD05b, SKCD07]. canadum [CKW^{+}01]. canal [BSG^{+}05]. canaliculus [KWKLG09]. cancer [BSTW08, CCB03]. candidate [VLK^{+}04]. Candidatus [MBD^{+}02]. canicula [GDP^{+}03]. canis [BF02a]. capacity [RAG^{+}07]. caprodes [CM04]. Capsalidae [EWCT05]. capsid [CAB^{+}07, HSK^{+}08, HLWH07, KJL07, KL07, KJL08, WCK^{+}09]. captive [BHRB04, JH01]. capture [CWS02, FDG^{+}06, SN00]. Carassius [AHB07, GRL04, IRA03, JBB^{+}07, MGB^{+}07, PB04, SMG08, WYWL01, ZWW^{+}06]. Carbohydrate [VLT^{+}07, MA04]. carbon [GWF^{+}02]. Carcharhinus [TBB^{+}08]. Carcharias [BHRB04, PGB^{+}06]. Cardiac [MIGR09, MALC00, PNA02, KLG^{+}06, PBIM09, SPK^{+}05, YIT05]. Cardicola [CMD01]. cards [SMK^{+}09]. carp [BCM^{+}02, DJW^{+}09, GDR^{+}07, GKVS04, LLZ08, PV08, RM04, STM^{+}01, SSK^{+}08, SHBJ^{+}09, UMIK09, VSNS08, VNSB00, WYWL01, ZWW^{+}06]. carpio [DJW^{+}09, GDR^{+}07, GKVS04, RM04, STM^{+}01, SSK^{+}08, SHBJ^{+}09, UMIK09, VSNS08, VNSB00]. carrier [CLVO^{+}05, GUE06, KM01, MGE04, MGAE06]. carriers [GME04]. cartilaginous [BSTW08]. case [BLGR09, BWCM04, GSMM09, HFI09, PTB07]. cases [SKR^{+}08, SI06, TBB^{+}08]. Caspase [CCL09, HHW05]. caspase-8 [HHW05]. Caspase-dependent [CCL09]. Castelnau [CMD01, DBJN05, HBN09, JDBN08, WMB02]. catalase [DRCA^{+}06]. Cataract [WHB^{+}03, BBBW01, BBC^{+}05]. catchments [CGP08]. Catfish [Rob05b, BNS03, BSW03, CUN09, CDT^{+}02, DM09, DMS09, Goo06, GM07, KS06, NM02, SKLY03, SBG08, SGF02, SGG03, TJG04a, Wan02, XK02, XKS02, XK03, XKS04, XKS05, XKP06, ZKG^{+}01]. catla [ACS^{+}09, KMJM06, ACS^{+}09, KMJM06]. caught [BHB03, BHRB04, SKO04]. causal [And06]. causative [BBLC00, CWJL07, GEC09, HMO^{+}08, KSKEM07, LLJC03, MAF^{+}00, MMA02a, MMA02b, SKR^{+}08, VTA06, WBP^{+}05, ZM01b]. cause [CLS^{+}02, SGEM03]. caused [BSD05, BB01, CIK^{+}04, Dig01, GAH^{+}08, GGCFB09, HWZ^{+}00, HO03, MMB02, MZF01, WYWL01, WCJJ08]. causes [AGSS^{+}01, SLC^{+}06]. causing [FZA06, YIT05]. CB2 [PWS^{+}08]. cDNA [SKR^{+}08]. Cell [FW01, ACS^{+}09, AST^{+}05, ATNF02b, AW02, BMFL^{+}07, BOB^{+}06, BBJ04, CYL^{+}06, CAB^{+}07, CLVO^{+}05, DSJPP08, GWJR08, HMO^{+}08, HHCBR^{+}07, HHW05, HGH^{+}08, HCFM04, HLWH07, IMAN01, KOK^{+}03, LMJ^{+}00, LMC^{+}01, LJL^{+}03, LCC^{+}08, LVW^{+}06, RM04, Shi02]. cell-culture [IMAN01]. cell-mediated [AW02]. cells [AVP^{+}09, BMPB^{+}00, BN04b, CCL09, DGA^{+}03, DG08, HLC00, IRA03, IA03, INM^{+}01, Jay09, MASS07, PWS^{+}08, TFN01]. cellular [BHP^{+}04, CMD02, VNSB00]. cent [LLJ^{+}02]. central [GDHM03, YWS^{+}08]. centres [AR03]. cephalus [CLS^{+}02, DGS^{+}02, GKNLN^{+}04]. Ceratomyxa [BB09, FYO08]. Ceratothoa [HO03]. cercariae [VSST09]. cerebralis [AAH^{+}07, EMS05, KSKEM07, SGEM03]. Cestoda [BSG^{+}05, DADB00, DPS^{+}07, SNNB01]. cestode [BD00]. cestodes [BC06]. chain [APYH00, DCTCB^{+}06, FPO^{+}09, KBH^{+}08, PTKA02, TSL^{+}01]. Challenge [AJ09, AHTM06, BNS03, BSD06, BSW03, BSE02, BLSE03, BMS^{+}00, GM07, GPS^{+}08, IKKS06, KKD^{+}09, LLJ^{+}02, NFR08, RLNB03, RRKM07, RM08, SKO04, SOF^{+}09, VMN06, WZJ^{+}08]. challenged [BCMN05, SDD^{+}00, TBLC07]. Changes [BMPB^{+}00, BSG^{+}05, DGLS09, GPB^{+}05, LGR08, BWD^{+}07, DPS^{+}07, GPS^{+}08, RSD04, SFPF01, TOW02, USR^{+}09]. changing [LBGR08]. Channa [PTKA02]. Channel [Rob05b, SGG03, BNS03, BSW03, CUN09, DM09, DMS09, Goo06, GM07, NM02, SKLY03, SBG08, SGF02, TJG04a, XK02, XKS02, XK03, XKS04, XKS05, XKP06]. char [Sta01]. characterisation [HAEM02]. Characteristics [JAM^{+}01, GWJR08, JDBN08, OBHF00, SHS05, TJG04b]. Characterization [CTB^{+}06, CLL01, EKG^{+}02, JST^{+}04, KSKEM07, LLJ^{+}02, LCC^{+}08, NUS^{+}06, PDOZ08, SWTC01, WWK03, XNS04, ACS^{+}09, BFR07, BAS^{+}09, BSQW05, CFA^{+}00, CPRL09, Cvi04a, DNC^{+}05, EBASW^{+}08, EKS09, FKA^{+}05, HOA09, HCJ08, JTR^{+}04, JTM^{+}01, KOK^{+}03, LMJ^{+}00, LZ09, LHP^{+}03, LWWS04, MHHK^{+}01, MWG^{+}02, NHA07, PWW^{+}06, SME^{+}01, TH08, VAH09, VJR^{+}09, ZSZB04, ZXZ^{+}08]. charr [MPP^{+}03]. chemical [BSM^{+}06, Dig01, EWCT05, GCWR07, SKLY03]. chemistry [DSBB05]. chemotactic [WKES08]. Chemotaxis [VSNS08]. chemotherapeutic [AW01]. Cherax [Edg00, RJ02, SHZ^{+}00]. Chile [BBFB07, BEL09, RHS01, SRAHJ^{+}08, SLC^{+}06, TLC^{+}02, VAH09, VJR^{+}09]. chilensis [HDP^{+}02]. China [CLW03, HWZ^{+}00, PV08, TXZ^{+}09, WHW^{+}02, WWH^{+}02, WGS^{+}07, YYY^{+}06, YWS^{+}08, ZWW^{+}06, ZB07, ZXZ^{+}08]. Chinese [PWS^{+}08, ZSZB04]. Chinook [FF07, JH01, JD02, KM01, KH06, LWWS04, LCH^{+}02, LMG^{+}03, LM03, LRH^{+}08, OR05, OB07, POHJ05, TPF09]. chloride [AW01]. Chloromyxum [AHB07]. chlororaphis [GGCFB09]. cholangiocarcinoma [BHB03]. cholesterol [MIGR09]. Chondroitin [STV06, SRYC02]. Chondrostoma [IPA^{+}01]. chronic [BBG^{+}05]. chrysops [BWCM04, MOG09]. chuatsi [HWZ^{+}00, SXSN07]. chungmuensis [ITK^{+}04]. ciliate [CPB^{+}07, SLB04]. ciliates [SHM00]. Ciliophora [BCB02, CPB^{+}07, Kha04a]. Cirrhinus [SMS02]. clade [KRKS06]. Clair [EFT^{+}06]. clam [DGMD^{+}09]. Clarias [KS06]. clarkii [LMX^{+}05, SWZ^{+}05]. class [XCD09]. clearance [BSW03]. Clinical [MNM^{+}02, MBD^{+}02, GSM^{+}04, PGB^{+}06, RJC04, SHRM07]. clinically [BMD^{+}06, WGM^{+}08]. clonality [NLO08]. Cloning [CWJL07, CFA^{+}00, HKV^{+}06, NHA07, NHA05, YTLN01]. Clostridium [PWS^{+}08]. ClpXP [HKV^{+}06]. Clupea [HGP^{+}07, JD02]. co [LVW^{+}06]. co-culture [LVW^{+}06]. coast [DCTCB^{+}06, GPRCH09, PMZD08, SN00, WHD^{+}08]. coastal [BC06, JO00, SYYO06]. coat [HTC^{+}01, YKR^{+}02]. cobalamin [BAS^{+}09]. cobalamin-binding [BAS^{+}09]. cobia [CKW^{+}01]. Coccidia [GA05]. coccidiosis [ACP01]. Cochennec [HDP^{+}02]. Cochennec-Laureau [HDP^{+}02]. cod [BBG^{+}05, BFR07, BTSH03, FCM^{+}04, GUE06, GKOF09, GBGB06, Kha04a, KKD^{+}09, MBG^{+}02, MGL^{+}06, MASS07, MOR^{+}09, ØSKH06, OMR^{+}06, ONI^{+}08, SBE03, SBM^{+}06, SWS07, SKGR05, UBB^{+}09]. coding [HLWH07]. codons [PV08]. cohabited [XK03]. Coho [CDL^{+}08, SLC^{+}06]. cold [RDT^{+}03, SC06]. coldwater [LLJC03]. Colisa [IGH01]. collected [AVN08, BC06, SC06]. Colonization [VTA06, BNN03]. Columbia [GPRCH09, SKCD07, JD02]. column [DHOM^{+}03, PD09]. columnare [AG01, BNS03, DM09, DMS09, FKA^{+}05, GLC04, LNZL08, MMB02, SC06, SKLY03, SN08, SRYC02, STV05, STV06, TJG04b, XNS04, ZASK06]. columnaris [MMB02, SBV09, TJG04a]. combined [IMAN01, ZWW^{+}06]. commercial [BBA08, GSMM09, GLMG03, MOG09, SSH^{+}07, SRT^{+}07]. commercially [CDL^{+}08, KWD^{+}01, SHRM07]. commercially-important [KWD^{+}01]. common [DJW^{+}09, GDR^{+}07, GKVS04, MCS08, SHBJ^{+}09, UMIK09, VSNS08, VNSB00]. Comparative [BNS03, CAC^{+}08, DBR^{+}05, MD00, RMG^{+}09, ASEK03, GBGB06]. compared [LSDB06, SSH^{+}07]. Comparison [BDSY07, BMŠH^{+}02, BT09, DJW^{+}09, SB01, ZASK06, BEL09, GLMG03, HRG03, ME08, SPL05, SNN^{+}08, SBB^{+}06, TPF09, XLZS08]. compendium [Edg00]. Competition [VLK^{+}04]. competitive [TSL^{+}01]. complement [PB04]. Complete [SMR01, SMS02]. complex [CLN05, LRH^{+}08, NM02]. complications [CHWK^{+}04]. components [ABC^{+}08, BHP^{+}04]. composition [SPK^{+}05, SKLY03]. compromises [MDB^{+}09]. concentration [HKK^{+}09a, YCR^{+}09]. concentrations [PSB06]. concept [Jön04]. concurrently [SSL^{+}08b]. condition [And06, CFA^{+}09, SRT^{+}07]. conditions [DRCMB03, KBEJ^{+}07, LAZ05, SBDPAP07, STV05]. conferred [YMN09]. confers [MBB^{+}06]. Confirmation [ZKG^{+}01, GDR^{+}07, GRF^{+}07]. confirmed [WCT^{+}01]. Congo [CK08]. conjugate [AW01]. conjugated [AST^{+}05]. Conjugation [AW01]. conservation [BKNS02]. Considerations [Bro09]. consumption [HFW01]. contain [MGAE06]. containing [GAMR^{+}04, PMZD08]. contamination [LKH08]. contemporary [GER^{+}06]. Continuous [BM09]. contrasting [SRT^{+}07]. Control [RJC04, And06, BS04, Bro09, BA05, BNFA07, Dig01, GER^{+}06, IA02b, IA03, JH01, NA09a, NA09b, SMM09, SBSR02]. controlling [BDSY07]. controls [IRA03]. convenient [CK08]. conversion [RSD04]. copepod [CAC^{+}08]. Copepoda [KPWK06, BB01, LB09, PSS09, RJC04]. copepodids [PD09, TSW00]. copepods [KPWK06]. copper [GM07]. coral [BB01]. coralliilyticus [KWKLG09]. coralliilyticus/neptunius [KWKLG09]. coralliilyticus/neptunius-like [KWKLG09]. Coregonidae [DPS^{+}07]. Coregonus [BWD^{+}07, DPS^{+}07, SKO04]. coronary [SPL05, SSL^{+}08a]. correlated [LB00]. correlations [TJG04b]. corruscans [EKS09]. cortisol [BSW03, VNSB00]. cost [Cos09]. Costa [SHFM09]. cotton [RJB00]. Cottus [FLHM04]. counting [Cvi04b]. coverage [KaVT06]. CpG [BBN03]. CpSBV [CCB03]. crab [CCB03, FHY^{+}08, NSK00, WGS^{+}07, ZSZB04, ZB07]. Crassostrea [ITK^{+}04, RLA01, WHD^{+}08, WP00]. crassum [BSG^{+}05]. crassus [GKB05]. crayfish [Edg00, LMX^{+}05, RJ02, SHZ^{+}00, SWZ^{+}05]. Cribb [CMD01]. croaker [CLW03, WYJ05]. crocea [CLW03, WYJ05]. Cromileptes [BTSH03, YKR^{+}02]. cross [BGBG05, LKH08, PTKA02, XKP06]. cross-contamination [LKH08]. cross-protection [XKP06]. crucian [WYWL01, ZWW^{+}06]. cruciata [GA05]. crude [OBHF00]. cruentum [DRCA^{+}08]. Crustacea [KPWK06, TSW06]. crustacean [HHCBR^{+}07, SWS07]. crustaceans [CZS^{+}01, YMNH04]. Cryptobia [AW01, AW02, CGW04, FW01, HFW01, Woo03, Woo04, ZW00]. cryptobiosis [Jön04, Woo03, Woo04]. Ctenopharyngodon [LLZ08]. Cultural [GWJR08]. Culture [CPB^{+}07, Rob05b, CLVO^{+}05, DRCMB03, IMAN01, LVW^{+}06, MGE04, SME^{+}01, WMA^{+}02]. cultured [AN03, AN04b, BM04, CFA^{+}09, CL02, CKW^{+}01, CDT^{+}02, CDI^{+}01, DCTCB^{+}06, DAMQ^{+}04, EKG^{+}02, FZA06, GSM^{+}04, GCN04, HWZ^{+}00, INKS04, Kha04a, KHA04b, Koz07, LVA^{+}06, MGGSREAG08, MKW^{+}09, MSD^{+}00, NDB^{+}04, PC04, RBA^{+}03, RJ02, SLC^{+}06, SHRM07, TKNM03, TMM01, TLC^{+}02, TOW02, WSKC03, WCSC07, WCW^{+}08, WHW^{+}02, WWH^{+}02, WGS^{+}07, ZB07]. cultures [AST^{+}05, LKH08]. cuneata [LNM^{+}09]. current [MO06, SRM^{+}03, WKB^{+}02]. Cutaneous [XKS02, XKS05, LLZ08, XK03]. cutlass [LSM^{+}02]. Cyathocephalus [DADB00]. cycle [AHB07, BB09, DSJPP08, KEEM^{+}05, LWWS04, SMR01]. CyHV [GSMM09, SMG08]. CyHV-2 [GSMM09]. CyHV-3 [SMG08]. Cymothoidae [HO03]. CYP1A [BWD^{+}07]. cyprinid [JBB^{+}07, MES^{+}02, UMIK09]. Cyprinidae [LFF05]. cyprinids [LFF05]. Cyprinus [DJW^{+}09, GDR^{+}07, GKVS04, RM04, STM^{+}01, SSK^{+}08, SHBJ^{+}09, UMIK09, VSNS08, VNSB00]. cyst [GDG00]. Cytotoxic [TFN01]. Czech [KBPD04].

D [GWF^{+}02]. Daintith [CMD01]. damage [KLG^{+}06]. damsela [BPV^{+}03, BVG^{+}03, BHP^{+}04, MMAP00]. damselae [AVP^{+}09, AOA^{+}09, ARC^{+}05, BAS^{+}09, CRBM05, DRCMB03, DRCA^{+}06, JTM^{+}01, KHA05, LVA^{+}06, MMTPDR^{+}08, NHA05, NHA07, PSLN^{+}09, RP04, TO02, LVA^{+}06, PSLN^{+}09]. danilewskyi [PB04]. Danio [CHWK^{+}04, HBK07, KBS03, RWSK09, TL09]. Danish [DM00, JHKB08, MD00, PSLN^{+}09]. data [MG09]. database [JGK^{+}09]. Davis [ABCRH09]. Day [EKG^{+}02]. DCH [RKG^{+}00]. dead [IA03, PWS^{+}08]. death [HGH^{+}08]. Decapoda [NSK00]. defective [HHCBR^{+}07]. deficiency [BB07, MST^{+}04, SHR04]. deficient [PC04]. deformation [SHRM07]. deformity [AGM^{+}08, PGB^{+}06, SGRM07, SRT^{+}07]. degeneration [RJB00]. degranulation [DG08]. dehydrogenase [DW05b]. deliver [OLR03]. Delivering [TL09]. delivery [SGMP^{+}05]. Delta [CDT^{+}02, FTS^{+}01]. dendriticum [DPS^{+}07]. Denmark [PSLN^{+}08]. densities [BBJ04]. density [BWMSB02, BLSE03]. dentex [MOO^{+}09, MKW^{+}09]. depend [BLSE03]. dependent [CCL09, RSSD01, HKV^{+}06]. depigmentation [CUN09]. deprivation [SKLY03]. derived [ACS^{+}09, AMI^{+}07, HMO^{+}08, KOK^{+}03, LVW^{+}06, PZS02]. derjavini [BNN03, OLB03]. dermatopathy [BBG^{+}05]. Dermocystidium [FLHM04]. describing [OB07]. Description [FYO08, BD00, BT09, FZA06, IPA^{+}01, SRAHJ^{+}08]. Descriptions [LFF05, PS02]. desiccation [EWCT05]. destructive [GME04, MGE04, ME08]. detect [TPF09, YDNM08]. detected [ILAW03, JDBN08, KRKS06]. Detecting [PV08, KBEJ^{+}07, MGE04]. Detection [CLQAGCRP08, CGMS03, CDGF07, FDG^{+}06, FPO^{+}09, GML^{+}02, GMW^{+}07, GKVS04, HKKO07, HTC^{+}01, IFA05, JTR^{+}04, JO09, KHA05, LLZ08, MMAP00, MBLK06, PTAR09, POHJ05, SMG08, Shi02, SSK^{+}08, XMM^{+}09, AVN08, AL09, CFA^{+}09, CABF08, CBDC^{+}03, CL05, CLVO^{+}05, DCTCB^{+}06, DADB00, DHCHN01, DLS^{+}07, FNE09, GME04, GEC09, GDR^{+}07, GSM^{+}04, GOS04, HKK^{+}09a, HWT^{+}06, IMAN01, IYS^{+}07, JLSA08, KMS^{+}08, KBH^{+}08, úMMUB08, MMA02a, MN08, OLP01, PGH^{+}07, PBW06, RA01, RB03, SKIS05, SB01, SRM^{+}03, SMM09, TXZ^{+}09, VCLS04, WDC^{+}03, WMA^{+}02, YSH^{+}06, YMNH04, YWBH05]. detections [AAH^{+}07]. Determination [OMLF^{+}06]. determine [CFA^{+}09]. Deuteromycotic [BTSH03]. developing [Bro09]. Development [ACS^{+}09, AOA^{+}09, AL09, BSE02, CPRL09, CAAN07, CDS^{+}00, Cvi04b, DZP^{+}00, DHCHN01, EMS05, ESMB00, FF07, FNE09, HKK^{+}09a, LLHF02, MMA02a, PGH^{+}07, SSL^{+}08a, SMM09, TPF09, YDNM08, And06, BSD06, BBBW01, BBC^{+}05, BNFA07, DRA00, FBP09, JKB08, KHK^{+}04, LFF05, MEC^{+}06, YSN^{+}06]. developmental [KSKEM07, SMR01]. dhufish [SCJ^{+}01, SCJ^{+}02]. diagnose [CN02]. Diagnosis [JLSA08, BMŠH^{+}02, EMS05]. diagnostic [CK08, DZP^{+}00, GEB^{+}08, MEC^{+}06, MO06, NDH^{+}05b, NDH05a, SRM^{+}03, YDNM08]. diatoms [IFA05]. dicentrarchi [CPB^{+}07, SPC^{+}08]. Dicentrarchus [ABPK03, BPV^{+}03, BVG^{+}03, BPBT02, CZP01, HO03, RTNA02, RNAA04, RP04, SBDPAP07]. Dicologoglossa [LNM^{+}09]. diet [GPS^{+}08]. Dietary [KS06, DRCA^{+}08, GWF^{+}02, MST^{+}04, NA09a, RP04, RDT^{+}03, SPK^{+}05, SMS02, WHB^{+}03]. diets [DRA00, GAMR^{+}04, Goo06, SRT^{+}07]. Differences [KH06, SKCD07, SRYC02, SSO^{+}09]. different [ATNF02a, ALAC09, DRCMB03, HMB07, KRJO02, MHD^{+}04, PKK09, SSH^{+}07, SBK^{+}05, SSL^{+}08a, SKGR05, SRT^{+}07, USR^{+}09, VSNS08, XKP06, ZASK06]. Differential [SBG08, BPV^{+}03]. Differentiating [RLL^{+}07]. Differentiation [KJB07, PGH^{+}07, RMSH09]. diffuse [BSG^{+}05, DGA^{+}03]. diffusion [TO02]. Digenea [CMD01]. digestibility [RKG^{+}00]. digestive [BMKG^{+}07, KHK^{+}04, LYO04]. dilation [FF07, LCH^{+}02, LMG^{+}03]. dilution [TO02]. dioxide [GWF^{+}02]. diphyllobothriosis [TLC^{+}02]. Diphyllobothrium [DPS^{+}07]. diplomonad [PS02]. Diplostomum [PSTV06, VSST09]. diquat [DM09]. direct [Cvi04b]. disc [TO02]. discovered [LWWS04]. discrimination [SGS01]. Disease [Hum07, Kha04a, AN01, AN03, AN04a, AEN04, AN04b, AVN08, AGSS^{+}01, ACGR^{+}05, BSDD02, BSD05, BM04, BM06, BSQW05, BML^{+}06, BN04a, BBN03, BCMN05, BWD^{+}07, BH02, CFA^{+}09, CLS^{+}02, CLW03, CMD02, DGMD^{+}09, DL01, EMH02, FPG^{+}02, FKT^{+}05, FZM00, FBP09, FZA06, FTE^{+}03, GEC09, GKCF04, GKNLN^{+}04, GCR^{+}02, GRW^{+}03, GCG^{+}04, HWZ^{+}00, HSK^{+}08, HKK^{+}09b, HCFM04, JLSA08, KSKEM07, KHA04b, KL07, KHSW09, KKT^{+}04, KS06, LLJC03, LHHP07, LLHF02, MD08, MFLC01, MNM^{+}02, MMB02, MAF^{+}00, MMA02a, MMA02b, MCN04, MN08, MZF01, NSK00, NFR08, OMR^{+}06, PWS^{+}08, PNS^{+}05, PBW06, PNA02, QSZ^{+}03, RJB00, RHS01, RP03, RM07, SGEM03, SKR^{+}08, SBV09, TOB^{+}07, TTD^{+}06, VMN06, WKB^{+}02, WBP^{+}05, WCSC07, WCW^{+}08, WHW^{+}02, WGS^{+}07, YDNM08, ZB07, ZM00, ZM01b, ZGM01]. disease-affected [MN08]. diseased [CDT^{+}02, GVGR00, KKD^{+}09, Koz07, SBE03, ZXZ^{+}08, ZAC^{+}03]. Diseases [MSCH03, BNFA07, KSW03, RJ02]. disinfectants [GCWR07]. dismutase [DRCA^{+}06]. disorders [MALC00]. disparate [Jön04]. dispersion [AM09]. display [ZWW^{+}06]. disruption [DW05a]. dissecting [PT09]. dissolved [GWF^{+}02]. Dissonidae [BB01]. Dissonus [BB01]. Distended [RSFS00]. distinct [GDG00, GDG01, ZS04]. Distribution [AN01, BMD^{+}06, INKS04, NLO08, UMIK09, WGM^{+}08, AAH^{+}07, DGA^{+}03, DHOM^{+}03, DHNB05, GER^{+}07, KBS03, PMZD08, SSM00, SSM01, Wan02]. divalent [ARC^{+}05]. diversicolor [CCTL06]. diversity [CDL^{+}08, SC06]. DNA [CWW^{+}02, GML^{+}02, HKKO07, PLC09, PWW^{+}06, RA01, SSL^{+}08b, SMK^{+}09, VCLS04, WCT^{+}01]. do [BLSE03]. Does [RAG^{+}07, WM04]. dogfish [BF02a, BF02b, CKK02, GDP^{+}03]. dogieli [MCS08]. domain [KL07]. Dominant [Koz07]. dose [BSDD02, SBE03, UMG^{+}08]. dot [DCTCB^{+}06, DHCHN01, GOS04, WDC^{+}03]. dot-blot [DCTCB^{+}06, GOS04, WDC^{+}03]. Dr [Aok08]. dragons [GAH^{+}08]. dramatically [CN02]. Drasche [HOA09]. dried [CRM^{+}03]. Drug [MHKA08]. drum [PWS^{+}08, WWH^{+}02, WCB08]. ductal [BF02b]. due [CHWK^{+}04, DBR04]. dumerili [WCT^{+}01]. Dunbar [LWWS04]. duodeno [BD00]. duodeno-spiral [BD00]. duration [SSS^{+}00]. during [ATT^{+}09, BWMSB02, BM09, CM04, CVGG06, DW05a, DW05b, DGLS09, GCR^{+}02, MSD^{+}00, PSB06, PSTV06, RDT^{+}03, RAG^{+}07, SHN^{+}00]. Dynamics [LYW03, GMS^{+}09, HDP^{+}02, OB07, SYYO06, Wan02]. dysgalactiae [NMJ^{+}04, NUS^{+}06].

E. [ASEK03]. E2 [FRW^{+}08]. Early [JKB08, BM09, DSJPP08, SWS07, YNH03]. East [LSDB06, WHD^{+}08]. eastern [AVN08, CWW^{+}02, GMB^{+}07, KWD^{+}01, RA08]. economic [Cos09]. ectasia [BF02b]. ectoparasite [CN02, VNSB00]. ectoparasites [CUHN08, MOO^{+}09]. Ecuador [RBA^{+}03, RJ02]. edible [CCB03, HOA09]. Edited [Ano02b]. Edition [Ano02b]. Editorial [Ano01c, Ano01d, Ano07a, RLB09, RW05a, RW05b, RW05c, RW06, Rob07, RW08]. Edwards [BWS^{+}08, DSBB05, Dov05, Jay09, ZB07]. Edwardsiella [ASEK03, BSW03, CTB^{+}06, CDT^{+}02, GM07, HKL^{+}06, NM02, PZD^{+}06, WKES08, XCD09]. edwardsiellosis [CL02]. edwardsii [Dig01, HSD07]. eel [CL02, JNS^{+}09, LG03, NEG06]. eels [GKB05, Ken07]. Effect [BWMSB02, DRCA^{+}08, FHY^{+}08, GAMR^{+}04, ONS07, RP04, SNNB01, SBDPAP07, SMS02, VCN09, BHP^{+}04, BCM^{+}02, BCMN05, BBK04, BNN03, DHNB05, GPS^{+}08, KMJM06, PDOZ08, SOF^{+}09, TSW00, XK03, ZM00, ZM01b]. Effectiveness [ARC^{+}05, OLR03]. effector [BF07]. Effects [AG01, BSDD02, BBJ04, CHWK^{+}04, EWCT05, KHSW09, MPEN05, VNSB00, BCB02, BB09, BGB^{+}04, CCTL06, FM00, GWF^{+}02, HDP^{+}02, HFW01, RDT^{+}03, STM^{+}01, SSM01, SPC^{+}08, YSN^{+}06]. Efficacy [GER^{+}06, JH01, MLB^{+}06, SGMP^{+}05, LBGR08, PEP^{+}05, RP03, SSS^{+}00, TJG04a, VSST09, WYWL01]. efficiency [BMŠH^{+}02]. effort [BGB^{+}04]. egg [EWCT05, ØGW04, TWS09]. eggs [BPBT02, MMD05, VTA06]. Egusa [Aok08]. eight [PSLN^{+}08]. EILO [ASEK03]. either [OL06]. elastolytic [CFA^{+}00]. electrolyte [PSB06]. electron [CZP01, MIS07, MA04]. electrophoresis [AMI^{+}07]. elevated [ØGW04]. elicits [NM02]. ELISA [CWS02, FDG^{+}06, HTC^{+}01, RB03, Shi02]. elongatus [ØSKH06, ØH07, RGTR02]. emamectin [GER^{+}06, LBGR08, SSS^{+}00]. embedded [CDGF07, KMS^{+}08, PTAR09]. Embryonal [IGH01]. embryonation [EWCT05]. Emergence [CDT^{+}07]. emerging [FLA06, PSLN^{+}09, SKIS05, SHFM09]. empA [XMM^{+}09]. emphasis [RBA^{+}03]. encephalitis [GKOF09]. encephalopathy [ATNF02a, ABPK03, JST^{+}04, JGS^{+}04]. encoding [CWJL07, KL07]. endemic [LBO^{+}01, WCJJ08]. endocrine [BSG^{+}05, DGA^{+}03]. endothelial [MASS07]. energy [RAG^{+}07]. England [CGP08, FPG^{+}02, FLHM04, PFL^{+}08]. Engulfed [HHL^{+}05]. Enhanced [CYL^{+}06, MSD^{+}00, RLNB03, LSDB06, PWS^{+}08, PSTV06]. Enhancement [FLGS00]. enhances [RP03]. enriched [CRM^{+}03]. enrofloxacin [FHY^{+}08]. enteric [DGA^{+}03]. enteritic [MBD^{+}02]. enteritis [BMPB^{+}00, KHK^{+}04]. Enterococcus [CL02]. enterocyte [MDB^{+}09]. Enteromyxum [SBDPAP07, YSN^{+}06]. entry [HSW03]. envelope [LMX^{+}05]. environment [BMD^{+}06, GMW^{+}07, MMD05, MMA02b]. Environmental [BM04, LAZ05, AM09, DHNB05, HBK07]. enzymatic [GHBW03, SMK^{+}09]. Enzyme [PB04, BMPB^{+}00, NHA07, RB03, SB01]. enzyme-linked [SB01]. epidemic [OR05, OB07]. epidemics [BWMSB02]. Epidemiological [RM07, HRG03, LGR08]. Epidemiology [ARKK08, FRW^{+}08, RGTR02, SCJ^{+}01, TSW06]. epidermal [HKK^{+}09b, KaVT06, KaVKT09, ONS07, TH08]. epidermidis [HLC00]. epidermis [VNSB00]. Epinephelus [CWT^{+}00, GKNLN^{+}04, LMJ^{+}00, LMC^{+}01, LCM^{+}01, LJL^{+}03, MWG^{+}02, TMA^{+}01, TTM04, YMKN09, YMN09]. epithelial [BBJ04, BN04b]. Epithelio [RM04]. Epithelio-reticular [RM04]. Epitheliocystis [CZP01, NL06, GPN05]. epitopes [CAB^{+}07, VLT^{+}07]. epizootic [AJ09, WSKC03, WCJJ08]. epizootics [HBN09]. Epizootiology [SHBB^{+}02]. Ergens [YWS^{+}08]. Eriocheir [ZSZB04, ZB07]. erosions [ONS07]. erosive [BBG^{+}05]. Erratum [Ano03b, Ano03c, Ano07b, Ano08a, Ano08b, Ano08c, Gar05, TCC10]. erythopterus [CWJL07]. Erythro [KM01]. Erythro-200 [KM01]. erythrocytes [GCR^{+}02, ZW00]. Erythrocytic [Rod07]. erythromycin [CR02, CRM^{+}03, KM01, MRM^{+}00]. erythrophthalmi [MEM^{+}09]. erythrophthalmus [MEM^{+}09]. Esox [EFT^{+}06, GCG^{+}04]. essential [MST^{+}04]. established [LMJ^{+}00, LCC^{+}08]. establishes [GUE06]. Establishment [KOK^{+}03, LJL^{+}03]. ester [SBE03]. estimate [Cvi04b]. estimated [KaVT06]. Estimating [GEB^{+}08]. Estimation [NDH^{+}05b, NDH05a, UMG^{+}08]. estuarine [JZKS04]. estuary [LGF04]. Eubothrium [BSG^{+}05, SNNB01]. euglenoid [ZRT04]. Euphrasen [DAMQ^{+}04]. Europe [BFR07, RA01]. European [AJ09, CL02, CZS^{+}01, FLA06, FRW^{+}08, GKB05, JNS^{+}09, MPP^{+}03, PGH^{+}07, PSTV06, SBDPAP07]. euryhaline [AG01]. Eustrongylides [MOG09]. Evaluation [CVGG06, DM09, DMS09, DLS^{+}07, FZM00, GOS04, OR05, PTT^{+}08, SSH^{+}07, XSK07, BGBG05, CL02, FNE09, GCWFB08, GER^{+}06, SGMP^{+}05, SRM^{+}03, SBSR02, TO02, ZAC^{+}03]. Evidence [DW05a, PSTV06, SGF02, AN04a, BCMN05, Cvi04a, DG08, MMA02b, NLO08]. evolving [KRKS06]. ex [MGAE06]. exacerbates [RWSK09]. examination [SBK^{+}05]. excised [LLZ08, XK02, XKS02]. excretory [Dov05]. exhibit [GKB05]. exhibiting [ZB07]. exitiosus [HDP^{+}02]. exogenous [DRCMB03]. Exophiala [MKW^{+}09]. exophthalmos [SCJ^{+}01]. exotoxins [GHBW03]. Expanded [AAH^{+}07]. Experimental [AN04a, ATNF02a, AW02, BGB^{+}04, BGBG05, BML^{+}06, BLSE03, BPBT02, CZS^{+}01, GDG01, HBK07, IKKS06, JKD06, KRJO02, LVDBB07, LMG^{+}03, SBM^{+}00, SKGR05, SAH^{+}07, UBB^{+}09, WBNG02, BMS^{+}00, CAC^{+}08, CLQAGCRP08, CN04, DM09, DMS09, EN03, GMS^{+}09, GPB^{+}05, GBGB06, KT09, LYO04, LYW03, OR05, OB07, RRKM07, SSM00, SHZ^{+}00, SBDPAP07, SKO04, SRM^{+}03, TSL^{+}01, TMA^{+}01, ZM00]. experimentally [MBG^{+}02, RSD04, RSFS00, SBM^{+}06, SGG03, VFAN03, YSN^{+}06, YNH03]. experiments [EMH02, JZKS04]. explant [LKH08]. explants [LLZ08]. exposed [GPB^{+}05, SMG08, SHBJ^{+}09]. exposure [BM09, EMH02, GM07, HBK07, LLHF02, MBB^{+}06, PDOZ08, RSSD01, ZGM01]. express [OLR03]. expressing [CYL^{+}06]. Expression [ALAC09, BAS^{+}09, KJL07, XLZS08, BPV^{+}03, BSTW08, CWJL07, DW05b, HLWH07, LSDB06, MGB^{+}07, OL06, SSM00, SMLZ05, SLB04, SDD^{+}00, SGG03, XKS06]. extirpation [EWBB02]. Extra [QSZ^{+}03, PBW06, SAH^{+}07, WCSC07, YWBH05]. extracellular [BHP^{+}04, CWJL07, CFA^{+}00, CTAR01, FAM^{+}06, OBHF00, PEP^{+}05, VCN09, WKES08, ZAC^{+}03]. extraction [GTBA09]. extractive [CHWK^{+}04]. extracts [BSM^{+}06].

Fabricius [KHA04b, MFLC01, TSL^{+}01, YNH03]. factor [BMKG^{+}07, CTAR01, DW05b, HGH^{+}08]. Factors [KaVT06, KaVKT09, LBGR08, And06, BM04, DHNB05, GER^{+}07, HD05b, JSG^{+}08, MPP^{+}03, OL06, PRG^{+}05, SBDPAP07, SGRM07, TWS09, WMB02]. faecium [CL02]. false [MMA02a]. families [CGW04, DJW^{+}09]. family [SXSN07]. farm [HFI09, YYY^{+}06]. farmed [AEN04, And06, ABPK03, BM06, BSG^{+}05, BB07, Bro09, CRBM05, CRA^{+}05, CGP08, CBDC^{+}03, CVGG06, CDT^{+}07, DCCF^{+}03, DBJN05, EWCT05, FTS^{+}01, FCM^{+}04, FKT^{+}05, GKOF09, HRG03, JTR^{+}04, JST^{+}04, KSW03, áKBH06, LGR08, LHW08, LVDBB07, LCH^{+}02, LM03, MMSM06, MHD^{+}04, MNM^{+}02, MALC00, MBD^{+}02, MG09, NUS^{+}06, ØH07, OMR^{+}06, ONI^{+}08, PNS^{+}05, PTB07, PBIM09, PT09, RRG^{+}05, RA08, RM07, RSFS00, RDT^{+}03, RMSH09, SPK^{+}05, SIM^{+}00, SKR^{+}08, SHM00, SFT^{+}08, TOB^{+}07, TTD^{+}06, TBLC07, VHM^{+}00, WBP^{+}05, YWS^{+}08, ZAC^{+}03]. farming [Cos09, HBN09, SOM05]. farms [BLGR09, DM00, GSMM09, GMW^{+}07, GER^{+}07, HOM^{+}09, LBGR08, MHHK^{+}01, PSLN^{+}08, PMZD08, PD09, SKCD07]. Fas [XKS06]. fasciata [IGH01]. fasciatus [JO00, KJL07, KJL08]. fathead [RFBD^{+}05]. fatty [ASEK03, MST^{+}04, RDT^{+}03]. features [BT09, Dov05, MNM^{+}02]. fecundity [BEL09]. fed [BMKG^{+}07, CRM^{+}03, Goo06, SSH^{+}07, SBK^{+}05, SSL^{+}08a]. Feed [SKLY03, RSD04]. Feeding [SAC^{+}09, ATT^{+}09, BSD05, CIK^{+}04, CRM^{+}03, HSD07, HO03, TL09]. female [MOG09]. fennica [VSST09]. feral [KSW03]. ferrugineus [PP02].

Fibrosarcoma [BAMG^{+}05]. fibrosis [BF02b]. Field [RRKM07, AMI^{+}07, EMH02, KBEJ^{+}07, PTB07]. fighting [PTKA02]. filamentous [MBD^{+}02]. filled [RSFS00]. fin [Goo06]. finding [LB09]. findings [CMD01, JGS^{+}04]. fingerlings [Goo06]. Finnish [MHHK^{+}01]. First [EKS09, FZA06, GDR^{+}07, GVGR00, GRW^{+}03, GCG^{+}04, GRF^{+}07, LVA^{+}06, LNM^{+}09, SRAHJ^{+}08, SKR^{+}08, AAH^{+}07, CRM^{+}03, CDI^{+}01, GKNLN^{+}04, SFT^{+}08, TL09]. first-feeding [CRM^{+}03, TL09]. Fish [BLL^{+}07, Hum07, AR03, ABR^{+}09, AG01, AOA^{+}09, AMI^{+}07, BOB^{+}06, BWMSB02, BMD^{+}06, BHO05, BSTW08, BBA08, CIK^{+}04, CPB^{+}07, CGMS03, CRBM05, CGP08, CUHN08, CDS^{+}00, CDGF07, CDT^{+}07, Cvi04b, DZP^{+}00, DSJPP08, DG08, DHCHN01, DHDCN02, EMS05, FH03, GSM^{+}04, GOS04, GML^{+}02, HHCBR^{+}07, HSK^{+}08, INKS04, INM^{+}01, JSBK09, JZKS04, KRKS06, KJC^{+}02, KJL08, KBEJ^{+}07, Koz07, LZ09, LHP^{+}03, LSM^{+}02, LYW03, LVDBB07, MHHK^{+}01, MHKA08, úMMUB08, MOG09, MES^{+}02, MAF^{+}00, MZF01, MKM02, NDPH03, NMJ^{+}04, NUS^{+}06, NDB^{+}04, NL06, OMLF^{+}06, OKK^{+}06, ØH07, OFKEA08, PD01, PTAR09, PS02, PTKA02, RBBC^{+}09, SKIS05, SC06, SNN^{+}08, SOM05, SGEM03, SME^{+}01, SI06, SHR04, SXSN07, STV06, TKNM03, TJG04b, TF04, TDH^{+}07, VCLS04, VLK^{+}04, WGM^{+}08, YYY^{+}06, ZFF^{+}08, ZXZ^{+}08, ZM00, ZGM01, Rob00, Rob05a]. Fish [Ano02b]. fish-pathogenic [LHP^{+}03, MHKA08]. fisheries [TSW06, TWS09, Pen04]. Fishes [Rob06]. fishing [BKNS02]. Fishnet [BH02]. FishPathogens.eu [JGK^{+}09]. FishPathogens.eu/ [JGK^{+}09]. Fitness [PKK09]. five [CGW04, CR02]. fixed [CDGF07, KMS^{+}08, MMAP00, PTAR09]. fjord [HOM^{+}09]. flagellates [PS02]. flat [HDP^{+}02]. flavescens [BAMG^{+}05]. Flavobacterium [AG01, AMI^{+}07, BNS03, CDL^{+}08, CDGF07, CK08, DM09, DMS09, DDL^{+}01, EN03, EEF06, FKA^{+}05, FNE09, GLC04, HPMW08, ILAW03, IFA05, IOK^{+}07, LLJ^{+}02, LLJC03, LNZL08, LHW08, MHHK^{+}01, MLB^{+}06, MD00, MMD05, MD08, MMB02, MEBD05, NDPH03, OBHF00, OL06, PLC09, RLL^{+}07, SC06, SKLY03, SN08, SRYC02, STV05, STV06, TJG04b, VAH09, VTA06, XNS04, ZASK06]. flesh [DCCF^{+}03]. Flexibacter [CGMS03]. flora [GCWFB08]. florfenicol [SBM^{+}06]. flounder [BCB02, HKL^{+}06, HKK^{+}09b, HCFM04, KOK^{+}03, KHK^{+}04, KBH^{+}08, SYYO06]. flow [AL09, SOF^{+}09]. fluids [SM08]. fluke [CMD01]. flumequine [RTNA02]. fluorescent [CYL^{+}06]. fluviatilis [AJ09, GGCFB09, GCWFB08, KPWK06, LHW08, WBP^{+}05]. fluxes [PSB06]. foliaceus [TWS09]. following [BSW03, FHY^{+}08, GLC04, GM07, GBGB06, KM01, LLJ^{+}02, LLHF02, RLNB03, SSM00, SSL^{+}08b, SRM^{+}03, SSS^{+}00, SLC08, WZJ^{+}08, YCR^{+}09]. forcing [AM09]. form [KJB07, ZS04]. formalin [BBJ04, CDGF07, IRA03, KMS^{+}08, PTAR09]. formalin-fixed [CDGF07, KMS^{+}08, PTAR09]. formalin-inactivated [IRA03]. formation [WHB^{+}03]. forms [CR02, Cvi04a, MMB02]. formulations [BPV^{+}03]. Forsskål [DW05a, DW05b, FHY^{+}08, GKNLN^{+}04, WCT^{+}09]. Forster [GDHM03]. forsteri [CMD01]. Fouquet [SB01, XK03]. four [BDSY07, BEL09, JZKS04]. Fowler [PTKA02]. fraction [HPMW08, LNZL08]. France [DGMD^{+}09, EWBB02, RLA01]. franciscana [MRM^{+}00, SAC^{+}09]. Francisella [BBFB07, MOR^{+}09, OMR^{+}06, ONI^{+}08, SHFM09]. Fraser [SHBB^{+}02]. Fredericella [MMBA05]. free [BF02a, BSTW08, BT09, BKSB09]. free-living [BF02a]. free-ranging [BT09, BKSB09]. freeze [CRM^{+}03]. freeze-dried [CRM^{+}03]. frequency [SGRM07]. fresh [ABPK03, BCB^{+}09, RP03, SOF^{+}09]. freshwater [AN04b, ABR^{+}09, AG01, BSM^{+}06, BSQW05, CDT^{+}02, DM00, Edg00, HWT^{+}06, HHL^{+}05, KRJO02, QSZ^{+}03, RB03, SHZ^{+}00, SBSR02, SGRM07, WDC^{+}03, WCB08, ZB07]. Freund [SMS02]. friendly [JGK^{+}09]. Frisian [WHD^{+}08]. frog [WHW^{+}02]. fry [BWMSB02, DRA00, GKCF04, LLJC03, LBN02, MMD05, VTA06]. FTA [SMK^{+}09]. fugu [HMO^{+}08]. fugu-derived [HMO^{+}08]. full [CGW04]. full-sib [CGW04]. fulvidraco [Wan02]. fulvidraconi [Wan02]. fumagillin [RKG^{+}00, WYWL01]. function [GMBN04]. functional [NHA05]. fungi [BTSH03]. fungus [LHP^{+}03]. furcatus [SBG08]. Further [FBP09, AN04a]. furunculosis [BGBG05, GLMG03, IA02b, IA03, OB07, RDT^{+}03, SGMP^{+}05]. Fusarium [KHA04b]. fuscoguttatus [GKNLN^{+}04]. fusion [CYL^{+}06, LMX^{+}05].

G [XNS04]. Gadus [BFR07, FCM^{+}04, GUE06, GKOF09, GBGB06, Kha04a, KKD^{+}09, MBG^{+}02, MGL^{+}06, MOR^{+}09, OMR^{+}06, ONI^{+}08, SBE03, SBM^{+}06, SWS07, SKGR05, UBB^{+}09]. Galicia [IPA^{+}01]. Galician [ACP01]. gallaicus [IPA^{+}01]. galloprovincialis [CABF08, DBR04]. gangliosides [SKI^{+}06]. garlic [NA09b]. garvieae [ABR^{+}09, APYH00, CLS^{+}02, EKS09, MOO^{+}09, MHKA08, SKI^{+}06]. Gasterosteus [JKD06]. Gastric [LCH^{+}02, FF07, LMG^{+}03, TBB^{+}08]. gastrointestinal [BC06, HBK07]. gavage [JH01]. gel [AMI^{+}07]. Gene [SXSN07, CWJL07, CFA^{+}00, DW05b, FRW^{+}08, HSK^{+}08, KBH^{+}08, NHA05, SGG03, WCK^{+}09, XMM^{+}09, YTLN01]. genera [DBR^{+}05]. genes [ALAC09, BF07, CYL^{+}06, HKV^{+}06, KHA05, KL07, NPF05, OLR03, RHS^{+}06, SLB04, WZJ^{+}08, XNS04]. Genetic [BGR08, CLN05, CDL^{+}08, NBB08, Pen04, RBBC^{+}09, SC06, SNN^{+}08, SMN^{+}09, VJR^{+}09, BFR07, NPF05, RM08, SBB^{+}06, SGRM07, TXZ^{+}09, TJG04b, VAH09, ZXZ^{+}08]. genetically [BMKG^{+}07, SSH^{+}07]. genogroup [CDT^{+}07, PGC^{+}09, ZS04]. genogroup-specific [PGC^{+}09]. Genome [WDC^{+}03, RHS^{+}06]. Genome-based [WDC^{+}03]. genomes [KHA05]. genomic [ASEK03]. genomovar [MMB02]. genomovars [ZASK06]. genotype [BLL^{+}07, BBC^{+}05, EFT^{+}06, HAEM02]. genotypes [IOK^{+}07, ØH07, SKGR05]. genotypic [MHHK^{+}01]. genus [ABCRH09, DNC^{+}05, HBCS^{+}06, KPWK06, OMR^{+}06]. geographic [CWW^{+}02]. geographical [AAH^{+}07]. Germany [JNS^{+}09, WHD^{+}08]. giant [BSQW05, HWT^{+}06, HHL^{+}05, IGH01, KRJO02, QSZ^{+}03, RB03, WDC^{+}03]. gibelio [WYWL01]. gigas [RLA01, WHD^{+}08]. Gilbert [IGH00, ILAW03, WWK03]. Gill [BB01, GRL04, MZF01, NDB^{+}04, Rob06, AN01, AN03, AN04a, AEN04, AN04b, AVN08, BSDD02, BSD05, BM04, BM06, BN04a, BBN03, BCMN05, FZM00, FBP09, HCFM04, IFA05, KHA04b, LVW^{+}06, LHHP07, MPEN05, MCN04, MN08, MO06, PNA02, RP03, SKR^{+}08, TBB^{+}08, VMN06, YDNM08, ZB07, ZM00, ZM01b, ZGM01]. gill-associated [MO06]. gill-derived [LVW^{+}06]. gills [AN01, CAAN07, DG08, GCR^{+}02, IPA^{+}01, MN08, ZGM01]. gilt [CFA^{+}09]. gilt-head [CFA^{+}09]. gilthead [AVP^{+}09, CRA^{+}05, CVGG06, MMTPDR^{+}08, MST^{+}04, PRG^{+}05, RKG^{+}00, RAT^{+}02, SBDPAP07]. ginger [NA09a]. Gingival [BHRB04]. GIV [MWG^{+}02]. glands [BT09]. glauca [BD00, BKNS02, BHB03, BHRB04, BSTW08, BT09, BKSB09]. Glaucosoma [SCJ^{+}01, SCJ^{+}02]. global [Cos09]. glomerata [SNOR09]. Glomerulonephritis [MST^{+}04, LRH^{+}08]. glucan [BCMN05, GMS06, GMS07, KMJM06, KS06]. glucose [DW05b]. glucose-6-phosphate [DW05b]. Glugea [LYO04]. glycans [KSKEM07]. glycoconjugates [FW01, TWOM06]. glycogen [TTD^{+}06]. glycoprotein [PV08]. gnathiids [HSD07]. GNNV [CLL01]. gobio [FLHM04]. goby [TL09]. Goitre [GDP^{+}03]. gold [NDH05a]. goldfish [AHB07, GSMM09, GRL04, IRA03, JBB^{+}07, MGB^{+}07, PB04, SMG08]. good [MD08]. Gorbunova [SMR01]. gorbuscha [JKB08]. gossypol [GAMR^{+}04]. Gould [WP00]. gourami [FLGS00, IGH01]. Goussia [GA05]. gradient [MN08]. gradient-2 [MN08]. Gram [AJA03, FH03]. Gram-negative [AJA03, FH03]. grani [CAC^{+}08]. granulomas [MIGR09]. granulomatous [KHH^{+}05, OMR^{+}06]. grass [LLZ08]. grayling [MPP^{+}03, PSTV06]. greater [LGF04]. Greece [ABPK03]. green [CYL^{+}06]. Greenshell^{T M} [KWKLG09]. grey [CLS^{+}02]. Gross [AEN04, JDBN08, PNS^{+}05]. groundwater [PC04]. group [NMJ^{+}04, NUS^{+}06, EKG^{+}02]. grouper [CWT^{+}00, CLL01, CCL09, GKNLN^{+}04, HLWH07, LMJ^{+}00, LMC^{+}01, LCM^{+}01, LJL^{+}03, LCC^{+}08, MWG^{+}02, TMA^{+}01, TKNM03, TTM04, YMKN09, YMN09, YKR^{+}02]. groups [MHD^{+}04, WMB02]. grow [TTM04]. grow-out [TTM04]. Growth [SSL^{+}08b, BHP^{+}04, BCB02, DW05b, DGLS09, GWF^{+}02, HKPKS06, LVW^{+}06, RSD04, RRG^{+}05, SNNB01, SN08, VSNS08]. grunniens [WCB08]. Guidelines [PS02]. Gulf [CLQAGCRP08]. Günther [AVN08]. guppies [LAZ05, LZ09]. gurnard [MIGR09]. gut [VSNS08]. gymnamoeba [CAAN07]. gyrA [RMHA08]. gyrB [IYS^{+}07]. Gyrodactylidae [áKBH06]. gyrodactylids [JHKB08]. gyrodactylosis [XSK07]. Gyrodactylus [ARKK08, BCB02, BNN03, HMB07, KJB07, áKBH06, LSDB06, OLB03, SGS01, YYY^{+}06, YWS^{+}08].

H [Jay09]. habits [TWS09]. haddock [TBLC07]. haem [NHA05]. haematology [PSB06]. haematopoietic [AST^{+}05, ACS^{+}08, AJ09, BMŠH^{+}02, GSMM09, LBO^{+}01, MBB^{+}06, MBLK06, OL06, PGC^{+}09, YSH^{+}06]. haemocytes [HHL^{+}05, Jay09, LB00, SWZ^{+}05]. haemoflagellate [AW01, CGW04, FW01, HFW01, ZW00]. Haemoglobin [SCJ^{+}02]. haemolymph [SMK^{+}09]. haemolysis [ZW00]. haemolytic [EKG^{+}02, NHA07, SLC^{+}06]. haemorrhage [PT09]. haemorrhagic [Bow03, EFT^{+}06, GMB^{+}07, HGP^{+}07, HO03, JGK^{+}09, KIIK04, LVDBB07, SKO04, SOM05, SKGR05, SFT^{+}08]. Hafnia [PRG^{+}05]. Hainan [TXZ^{+}09]. halibut [Bow03, GME04, GLMG03, GBGB06, JGS^{+}04, KKD^{+}09, ONS07, SKGR05, SIM^{+}00]. Haliotis [CCTL06]. Haliphthoros [Dig01]. Halocynthia [HOA09]. Hamilton [ACS^{+}09, CHWK^{+}04, GDR^{+}07, GPS^{+}08, HBK07, KMJM06, KBS03, RWSK09, SMS02, TL09]. Haná [KBPD04]. Haploblepharus [HSD07]. hapten [SLC08]. Harbour [GDR^{+}07]. Hardanger [HOM^{+}09]. hardness [RP03]. harvest [AGM^{+}08]. harvest-sized [AGM^{+}08]. harvested [ZGM01]. harveyi [AGSS^{+}01, ARC^{+}05, ABC^{+}08, APR03, CRA^{+}05, CN04, WZJ^{+}08, ZA00]. hatcheries [MMD05]. hatchery [BCB02, RA01]. hatchery-reared [BCB02, RA01]. hatching [EWCT05]. Hawaii [PWW^{+}06]. Hay [WCB08]. head [CFA^{+}09, CLQAGCRP08, CDS^{+}00, DCTCB^{+}06, PZS02]. healing [ONS07]. health [GWF^{+}02, NDB^{+}04, PSLN^{+}08, TF04]. healthy [BMD^{+}06, IGH00, Koz07]. Heart [KKT^{+}04, ATT^{+}09, FKT^{+}05, KHTF06, KT09, MCS08]. heat [PLC09, SAC^{+}09]. hebraicum [SCJ^{+}01, SCJ^{+}02]. held [SBE03]. helminth [Ken07]. helminths [DGA^{+}03]. Henneguya [KEEM^{+}05, YIT05]. henneguyosis [YIT05]. Hepatic [ACP01, BHB03]. hepatitis [PFL^{+}08]. hepatotoxin [SGF02]. Heritability [NFR08]. Herpes [RA01, GCG^{+}04, RLA01, SMG08]. Herpes-like [RA01, GCG^{+}04, RLA01]. Herpesviral [GSMM09]. Herpesvirus [WCJJ08, DJW^{+}09, GML^{+}02, GKVS04, HKKO07, HKK^{+}09a, JBB^{+}07, MIS07, SHBJ^{+}09, UMIK09, JNS^{+}09]. herring [HGP^{+}07, JD02]. heterogeneity [SMN^{+}09]. High [HLWH07, LVW^{+}06, GCR^{+}02, JO00, TWOM06]. high-molecular [TWOM06]. highly [HGP^{+}07, LSDB06, VFT^{+}03]. Hilgendorf [KL07]. Hine [HDP^{+}02]. Hippocampus [AGSS^{+}01]. Hippoglossoides [MFLC01]. Hippoglossus [Bow03, GME04, GLMG03, GBGB06, JGS^{+}04, KKD^{+}09, ONS07, SKGR05, SIM^{+}00]. hirame [SYYO06]. Hirudinea [WCB08]. Histidine [BBC^{+}05]. histochemical [BMPB^{+}00]. histochemistry [MA04]. Histological [BMKG^{+}07, CIK^{+}04, BT09]. histologically [TPF09]. Histopathological [KIIK04, RJ02, TTM04, BT09, GPS^{+}08, JDBN08, MNM^{+}02, RJC04, SBM^{+}00]. Histopathology [DGS^{+}02, LGF04, AEN04, KM01, NDB^{+}04, PNS^{+}05, RBA^{+}03, SBB^{+}06]. histophagous [CPB^{+}07]. History [Rob06, PNA02]. hoferi [CWW^{+}02, JD02]. Hoffman [LWWS04]. Homarus [BWS^{+}08, DSBB05, Dov05]. Homogeneity [MOO^{+}09]. hooks [BKNS02]. horizontal [BS04, SBDPAP07]. hormonal [BMKG^{+}07]. horse [ACP01, GA05]. Host [ØSKH06, BCB02, BBA08, CABF08, Cvi04a, ESMB00, HSW03, LB09, OLB03, SBDPAP07, XKS04]. hosts [ABCRH09, LYW03, NPF05, OW02]. hovorkai [LYW03]. hsp70 [DW05a]. Humoral [CGW04, AW02, BCM^{+}02, WMB02]. humpback [YKR^{+}02]. Husbandry [RWSK09, SRT^{+}07]. Hutton [Dig01]. HVA [JNS^{+}09]. HY9901 [CWJL07]. hybrid [AL09, BWCM04]. hybridization [CABF08, DCTCB^{+}06, HSW03, HWT^{+}06, úMMUB08, PTKA02, VCLS04, WZJ^{+}08, WDC^{+}03]. hybrids [SBG08]. hydrogen [DRCMB03, MPEN05]. hydrophila [CFA^{+}00, CL05, FLGS00, GPS^{+}08, IRA03, KMJM06, NM02, NA09a, NA09b, VSNS08]. hyperinfection [CZP01, GPN05]. hyperoxygenation [SOF^{+}09]. hyperparasitic [FBS03]. hyperplastic [BSTW08]. hypochlorite [HPG08]. hypodermal [MBB^{+}06, YSH^{+}06]. hypophthalmus [CDT^{+}02, FTS^{+}01]. hypoxia [GKB05].

I. [SBG08]. Iberian [CDT^{+}07, IPA^{+}01]. Ibero [GA05]. Ibero-Atlantic [GA05]. Iceland [áKBH06]. Ich [XKS05]. Ichthyobodo [CN02, CLN05]. Ichthyophonus [CWW^{+}02, HPG08, JD02, KWD^{+}01, KH06, KLG^{+}06, KHSW09, LKH08]. Ichthyophonus-induced [KLG^{+}06]. Ichthyophonus-infected [KHSW09, LKH08]. Ichthyophthirius [RF01, SB01, SLB04, XK02, XKS02, XK03, XKS04, XKS05, XKS06, XKP06]. ichthyotoxin [ZRT04]. ictaluri [BSW03, ASEK03, CDT^{+}02, GM07, NM02]. Ictalurus [BNS03, BSW03, CUN09, DM09, DMS09, Goo06, GM07, NM02, SKLY03, SBG08, SGF02, SGG03, TJG04a, XK02, XKS02, XK03, XKS04, XKS05, XKP06, ZKG^{+}01]. Idaho [PGC^{+}09]. idella [LLZ08]. Identification [BBFB07, CAB^{+}07, CDT^{+}02, GCWFB08, IYS^{+}07, Jay09, OBAA09, PTKA02, Rob05a, SLC08, WZJ^{+}08, ZXZ^{+}08, ZRT04, AOA^{+}09, CCTL06, IFA05, KFM^{+}00, PTT^{+}08, PTAR09, SBB^{+}06, TYH^{+}01]. identified [RHS^{+}06]. Idiopathic [PBIM09, RST^{+}07, Edg00, PNS^{+}05]. ignotus [MOG09]. IHNV [ACS^{+}08, KBEJ^{+}07, LBO^{+}01, OLP01]. II [XCD09]. III [BF07]. IL-1 [LSDB06, PZS02]. immersion [CN04, SGMP^{+}05]. immobilization [SB01, XKP06]. Immune [BNN03, CN04, DMG09, PSP^{+}09, PWS^{+}08, XKS04, YKR^{+}02, AW02, ABC^{+}08, BMPB^{+}00, BMKG^{+}07, BPV^{+}03, BCM^{+}02, FM00, GPB^{+}05, Jay09, LLZ08, LRH^{+}08, MFLC01, NEG06, SSH^{+}07, SLB04, SLC08, VFAN03, WZJ^{+}08, WMB02, WM04, XK02, XKS02, XK03, XKS05]. immune-complex [LRH^{+}08]. immune-related [WZJ^{+}08]. immune-response [SSH^{+}07]. immunity [FLGS00, GPS^{+}08, KS06, TMA^{+}01]. Immunization [HPMW08, GLC04, LLJ^{+}02, LLJC03, SKSE02, SLC08, XKS04, ZWZS02]. immunized [YMKN09]. immuno [AW01, BCM^{+}02]. immuno-adjuvant [BCM^{+}02]. immuno-chemotherapeutic [AW01]. immunocytochemical [BSTW08]. immunodominant [VLT^{+}07]. immunogenic [XLZS08]. immunogold [MA04]. Immunohistochemical [DADB00, MAF^{+}00, MN08, DPS^{+}07]. immunohistochemistry [DGS^{+}02, EN03, FTE^{+}03, JTM^{+}01, MMAP00, SBB^{+}06]. immunolocalization [LCM^{+}01]. Immunological [CBDC^{+}03, CVGG06, OL06]. immunomodulatory [FM00]. immunoperoxidase [GJR^{+}03, GJR^{+}05]. immunoperoxidase-based [GJR^{+}03, GJR^{+}05]. immunosorbent [RB03, SB01]. Immunostimulant [STM^{+}01, NA09a]. Immunostimulants [SBV09]. Immunostimulation [MGL^{+}06]. Immunostimulatory [BBN03]. immunosuppression [MST^{+}04]. impact [TWS09]. implications [BKNS02, CLN05, ØSKH06, OLB03, SOM05, WMB02]. important [KWD^{+}01, WM04]. imported [MMB02]. imprints [WMA^{+}02]. Improved [BLT01, úMMUB08, MMA02a]. Inactivated [ACS^{+}08, IRA03, MBB^{+}06, PSP^{+}09, YMKN09, YMN09]. Inactivation [HPG08, BSM^{+}06, CFA^{+}00]. incarnatum [KHA04b]. incidence [GBB^{+}06, SRT^{+}07]. inclusion [Rod07, USR^{+}09]. increase [BBN03]. Increased [ATT^{+}09, LHHP07]. increases [PB04]. incubation [ØGW04]. index [Ano01b, Ano01e, Ano02a, Ano02c, Ano03a, Ano04a, Ano04b, Ano04c, Ano05a, Ano05b, Ano05c, Ano03d]. indicate [CWW^{+}02]. indicates [GDG00]. indices [SKLY03]. indicus [Jay09, RJB00]. indigenous [BDSY07]. induce [ABC^{+}08, XKS05]. Induced [XKP06, ATNF02a, AGM^{+}08, BMPB^{+}00, BGB^{+}04, BGBG05, BML^{+}06, GBGB06, HHL^{+}05, KHA05, KLG^{+}06, LCC^{+}08, SBM^{+}06, TOW02]. induces [HHW05, SLB04]. Induction [HCFM04, HLC00, CCL09, IKKS06, MCN04]. industry [Cos09]. infected [BSG^{+}05, CYL^{+}06, CMD01, DGF02, DADB00, DGS^{+}02, DG08, GKB05, GDG00, GML^{+}02, GMBN04, KHSW09, LKH08, LZ09, LLHF02, MBG^{+}02, MMA02b, NM02, RSD04, SMG08, SWZ^{+}05, SGG03, VFAN03, WGM^{+}08, YSN^{+}06, YNH03, ZGM01, GMW^{+}07]. infecting [BTSH03, BBA08, FBS03, GA05, IPA^{+}01, KJC^{+}02, MOG09, RLA01, ZWW^{+}06]. Infection [JZKS04, MCS08, SYYO06, AG01, ATNF02a, AHTM06, BLL^{+}07, BCB^{+}09, BGB^{+}04, BGBG05, BWCM04, CFA^{+}09, CAC^{+}08, CMD02, CHWK^{+}04, CZS^{+}01, CDGF07, CDI^{+}01, Cvi04a, DM09, DMS09, DPS^{+}07, DBR^{+}05, DLS^{+}07, EN03, FCM^{+}04, GSMM09, GRF^{+}07, GMS^{+}09, GMBN04, GPB^{+}05, GBGB06, HBK07, HGH^{+}08, HCFM04, IRA03, ITK^{+}04, JO00, KH06, áKBH06, LCM^{+}01, LYO04, LAZ05, LZ09, LSDB06, LVDBB07, LWWS04, MBG^{+}02, MKW^{+}09, MZF01, MSD^{+}00, NMJ^{+}04, NA09b, OL06, PZD^{+}06, PSB06, PSTV06, RBA^{+}03, SSM00, SHZ^{+}00, SBM^{+}00, SNOR09, SKO04, SRM^{+}03, SHN^{+}00, SHM00, SAC^{+}09, STV05, TSL^{+}01, TJG04a, TBLC07, TPF09, UBB^{+}09, VNSB00, WSKC03, WCW^{+}08, WCT^{+}09, YMKN09, YWS^{+}08, YCR^{+}09, ZM00]. Infections [WCT^{+}01, AW02, CN02, CM04, FLHM04, HRG03, HDP^{+}02, HKPKS06, JNS^{+}09, JKD06, KWD^{+}01, KBPD04, LGF04, MFLC01, MG07, MMB02, MKM02, NDPH03, NA09a, OLB03, RWSK09, RA01, SBSR02, TSW06, TDH^{+}07, UCD04, Woo04, YYY^{+}06]. Infectious [GUE06, PGC^{+}09, RP05, SBB^{+}06, SLC^{+}06, AST^{+}05, ACS^{+}08, AL09, BGR08, BMŠH^{+}02, BWMSB02, BM09, BLL^{+}07, BB09, BSE02, GMW^{+}07, GMS^{+}09, GER^{+}07, GEB^{+}08, GBB^{+}06, HD05a, HD05b, JO09, KKT^{+}04, LBO^{+}01, MASS07, MEC^{+}06, MHD^{+}04, MBB^{+}06, MBLK06, MGAE06, NDH^{+}05b, NDH05a, OL06, RW03, RBBC^{+}09, RAG^{+}07, RMSH09, SBM^{+}00, SFPF01, SM08, SRM^{+}03, SMM09, SOF^{+}09, UMG^{+}08, VLT^{+}07, WGM^{+}08, WMA^{+}02, YSH^{+}06, RMG^{+}09]. infective [CABF08, DSJPP08]. infectivity [DGF02, JD02]. infestation [CKW^{+}01, JH01, LGR08, RFBD^{+}05]. infestations [BB01, GER^{+}06, LBGR08, SSS^{+}00]. inflammation [ATT^{+}09, FKT^{+}05, KKT^{+}04, KHTF06, KT09, KHH^{+}05, YCR^{+}09]. inflammatory [OMR^{+}06, SMS02]. inflation [PDOZ08]. Influence [BSD05, DRA00, DGA^{+}03, EMH02, SPK^{+}05, STV05, GSW^{+}07, GWJR08, RM08, SN08, SGRM07, SRT^{+}07, TWS09]. influences [BLL^{+}07, SKLY03]. ingestion [SGF02]. inherent [BCMN05]. Inhibition [AST^{+}05, CK08, SDD^{+}00]. Inhibitory [HFW01, GCWFB08, HBB09]. iniae [NBB08, SKSE02, ZXZ^{+}08]. initial [AN04b, CABF08, JTM^{+}01]. initiated [LKH08]. injected [YKR^{+}02]. injection [CN04, KM01, MLB^{+}06, RTNA02, SGMP^{+}05]. innate [GPB^{+}05, KS06]. innesi [MMB02]. INNO [PTT^{+}08]. INNO-LiPA [PTT^{+}08]. inoculation [YMN09]. insertional [CFA^{+}00]. insoluble [LNZL08]. insulin [DW05b]. insulin-like [DW05b]. intact [TWOM06]. intake [RSD04]. integrated [Bro09]. integrity [MDB^{+}09, SOF^{+}09]. intensity [KaVT06]. interacting [Jay09]. Interaction [BSD06, BCB02, CRBM05, LB00]. Interactions [CRA^{+}05, BN04b, DSJPP08, NDB^{+}04]. interest [BH02]. interleukin [BCM^{+}02]. interleukin-1 [BCM^{+}02]. internal [MEM^{+}09, NPF05]. intestinal [BMPB^{+}00, BD00, DGA^{+}03, MDB^{+}09, SBK^{+}05, SKI^{+}06, SOF^{+}09, USR^{+}09, VLK^{+}04]. intestine [BMFL^{+}07, DADB00, DGS^{+}02, JOG^{+}06, JSG^{+}08, TOW02]. intestines [DG08]. intimal [SSL^{+}08a]. intracellular [DDL^{+}01, FH03, MASS07, WP00]. intraerythrocytic [CKK02]. intramuscular [MES^{+}02]. Intramyocardial [PT09]. intranuclear [EWBB02]. Intraperitoneal [GMS06, CN04, GMS07, KKD^{+}09]. intravascular [RTNA02]. introduced [UMIK09]. Introduction [SPL05]. intubation [CN04]. invadans [JZKS04, LHP^{+}03]. Invasion [AVP^{+}09, PSTV06]. Investigation [BPV^{+}03, SKO04, WMB02, CGP08, GMS^{+}09]. involvement [JSG^{+}08]. involves [AHB07, KEEM^{+}05]. iodine [HPG08]. ion [PSB06]. Ionophore [SDD^{+}00]. Ionophore-mediated [SDD^{+}00]. IPN [RRKM07, RM08]. IPNV [BLSE03, CLVO^{+}05, GME04, GMW^{+}07, MGE04, ME08, UBB^{+}09, WGM^{+}08]. IPNV-carriers [GME04]. IPNV-infected [GMW^{+}07]. IPNV-like [CLVO^{+}05]. Ireland [BM04, GJR^{+}03, GCG^{+}04, NFR08, RM07, RMSH09]. irido [KJC^{+}02, LJL^{+}03]. irido- [LJL^{+}03]. irido-like [KJC^{+}02]. iridoviral [GKNLN^{+}04]. Iridovirus [JO00, CLW03, CCL09, DLS^{+}07, HWZ^{+}00, KJL07, KJL08, LMJ^{+}00, LCC^{+}08, MWG^{+}02, OKK^{+}06, WSKC03, WHW^{+}02, WWH^{+}02]. Iridovirus-like [JO00, HWZ^{+}00]. iridovirus-susceptible [LMJ^{+}00]. Iron [MEBD05, RP04, RDT^{+}03]. ISAV [GMS^{+}09, SRM^{+}03, SMM09, WMA^{+}02].

Island [BWS^{+}08]. Isles [SBB^{+}06]. isogenic [SSH^{+}07]. isolate [ABR^{+}09, ILAW03, JGK^{+}09, KWKLG09, KKD^{+}09]. isolated [BFR07, Bow03, CTB^{+}06, CRBM05, CRA^{+}05, CWW^{+}02, HPMW08, IGH00, KHA04b, KJL07, KL07, Koz07, LVDBB07, MPP^{+}03, MMTPDR^{+}08, MWG^{+}02, NLO08, NUS^{+}06, OKK^{+}06, PTKA02, RBBC^{+}09, STM^{+}01, SBE03, SC06, SNN^{+}08, SRAHJ^{+}08, SME^{+}01, TH08, VAH09, VJR^{+}09, WCK^{+}09, ZS04, ZXZ^{+}08, ZAC^{+}03]. isolates [AMI^{+}07, ASEK03, GDG00, GDG01, GHBW03, IYS^{+}07, IOK^{+}07, JTR^{+}04, KFM^{+}00, MOO^{+}09, MD00, NBB08, OFKEA08, PV08, RMG^{+}09, RMSH09, SBB^{+}06, SRYC02, TJG04b]. Isolation [ACGR^{+}05, BHO05, CCTL06, CLVO^{+}05, DAMQ^{+}04, EEF06, EFT^{+}06, FKA^{+}05, GMB^{+}07, HBB09, JPPS06, JBB^{+}07, SM08, SIM^{+}00, Sta01, TWOM06, VFT^{+}03, BMŠH^{+}02, CDS^{+}00, EKS09, GVGR00, GRW^{+}03, KBEJ^{+}07, KFM^{+}00, LVA^{+}06, LNM^{+}09]. isometamidium [AW01]. Isopoda [HO03]. isothermal [AL09, EMS05, FNE09, GEC09, GKVS04, PBW06, SKIS05, SSK^{+}08]. Israeli [UCD04]. issue [RLB09]. Isurus [BT09, BKSB09]. Italy [GRF^{+}07, ZFF^{+}08]. ITS-1 [NPF05]. ivermectin [JH01]. Iwagaki [ITK^{+}04].

J [Ano02b]. jack [INM^{+}01, MOO^{+}09, MKW^{+}09]. Janeiro [LSM^{+}02]. Japan [HKKO07, INKS04, ILAW03, MKW^{+}09, NMJ^{+}04, SNN^{+}08, SYYO06, SSO^{+}09, WCT^{+}01]. Japanese [AMI^{+}07, FYO08]. japonicus [CLL^{+}00, SMK^{+}09, VNSB00, WM04]. jarbua [WCT^{+}09]. Jasus [Dig01]. jaundice [SLC^{+}06]. JF3835 [GGCFB09]. Jordan [FYO08, IGH00, ILAW03]. junction [BD00]. juvenile [BLL^{+}07, CKW^{+}01, CDI^{+}01, Dig01, FPG^{+}02, FDG^{+}06, GUE06, GGCFB09, GPRCH09, HSD07, JKD06, JKB08, LFF05, PC04, PSB06, SYYO06, SMN^{+}09, TPF09, WYWL01]. juveniles [GPS^{+}08, HGP^{+}07, MGL^{+}06, MST^{+}04, TKNM03].

Kabata [BB01]. Kagawa [INKS04]. Katz [AW01, HFW01, ZW00]. Kaup [ACGR^{+}05, ARC^{+}05, CRBM05, CRA^{+}05, DRCA^{+}08, MMTPDR^{+}08, ZAC^{+}03]. kDa [TH08]. Kellogg [SAC^{+}09]. Keyword [Ano04b, Ano05b]. kidney [BH02, CMD02, CDS^{+}00, Cvi04b, DL01, EMH02, FPG^{+}02, GEC09, GUE06, GCR^{+}02, JLSA08, KM01, LMJ^{+}00, LLHF02, MAF^{+}00, MMA02a, MMA02b, MGAE06, PZS02, WKB^{+}02, WZJ^{+}08]. KilC [TH08]. killing [Jay09, TH08]. kinase [HGH^{+}08]. Kinetoplastida [AW02, FW01]. Kingdom [GRW^{+}03, SFT^{+}08]. kingfish [MPEN05]. kisutch [CDL^{+}08, SLC^{+}06]. kisutchi [HCJ08]. Klamath [TPF09]. klunzingeri [EKG^{+}02]. knowledge [NEG06, WKB^{+}02]. known [HBCS^{+}06]. koi [DJW^{+}09, GML^{+}02, GKVS04, HKKO07, HKK^{+}09a, MIS07, SMG08, SSK^{+}08, SHBJ^{+}09]. Korea [KJC^{+}02]. Korean [JO00]. Krøyer [JKD06, JKB08, PMZD08, PD09, SSS^{+}00, SBSR02, TSW00]. kuchijirosho [HMO^{+}08]. Kudoa [BBA08, CBDC^{+}03, DCCF^{+}03, HBCS^{+}06]. kudoosis [LJM08]. Kuwait [DAMQ^{+}04, EKG^{+}02].

L. [ACP01, AVP^{+}09, AN01, AN03, AN04a, AEN04, AN04b, AVN08, ATT^{+}09, ATNF02a, AW02, AJ09, ABPK03, AHB07, AGM^{+}08, AHTM06, BLGR09, BMPB^{+}00, BMKG^{+}07, BMFL^{+}07, BPV^{+}03, BVG^{+}03, BFR07, BM04, BM06, BCB^{+}09, BBFB07, BBBW01, BGB^{+}04, BGBG05, BD00, BF02a, BKNS02, BF02b, BHB03, BHRB04, BSTW08, BT09, BKSB09, BSE02, Bow03, BMS^{+}00, BCM^{+}00, BBC^{+}05, BPBT02, BBN03, BCMN05, BBK04, BWD^{+}07, BLT01, CIK^{+}04, CFA^{+}09, CTB^{+}06, CRA^{+}05, CL02, CBDC^{+}03, CKW^{+}01, CLS^{+}02, CGW04, CKK02, CVGG06, CR02, CCB03, CZP01, CAAN07, CLVO^{+}05, DCCF^{+}03, DCTCB^{+}06, DADB00, DGS^{+}02, DGA^{+}03, DPS^{+}07, DMG09, DJW^{+}09, EEF06, EKG^{+}02, EKS09, FLHM04, FCM^{+}04, FKT^{+}05, FZM00, FM00, FBP09, FTE^{+}03, GME04, GUE06, GDR^{+}07, GA05, GKNLN^{+}04, GKOF09, GGCFB09, GCWFB08, GKB05]. L. [GPN05, GRL04, GLC04, GWF^{+}02, GCR^{+}02, GCG^{+}04, GJR^{+}05, GDG00, GDG01, GMS^{+}09, GDP^{+}03, GMBN04, GCN04, GPB^{+}05, GLMG03, GBGB06, GKVS04, GER^{+}06, GER^{+}07, GBB^{+}06, HRG03, HSW03, HO03, IRA03, JNS^{+}09, JBB^{+}07, JTR^{+}04, JST^{+}04, JGS^{+}04, JKD06, JHKB08, KPWK06, Kha04a, KKT^{+}04, KHTF06, KT09, KFM^{+}00, KHH^{+}05, KaVT06, KaVKT09, KKD^{+}09, KBPD04, KS06, LHHP07, LGR08, LJM08, LSM^{+}02, LSDB06, LGF04, LHW08, LVDBB07, LBN02, MPP^{+}03, MIGR09, MBG^{+}02, MGL^{+}06, MMSM06, MGB^{+}07, MMTPDR^{+}08, MHD^{+}04, MNM^{+}02, MALC00, MOR^{+}09, MCS08, MEM^{+}09, MST^{+}04, MN08, MGE04, MGAE06, ME08, NLO08, NFR08, NDB^{+}04, OKK^{+}06, OLB03, OMR^{+}06, ØGW04, ONI^{+}08, ONS07, PRG^{+}05, PZD^{+}06, PFL^{+}08, PMZD08, PSS09, PB04, PTB07, PNA02, PSTV06, RM08, RGTR02, RRG^{+}05, RST^{+}07, RKG^{+}00, RAT^{+}02, RTNA02, RNAA04]. L. [RMG^{+}09, RHS01, RP03, RP05, RA08, Rod07, RM07, RP04, RM04, RDT^{+}03, RAG^{+}07, SMG08, SSH^{+}07, STM^{+}01, SNNB01, SBE03, SBK^{+}05, SPC^{+}08, SGMP^{+}05, SPL05, SPK^{+}05, SSL^{+}08a, SBM^{+}06, SKSE02, SSK^{+}08, SBM^{+}00, SFPF01, SBDPAP07, SKO04, SSL^{+}08b, SBB^{+}06, SM08, SWS07, SRM^{+}03, SKGR05, SHFM09, SHBJ^{+}09, SIM^{+}00, Sta01, SKR^{+}08, SN00, SHM00, SSS^{+}00, SBSR02, SHS05, SGRM07, SRT^{+}07, SHRM07, SOF^{+}09, TFN01, TOB^{+}07, TBLC07, TSW00, TH08, UMIK09, USR^{+}09, UMG^{+}08, UBB^{+}09, VJR^{+}09, VHM^{+}00, VSNS08, VNSB00, VCN09, VFT^{+}03, VFAN03, VMN06, WHB^{+}03, WBP^{+}05, WGM^{+}08, WZJ^{+}08, WWH^{+}02, XSK07, XCD09, ZWW^{+}06, ZM00, ZM01a]. Labeo [ACS^{+}09, GPS^{+}08]. Labile [OLB03]. laboratory [GRF^{+}07, MOR^{+}09, MCN04, ØSKH06, TBLC07]. laboratory-based [MCN04]. laboratory-challenged [TBLC07]. labrax [ABPK03, BPV^{+}03, BVG^{+}03, BPBT02, CZP01, HO03, RTNA02, RNAA04, RP04, SBDPAP07]. Lacépède [BB01, GAH^{+}08]. lactococcosis [BA05]. Lactococcus [ABR^{+}09, APYH00, CLS^{+}02, EKS09, MOO^{+}09, MHKA08, SKI^{+}06]. lactoferrin [KMJM06]. ladderlectin [YCR^{+}09]. laevis [DGS^{+}02]. Lake [GDR^{+}07, PGC^{+}09, WCB08, KaVKT09, EEF06, EFT^{+}06]. lalandi [MPEN05]. LAMP [GEC09, SSK^{+}08]. lamprey [EEF06]. Lampreys [Rob06]. Lancefield [NMJ^{+}04, NUS^{+}06]. large [CLW03, GDP^{+}03, WYJ05]. largemouth [GPN05, WBNG02]. Larimichthys [CLW03, WYJ05]. Larvae [YNH03, BPBT02, CCTL06, DMG09, JPPS06, KWKLG09, KBPD04, MGL^{+}06, MBB^{+}06, PMZD08, RA01, RLA01, SAC^{+}09, TL09, VFT^{+}03]. Larval [HGP^{+}07, BWD^{+}07, KHK^{+}04]. latently [NM02]. Lateolabrax [CLL^{+}00]. Lateral [BB07, CUN09, AL09]. Lates [BB07, BCM^{+}02, CCL09, CN04, FDG^{+}06, GKCF04, HTC^{+}01, LCC^{+}08, PSP^{+}09, PC04]. Latreille [MRM^{+}00]. Latris [GDHM03]. Laureau [HDP^{+}02]. lavaretus [BWD^{+}07, DPS^{+}07, SKO04]. Laveran [PB04]. Lawrence [CM04]. laying [TWS09]. LD [KM01]. leading [HGH^{+}08]. lectin [BHO05, MA04]. Leeches [Rob06]. leei [SBDPAP07]. lenok [YWS^{+}08]. lentus [FAM^{+}06]. leopardus [BB01]. Lepeophtheirus [BLGR09, BMS^{+}00, BBK04, FBS03, GPRCH09, HRG03, JKD06, JKB08, LBGR08, LB09, MSD^{+}00, PMZD08, PD09, PSS09, RRG^{+}05, SKCD07, SSS^{+}00, SBSR02, TSW00]. lepturus [LSM^{+}02]. Lereboullet [EWBB02]. Lernaeapodidae [RJC04]. Lernaeocera [SWS07]. Lesions [BD00, AN01, AGM^{+}08, BBLC00, BF02a, BC06, BT09, Edg00, EEF06, KHA05, SPL05, WP00]. Lesson [PP02]. lethal [DLS^{+}07, FAM^{+}06]. Leuciscus [DGS^{+}02]. leucocytes [CTAR01, CLVO^{+}05, GUE06, MGE04, PZS02]. levamisole [FZM00, FM00]. levan [GPS^{+}08]. level [GER^{+}07, MGAE06, WHB^{+}03]. levels [BWD^{+}07, RDT^{+}03, SSH^{+}07, SKCD07, USR^{+}09, YMKN09]. library [ZWW^{+}06]. Lice [Rob06, AM09, BLGR09, BMS^{+}00, BEL09, Bro09, Cos09, GPRCH09, GER^{+}06, HBN09, HRG03, HOM^{+}09, LGR08, RGTR02, RRG^{+}05, SSS^{+}00, SBSR02]. licheniformis [RLNB03]. life [AHB07, BM09, KEEM^{+}05, KRJO02, LWWS04]. life-stages [KRJO02]. lifetime [SSL^{+}08a]. light [BS04, BBK04, CZP01]. like [BMFL^{+}07, CWT^{+}00, CKW^{+}01, DGMD^{+}09, DW05b, GCG^{+}04, HWZ^{+}00, JO00, KWD^{+}01, KWKLG09, KJC^{+}02, QSZ^{+}03, RA01, RLA01, SHM00, ZB07, CLVO^{+}05]. Lile [áKBH06]. limited [MMB02]. line [ATNF02b, BOB^{+}06, CYL^{+}06, CUN09, GWJR08, HMO^{+}08, HHCBR^{+}07, HHW05, HCFM04, HLWH07, LMC^{+}01, LVW^{+}06, SSH^{+}07]. lineata [GDHM03]. lines [ACS^{+}09, KOK^{+}03, LMJ^{+}00, LJL^{+}03, LCC^{+}08]. link [XSK07]. linked [RB03, SB01, SGF02]. LiPA [PTT^{+}08]. lipid [CIK^{+}04, SPK^{+}05, SSL^{+}08a, TL09, WHB^{+}03]. lipids [DRA00]. Lipoma [MMSM06]. lipomas [JDBN08]. Lipopolysaccharide [GTBA09, ZASK06]. lipopolysaccharides [ASEK03]. liquefaciens [Sta01]. Listonella [SBM^{+}06]. Litopenaeus [DCTCB^{+}06, MGGSREAG08]. litulon [FYO08]. live [CRM^{+}03, PWS^{+}08, TRY^{+}05, XKS04]. liver [CIK^{+}04, DMG09, GA05, LMJ^{+}00]. living [BF02a]. Liza [EKG^{+}02]. LLD [CUN09]. loach [KBPD04]. loads [BSW03]. lobster [Dig01, DSBB05, Dov05, RST^{+}07, SN00]. lobsters [BWS^{+}08]. local [KHA05]. Localization [MA04]. localized [SLC08]. location [HBCS^{+}06, SHN^{+}00]. loch [AM09, PMZD08, PD09]. locus [HSW03]. Loma [BS04, BSD05, BSD06, BHO05, GMS06, GMS07, LWWS04, MSD^{+}00, PSB06, RSSD01, RSD04, SSM00, SSM01, SDD^{+}00]. long [CIK^{+}04, GDG00, GDG01, HBN09, BWS^{+}08]. long-spined [GDG00, GDG01]. long-term [CIK^{+}04, HBN09]. Longitudinal [GJR^{+}05, KHTF06]. Loop [SKIS05, AL09, EMS05, FNE09, GEC09, GKVS04, PBW06, SSK^{+}08]. Loop-mediated [SKIS05, AL09, EMS05, FNE09, GEC09, GKVS04, PBW06, SSK^{+}08]. lophii [FYO08]. Lophius [FYO08]. Loricariidae [PD01]. loss [Goo06]. louse [BBK04, FBS03, JKB08, LB09, MSD^{+}00]. low [AG01, HPMW08, HKK^{+}09b, LSDB06, MGAE06, SI06]. LR [SGF02]. lucioperca [KPWK06]. lucius [GCG^{+}04]. Lutjanus [CWJL07]. LuxS [YML^{+}08]. lyase [SRYC02, STV06]. Lymphocystis [HKK^{+}09b, ACGR^{+}05, CFA^{+}09, HSK^{+}08, HCFM04, KL07, MFLC01]. lymphokine [CTAR01]. lyra [MIGR09]. lysis [JO09, PB04].

M. [WCW^{+}08]. M74 [LBN02]. MAb [HFW01]. MAb-001 [HFW01]. maccoyii [CMD01, DBJN05, HBN09, JDBN08, WMB02]. Maccullochella [BBG^{+}05]. mackerel [ACP01, GA05, LJM08]. Macrobrachium [BSQW05, HWT^{+}06, HHL^{+}05, JPPS06, KRJO02, PNS^{+}05, PBW06, QSZ^{+}03, RB03, SAH^{+}07, WCSC07, WCW^{+}08, WDC^{+}03, YWBH05]. macrophage [AR03, BCMN05, CTAR01, JO09, WKES08]. macrophages [CDS^{+}00, MGAE06]. Maine [GER^{+}07, GEB^{+}08]. mainland [SBB^{+}06]. maintenance [ZM00]. maize [SSH^{+}07, SBK^{+}05]. major [CFA^{+}00, FLGS00, HSK^{+}08, KL07, KJL08, MGGSREAG08, WCK^{+}09, WSKC03, YIT05, ZWZS02]. mako [BT09, BKSB09]. Malacosporea [MMBA05]. male [MOG09]. Malmberg [HMB07, SGS01]. Man [HWT^{+}06, JPPS06, KRJO02, PNS^{+}05, PBW06, RB03, WCSC07, WCW^{+}08]. management [Bro09, CLN05, MD08]. Manchurian [YWS^{+}08]. mandarin [SXSN07]. mandarinfish [HWZ^{+}00]. Manila [DGMD^{+}09]. manteri [BB01]. manual [JH01, Rob05a]. mariculture [PSLN^{+}09, UCD04]. maricultured [CLW03]. marina [BBLC00]. marine [BHP^{+}04, BSM^{+}06, BDSY07, CKW^{+}01, GSM^{+}04, HBB09, INKS04, JSBK09, LVDBB07, MZF01, NDB^{+}04, OKK^{+}06, PSLN^{+}08, RGTR02, SNNB01, SOM05, TKNM03, WGM^{+}08, XLZS08, ZS04]. marinum [HBK07]. marinus [BBLC00, EEF06]. maritimum [AHTM06, LNM^{+}09, VCN09]. maritimus [CGMS03]. markers [BSTW08, SNOR09]. Marsupenaeus [DBR^{+}05, SMK^{+}09]. Marteilia [CAC^{+}08, CABF08, NPF05, SNOR09]. Marteilioides [ITK^{+}04]. Martens [Edg00, RJ02]. masou [ILAW03]. masquinongy [EFT^{+}06]. Mass [CKW^{+}01, HPMW08, Kha04a, VFT^{+}03, WHD^{+}08]. Mast [DG08]. mathematical [RRG^{+}05]. matrix [SMK^{+}09]. maturation [ME08]. maurini [CAC^{+}08]. maxima [CTB^{+}06, LVDBB07, NLO08, SPC^{+}08, WZJ^{+}08, XCD09]. maximus [AHTM06, BGB^{+}04, BGBG05, Bow03, CLVO^{+}05, JST^{+}04, LLZ08, OKK^{+}06, PZD^{+}06, SGMP^{+}05, SHM00, TFN01, VFT^{+}03, VFAN03]. may [XKS05]. MBV [SBSH08]. meal [BMPB^{+}00, MDB^{+}09, USR^{+}09]. meal-induced [BMPB^{+}00]. measure [BLGR09]. measurement [DL02]. measures [And06]. mechanism [And06, Jay09, KLG^{+}06, MHKA08]. mechanisms [MEBD05, PGC^{+}09]. media [BPV^{+}03, LG03]. mediated [AL09, AW02, EMS05, FNE09, GEC09, GKVS04, PBW06, SKIS05, SSK^{+}08, SDD^{+}00, XK02]. medicated [CRM^{+}03]. Mediterranean [ABCRH09]. megalocytivirus [OKK^{+}06]. megalocytiviruses [WCK^{+}09]. Mekong [CDT^{+}02, FTS^{+}01]. melanin [KHH^{+}05]. Melano [AR03]. Melano-macrophage [AR03]. Melanogrammus [TBLC07]. Melanomacrophages [BKSB09]. melanostigma [CWT^{+}00]. members [CLN05, KRKS06]. membrane [ASEK03, FW01, LNZL08, XNS04]. membrane-associated [XNS04]. Meserve [MPEN05]. Mesnil [PB04]. mesophilic [Koz07]. mesothelioma [BHB03]. metabolic [BMKG^{+}07, TF04]. metallo [ZW00]. metallo-protease [ZW00]. metalloprotease [TH08]. metamorphosis [HGP^{+}07]. metaplasia [GRL04]. method [AL09, CDS^{+}00, Cvi04b, GEC09, IMAN01, JO09, KBEJ^{+}07, úMMUB08, MGE04, RSSD01, SMK^{+}09, WMA^{+}02]. methods [BMŠH^{+}02, Dig01, DLS^{+}07, GHBW03, HKK^{+}09a, MD08, SPL05, WDC^{+}03]. methylene [PDOZ08]. Mexico [DCTCB^{+}06, MGGSREAG08]. MHC [XCD09]. Miamiensis [SSO^{+}09]. mice [MOR^{+}09]. Michigan [EFT^{+}06]. micro [BF02a, Cvi04b]. micro-organisms [BF02a, Cvi04b]. Microbial [GPS^{+}08, MDB^{+}09]. microcystin [SGF02]. microcystin-LR [SGF02]. Microcystis [ZKG^{+}01]. Microfauna [BM06]. Microhabitat [HMB07]. Micromesistius [ACP01]. microorganism [WP00]. microorganisms [CRBM05, CRA^{+}05]. Micropterus [GPN05]. microscope [CZP01]. microscopic [Cvi04b]. microscopy [MIS07, MA04]. Microspora [BSD06, LYO04, RSSD01, RSD04, SSM00, SDD^{+}00]. microsporidia [MMB02, KBS03]. microsporidial [BSDD02, BSD05]. microsporidian [BHO05, FBS03, GMS07, MSD^{+}00]. microvilli [MDB^{+}09]. migration [DG08]. Miichthys [PWS^{+}08]. miiuy [PWS^{+}08]. Mikailov [BNN03, OLB03]. Milne [BWS^{+}08, DSBB05, Dov05, Jay09, ZB07]. Milne-Edwards [BWS^{+}08, DSBB05, Dov05]. minerals [GWF^{+}02]. minibicornis [SHBB^{+}02, TPF09]. minnow [DMG09]. minnows [RFBD^{+}05]. minor [áKBH06]. Mitchell [BBG^{+}05]. Mitchill [BF02a, BAMG^{+}05, EFT^{+}06]. mitten [ZSZB04]. mixed [CM04, ZWZS02]. mixtures [BVG^{+}03]. Mo [áKBH06]. Mode [BC06]. model [AHTM06, BNS03, BSD06, BSE02, HBK07, LYO04, RRG^{+}05]. Modelling [AM09, DGLS09, OB07]. models [ATNF02a, BSM^{+}06, RRKM07]. Modes [LYO04]. modified [BMKG^{+}07, SSH^{+}07]. Modulation [DW05b]. Molecular [AMI^{+}07, BSTW08, HAEM02, LHP^{+}03, MMA02b, MWG^{+}02, NHA05, OFKEA08, PWW^{+}06, RMSH09, SME^{+}01, TXZ^{+}09, YTLN01, ABCRH09, BMŠH^{+}02, DNC^{+}05, EBASW^{+}08, FBS03, FRW^{+}08, HPMW08, HCJ08, HBCS^{+}06, MBLK06, MES^{+}02, TWOM06]. molluscs [LB00]. monensin [BSDD02]. Monitoring [JNS^{+}09]. monoclonal [CPRL09, HFW01, JAM^{+}01, JTM^{+}01, LCM^{+}01, LLWW02, SWTC01]. monodon [SBSH08, KHA04b, MO06, SBSH08, TSL^{+}01, YNH03]. Monogenea [ARKK08, BCB02, EWCT05, áKBH06, SGS01, SYYO06, WCT^{+}01, YYY^{+}06]. monogenean [MPEN05]. Moreau [LNM^{+}09]. morhua [BFR07, FCM^{+}04, GUE06, GKOF09, GBGB06, Kha04a, KKD^{+}09, MBG^{+}02, MGL^{+}06, MOR^{+}09, OMR^{+}06, ONI^{+}08, SBE03, SBM^{+}06, SWS07, SKGR05, UBB^{+}09]. Moritella [BBLC00, BGB^{+}04, GBGB06]. Morone [BWCM04, MOG09, TWOM06]. moronecidins [SXSN07]. Morphogenesis [MIS07, EBASW^{+}08]. morpholino [AST^{+}05]. Morphological [ABCRH09, HOA09, LWWS04, MES^{+}02, SHS05, TJG04b, DNC^{+}05]. morphology [BD00, CIK^{+}04, EBASW^{+}08, HBCS^{+}06, MMBA05, MDB^{+}09, PNA02, USR^{+}09, WCT^{+}01]. Morphometric [SGS01]. morphotypes [GDG00, GDG01]. mortalities [BLSE03, HD05b, JO00, NMJ^{+}04, VFT^{+}03, ZB07]. Mortality [HD05a, CKW^{+}01, DBR04, GGCFB09, GCR^{+}02, GBB^{+}06, Kha04a, LHW08, NFR08, SGF02, SI06, WHD^{+}08, WCJJ08, YNH03, ZKG^{+}01]. motile [FZA06]. Mountain [YYY^{+}06]. movement [LBO^{+}01, MG09]. Mr [BSQW05, PBW06, RB03, SAH^{+}07]. mrigal [SMS02]. mrigala [SMS02]. mRNA [BWD^{+}07]. mucosa [BMPB^{+}00]. mucosal [GLC04, LLJ^{+}02, SLC08]. mucous [BBJ04]. mucus [CRBM05, SN08, TWOM06, TH08, VSNS08, VLK^{+}04]. mud [FHY^{+}08, WGS^{+}07]. Mugil [CLS^{+}02, GKNLN^{+}04]. mullet [CLS^{+}02, EKG^{+}02, GKNLN^{+}04]. Mulsow [JD02]. multifiliis [RF01, SB01, SLB04, XK02, XKS02, XK03, XKS04, XKS05, XKS06, XKP06]. multiple [BNFA07]. Multiplex [CL05, AOA^{+}09, YSH^{+}06, YWBH05]. multiplication [DDL^{+}01, HFW01]. Multivalvulida [GDHM03]. multivariate [GWF^{+}02]. mummichog [GMB^{+}07]. Munday [CMD01]. murmanica [BCB02, Kha04a]. Murray [BBG^{+}05]. Muscle [MMB02, RJB00, ATT^{+}09, CBDC^{+}03, CCL09, DGMD^{+}09, FKT^{+}05, KKT^{+}04, KHTF06, KT09, KHH^{+}05, QSZ^{+}03, RST^{+}07, SN00]. muscular [BB07]. mushroom [KMJM06]. muskellunge [EFT^{+}06]. mussel [KWKLG09]. Mustelus [BF02a]. mutant [ASEK03, TRY^{+}05]. myc [BSTW08]. mycobacteria [BMD^{+}06, PTT^{+}08, PTAR09]. mycobacterial [HBK07, RWSK09]. mycobacteriosis [JSBK09]. Mycobacterium [CTAR01, HBK07, KRKS06, KWD^{+}01, PWW^{+}06, PTKA02, WWK03, ZFF^{+}08]. Mycobacterium-like [KWD^{+}01]. mycosis [Dig01]. Myers [MMB02]. mykiss [MBD^{+}02, RJC04, AJ09, ABC^{+}08, AJA03, BWMSB02, BM09, BSDD02, BSD05, BSD06, BLL^{+}07, BML^{+}06, BSG^{+}05, BA05, BNFA07, BNN03, BBJ04, CGP08, CTAR01, CDL^{+}08, CMD02, CDGF07, DM00, DL01, DGF02, DDL^{+}01, DGLS09, FPO^{+}09, FZA06, GVGR00, GRW^{+}03, GDG01, GMS06, GMS07, HPMW08, IA02b, IA03, JTR^{+}04, JO09, JOG^{+}06, JSG^{+}08, KHSW09, LLJ^{+}02, LLJC03, LBO^{+}01, LKH08, LYO04, LLHF02, MLB^{+}06, MMD05, MDB^{+}09, MSD^{+}00, NA09a, NA09b, OBAA09, OBHF00, OLR03, OL06, PZS02, PSLN^{+}08, PSLN^{+}09, PLC09, PBIM09, PT09, PSB06, RLNB03, RSSD01, RSD04, RF01, RSFS00, SSM00, SLB04, SHN^{+}00, SFT^{+}08, STV05, SBV09, SLC08, TOB^{+}07, TLC^{+}02, TTD^{+}06, VTA06, YYY^{+}06, YWS^{+}08]. Myocardial [TTD^{+}06, FKT^{+}05, PTB07]. myoliquefactive [LJM08]. myonecrosis [AL09]. myopathy [BB07, PC04]. myositis [OBHF00]. Mytilus [CABF08, DBR04]. Myxobolid [IPA^{+}01]. Myxobolidae [ESMB00, MCS08, MEM^{+}09, SMR01]. Myxobolus [AAH^{+}07, CM04, EMS05, ESMB00, HCJ08, IPA^{+}01, KSKEM07, LLWW02, MES^{+}02, MCS08, MEM^{+}09, SGEM03, SMR01, ZWW^{+}06]. Myxosporea [ESMB00, LYW03, LFF05, LLWW02, SMR01, WYWL01, YIT05]. myxosporeans [TOW02]. myxosporidean [CKW^{+}01]. myxosporidiosis [FPG^{+}02, PFL^{+}08]. Myxozoa [AHB07, CM04, HAEM02, KEEM^{+}05, KSKEM07, LYW03, LFF05, LLWW02, MMBA05, MCS08, MEM^{+}09, SBDPAP07, YSN^{+}06, YIT05]. myxozoan [ABCRH09, LFF05]. myxozoanosis [GAH^{+}08].

n [DNC^{+}05, FYO08, GAH^{+}08, HBCS^{+}06, MEM^{+}09, YIT05]. Nant [EWBB02]. nase [IPA^{+}01]. native [PWW^{+}06]. Natural [Rob06, ATT^{+}09, CDGF07, KHTF06, LWWS04, NFR08, RBBC^{+}09, RDT^{+}03, RAG^{+}07, SSM01]. Naturally [BWS^{+}08, BF02a, BSG^{+}05, DADB00, DGS^{+}02, DG08, FTE^{+}03, MBG^{+}02, MNM^{+}02, TBLC07]. naturally- [TBLC07]. naturally-occurring [FTE^{+}03]. NCIMB [BBLC00]. near [EWBB02, SSH^{+}07]. near-isogenic [SSH^{+}07]. Nebrius [PP02]. necator [CN02, CLN05]. necrosis [AST^{+}05, ACS^{+}08, AJ09, BGR08, BMŠH^{+}02, BWMSB02, BM09, BSE02, CYL^{+}06, CLL01, CDI^{+}01, DZP^{+}00, FTS^{+}01, GUE06, GSMM09, GMW^{+}07, GBB^{+}06, HLWH07, INM^{+}01, JO09, KKD^{+}09, LBO^{+}01, LMC^{+}01, LCM^{+}01, MASS07, MBB^{+}06, MBLK06, MGAE06, OL06, PTB07, PGC^{+}09, RST^{+}07, RP05, RBBC^{+}09, RAG^{+}07, RMSH09, SBB^{+}06, SM08, SN00, SOF^{+}09, TMA^{+}01, TKNM03, TTM04, UMG^{+}08, WGM^{+}08, YMN09, YSH^{+}06]. Necrotic [OBHF00, JPPS06]. need [NEG06]. negative [AJA03, FH03]. negatives [MMA02a]. Negligible [LBO^{+}01]. nematode [BF02b, DMG09]. Neoheterobothrium [SYYO06]. neon [MMB02]. Neoparamoeba [AN04a, BN04b, CAAN07, DHDCN02, DHNB05, DNC^{+}05, GMBN04, GCN04, GPB^{+}05, LVW^{+}06, MZF01, VLT^{+}07, YDNM08]. neopilchardus [WCJJ08]. Neoplasms [TMM01, BHRB04, LM03]. neoplastic [BSTW08]. Nephrops [RST^{+}07, SN00]. neptunius-like [KWKLG09]. nerka [CRM^{+}03, PGC^{+}09, SHBB^{+}02, TF04]. nervous [CYL^{+}06, CLL01, CDI^{+}01, DZP^{+}00, GDHM03, HLWH07, INM^{+}01, KKD^{+}09, LMC^{+}01, LCM^{+}01, TMA^{+}01, TKNM03, TTM04, YMN09]. nested [CGMS03, DZP^{+}00]. net [PSB06]. network [MG09]. neurofibrosarcoma [JDBN08]. neuromodulators [BSG^{+}05, DADB00]. neurophila [GDHM03]. neurophilia [KBS03]. neutralization [GJR^{+}03, GJR^{+}05, LCM^{+}01]. Neutralizing [YMKN09]. newly [LCC^{+}08, LWWS04]. NF [HGH^{+}08]. NF- [HGH^{+}08]. Nile [EKS09, XSK07]. niloticus [BCB^{+}09, EKS09, GLC04, SKSE02, SHFM09, XSK07]. nippona [ITK^{+}04]. Niteroi [LSM^{+}02]. Nitzsch [DPS^{+}07]. NNV [KKD^{+}09]. no [BNN03, GSM^{+}04, YNH03]. noatunensis [MOR^{+}09]. nobilis [CDI^{+}01]. Nocardia [SNN^{+}08, WCT^{+}09]. Nocardiosis [BCM^{+}00, CLL^{+}00, WYJ05, IKKS06]. Nodavirus [ATNF02b, CDI^{+}01, GKOF09, UCD04, ATNF02a, BSQW05, BPBT02, HWT^{+}06, HTC^{+}01, JST^{+}04, LMC^{+}01, PBW06, QSZ^{+}03, SIM^{+}00, SAH^{+}07, WCW^{+}08, WDC^{+}03, YWBH05]. nodavirus-susceptible [LMC^{+}01]. nodaviruses [IMAN01, LJL^{+}03]. Non [KFM^{+}00, MFLC01, VFAN03, AVP^{+}09, DLS^{+}07, FZA06, GME04, GPS^{+}08, GER^{+}07, HLWH07, KJB07, MGE04, MGAE06, ME08, PDOZ08, PWW^{+}06, SMK^{+}09, XK03]. non-coding [HLWH07]. non-destructive [GME04, MGE04, ME08]. non-enzymatic [SMK^{+}09]. non-immune [XK03]. non-inflation [PDOZ08]. non-lethal [DLS^{+}07]. non-motile [FZA06]. non-native [PWW^{+}06]. non-pathogenic [KJB07]. non-phagocytic [AVP^{+}09]. Non-pigment-producing [KFM^{+}00]. non-replicating [MGAE06]. non-spatial [GER^{+}07]. Non-specific [MFLC01, VFAN03, GPS^{+}08]. nonspecific [FM00]. Nordmann [ØSKH06, ØH07, RGTR02]. Normal [Hum07, SSM00, CIK^{+}04, GCWFB08]. north [BF02a, BF02b, CWW^{+}02, GPRCH09, PMZD08, CDI^{+}01, EFT^{+}06, KWD^{+}01, LJM08]. north-eastern [CWW^{+}02]. north-west [PMZD08]. north-western [BF02a, BF02b]. northern [BFR07, GCG^{+}04, JNS^{+}09, MMSM06, RA08, GJR^{+}03]. Northwest [CKK02, DCTCB^{+}06, GVGR00]. northwestern [BT09, BKSB09]. norvegicus [RST^{+}07, SN00]. Norway [HRG03, ONI^{+}08, RST^{+}07, SN00, TOB^{+}07]. Norwegian [SKR^{+}08]. notes [LWWS04]. Notropis [IGH00]. nov [BF02b, IPA^{+}01]. novel [BVG^{+}03, GKCF04, HBB09, ILAW03, OMR^{+}06, SGS01]. nsP3 [FRW^{+}08]. nucleic [AL09]. Nucleospora [FPO^{+}09]. Nucleotide [MGGSREAG08, FRW^{+}08, STM^{+}01]. nuesslini [KEEM^{+}05]. number [BWMSB02]. numbers [Cvi04b, DCCF^{+}03, HOM^{+}09]. nurse [PP02]. nutrition [BBC^{+}05]. NV [BSQW05, PBW06, RB03, SAH^{+}07]. NW [IPA^{+}01, NLO08].

O. [CDL^{+}08]. O1 [TRY^{+}05]. O3 [SRAHJ^{+}08]. Observation [GCR^{+}02, GCG^{+}04, WYWL01]. Observations [FLHM04, CZP01, FKT^{+}05, JTM^{+}01, RM07]. observed [Edg00, MIS07]. obtained [GDG01, JTR^{+}04, XNS04]. Occurrence [JHKB08, LJM08, ONI^{+}08, WYWL01, ZFF^{+}08, Cvi04a, DL01, KaVKT09, MMB02, RBBC^{+}09]. occurring [BWS^{+}08, BF02a, FTE^{+}03, MNM^{+}02]. Ocean [CWW^{+}02, KWD^{+}01]. ocellata [WWH^{+}02]. oestradiol [CHWK^{+}04]. oestroides [HO03]. officinale [NA09a]. oil [CIK^{+}04, SSL^{+}08b]. oil-adjuvanted [SSL^{+}08b]. oils [CIK^{+}04]. Olafsen [áKBH06]. old [BML^{+}06]. oligochaete [ESMB00, LYW03]. oligochaetes [EMS05]. oligodeoxynucleotides [BBN03]. oligomers [AST^{+}05]. olivaceus [HKL^{+}06, HKK^{+}09b, KOK^{+}03, KHK^{+}04, SYYO06]. olive [HKL^{+}06, HKK^{+}09b, KHK^{+}04, SYYO06]. olseni [ALAC09]. omega [RDT^{+}03]. omega-3 [RDT^{+}03]. Oncorhynchus [AJ09, ABC^{+}08, AJA03, BWMSB02, BM09, BSDD02, BSD05, BSD06, BLL^{+}07, BML^{+}06, BSG^{+}05, BA05, BNFA07, BNN03, BBJ04, CGP08, CTAR01, CDL^{+}08, CMD02, CRM^{+}03, CDGF07, DM00, DL01, DGF02, DDL^{+}01, DGLS09, FPO^{+}09, FF07, FZA06, GVGR00, GRW^{+}03, GDG01, GMS06, GMS07, HPMW08, IA02b, IA03, ILAW03, JTR^{+}04, JO09, JH01, JD02, JKD06, JKB08, JOG^{+}06, JSG^{+}08, KM01, KH06, KHSW09, LLJ^{+}02, LLJC03, LBO^{+}01, LKH08, LYO04, LLHF02, LWWS04, LCH^{+}02, LMG^{+}03, LM03, LRH^{+}08, MLB^{+}06, MMD05, MDB^{+}09, MBD^{+}02, MSD^{+}00, NA09a, NA09b, OR05, OB07, OBAA09, OBHF00, OLR03, OL06, PZS02, PSLN^{+}08, PSLN^{+}09, PLC09, PBIM09, PT09, POHJ05, PSB06, PGC^{+}09, RLNB03, RSSD01, RSD04, RJC04, RF01, RSFS00, SSM00, SLB04, SLC^{+}06, SHN^{+}00, SHBB^{+}02, SFT^{+}08, STV05, SBV09, SLC08, TOB^{+}07]. Oncorhynchus [TF04, TLC^{+}02, TTD^{+}06, TPF09, VTA06, YYY^{+}06, YWS^{+}08]. one [BML^{+}06, HAEM02, JGS^{+}04, SXSN07, YWBH05]. one-step [YWBH05]. one-year-old [BML^{+}06]. onset [SN00]. Ontario [GDR^{+}07, EEF06, WCB08]. Oplegnathus [JO00, KJL07, KJL08]. opportunistic [AVN08, LG03]. optimization [MEC^{+}06]. optimized [HSW03]. Oral [IRA03, FHY^{+}08, PWS^{+}08, RLNB03, RF01, SGMP^{+}05, SSS^{+}00, WBNG02]. Oreochromis [BCB^{+}09, EKS09, GAMR^{+}04, GLC04, HFI09, HLC00, SKSE02, SHFM09, XSK07]. organ [HSW03]. organism [CWT^{+}00, MAF^{+}00, MMA02b, OW02]. organisms [BF02a, Cvi04b, ZS04]. organosomatic [SKLY03]. organs [MEM^{+}09]. orientalis [SMN^{+}09]. ornamental [BMD^{+}06, HSK^{+}08, ZFF^{+}08]. ornamentation [GDG00]. osmoregulatory [LMG^{+}03, RSFS00]. Osteoma [PP02]. Osteomas [LSM^{+}02]. Ostrea [HDP^{+}02]. other [MOG09, SCJ^{+}02, ZS04, Rob05a]. out-membrane [XNS04]. Outbreak [BCB^{+}09, ATT^{+}09, EMH02, FKT^{+}05, GCR^{+}02, JGS^{+}04, KHTF06]. Outbreaks [CLW03, WHW^{+}02, WWH^{+}02, CVGG06, HD05a, Kha04a, MD08, RDT^{+}03, RAG^{+}07, SLC^{+}06]. outer [ASEK03, LNZL08]. ovarian [ITK^{+}04, SM08]. over-expressing [CYL^{+}06]. overload [RP04]. oxidized [DRA00]. oxolinic [RAT^{+}02, SBE03]. oxygen [HFW01, SCJ^{+}02]. oxyrhinchus [BT09, BKSB09]. oxytetracycline [RNAA04, TL09]. oyster [ITK^{+}04, WP00]. oysters [HDP^{+}02, SNOR09, WHD^{+}08].

p [BSTW08]. p-53 [BSTW08]. p53 [BWD^{+}07]. Pacific [KWD^{+}01, Bro09, CLQAGCRP08, CKK02, CWW^{+}02, HGP^{+}07, JD02, JKD06, RW03, SMN^{+}09, WHD^{+}08]. padlock [MBLK06]. PAGE [SGEM03, AN04a, DHCHN01, DHDCN02, DNC^{+}05]. Pagellus [ACGR^{+}05]. pagri [YIT05]. Pagrus [LVA^{+}06, WSKC03, YIT05, ZWZS02]. pagurus [CCB03]. Palaemon [DBR^{+}05]. palaemonid [DBR04]. Pallas [DADB00, FLGS00, YWS^{+}08]. pallasi [JD02]. pallasii [HGP^{+}07]. pallipes [EWBB02]. Pampus [DAMQ^{+}04]. Pan [RA08]. Pan-steatitis [RA08]. Pancreas [TOB^{+}07, DMG09, MNM^{+}02, NFR08, RM07]. Pancreatic [BF02b, BGR08, BMŠH^{+}02, BWMSB02, BM09, BSE02, GUE06, GMW^{+}07, GBB^{+}06, JO09, MASS07, MGAE06, RP05, RBBC^{+}09, RAG^{+}07, RMSH09, SBB^{+}06, SM08, SOF^{+}09, UMG^{+}08, WGM^{+}08]. Pancreatonema [BF02b]. Pangasius [CDT^{+}02, FTS^{+}01]. panophthalmitis [VHM^{+}00]. papillomas [ONS07]. papillomatosis [KaVT06, KaVKT09]. Paracartia [CAC^{+}08]. Paracheirodon [MMB02]. paraffin [KMS^{+}08, PTAR09]. paraffin-embedded [KMS^{+}08, PTAR09]. Paralichthys [HKL^{+}06, HKK^{+}09b, KOK^{+}03, KHK^{+}04, SYYO06]. parameters [CVGG06, GAMR^{+}04, SGS01]. Paramoeba [DHCHN01, ZM01b, ZGM01, ZM01a]. paramoebae [DHOM^{+}03]. Paramyxea [CABF08]. parasite [AW02, BCB02, BHO05, CKK02, DBJN05, EWCT05, ITK^{+}04, LZ09, SPC^{+}08, SWS07]. Parasites [Rob06, BBA08, IPA^{+}01, Ken07, LFF05]. parasitic [CMD02, HSD07, LGF04, LB09, SLB04]. parasitising [SGS01]. parasitizing [BCB02, KPWK06]. parC [RMHA08]. parent [ASEK03]. parental [SSH^{+}07]. parr [SBK^{+}05, SRT^{+}07]. Partial [HHCBR^{+}07, FRW^{+}08, LMJ^{+}00]. partially [HGP^{+}07]. particles [DGMD^{+}09, GCR^{+}02, GCG^{+}04, QSZ^{+}03]. Parvicapsula [SHBB^{+}02, TPF09]. Passive [LLJC03, SKSE02]. Pasteurellosis [FTE^{+}03]. Pathogen [BSW03, BHP^{+}04, CRA^{+}05, FLA06, FH03, HHL^{+}05, LG03, MGGSREAG08, PSLN^{+}09, SHFM09, Sta01, WDC^{+}03]. pathogen-induced [HHL^{+}05]. Pathogenesis [SCJ^{+}01, EBASW^{+}08]. Pathogenic [YYY^{+}06, AN04a, AW02, BDSY07, CCTL06, CL05, CDT^{+}07, HFW01, JPPS06, KJB07, Ken07, Koz07, LHP^{+}03, MHKA08, PBW06, SHS05, STV06, UBB^{+}09, VFT^{+}03, VFAN03, VLK^{+}04, ZW00]. Pathogenicity [HKL^{+}06, ZA00, GDG01, KFM^{+}00, LFF05, MOR^{+}09, MWG^{+}02, OBAA09, PRG^{+}05, RP04, VSNS08]. pathogens [BWCM04, DM00, HBB09, KWKLG09, LB00, úMMUB08, RJ02, SKIS05]. Pathological [CMD01, TOW02, BWD^{+}07, DPS^{+}07, JGS^{+}04, PGB^{+}06]. Pathology [BKNS02, EN03, HSD07, Hum07, NSK00, PZD^{+}06, WHD^{+}08, ZM00, AN03, AEN04, AN04b, AR03, BBG^{+}05, BB01, GBGB06, LZ09, LBN02, PBIM09, RST^{+}07, SPK^{+}05, SWS07, SHR04, WCB08, Ano02b]. pathophysiology [MPEN05]. pathway [PB04]. pattern [ALAC09]. patterns [HD05a, HRG03, LGR08]. PCNA [BSTW08]. PCR [AOA^{+}09, CGMS03, CL05, CLVO^{+}05, DZP^{+}00, DL02, GSM^{+}04, HWT^{+}06, HTC^{+}01, IMAN01, IFA05, IYS^{+}07, IOK^{+}07, JLSA08, KJB07, KMS^{+}08, KJC^{+}02, KBEJ^{+}07, MAF^{+}00, MMA02a, OLP01, PGH^{+}07, POHJ05, PTKA02, RLL^{+}07, RA01, SMM09, TYH^{+}01, TPF09, WCSC07, WCK^{+}09, WDC^{+}03, XMM^{+}09, YSH^{+}06, YWBH05, YDNM08]. PCR-based [CLVO^{+}05, GSM^{+}04, IFA05]. PCR-RFLP [KJB07, IOK^{+}07]. pediculata [WCB08]. pedigreed [GBB^{+}06]. peelii [BBG^{+}05, BBG^{+}05]. pelagius [VFT^{+}03, VFAN03]. pellets [CRM^{+}03]. pemaquidensis [DHCHN01, DHDCN02, DHNB05, LVW^{+}06, MZF01]. penaeid [HBB09, Jay09, YSH^{+}06]. Penaeus [CLQAGCRP08, Jay09, KHA04b, MBB^{+}06, MO06, RJB00, RBA^{+}03, SBSH08, TSL^{+}01, WM04, YNH03]. peninsula [JO00, CDT^{+}07]. Pentacapsula [GDHM03]. peptide [AST^{+}05, DBR04, PZS02, SXSN07]. peptide-conjugated [AST^{+}05]. Perca [AJ09, BAMG^{+}05, GGCFB09, GCWFB08, KPWK06, LHW08, WBP^{+}05]. perch [AJ09, BAMG^{+}05, GGCFB09, GCWFB08, KPWK06, LHW08, WBP^{+}05]. Percidae [CM04]. Percina [CM04]. peregrinum [HBK07]. performance [GWF^{+}02, SSL^{+}08b, TF04]. periplasmic [BAS^{+}09]. perkinsosis [ALAC09]. Perkinsus [ALAC09]. permanganate [DMS09]. Perna [KWKLG09]. peroxide [DRCMB03, MPEN05]. persistence [DGF02]. persists [HKK^{+}09b]. perurans [YDNM08]. pest [Bro09]. Peters [LAZ05, LZ09]. Petromyzon [EEF06]. pflugfelderi [ABCRH09]. pH [GSW^{+}07, SI06]. phage [ZWW^{+}06]. phagocyte [GMBN04]. phagocytes [DDL^{+}01, DRCA^{+}08]. phagocytic [AVP^{+}09, Jay09]. phagocytosis [DDL^{+}01, PZS02]. pharmacokinetic [RTNA02, SBE03]. Pharmacokinetics [RAT^{+}02, FHY^{+}08]. pharmacological [RF01]. Pharmacology [Rob00]. phenol [GTBA09]. phenotypes [BMPB^{+}00, HAEM02]. Phenotypic [BFR07, MHHK^{+}01, VAH09, SNN^{+}08]. Philasterides [CPB^{+}07, SPC^{+}08]. Philippi [HDP^{+}02]. philippinarum [DGMD^{+}09, RLA01]. Philometridae [Wan02]. Philometroides [Wan02]. philomiragia [MOR^{+}09]. phocae [VJR^{+}09]. phosphatase [WMA^{+}02]. phosphatase-based [WMA^{+}02]. phosphate [DW05b]. phospholipase [NHA07]. phosphorus [SHR04]. Photobacterium [AVP^{+}09, AOA^{+}09, ARC^{+}05, BPV^{+}03, BVG^{+}03, BHP^{+}04, BAS^{+}09, CRBM05, DRCMB03, DRCA^{+}06, JTM^{+}01, KHA05, LVA^{+}06, MMAP00, MMTPDR^{+}08, NHA05, NHA07, PSLN^{+}09, RP04, TO02]. Phoxinus [DMG09]. Phylactolaemata [MMBA05]. Phyllopteryx [GAH^{+}08]. phyllopteryxa [GAH^{+}08]. Phylogenetic [FRW^{+}08, HSK^{+}08, FBS03]. phylogeny [FYO08, HBCS^{+}06, SKR^{+}08]. physical [BSM^{+}06]. physiological [BMS^{+}00, LAZ05, SHS05]. physiology [FF07]. pigment [KFM^{+}00]. pike [GCG^{+}04]. pilchard [WCJJ08]. pilot [DHOM^{+}03]. Pimephales [RFBD^{+}05]. pink [JKB08]. pintado [EKS09]. pipefish [LGF04]. piper [MIGR09]. Pisces [LFF05]. piscicida [AVP^{+}09, AOA^{+}09, ARC^{+}05, BPV^{+}03, BVG^{+}03, BHP^{+}04, BAS^{+}09, CRBM05, DRCMB03, DRCA^{+}06, JTM^{+}01, KHA05, MMAP00, MMTPDR^{+}08, NHA05, NHA07, ONI^{+}08, RP04, TO02]. piscidin [CUHN08]. piscidins [SXSN07]. piscine [IMAN01, ZW00]. Piscirickettsia [CWT^{+}00, FH03, JAM^{+}01, KMS^{+}08, VCLS04]. PKD [DL01]. PKX [MAF^{+}00, OW02]. plaice [MFLC01, TH08]. Plasma [SFPF01, YCR^{+}09, BSW03, PSB06]. plasmid [MD00]. plasmids [NLO08]. plasmodia [DCCF^{+}03]. plastic [MGE04]. platessa [TH08]. platessoides [MFLC01]. play [WM04]. plecoglossi [LYO04]. plecoglossicida [IYS^{+}07]. Plecoglossus [KIIK04]. Plectropomus [BB01]. Plehn [JD02]. Pleistophora [RFBD^{+}05]. plerocercoids [DPS^{+}07]. pleurocidins [SXSN07]. Pleuronectes [TH08]. pleuronecti [BCB02]. plus [DZP^{+}00]. Poecilia [LAZ05, LZ09]. poeciliids [PWW^{+}06]. Poland [Koz07]. polyclonal [LMX^{+}05, MEC^{+}06, YMNH04]. polylepis [IPA^{+}01]. polymerase [APYH00, DCTCB^{+}06, FPO^{+}09, KL07, KBH^{+}08, PTKA02, TSL^{+}01]. Polymicrobic [BWCM04]. Polymorphisms [NPF05]. polypeptide [CUHN08]. polyvalent [SSL^{+}08b]. polyvinyl [HPG08]. pomfret [DAMQ^{+}04]. Pomphorhynchus [DGS^{+}02]. pond [CDS^{+}00]. pond-side [CDS^{+}00]. ponds [TJG04a]. population [BLSE03, UMIK09]. populations [CGP08, CWW^{+}02, GBB^{+}06, KaVT06, MDB^{+}09, RRG^{+}05, SHBJ^{+}09]. Porphyridium [DRCA^{+}08]. Portugal [NLO08]. Portuguese [SME^{+}01]. Portunidae [NSK00]. position [KPWK06]. positive [SMM09]. positively [PV08]. positives [MMA02a]. possible [Dig01]. Post [HO03, BLSE03, CCTL06, Jay09, MBB^{+}06, MGAE06, NFR08, RAG^{+}07, SBB^{+}06, SN00, UMG^{+}08, YNH03]. post-capture [SN00]. Post-haemorrhagic [HO03]. post-larvae [CCTL06, MBB^{+}06, YNH03]. post-phagocytic [Jay09]. post-smolt [RAG^{+}07]. post-smolts [BLSE03, MGAE06, NFR08, SBB^{+}06, UMG^{+}08]. postmortem [LJM08]. pot [GKCF04]. pot-belly [GKCF04]. potassium [DMS09, PC04]. potassium-deficient [PC04]. Potential [LKH08, TRY^{+}05, GCWFB08, HBK07, MD08, NDB^{+}04, RSSD01, SMLZ05, ZAC^{+}03]. potentiates [KS06]. poutassou [ACP01]. Practical [Rob05a, TJG04a]. prawn [BSQW05, HWT^{+}06, HHL^{+}05, KBH^{+}08, KRJO02, QSZ^{+}03, RB03, WDC^{+}03]. praziquantel [VSST09]. Pre [MBB^{+}06, ZM00]. Pre-exposure [MBB^{+}06]. pre-maintenance [ZM00]. preadult [BB01]. preclinical [SRT^{+}07]. predominance [MMB02]. Prefecture [INKS04]. preference [ØSKH06]. prelabelled [TFN01]. preliminary [DJW^{+}09]. preparation [OBHF00, SMK^{+}09]. Presence [FAM^{+}06, GCN04]. present [NEG06, RJ02]. Presumptive [DCTCB^{+}06]. prevalence [BLGR09, CGP08, KH06, MHD^{+}04, PFL^{+}08]. prevent [MMA02a]. prevention [SBV09]. prevents [DL01]. previously [GMBN04]. primary [AN04a, HBK07, MSD^{+}00]. primers [APYH00, IYS^{+}07]. Prionacae [BKSB09]. Prionace [BD00, BKNS02, BHB03, BHRB04, BSTW08, BT09]. pro [WHB^{+}03]. pro- [WHB^{+}03]. probiotic [BA05, CL02, IA03, VLK^{+}04]. Probiotics [IA02a, ABC^{+}08, BNFA07, IA02b]. Procambarus [SWZ^{+}05]. processed [LBO^{+}01]. processing [MASS07]. produce [ZRT04]. produced [GHBW03, PLC09, XK03]. producing [KFM^{+}00, WGM^{+}08]. product [CYL^{+}06]. Production [YMNH04, ZGM01, AVN08, CTAR01, JTM^{+}01, LYW03, LMG^{+}03, SPC^{+}08, SMN^{+}09]. products [CTAR01, FAM^{+}06, PEP^{+}05, VCN09, WKES08, ZAC^{+}03]. Professor [Aok08]. profiles [DGLS09, MD00, ZASK06]. progeny [SAH^{+}07]. programme [Bro09]. progression [KHSW09, ONS07, RST^{+}07]. project [BH02]. prokaryotic [WP00]. prolapse [TBB^{+}08]. Proliferation [HMO^{+}08, SHN^{+}00, VNSB00]. Proliferative [BH02, CMD02, FPG^{+}02, WKB^{+}02, DL01, EMH02, LLHF02, MAF^{+}00, MMA02a, MMA02b]. promelas [RFBD^{+}05]. promoting [BPV^{+}03]. pronounced [GKB05]. Propagation [LMC^{+}01]. properties [GSW^{+}07]. protease [CWJL07, CFA^{+}00, FAM^{+}06, HKV^{+}06, XNS04, ZW00]. protected [HGP^{+}07]. Protection [GLMG03, LMX^{+}05, YMN09, BGBG05, GMS07, MBB^{+}06, SPC^{+}08, XKS04, XKP06, YMKN09]. Protective [TMA^{+}01, XK03, FLGS00]. protects [GMS06, SAC^{+}09]. Protein [SNOR09, BAS^{+}09, CYL^{+}06, CAB^{+}07, HSK^{+}08, HTC^{+}01, HLWH07, KJL07, KL07, KJL08, LMX^{+}05, SFPF01, SAC^{+}09, SLC08, WCK^{+}09, XNS04, YMNH04, YKR^{+}02, ZASK06]. protein-based [HTC^{+}01]. protein-hapten [SLC08]. proteins [ASEK03, MGGSREAG08, PLC09, SMLZ05]. proteolytic [OBHF00]. Proteomic [LNZL08]. protozoan [CN02, ITK^{+}04, SPC^{+}08]. Province [TXZ^{+}09]. provoked [RSFS00]. provokes [GKOF09]. Psetta [CTB^{+}06, LVDBB07, NLO08, OKK^{+}06, SPC^{+}08, WZJ^{+}08, XCD09]. Pseudobagrus [Wan02]. Pseudocaranx [MOO^{+}09, MKW^{+}09]. pseudodispar [SMR01]. Pseudoloma [KBS03]. Pseudomonas [BFR07, FCM^{+}04, GGCFB09, IYS^{+}07]. Pseudophyllidea [SNNB01]. Pseudoplathystoma [EKS09]. Pseudopleuronectes [BCB02]. psychrophilum [AMI^{+}07, CDL^{+}08, CDGF07, CK08, DDL^{+}01, EN03, EEF06, FNE09, HPMW08, ILAW03, IFA05, IOK^{+}07, LLJ^{+}02, LLJC03, LHW08, MHHK^{+}01, MLB^{+}06, MD00, MMD05, MD08, MEBD05, NDPH03, OBHF00, OL06, PLC09, RLL^{+}07, VAH09, VTA06]. Pterophyllum [PDOZ08]. puffadder [HSD07]. puffer [TOW02, YSN^{+}06]. pulsed [AMI^{+}07]. pulsed-field [AMI^{+}07]. pumilus [HBB09]. punctatus [BNS03, BSW03, CUN09, DM09, DMS09, Goo06, GM07, NM02, SKLY03, SBG08, SGF02, SGG03, TJG04a, XK02, XKS02, XK03, XKS04, XKS05, XKP06, ZKG^{+}01]. Purification [TH08, ZSZB04]. purified [ZW00]. Putative [MPP^{+}03, RHS^{+}06]. Putz [LWWS04]. pyrrolidone [HPG08].

Qinling [YYY^{+}06]. qPCR [FPO^{+}09]. quadricarinatus [Edg00, RJ02, SHZ^{+}00]. quality [DCCF^{+}03, EMH02]. quantification [CWS02, SM08]. Quantitative [DL02, OL06, TSL^{+}01, FPO^{+}09, OLP01, POHJ05, TPF09]. Quebec [CM04]. quinolone [IOK^{+}07, KHA05, RMHA08]. quinolone-resistance [KHA05]. quinqueradiata [IKKS06, SKI^{+}06, WCT^{+}01]. quorum [TYH^{+}01, YML^{+}08]. quorum-sensing [YML^{+}08].

R. [Cvi04a, VSST09]. Rachycentron [CKW^{+}01]. radiological [SHRM07].

Rafinesque [BNS03, BSW03, BHRB04, BT09, BKSB09, CM04, CUN09, DM09, DMS09, Goo06, GM07, NM02, PGB^{+}06, RFBD^{+}05, SKLY03, SBG08, SGF02, SGG03, TJG04a, TBB^{+}08, WCB08, XK02, XKS02, XK03, XKS04, XKS05, XKP06, ZKG^{+}01]. rainbow [AJ09, ABC^{+}08, AJA03, BWMSB02, BM09, BSDD02, BSD05, BSD06, BLL^{+}07, BML^{+}06, BSG^{+}05, BA05, BNFA07, BNN03, BBJ04, CGP08, CTAR01, CDL^{+}08, CMD02, CDGF07, DM00, DRA00, DL01, DGF02, DDL^{+}01, DGLS09, FZA06, GVGR00, GRW^{+}03, GDG01, GMS06, GMS07, HPMW08, IA02b, IA03, JTR^{+}04, JO09, JOG^{+}06, JSG^{+}08, KHSW09, LLJ^{+}02, LLJC03, LBO^{+}01, LKH08, LYO04, LVW^{+}06, LLHF02, MLB^{+}06, MMD05, MDB^{+}09, MBD^{+}02, MSD^{+}00, NA09a, NA09b, OBAA09, OBHF00, OLR03, OL06, PZS02, PSLN^{+}08, PSLN^{+}09, PLC09, PBIM09, PT09, PSB06, RLNB03, RSSD01, RSD04, RJC04, RF01, RSFS00, SSM00, SB01, SLB04, SHN^{+}00, SFT^{+}08, STV05, SBV09, SLC08, TOB^{+}07, TLC^{+}02, TTD^{+}06, VTA06, YWS^{+}08]. raised [CDL^{+}08]. Rana [WHW^{+}02]. ranaviruses [PGH^{+}07]. ranched [BLT01]. random [RHS^{+}06]. range [GCWR07]. ranging [BSTW08, BT09, BKSB09]. ranked [TPF09]. Raphidascaris [KBPD04]. Rapid [MEC^{+}06, PBW06, TYH^{+}01, EMS05, FNE09, GJR^{+}03, GJR^{+}05, IMAN01, LVW^{+}06, SN00]. ras [BSTW08]. rate [BLT01, RSD04, TF04]. rates [BSD06, UMG^{+}08]. Rathbun [NSK00]. ratio [BSD05]. RCBH [PTKA02]. rDNA [FBS03, WWK03]. rDNA-based [FBS03]. re [GMBN04, ASEK03]. RE-33 [ASEK03]. re-infection [GMBN04]. reaction [APYH00, DCTCB^{+}06, FPO^{+}09, KBH^{+}08, PTKA02, TSL^{+}01]. reactivity [BMFL^{+}07]. Real [KMS^{+}08, OLP01, DL02, JLSA08, PGH^{+}07, SMM09]. Real-time [OLP01, JLSA08, PGH^{+}07, SMM09]. reared [BCB02, BCM^{+}00, DM00, DGLS09, GAMR^{+}04, RA01]. rearing [STV05]. reassortment [RBBC^{+}09]. recapture [BLT01]. receptor [NHA05, XKS06]. recirculating [GAMR^{+}04]. recirculation [MD08]. recognition [BPV^{+}03]. recombinant [HTC^{+}01, KJL08, PLC09, SMLZ05, YMNH04, YKR^{+}02]. record [CDI^{+}01]. recovered [CDL^{+}08]. Recovery [AJA03, TF04, CIK^{+}04]. red [CK08, HFI09, HLWH07, WSKC03, WWH^{+}02, YIT05]. red-spotted [HLWH07]. redbanded [LVA^{+}06]. redclaw [Edg00, RJ02]. Redescription [HCJ08, ABCRH09, BBA08]. redfin [AJ09]. Redfish [PGC^{+}09]. Reduced [RP03, KLG^{+}06, SOF^{+}09]. reduces [CN02, DBR04, GGCFB09]. reducing [CL02]. reduction [BSDD02]. reedi [WWK03]. refringens [CAC^{+}08, CABF08]. Regan [PTKA02]. regeneration [FKT^{+}05]. regime [DL01]. regimes [BBBW01, SRT^{+}07]. region [IYS^{+}07, NPF05, YYY^{+}06]. regions [HLWH07]. Regulation [YML^{+}08]. Reinboth [ABCRH09]. related [DNC^{+}05, FAM^{+}06, NPF05, RHS^{+}06, STV06, USR^{+}09, WZJ^{+}08]. relation [PNA02]. relationship [AEN04, BBLC00, BN04a, DCCF^{+}03]. relationships [CLN05, TF04]. Relative [RW03, LLJ^{+}02, WHB^{+}03]. release [MMA02b]. released [MMBA05]. relevant [SLB04]. remarks [HOA09]. Remodelling [CM04]. removal [JH01]. Renal [GAH^{+}08, FPG^{+}02, PFL^{+}08]. Renibacterium [CGP08, Cvi04a, GEC09, HKPKS06, JLSA08, POHJ05]. reovirus [WGS^{+}07, ZSZB04]. repeatability [NDH^{+}05b]. replicating [MGAE06]. replication [ATNF02b, HHCBR^{+}07]. report [BWCM04, GKNLN^{+}04, SFT^{+}08, YWS^{+}08]. reproducibility [NDH^{+}05b]. reproducible [BSE02]. reproduction [CHWK^{+}04]. Republic [KBPD04]. requires [HLWH07]. rerio [CHWK^{+}04, HBK07, KBS03, RWSK09, TL09]. research [NEG06]. resembling [FKT^{+}05]. reservoir [DHDCN02]. reservoirs [BBA08]. resistance [BBN03, BCMN05, DJW^{+}09, IOK^{+}07, JKB08, KHA05, KS06, LHHP07, MHKA08, OLB03, PWS^{+}08, RLNB03, RMHA08, RW03, SSM01, WM04, XCD09]. resistant [GMBN04, VMN06]. respect [WGM^{+}08]. respiratory [BCMN05, DRCA^{+}08]. Response [BMFL^{+}07, Hum07, SWZ^{+}05, ALAC09, BSW03, BMS^{+}00, BCM^{+}02, CMD02, CGW04, Cvi04a, DMG09, GKB05, GLC04, GPB^{+}05, MFLC01, NFR08, SSH^{+}07, SMS02, SHBJ^{+}09, SLC08, VFAN03, XK02, YKR^{+}02]. Responses [ZM01a, AW02, ABC^{+}08, BMKG^{+}07, BNN03, CN04, LLJ^{+}02, PSP^{+}09, PWS^{+}08, XKS04]. responsible [NMJ^{+}04]. retained [BKNS02]. reticular [RM04]. reticulata [LAZ05, LZ09]. retinochoroiditis [GKOF09]. retinopathy [ATNF02a, ABPK03, JST^{+}04, JGS^{+}04]. retrieval [MMAP00]. return [TF04]. return-run [TF04]. reveals [EFT^{+}06, HAEM02, HBK07, RBBC^{+}09]. reverse [AL09, DCTCB^{+}06, PTKA02]. reverse-transcription [AL09]. Review [Ano02b, ABCRH09, And06, EBASW^{+}08, JSBK09, MG07, LB09, MKM02, SOM05, SHR04]. reviews [Rob02a, Rob02b, Bru00, Tac00]. revisited [MCN04]. RFLP [IOK^{+}07, KJB07]. rhabdomyosarcoma [IGH01]. Rhabdosargus [DW05b, DW05a]. rhabdovirus [LLZ08, SSL^{+}08b]. rhamnose [BHO05]. rhamnose-binding [BHO05]. Rhipidocotyle [VSST09]. rhodurus [ILAW03]. Ribosomal [CWW^{+}02, NPF05, WCT^{+}01]. ribotypes [MD00]. Rica [SHFM09]. rich [SI06]. Richardson [CLW03, DLS^{+}07, LG03, SCJ^{+}01, SCJ^{+}02, Wan02, WYJ05]. rifampicin [ASEK03]. rifampicin-mutant [ASEK03]. Rio [LSM^{+}02]. Risk [HD05b, BM04, GER^{+}07, LBO^{+}01, TWS09]. Risso [ACP01, WCT^{+}01]. River [CM04, TPF09, ARKK08, CGP08, FLHM04, GDG00, GDG01, HKKO07, ARKK08, KH06, KBPD04, SHBB^{+}02]. rivers [FPG^{+}02, JHKB08, PFL^{+}08]. RNA [INM^{+}01, IGH00, KL07, NPF05, STM^{+}01]. RNA2 [HLWH07]. roach [KaVT06, KaVKT09]. Roberts [Ano02b]. rock [Dig01, KJL07, KJL08, SNOR09]. rockfish [KWD^{+}01, KL07, WWK03]. rogercresseyi [BEL09]. rohita [ACS^{+}09, GPS^{+}08]. rohu [ACS^{+}09]. role [AN04a, AR03, WM04, ZAC^{+}03]. Ronald [Ano02b]. roni [ZB07]. roni-like [ZB07]. roretzi [HOA09]. Roscoe [NA09a]. rosenbergii [BSQW05, HWT^{+}06, HHL^{+}05, JPPS06, KRJO02, PNS^{+}05, PBW06, QSZ^{+}03, RB03, SAH^{+}07, WCSC07, WCW^{+}08, WDC^{+}03, YWBH05]. rotundus [LLWW02, ZWW^{+}06]. route [HBK07, HSW03]. routes [DBR^{+}05, ZM01a]. routine [BMŠH^{+}02]. rpoS [KBH^{+}08]. rRNA [RLL^{+}07]. RT [DZP^{+}00, HWT^{+}06, HTC^{+}01, IMAN01, KBEJ^{+}07, WCSC07, WDC^{+}03, YWBH05]. RT-PCR [DZP^{+}00, HWT^{+}06, KBEJ^{+}07, WDC^{+}03, YWBH05, WCSC07]. RTG [TFN01]. RTG-2 [TFN01]. RTgill [LVW^{+}06]. RTgill-W1 [LVW^{+}06]. rubripes [TOW02, YSN^{+}06]. ruckeri [BNN03, FZA06, RLNB03, SME^{+}01, TYH^{+}01, TRY^{+}05, TDH^{+}07, YTLN01]. rudd [MEM^{+}09]. Ruditapes [DGMD^{+}09, RLA01]. rugulosa [WHW^{+}02]. run [TF04]. Rüppell [TL09]. Rutilus [KaVT06, KaVKT09].

s [ZRT04, MGB^{+}07, RB03]. S-100 [MGB^{+}07]. S-ELISA [RB03]. S. [WCT^{+}01]. s.l. [MCS08]. S0 [ATT^{+}09]. sac [BWD^{+}07, LBN02]. Saccostrea [SNOR09]. sacculitis [FF07, LCH^{+}02]. SAF [BOB^{+}06]. SAF-1 [BOB^{+}06]. sagax [WCJJ08]. salar [AN01, AN03, AN04a, AEN04, AN04b, AVN08, ATT^{+}09, AW02, AGM^{+}08, BLGR09, BMPB^{+}00, BMKG^{+}07, BMFL^{+}07, BM04, BM06, BBFB07, BBBW01, BSE02, BMS^{+}00, BCM^{+}00, BBC^{+}05, BBN03, BCMN05, BBK04, BLT01, CBDC^{+}03, CGW04, CAAN07, DCCF^{+}03, FKT^{+}05, FZM00, FM00, FBP09, FTE^{+}03, GWF^{+}02, GCR^{+}02, GJR^{+}05, GMS^{+}09, GMBN04, GCN04, GPB^{+}05, GER^{+}06, GER^{+}07, GBB^{+}06, HRG03, HSW03, JTR^{+}04, JHKB08, KKT^{+}04, KHTF06, KT09, KFM^{+}00, KHH^{+}05, LHHP07, LGR08, LSDB06, LBN02, MHD^{+}04, MNM^{+}02, MN08, MGE04, MGAE06, ME08, NFR08, NDB^{+}04, OLB03, ØGW04, PMZD08, PSS09, PTB07, PNA02, RM08, RGTR02, RRG^{+}05, RMG^{+}09, RHS01, RP03, RP05, Rod07, RM07, RDT^{+}03, RAG^{+}07, SSH^{+}07, SNNB01, SBK^{+}05, SPL05, SPK^{+}05, SSL^{+}08a, SBM^{+}00, SFPF01, SSL^{+}08b, SBB^{+}06, SM08, SRM^{+}03, SKR^{+}08, SSS^{+}00, SBSR02]. salar [SHS05, SGRM07, SRT^{+}07, SHRM07, SOF^{+}09, TOB^{+}07, TSW00, USR^{+}09, UMG^{+}08, UBB^{+}09, VJR^{+}09, VHM^{+}00, VCN09, VMN06, WHB^{+}03, WGM^{+}08, ZM00, ZM01a]. salaris [ARKK08, HMB07, KJB07, LSDB06, SGS01]. salina [CR02]. saline [PC04]. salinities [AG01]. salinity [EWCT05, TSW00]. Salmincola [JH01, RJC04]. Salmo [AN01, AN03, AN04a, AEN04, AN04b, AVN08, ATT^{+}09, AW02, AGM^{+}08, BLGR09, BMPB^{+}00, BMKG^{+}07, BMFL^{+}07, BM04, BM06, BBFB07, BBBW01, BSE02, BMS^{+}00, BCM^{+}00, BBC^{+}05, BBN03, BCMN05, BBK04, BLT01, CBDC^{+}03, CGW04, CAAN07, DCCF^{+}03, DADB00, DGA^{+}03, FKT^{+}05, FZM00, FM00, FBP09, FTE^{+}03, GWF^{+}02, GCR^{+}02, GJR^{+}05, GDG00, GDG01, GMS^{+}09, GMBN04, GCN04, GPB^{+}05, GER^{+}06, GER^{+}07, GBB^{+}06, HRG03, HSW03, JTR^{+}04, JHKB08, KKT^{+}04, KHTF06, KT09, KFM^{+}00, KHH^{+}05, LHHP07, LGR08, LSDB06, LBN02, MHD^{+}04, MNM^{+}02, MALC00, MN08, MGE04, MGAE06, ME08, NFR08, NDB^{+}04, OLB03, ØGW04, PFL^{+}08, PMZD08, PSS09, PTB07, PNA02, RM08, RGTR02, RRG^{+}05, RMG^{+}09, RHS01, RP03, RP05, Rod07, RM07, RDT^{+}03, RAG^{+}07, SSH^{+}07, SNNB01, SBK^{+}05, SPL05, SPK^{+}05, SSL^{+}08a, SBM^{+}00, SFPF01]. Salmo [SSL^{+}08b, SBB^{+}06, SM08, SRM^{+}03, SKR^{+}08, SSS^{+}00, SBSR02, SHS05, SGRM07, SRT^{+}07, SHRM07, SOF^{+}09, TOB^{+}07, TSW00, USR^{+}09, UMG^{+}08, UBB^{+}09, VJR^{+}09, VHM^{+}00, VCN09, VMN06, WHB^{+}03, WGM^{+}08, ZM00, ZM01a]. salmoides [GPN05]. salmon [AN01, AN03, AN04a, AEN04, AN04b, AVN08, ATT^{+}09, ARKK08, AW02, AGM^{+}08, BLGR09, BMPB^{+}00, BMKG^{+}07, BMFL^{+}07, BM04, BM06, BLL^{+}07, BBLC00, BBFB07, BBBW01, BSE02, BMS^{+}00, BCM^{+}00, BBC^{+}05, BBN03, BCMN05, Bro09, BBK04, BLT01, BN04b, CBDC^{+}03, CDL^{+}08, CGW04, CRM^{+}03, CAAN07, DCCF^{+}03, FKT^{+}05, FZM00, FM00, FBP09, FF07, FTE^{+}03, FBS03, GPRCH09, GWF^{+}02, GCR^{+}02, GJR^{+}05, GMW^{+}07, GMS^{+}09, GMBN04, GCN04, GPB^{+}05, GER^{+}06, GER^{+}07, GEB^{+}08, GBB^{+}06, HD05a, HD05b, HRG03, HOM^{+}09, JTR^{+}04, JH01, JD02, JKD06, JKB08, JHKB08, KM01, KH06, KKT^{+}04, KHTF06, KT09, KFM^{+}00, KHH^{+}05, LHHP07, LGR08, LBGR08, LSDB06, LWWS04, LCH^{+}02, LMG^{+}03, LM03, LRH^{+}08, LBN02, MEC^{+}06, MHD^{+}04, MNM^{+}02, MBLK06, MN08, MGE04, MGAE06, ME08, NDH^{+}05b, NDH05a, NFR08, NDB^{+}04, OR05, OB07]. salmon [OLB03, ØGW04, PMZD08, PD09, PSS09, PTB07, PNA02, POHJ05, PGC^{+}09, RM08, RGTR02, RRG^{+}05, RMG^{+}09, RHS01, RP03, RP05, Rod07, RM07, RW03, RDT^{+}03, RAG^{+}07, SSH^{+}07, SKCD07, SNNB01, SBK^{+}05, SPL05, SPK^{+}05, SSL^{+}08a, SBM^{+}00, SFPF01, SSL^{+}08b, SBB^{+}06, SM08, SLC^{+}06, SRM^{+}03, SMM09, SHBB^{+}02, SN08, SKR^{+}08, SSS^{+}00, SBSR02, SHS05, SGRM07, SRT^{+}07, SHRM07, SOF^{+}09, TOB^{+}07, TF04, TPF09, TSW00, USR^{+}09, UMG^{+}08, UBB^{+}09, VJR^{+}09, VHM^{+}00, VCN09, VMN06, WHB^{+}03, WGM^{+}08, WMA^{+}02, ZM00, ZM01a]. salmonae [BS04, BSD05, BSD06, BHO05, GMS06, GMS07, LWWS04, MSD^{+}00, PSB06, RSSD01, RSD04, SSM00, SSM01, SDD^{+}00]. salmonicida [BGBG05, NLO08, BF07, GHBW03, IRA03, JOG^{+}06, JSG^{+}08, KFM^{+}00, MPP^{+}03, MBG^{+}02, NLO08, OR05, OL06, TBLC07, YCR^{+}09]. Salmonid [BN04a, DSJPP08, PSS09, BHO05, Cos09, FRW^{+}08, Goo06, GJR^{+}03, GJR^{+}05, GSW^{+}07, GRF^{+}07, GCWR07, GWJR08, JTR^{+}04, MG09, NDPH03, SGEM03, TDH^{+}07, VCLS04, Woo03, ZGM01]. salmonids [And06, EN03, FPG^{+}02, GEC09, HMB07, KSW03, KLG^{+}06, MG07, MMA02b, OLP01, PD09, RMSH09, SGS01, SRAHJ^{+}08, VAH09, ZA00]. salmoninarum [CGP08, Cvi04a, GEC09, HKPKS06, JLSA08, POHJ05]. salmonis [BLGR09, BMS^{+}00, BBK04, CWT^{+}00, FPO^{+}09, FBS03, FH03, GPRCH09, HRG03, JAM^{+}01, JKD06, JKB08, KMS^{+}08, LBGR08, LB09, MSD^{+}00, PMZD08, PD09, PSS09, RRG^{+}05, SKCD07, SSS^{+}00, SBSR02, TSW00, VCLS04]. salmonis-like [CWT^{+}00]. salmositica [AW01, AW02, CGW04, FW01, HFW01, Woo03, ZW00]. Salvelinus [MPP^{+}03, Sta01]. sampled [KBEJ^{+}07]. samples [AOA^{+}09]. sampling [CDS^{+}00, DLS^{+}07]. sand [BHRB04, TBB^{+}08]. sandbar [IGH00]. Sander [KPWK06]. sandtiger [PGB^{+}06]. sandwich [RB03]. Sanguinicolidae [CMD01]. sapidus [NSK00]. Saprolegnia [GDG00, GDG01, LG03, SHS05]. sarba [DW05a, DW05b]. sarcosine [LNZL08]. sarcosine-insoluble [LNZL08]. Sardinops [WCJJ08]. Sarig [She09]. sativum [NA09b]. Sauvage [CDT^{+}02, FTS^{+}01]. SAV [FRW^{+}08]. saxatilis [BWCM04, MOG09, TWOM06]. SBN [PDOZ08]. scalare [PDOZ08]. scale [KaVT06]. Scardinius [MEM^{+}09]. scavenger [MASS07]. scepticus [IGH00]. ScFv [ZWW^{+}06]. Schlegel [HKL^{+}06, HKK^{+}09b, IKKS06, JO00, KOK^{+}03, KHK^{+}04, KJL07, KL07, KJL08, KIIK04, LMJ^{+}00, LMC^{+}01, LCM^{+}01, LJL^{+}03, MWG^{+}02, SKI^{+}06, SYYO06, SMN^{+}09, TOW02, WSKC03, WCT^{+}01, YSN^{+}06, YIT05, ZWZS02]. schlegelii [KL07]. Schneider [IGH01, MOO^{+}09, MKW^{+}09]. Schröder [KEEM^{+}05]. Schuberg [KEEM^{+}05]. Schubert [ABCRH09]. Schultz [PDOZ08]. schwannoma [MGB^{+}07]. Sciaenops [WWH^{+}02]. scleroperca [CM04]. Scomber [LJM08]. scombrus [LJM08]. Scophthalmus [AHTM06, BGB^{+}04, BGBG05, Bow03, CLVO^{+}05, JST^{+}04, LLZ08, PZD^{+}06, SGMP^{+}05, SHM00, TFN01, VFT^{+}03, VFAN03]. Scotland [AJA03, BGR08, GMW^{+}07, HRG03, LGR08, MG09, PMZD08, PD09, RGTR02, RRG^{+}05, SBB^{+}06, SN00]. Scottish [AM09, BLGR09, FKT^{+}05, LBGR08, SBB^{+}06, SIM^{+}00, WGM^{+}08]. Screamer [RHS01]. screening [XNS04]. scuticociliate [SSO^{+}09]. Scuticociliatia [CPB^{+}07]. Scyliorhinus [GDP^{+}03]. Scylla [FHY^{+}08, WGS^{+}07]. SDS [SGEM03]. SDS-PAGE [SGEM03]. Sea [GPRCH09, AVP^{+}09, ACGR^{+}05, AM09, ATNF02a, BPV^{+}03, BVG^{+}03, BM06, BMS^{+}00, BEL09, BPBT02, Bro09, BLT01, CIK^{+}04, CRA^{+}05, CLL^{+}00, CVGG06, Cos09, CZP01, DW05a, DW05b, EEF06, GAH^{+}08, GER^{+}06, HBN09, HO03, LGR08, MST^{+}04, LB09, MSD^{+}00, NFR08, PSP^{+}09, PMZD08, RGTR02, RRG^{+}05, RKG^{+}00, RAT^{+}02, RTNA02, RNAA04, RP04, SBE03, SBDPAP07, SSS^{+}00, SBSR02, SOF^{+}09, WSKC03, YIT05, ZWZS02, ZM00, LJM08]. sea-farmed [BM06]. sea-ranched [BLT01]. seabass [CDI^{+}01, GKCF04]. seabream [CFA^{+}09, EKG^{+}02, LVA^{+}06, MMTPDR^{+}08, PRG^{+}05]. seahorse [AGSS^{+}01]. Seasonal [HBN09, Wan02]. Seasonality [YWS^{+}08]. seawater [GMS^{+}09, MSD^{+}00, PBIM09, RSFS00, SLC^{+}06, SKR^{+}08]. seawater-cultured [SLC^{+}06]. seawater-farmed [PBIM09]. seaweed [BSM^{+}06]. Sebastes [KWD^{+}01, KL07, WWK03]. secondary [GDG00]. Secretory [SMLZ05]. sections [MMAP00]. segmented [MBD^{+}02]. Selected [DPS^{+}07, ABR^{+}09, CGP08, DJW^{+}09, DNC^{+}05, PV08, WMB02]. selection [HMB07]. seminal [SM08]. semipunctata [TL09]. senegalensis [ACGR^{+}05, ARC^{+}05, CRBM05, CRA^{+}05, DRCA^{+}08, MMTPDR^{+}08, ZAC^{+}03]. Senegalese [CRA^{+}05, MMTPDR^{+}08]. sensing [TYH^{+}01, YML^{+}08]. sensitive [DZP^{+}00, GEC09, MGE04]. Sensitivity [MO06, BMŠH^{+}02, NDH05a]. separation [BSQW05]. septemfasciatus [TMA^{+}01, TTM04, YMKN09, YMN09]. septicaemia [Bow03, EFT^{+}06, GMB^{+}07, HGP^{+}07, JGK^{+}09, LVDBB07, SKO04, SOM05, SKGR05, SFT^{+}08]. Sequence [KL07, JGK^{+}09, KJC^{+}02, MGGSREAG08, WCSC07, WCK^{+}09]. sequences [CWW^{+}02, FRW^{+}08, NPF05, PWW^{+}06, SKR^{+}08, WWK03]. sequencing [IFA05, RHS^{+}06]. Sequential [AN04b, AN03, JGS^{+}04]. sera [BPV^{+}03, JTR^{+}04, SKSE02]. serine [CWJL07, CFA^{+}00]. Seriola [EWCT05, IKKS06, MPEN05, SKI^{+}06, WCT^{+}01]. seriolae [EWCT05, MPEN05, SNN^{+}08, WCT^{+}09, WCT^{+}01]. Seriolella [AVN08]. serological [BFR07, GJR^{+}03, GJR^{+}05, GHBW03, KJL07, VAH09]. serotype [ILAW03, SRAHJ^{+}08]. serotypes [MD00, XKP06]. serovar [FZA06, FLA06]. Serranidae [BB01]. serrata [FHY^{+}08, WGS^{+}07]. Serratia [Sta01]. serum [DSBB05, LLJ^{+}02]. settlement [BBK04, TSW00]. sevenband [TMA^{+}01, TKNM03, TTM04, YMKN09, YMN09]. several [MMTPDR^{+}08, TKNM03]. severe [GKB05, MMB02]. severity [KH06, SGRM07, SRT^{+}07]. SF68 [CL02]. shaharomae [MEM^{+}09]. shark [BD00, BHB03, BHRB04, BT09, BKSB09, PP02, PGB^{+}06]. sharks [BF02a, BKNS02, BC06, BSTW08, BT09, BKSB09, TBB^{+}08]. shasta [BB09]. shedding [NM02, UMG^{+}08]. shell [NSK00]. shellfish [SKIS05]. Shetland [SBB^{+}06]. Shewanella [CCTL06]. shiners [IGH00]. Shmuel [She09]. shock [PLC09, SAC^{+}09]. short [CIK^{+}04, LLHF02]. short- [CIK^{+}04]. short-term [LLHF02]. shortfin [BT09, BKSB09]. show [YNH03]. shows [NHA07]. shrimp [BSM^{+}06, BDSY07, CLQAGCRP08, DBR^{+}05, DBR04, DL02, HBB09, Jay09, KHA04b, MRM^{+}00, RJB00, TXZ^{+}09, YSH^{+}06]. shyshark [HSD07]. Siamese [PTKA02]. sib [CGW04]. side [BGB^{+}04, CDS^{+}00]. side-effects [BGB^{+}04]. sighs [ZB07]. signalling [HGH^{+}08]. significant [DHDCN02, YNH03]. significantly [BLL^{+}07]. signs [GSM^{+}04, PNS^{+}05]. silver [DAMQ^{+}04, WYWL01]. simple [FNE09, RLL^{+}07, SMK^{+}09]. Simultaneous [YWBH05, YMN09, YSH^{+}06]. sinensis [ZB07, ZSZB04]. single [BMS^{+}00, RTNA02, SBE03, YSH^{+}06, YWBH05]. single-dose [SBE03]. single-step [YSH^{+}06]. Siniperca [HWZ^{+}00, SXSN07]. Sinuolinea [GAH^{+}08]. Siphonostomatoida [KPWK06]. site [NFR08]. sites [WGM^{+}08]. situ [CFA^{+}09, CABF08, HSW03, HWT^{+}06, úMMUB08, WDC^{+}03]. size [BSD05]. sized [AGM^{+}08]. SJNNV [CDT^{+}07, INM^{+}01]. Skeletal [PC04, ATT^{+}09, FKT^{+}05, KKT^{+}04, KHTF06, KT09]. Skin [LSDB06, Rob06, BBLC00, BBJ04, EEF06, Goo06, LYO04, LLZ08, ONS07, PP02, TWOM06, VNSB00, XK02, XKS02]. skinny [GKCF04]. sleeping [BML^{+}06, GRW^{+}03]. slit [TBB^{+}08]. small [GDP^{+}03, IGH00, NLO08, PBW06, QSZ^{+}03, SAH^{+}07, WCSC07, YWBH05]. smears [Cvi04b]. smolt [RAG^{+}07, WHB^{+}03]. smolting [BBBW01, GWF^{+}02]. smolts [BM04, BM06, BLSE03, MGAE06, NFR08, SBB^{+}06, UMG^{+}08]. smooth [BF02a]. snakehead [PTKA02]. sobria [GGCFB09, WBP^{+}05]. sockeye [CRM^{+}03, PGC^{+}09, SHBB^{+}02, TF04]. sodium [HPG08]. soft [HOA09]. sole [ACGR^{+}05, ARC^{+}05, CIK^{+}04, CRBM05, CRA^{+}05, DRCA^{+}08, LNM^{+}09, MMTPDR^{+}08, ZAC^{+}03]. Solea [ACGR^{+}05, ARC^{+}05, CRBM05, CRA^{+}05, DRCA^{+}08, MMTPDR^{+}08, ZAC^{+}03]. some [BMKG^{+}07, LB00, OBAA09]. sonicated [XKS04]. Sound [BWS^{+}08]. source [CIK^{+}04]. sources [PD09, SSL^{+}08a]. south [AVN08, JDBN08, PFL^{+}08]. South-West [PFL^{+}08]. southeastern [MOG09]. southern [BEL09, CMD01, CDI^{+}01, FLHM04, HBN09, JDBN08, JO00, WMB02, WHW^{+}02, WWH^{+}02, WGS^{+}07, DBJN05]. soy [BMFL^{+}07, SBK^{+}05]. Soybean [MDB^{+}09, BMPB^{+}00, USR^{+}09]. soybeans [BMKG^{+}07]. sp [AN04a, AGSS^{+}01, BBLC00, BBFB07, BD00, BF02b, BN04b, CTAR01, CAAN07, Dig01, DNC^{+}05, FLHM04, FYO08, GAMR^{+}04, GAH^{+}08, GMBN04, GCN04, GPB^{+}05, HFI09, HBCS^{+}06, IPA^{+}01, LG03, LAZ05, LZ09, MEM^{+}09, PD01, SNNB01, SHFM09, VSST09, WWK03, YIT05, ZM01b, ZGM01]. sp.n [GDHM03]. spacer [NPF05]. Spain [FZA06, GVGR00, GRF^{+}07, IPA^{+}01, LVA^{+}06, NLO08, NPF05]. Sparus [AVP^{+}09, ATNF02a, CIK^{+}04, CFA^{+}09, CRA^{+}05, CVGG06, DW05a, DW05b, EKG^{+}02, MMTPDR^{+}08, MST^{+}04, PRG^{+}05, RKG^{+}00, RAT^{+}02, SBDPAP07]. spathaceum [PSTV06]. Spatial [GER^{+}07, PMZD08, DHOM^{+}03, HOM^{+}09]. Spatio [DSBB05]. Spatio-temporal [DSBB05]. spawning [CDL^{+}08]. SPDV [NFR08]. special [RLB09]. Species [APYH00, LB00, ABR^{+}09, BF02a, BC06, BT09, BKSB09, BBA08, CDT^{+}07, DNC^{+}05, EN03, GCWFB08, HSK^{+}08, INKS04, JZKS04, Koz07, OKK^{+}06, PWW^{+}06, PS02, SGS01, SCJ^{+}02, ZAC^{+}03]. Species-specific [APYH00]. Specific [XCD09, APYH00, GPS^{+}08, HSW03, MFLC01, PGC^{+}09, RA01, VFAN03]. specificity [BBA08, LB00, MO06, NDH05a, OLB03]. Sphaerospora [CKW^{+}01, HSW03]. Sphaerospora-like [CKW^{+}01]. Spinal [PGB^{+}06, SHRM07, AGM^{+}08, SGRM07, SRT^{+}07]. spined [GDG00, GDG01]. spiny [BF02a, BF02b, CKK02, Dig01]. spiral [BD00]. Spix [EKS09]. spleen [DDL^{+}01, GAMR^{+}04, GCR^{+}02, WZJ^{+}08]. splendens [PTKA02]. splendidus [FAM^{+}06, KWKLG09]. spontaneous [SCJ^{+}01]. Sporadic [LM03]. spore [HBCS^{+}06]. spores [BHO05, CBDC^{+}03, DCCF^{+}03, HPG08, MMBA05, SGEM03]. sporulating [MBD^{+}02]. spot [BSM^{+}06, CWS02, CZS^{+}01, DBR04, DL02, EBASW^{+}08, KRJO02, LMX^{+}05, MBB^{+}06, MGGSREAG08, PKK09, RBA^{+}03, SHZ^{+}00, SWZ^{+}05, SWTC01, Shi02, SMK^{+}09, TSL^{+}01, TXZ^{+}09, WM04, YSH^{+}06, YNH03, YMNH04]. spotted [GDP^{+}03, HLWH07, áKBH06]. spp [BDSY07, EWCT05, HBN09, Jay09, JKD06, KWD^{+}01, MES^{+}02, MSCH03, PTKA02, SHS05, TSW06, VLT^{+}07, YSN^{+}06, ZFF^{+}08]. spp. [BPV^{+}03, DRCA^{+}06, JTM^{+}01, LVA^{+}06, NHA07, RP04, TO02]. Sprague [ABCRH09]. spray [TL09]. spring [GDR^{+}07, KM01, PV08, SSK^{+}08]. Squalus [BF02a, BF02b, CKK02]. SSN [HHCBR^{+}07]. SSN-1 [HHCBR^{+}07]. St [EFT^{+}06, CM04]. stable [CYL^{+}06]. stage [TTM04]. stages [BM09, CABF08, KSKEM07, KRJO02, SWS07, SGRM07]. stamina [KLG^{+}06, KHSW09]. standard [NDH05a]. Staphylococcus [GVGR00, HLC00]. state [GUE06, LSM^{+}02, WKB^{+}02]. status [GWF^{+}02, ØGW04, PSLN^{+}08]. Steatitis [Goo06, RA08]. steelhead [FPO^{+}09]. Steindachner [HLC00, IPA^{+}01, WCJJ08]. stellaris [GDP^{+}03]. stenohaline [AG01]. step [YSH^{+}06, YWBH05]. steps [DSJPP08]. sterile [YCR^{+}09]. stickleback [GMB^{+}07]. sticklebacks [JKD06]. still [ME08]. stillwater [TSW06, TWS09]. stimulated [CTAR01]. stimulates [PZS02]. STL2 [BHO05]. stochastic [OB07]. stock [LSDB06]. stocks [AJ09, CLN05, JD02, KH06]. stomach [RSFS00]. stone [KBPD04]. storage [TTD^{+}06]. story [KRKS06]. strain [CWJL07, GGCFB09, HBB09, JPPS06, LG03, SMLZ05, UBB^{+}09, VFT^{+}03]. strains [AW02, BFR07, CTB^{+}06, CL02, DNC^{+}05, FAM^{+}06, FW01, FKA^{+}05, FZA06, KHA05, LVDBB07, MMTPDR^{+}08, NLO08, PKK09, SBE03, SRAHJ^{+}08, SME^{+}01, SHS05, VAH09, VJR^{+}09, VSNS08, ZASK06, ZXZ^{+}08]. strategies [RRG^{+}05]. strategy [AW01, SMM09]. Streptococcosis [HFI09, BA05, XSK07]. Streptococcus [DAMQ^{+}04, EKG^{+}02, NBB08, NMJ^{+}04, NUS^{+}06, OFKEA08, PEP^{+}05, SKSE02, VJR^{+}09, ZXZ^{+}08]. stress [GKB05, LMG^{+}03, RWSK09, RSFS00, SSH^{+}07]. stress- [SSH^{+}07]. stressors [HDP^{+}02]. striata [PTKA02]. striped [BWCM04, GMB^{+}07, GDHM03, INM^{+}01, JO00, MOO^{+}09, MKW^{+}09, WCT^{+}09]. Strong [RM08]. structural [MGGSREAG08]. structure [AN01, BBJ04, DGLS09, GAMR^{+}04, SXSN07]. Studies [BSM^{+}06, KT09, MASS07, WSKC03, AJ09, EMH02, GSMM09, KIIK04, MD00, MMD05, MES^{+}02, MAF^{+}00, TTM04]. study [ASEK03, BLGR09, BSTW08, Bow03, DDL^{+}01, DJW^{+}09, DHOM^{+}03, FBS03, GKNLN^{+}04, GWF^{+}02, GBGB06, HHCBR^{+}07, HFI09, JGS^{+}04, KHTF06, LLWW02, PGB^{+}06, RTNA02, RJC04, RM04, SBE03, SBM^{+}00, SHRM07, UBB^{+}09]. sturgeon [DLS^{+}07]. stylirostris [CLQAGCRP08, DCTCB^{+}06]. subacute [DMS09]. Subcellular [ABC^{+}08]. subclinical [GKOF09, HKPKS06]. Subject [Ano01e, Ano02c, Ano03d]. sublineage [EFT^{+}06]. subsequent [SOF^{+}09, VMN06]. subsp [YCR^{+}09, BVG^{+}03]. subsp. [AVP^{+}09, AOA^{+}09, ARC^{+}05, BHP^{+}04, BGBG05, BAS^{+}09, CRBM05, DRCMB03, KHA05, MMTPDR^{+}08, MOR^{+}09, NLO08, NHA05, PSLN^{+}09]. subspecies [KFM^{+}00, MMAP00]. substrate [ONS07]. subtilis [RLNB03]. subtractive [WZJ^{+}08]. success [EWCT05]. Successful [GPN05, KBEJ^{+}07]. suggested [And06]. suggests [HBCS^{+}06, SXSN07]. sulphate [GM07]. sultana [MMBA05]. summer [MBD^{+}02]. sunshine [TMM01]. Superoxide [DRCA^{+}06]. supertexta [CCTL06]. support [SMM09]. suppressed [GMBN04]. suppression [WZJ^{+}08]. suprex [SSH^{+}07]. surface [OMLF^{+}06, SN08, VLT^{+}07]. Surveillance [PSLN^{+}08]. survey [DBJN05, GJR^{+}03, KSW03, RJ02]. surveys [GJR^{+}05]. Survival [DRCMB03, AVP^{+}09, ATT^{+}09, BWMSB02, DDL^{+}01, GSW^{+}07, LLJ^{+}02, RDT^{+}03, RAG^{+}07, TSW00]. survivors [HGP^{+}07]. Susceptibility [ABR^{+}09, BOB^{+}06, GCWR07, GM07, MMTPDR^{+}08, OKK^{+}06, TKNM03, ALAC09, BSD05, BLL^{+}07, Bow03, CGW04, GBGB06, HHCBR^{+}07, KRJO02, LJL^{+}03, LSDB06, MSD^{+}00, PB04, SKLY03, SBG08, SKGR05, TO02, UBB^{+}09]. susceptible [HGP^{+}07, LMJ^{+}00, LMC^{+}01, LSDB06]. Suspected [PTB07]. SVCV [PV08, SSK^{+}08]. swim [PDOZ08]. swimming [KLG^{+}06, KHSW09, TF04]. Swiss [BH02]. Switzerland [KSW03, WKB^{+}02]. Sydney [SNOR09]. sydneyi [SNOR09]. sympatric [GPRCH09]. sympatry [CWW^{+}02]. syndrome [BSM^{+}06, CWS02, CVGG06, CZS^{+}01, CPRL09, DRA00, DBR04, DL02, EBASW^{+}08, HOA09, KRJO02, LLJC03, LMX^{+}05, LBN02, MBB^{+}06, MBD^{+}02, MGGSREAG08, PKK09, Rod07, RBA^{+}03, SHZ^{+}00, SWZ^{+}05, SWTC01, Shi02, SMK^{+}09, TSL^{+}01, TXZ^{+}09, VTA06, WM04, YSH^{+}06, YNH03, YMNH04, ZB07]. Synergistic [RDT^{+}03]. Syngnathus [LGF04]. synthetic [DBR04]. system [BSG^{+}05, DGA^{+}03, FM00, GAMR^{+}04, GDHM03, NEG06, YML^{+}08]. Systematic [KPWK06]. Systemic [GLC04, HWZ^{+}00, Hum07, SHM00, CFA^{+}09, CCB03, GKNLN^{+}04, LHHP07, OMR^{+}06]. systems [SME^{+}01]. Syuzo [Aok08].

T [BMFL^{+}07]. T-cell-like [BMFL^{+}07]. taeniolatus [GAH^{+}08]. tail [BSQW05, PBW06, WCSC07, WCW^{+}08]. Taiwan [CLL^{+}00, CWT^{+}00, CKW^{+}01, CLS^{+}02, HWT^{+}06, WSKC03, WCSC07, WCW^{+}08, WCK^{+}09]. Takifugu [TOW02, YSN^{+}06]. Tanana [KH06]. tank [BCM^{+}00]. tank-reared [BCM^{+}00]. Taqman [SMM09]. tarda [CTB^{+}06, HKL^{+}06, PZD^{+}06, WKES08, XCD09]. target [HSW03]. targeting [IYS^{+}07, KHA05]. targeting-induced [KHA05]. Tasmania [AVN08, DHNB05, NDB^{+}04]. Tasmanian [GCN04]. Taura [CPRL09]. taurus [BHRB04, PGB^{+}06, TBB^{+}08]. tawny [PP02]. taxonomic [YTLN01]. technique [HSW03]. techniques [CFA^{+}09]. technology [CLVO^{+}05, SKIS05]. teleost [ABCRH09, BBA08, MKM02]. Teleosts [Hum07]. Temminch [HKK^{+}09b]. Temminck [HKL^{+}06, IKKS06, JO00, KOK^{+}03, KHK^{+}04, KJL07, KJL08, KIIK04, LMJ^{+}00, LMC^{+}01, LCM^{+}01, LJL^{+}03, MWG^{+}02, SKI^{+}06, SYYO06, SMN^{+}09, TOW02, WSKC03, WCT^{+}01, YSN^{+}06, YIT05, ZWZS02]. Temperature [YSN^{+}06, BSD06, BBBW01, DL01, EWCT05, FHY^{+}08, GSW^{+}07, GWJR08, KHSW09, ØGW04, SSM00, TSW00]. temperatures [HKK^{+}09b, SSM00]. Temporal [DHOM^{+}03, HOM^{+}09, DSBB05, PMZD08]. Tenacibaculum [AHTM06, LNM^{+}09, VCN09]. Tentacularia [BD00]. Terapon [WCT^{+}09]. Teratogenicity [ØGW04]. term [CIK^{+}04, HBN09, LLHF02]. terminals [MA04]. terminological [Jön04]. test [DHCHN01, GME04, GEB^{+}08]. tested [BSM^{+}06]. testicular [BHB03]. testing [ME08, TO02]. tests [MO06, NDH^{+}05b, NDH05a, SRM^{+}03]. tetra [MMB02]. Tetracapsula [DGF02, LLHF02, MAF^{+}00, MMA02a, MMA02b, OW02]. Tetracapsuloides [MMBA05, MA04]. tetradecylthioacetic [ATT^{+}09]. Tetrahymena [LAZ05, LZ09]. Text [Hum07]. their [AR03, BMD^{+}06, BKNS02, BN04a, CRBM05, GCWFB08, HGP^{+}07, Hum07, LJL^{+}03, LB00, ZAC^{+}03]. thelohanellosis [WYWL01]. Thelohanellus [LYW03, WYWL01]. therapeutant [WYWL01]. therapy [FBP09]. theronts [XKS04, XKS05, XKS06]. thickening [SSL^{+}08a]. Thiele [VNSB00]. Third [Ano02b]. Three [TBB^{+}08, BT09, BKSB09, BWCM04, EN03, GER^{+}07, NDH^{+}05b, NDH05a, TXZ^{+}09, TOW02, WCT^{+}09]. threespine [JKD06]. thresher [BT09, BKSB09]. throughout [JGS^{+}04]. Thunberg [TMA^{+}01, TTM04, WHD^{+}08, YMKN09, YMN09]. Thunnus [CMD01, DBJN05, HBN09, JDBN08, MMSM06, MSCH03, RA08, SMN^{+}09, WMB02]. Thymallus [MPP^{+}03, PSTV06]. thymoma [RM04]. thynnus [MMSM06, RA08]. thyroid [BT09]. thyrsites [CBDC^{+}03, DCCF^{+}03]. tiger [BHRB04, GKNLN^{+}04, KHA04b, TOW02, TBB^{+}08, WHW^{+}02, YSN^{+}06]. tigerfish [WCT^{+}09]. tigrina [WHW^{+}02]. tilapia [BCB^{+}09, EKS09, GAMR^{+}04, GLC04, HFI09, HLC00, SKSE02, SHFM09, XSK07]. Time [USR^{+}09, BSDD02, DL02, JLSA08, KMS^{+}08, ME08, OLP01, PGH^{+}07, SMM09]. Time-related [USR^{+}09]. Timing [GMS07, RSD04, RSSD01]. Tissue [KKD^{+}09, CGMS03, HBCS^{+}06, KBS03, KBEJ^{+}07, LMC^{+}01, MMAP00, SSM00, SSM01, WMA^{+}02]. Tissues [Hum07, BSTW08, CPB^{+}07, CBDC^{+}03, CDGF07, Cvi04b, GOS04, GML^{+}02, HKK^{+}09b, KMS^{+}08, PTAR09]. titration [Shi02]. titres [SB01]. TO-2 [HLC00]. tool [MEC^{+}06, YTLN01]. Tornionjoki [ARKK08]. Torridon [PD09]. total [Cvi04b, RP03]. Toxicity [RKG^{+}00]. toxins [ZKG^{+}01]. toyoi [CL02]. Trachurus [ACP01, GA05]. Tracing [HSW03]. tract [HBK07, KHK^{+}04, LYO04, SBK^{+}05]. tracts [BC06]. transcribed [NPF05]. transcriptase [DCTCB^{+}06]. transcriptase-polymerase [DCTCB^{+}06]. transcription [AL09, HGH^{+}08]. Transfection [INM^{+}01]. transfer [GPRCH09]. Translocation [JOG^{+}06, JSG^{+}08]. Transmission [KBS03, ATNF02a, BS04, BML^{+}06, BPBT02, CLQAGCRP08, KT09, KFM^{+}00, LYO04, RSSD01, SBSH08, SBDPAP07, SAH^{+}07, WBNG02]. transmontanus [DLS^{+}07]. transport [SCJ^{+}02]. Treating [SBM^{+}06]. treatment [AN04b, BSDD02, BBJ04, FZM00, FBP09, GPN05, MPEN05, PB04, RRG^{+}05, RP03, RF01, SSS^{+}00]. treatments [BSM^{+}06, EWCT05, SBSR02, TJG04a]. triactinomyxon [KEEM^{+}05, SGEM03]. triactinomyxon-type [KEEM^{+}05]. trial [HBN09]. trials [BVG^{+}03, RM08, SMN^{+}09]. Tricaine [CN02]. Trichiurus [LSM^{+}02]. Trichodina [BCB02, Kha04a]. Trichogaster [FLGS00]. trichopterus [FLGS00]. trifolia [HBCS^{+}06]. Trigla [MIGR09]. trophonts [XKS04]. tropical [HSK^{+}08, LJL^{+}03]. tropism [KKD^{+}09]. trout [AJ09, ABC^{+}08, AJA03, BWMSB02, BM09, BSDD02, BSD05, BSD06, BB01, BLL^{+}07, BML^{+}06, BSG^{+}05, BA05, BNFA07, BNN03, BBJ04, CGP08, CTAR01, CDL^{+}08, CMD02, CDGF07, DM00, DRA00, DL01, DGF02, DDL^{+}01, DGLS09, DGA^{+}03, FPO^{+}09, FZA06, GMB^{+}07, GVGR00, GRW^{+}03, GDG00, GDG01, GMS06, GMS07, HPMW08, IA02b, IA03, JTR^{+}04, JO09, JOG^{+}06, JSG^{+}08, KHSW09, LLJ^{+}02, LLJC03, LBO^{+}01, LKH08, LYO04, LVW^{+}06, LLHF02, MLB^{+}06, MMD05, MALC00, MDB^{+}09, MBD^{+}02, MSD^{+}00, NA09a, NA09b, OBAA09, OBHF00, OLR03, OL06, PZS02, PSLN^{+}08, PSLN^{+}09, PFL^{+}08, PLC09, PBIM09, PT09, PSB06, RLNB03, RSSD01, RSD04, RJC04, RF01, RSFS00, SSM00, SB01, SLB04, SHN^{+}00, SDD^{+}00, SFT^{+}08, STV05, SBV09, SLC08, TOB^{+}07, TSW06, TWS09, TLC^{+}02, TTD^{+}06, VTA06, YWS^{+}08]. trumpeter [GDHM03]. truncatus [DADB00]. trutta [DADB00, DGA^{+}03, GDG00, GDG01, MALC00, PFL^{+}08]. truttae [HSW03]. Trypanoplasma [Woo03]. trypanorhynch [BC06]. Trypanosoma [PB04]. tshawytscha [FF07, JH01, JD02, KM01, KH06, LWWS04, LCH^{+}02, LMG^{+}03, LM03, LRH^{+}08, OR05, OB07, POHJ05, TPF09]. Tsuyuki [WWK03]. tube [YWBH05]. tuberculosis [KRKS06]. Tubifex [ESMB00]. tuna [CMD01, DBJN05, HBN09, JDBN08, MMSM06, RA08, SMN^{+}09, WMB02]. tunas [MSCH03]. tunic [HOA09]. turbot [ATNF02b, AHTM06, BGB^{+}04, BGBG05, Bow03, CTB^{+}06, CLVO^{+}05, JST^{+}04, LVDBB07, NLO08, OKK^{+}06, PZD^{+}06, SPC^{+}08, SGMP^{+}05, SHM00, TFN01, VFT^{+}03, VFAN03, WZJ^{+}08, XCD09]. Two [KWKLG09, LMJ^{+}00, SI06, ABCRH09, BF02a, BC06, CL02, GDG00, GDG01, HAEM02, JDBN08, KOK^{+}03, LCC^{+}08, SPL05, SHBJ^{+}09, SRT^{+}07, USR^{+}09, VSNS08, XNS04]. type [BSQW05, BF07, KEEM^{+}05]. types [NPF05]. typical [BF07]. typing [AMI^{+}07, OFKEA08]. tyrosine [HGH^{+}08].

UK [ABR^{+}09, JBB^{+}07, LGF04]. ulcerated [AJA03]. ulcers [Goo06]. Ultrasonic [ZWZS02]. Ultrastructural [Dov05, GDG00, LBN02, DPS^{+}07, RM04]. Ultrastructure [GA05, DGS^{+}02, RBA^{+}03]. Ultraviolet [BS04]. unabraded [BNS03]. Unicapsula [ABCRH09]. uninfected [GKB05]. United [GRW^{+}03, SFT^{+}08]. untreated [GER^{+}06]. unusual [AJA03, HBCS^{+}06]. upon [KMJM06, RSSD01]. upper [KH06]. Uptake [CRM^{+}03, MASS07]. Uronema [SHM00]. Uronema-like [SHM00]. USA [EFT^{+}06, GER^{+}07, GEB^{+}08, MOG09, PV08, TPF09]. Use [BA05, IA02b, IA03, NA09a, NA09b, AW01, BLGR09, Jön04, RRG^{+}05, SMLZ05, TRY^{+}05]. used [GJR^{+}03]. useful [YTLN01]. user [JGK^{+}09]. user-friendly [JGK^{+}09]. using [ATNF02a, ASEK03, BNS03, BVG^{+}03, CGMS03, CWS02, CDGF07, DGLS09, EMS05, FPO^{+}09, GJR^{+}05, HPG08, HSW03, HTC^{+}01, ILAW03, JO09, JTM^{+}01, KJB07, KJL08, LYO04, LLWW02, MRM^{+}00, MMAP00, MA04, PWW^{+}06, POHJ05, PTKA02, RLL^{+}07, Shi02, SGS01, SMM09, SKR^{+}08, SMK^{+}09, TSL^{+}01, TL09, WWK03, XK02]. utilization [RAG^{+}07].

V. [ABC^{+}08, KWKLG09, ZWZS02]. v2 [PTT^{+}08]. Va [GOS04]. vaccinated [BLT01]. Vaccination [BVG^{+}03, KJL08, PLC09, SPC^{+}08, BGB^{+}04, KMJM06, RM08]. vaccination-and-challenge [RM08]. Vaccine [KHH^{+}05, AW02, ARC^{+}05, AGM^{+}08, BVG^{+}03, BGB^{+}04, BGBG05, GLMG03, PSP^{+}09, PEP^{+}05, SSL^{+}08b, TRY^{+}05, YMKN09, ZWZS02]. Vaccine-associated [KHH^{+}05]. vaccine-induced [AGM^{+}08, BGB^{+}04]. vaccines [ACS^{+}08, MLB^{+}06, PLC09, RRKM07, SGMP^{+}05, SMLZ05].

Valenciennes [BTSH03, HGP^{+}07, JD02, LVA^{+}06, LLZ08, MPEN05, SBG08, YKR^{+}02]. validation [DHCHN01, RRKM07, TPF09]. vannamei [CLQAGCRP08, DCTCB^{+}06, MBB^{+}06, MGGSREAG08, RBA^{+}03]. variability [NBB08]. variant [LWWS04]. variation [DSBB05, KL07, KaVKT09, PFL^{+}08, TXZ^{+}09]. variations [HOM^{+}09, MGGSREAG08]. varieties [SBK^{+}05]. variety [ZM01a]. Varracalbmi [VHM^{+}00]. varying [AM09]. vascular [LHHP07]. vector [SBSH08]. vectors [OLR03]. vegetable [CIK^{+}04]. velocity [BB09]. veronii [HKV^{+}06]. vertebral [DGLS09]. vertical [BPBT02, SAH^{+}07]. vetoquinol [SBE03]. VHSV [KBEJ^{+}07, SKO04, JGK^{+}09]. via [HGH^{+}08, LYO04]. viability [CYL^{+}06]. viable [JOG^{+}06]. Vibrio [AGSS^{+}01, ARC^{+}05, ABC^{+}08, APR03, BBLC00, CWJL07, CRA^{+}05, CN04, FLGS00, FAM^{+}06, FLA06, GOS04, JPPS06, Jay09, KWKLG09, KBH^{+}08, RHS^{+}06, RMHA08, SBE03, SMLZ05, SRAHJ^{+}08, SHN^{+}00, SAC^{+}09, VFT^{+}03, VFAN03, WZJ^{+}08, XMM^{+}09, YML^{+}08, ZA00, ZWZS02, ZAC^{+}03]. vibriosis [CWJL07, DW05a, DW05b, RDT^{+}03, SBM^{+}06]. vicinity [GMW^{+}07]. Vietnam [CDT^{+}02, FTS^{+}01, KHA04b]. viraemia [GDR^{+}07, PV08, SSK^{+}08]. Viral [ABPK03, SOM05, Bow03, CDI^{+}01, EFT^{+}06, GMB^{+}07, HGP^{+}07, HLWH07, JST^{+}04, JGS^{+}04, JGK^{+}09, KSW03, LMX^{+}05, LVDBB07, RA01, SKO04, SKGR05, SFT^{+}08, TMA^{+}01, TTM04, UMG^{+}08, YMN09]. virions [BSQW05]. Virulence [MOR^{+}09, PRG^{+}05, APR03, JSG^{+}08, MPP^{+}03, MD00, PKK09, PGC^{+}09, RMG^{+}09, RHS^{+}06, STV06, TJG04b, YML^{+}08, ZAC^{+}03]. virulence-related [RHS^{+}06]. virulent [FW01]. Virus [DGMD^{+}09, KBEJ^{+}07, AST^{+}05, ACGR^{+}05, ACS^{+}08, AL09, AJ09, BGR08, BSM^{+}06, BMŠH^{+}02, BM09, BLL^{+}07, BML^{+}06, BSE02, Bow03, CFA^{+}09, CLQAGCRP08, CWS02, CYL^{+}06, CLL01, CZS^{+}01, CCB03, CPRL09, CLVO^{+}05, DZP^{+}00, DCTCB^{+}06, DBR04, DL02, EWBB02, EFT^{+}06, EBASW^{+}08, GMB^{+}07, GUE06, GDR^{+}07, GSMM09, GCR^{+}02, GRW^{+}03, GJR^{+}03, GCG^{+}04, GSW^{+}07, GMW^{+}07, GMS^{+}09, GEB^{+}08, HD05a, HD05b, HKK^{+}09a, HHCBR^{+}07, HSK^{+}08, HKK^{+}09b, HCFM04, HLWH07, INM^{+}01, IGH00, JO09, JGK^{+}09, KJC^{+}02, KL07, KRJO02, KKD^{+}09, LBO^{+}01, LMC^{+}01, LCM^{+}01, LMX^{+}05, LVDBB07, MFLC01, MASS07, MEC^{+}06, MHD^{+}04, MBB^{+}06, MBLK06, MGGSREAG08, MGE04, MGAE06, MO06, NDH^{+}05b, NDH05a, OL06, PV08, PBW06, PKK09, PGC^{+}09, QSZ^{+}03, RA01, RLA01, RMG^{+}09, Rod07, RBA^{+}03, RW03, RBBC^{+}09, RMSH09, SMG08, SHZ^{+}00, SWZ^{+}05, SWTC01, Shi02, SSK^{+}08, SBG08]. virus [SBM^{+}00, SKO04, SOM05, SBB^{+}06, SM08, SRM^{+}03, SKGR05, SMM09, SGG03, SAH^{+}07, SMK^{+}09, SOF^{+}09, TSL^{+}01, TXZ^{+}09, TKNM03, UMG^{+}08, WGM^{+}08, WCSC07, WMA^{+}02, WBNG02, WM04, YMKN09, YSH^{+}06, YNH03, YMNH04, YWBH05, ZB07]. virus-infected [CYL^{+}06]. Virus-like [DGMD^{+}09, QSZ^{+}03]. viruses [DSJPP08, TXZ^{+}09]. Visceral [TLC^{+}02]. viscosa [BGB^{+}04, GBGB06]. viscosus [BBLC00]. vitamin [BB07, GWF^{+}02, ØGW04, SMS02]. vitro [BN04b, HFW01, Jay09, LKH08, LCM^{+}01, PZS02, SBE03, TO02, VSST09, ZW00]. vivo [BHP^{+}04, CAAN07, DDL^{+}01, MGAE06]. VNN [CDI^{+}01]. Voight [HSD07]. Volume [Ano04c, Ano05c]. VP15 [MGGSREAG08]. VP19 [MGGSREAG08]. VP2 [XLZS08]. VP26 [MGGSREAG08]. VP28 [MGGSREAG08, YMNH04]. VP3 [XLZS08]. vs [KBEJ^{+}07]. vulnificus [FLA06]. vulpinus [BT09, BKSB09].

W1 [LVW^{+}06]. Walbaum [AJ09, ABC^{+}08, AJA03, BCB02, BWMSB02, BM09, BSDD02, BSD05, BSD06, BLL^{+}07, BML^{+}06, BSG^{+}05, BA05, BNFA07, BNN03, BBJ04, CGP08, CTAR01, CDL^{+}08, CMD02, CRM^{+}03, CDGF07, DM00, DL01, DGF02, DDL^{+}01, DGLS09, FPO^{+}09, FF07, FZA06, GVGR00, GRW^{+}03, GDG01, GMS06, GMS07, HPMW08, IA02b, IA03, JTR^{+}04, JO09, JH01, JD02, JKB08, JOG^{+}06, JSG^{+}08, KM01, KH06, KHSW09, LLJ^{+}02, LLJC03, LBO^{+}01, LKH08, LYO04, LLHF02, LWWS04, LCH^{+}02, LMG^{+}03, LM03, LRH^{+}08, MLB^{+}06, MMD05, MDB^{+}09, MBD^{+}02, MSD^{+}00, NA09a, NA09b, OR05, OB07, OBAA09, OBHF00, OLR03, OL06, PZS02, PSLN^{+}08, PSLN^{+}09, PLC09, PBIM09, PT09, POHJ05, PSB06, PGC^{+}09, RLNB03, RSSD01, RSD04, RJC04, RF01, RSFS00, SSM00, SLB04, SLC^{+}06, SHN^{+}00, SHBB^{+}02, SFT^{+}08, STV05, SBV09, SLC08, TOB^{+}07, TF04]. Walbaum [TWOM06, TLC^{+}02, TTD^{+}06, TPF09, VTA06, YYY^{+}06, YWS^{+}08]. Wales [CGP08, FPG^{+}02]. warehou [AVN08]. warm [SC06]. warmwater [TJG04a]. warneri [GVGR00]. washings [IFA05]. Water [BBBW01, MD08, ABPK03, BSD06, BCB^{+}09, BB09, DL01, DGF02, DHOM^{+}03, EMH02, GDG00, GDG01, HKKO07, HKK^{+}09a, PD09, RP03, RSFS00, RDT^{+}03, SBE03, SSM00, SC06, SRM^{+}03, SI06, SOF^{+}09, ZM00]. water-borne [SRM^{+}03]. water-filled [RSFS00]. waterborne [BLL^{+}07]. waters [ACP01, BC06, GA05, SYYO06, WGM^{+}08]. watershed [EWBB02]. wax [CDGF07]. wax-embedded [CDGF07]. wedge [LNM^{+}09]. weedy [GAH^{+}08]. weight [TWOM06]. well [HBCS^{+}06]. well-known [HBCS^{+}06]. west [PMZD08, NFR08, PFL^{+}08, SN00, SCJ^{+}01, SCJ^{+}02]. western [BF02a, BF02b, SGEM03]. Westrheim [WWK03]. which [DL01, KBEJ^{+}07]. whirling [KSKEM07, SGEM03]. White [BSQW05, RBA^{+}03, SHZ^{+}00, BSM^{+}06, CLQAGCRP08, CWS02, CZS^{+}01, CDI^{+}01, DLS^{+}07, DBR04, DL02, EBASW^{+}08, KRJO02, KHH^{+}05, LMX^{+}05, MBB^{+}06, MGGSREAG08, PBW06, PKK09, SWZ^{+}05, SWTC01, Shi02, SMK^{+}09, TSL^{+}01, TXZ^{+}09, WCSC07, WCW^{+}08, WM04, YSH^{+}06, YNH03, YMNH04]. whitefish [BWD^{+}07, SKO04]. whiting [ACP01]. whitish [QSZ^{+}03]. Whole [PSB06, SKSE02]. widely [SMM09]. Wild [DHDCN02, ARKK08, BHB03, BHRB04, CLQAGCRP08, CGP08, EKG^{+}02, GSM^{+}04, GPRCH09, GDG00, GDG01, IGH00, JHKB08, LVDBB07, MAF^{+}00, NDB^{+}04, ØH07, ONI^{+}08, PFL^{+}08, PD09, Rod07, RBBC^{+}09, RMSH09, SKO04, UMIK09, WGM^{+}08, YWS^{+}08]. wild-caught [BHB03, BHRB04]. Wildlife [BBA08]. Wilson [JH01]. winter [BCB02, CVGG06]. within [PFL^{+}08, SSM00]. wolffish [áKBH06]. wood [CHWK^{+}04, HCJ08]. WSSV [CZS^{+}01, DBR^{+}05, KRJO02, MBB^{+}06, MGGSREAG08, SHZ^{+}00, SMK^{+}09, TSL^{+}01, YNH03]. wuhanensis [WYWL01].

xenobiotica [MKW^{+}09]. xenoma [BSDD02, BSD06, SSM00, SDD^{+}00]. xenoma-expression [SSM00, SDD^{+}00]. XSV [PBW06, QSZ^{+}03, SAH^{+}07].

Yasutake [HCJ08]. year [BML^{+}06, JGS^{+}04]. yeast [STM^{+}01]. yellow [BAMG^{+}05, CLQAGCRP08, CLW03, DCTCB^{+}06, LMC^{+}01, LCM^{+}01, MWG^{+}02, WYJ05]. yellow-head [CLQAGCRP08]. yellowtail [IKKS06, SKI^{+}06]. Yersinia [BNN03, FZA06, RLNB03, SME^{+}01, TYH^{+}01, TRY^{+}05, TDH^{+}07, YTLN01]. YGNNV [LMC^{+}01, LCM^{+}01]. YHV [CLQAGCRP08]. yield [LVW^{+}06]. yolk [LBN02]. yruI [TYH^{+}01]. yruR [TYH^{+}01]. Yukon [KH06].

zander [KPWK06]. Zealand [LM03]. zebrafish [CHWK^{+}04, HBK07, HHW05, KBS03, RWSK09, TL09]. Zeuxapta [MPEN05]. zfBcl [CYL^{+}06]. zfBcl- [CYL^{+}06]. Zingiber [NA09a]. zoospore [GDG00]. Zschokkella [FYO08].

References

Arsan:2007:EGD

[AAH^{+}07] E. L. Arsan, S. D. Atkinson, S. L. Hallett, T. Meyers, and J. L. Bartholomew. Expanded geographical distribution of Myxobolus cerebralis: first detections from Alaska. Journal of Fish Diseases, 30(8):483–491, August 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Arijo:2008:SCV

[ABC^{+}08] S. Arijo, J. Brunt, M. Chabrillón, P. Díaz-Rosales, and B. Austin. Subcellular components of Vibrio harveyi and probiotics induce immune responses in rainbow trout, Oncorhynchus mykiss (Walbaum), against V. harveyi. Journal of Fish Diseases, 31(8):579–590, August 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Alama-Bermejo:2009:MMR

[ABCRH09] G. Alama-Bermejo, M. Cuadrado, J. A. Raga, and A. S. Holzer. Morphological and molecular redescription of the myxozoan Unicapsula pflugfelderi Schubert, Sprague & Reinboth 1975 from two teleost hosts in the Mediterranean. A review of the genus Unicapsula Davis 1924. Journal of Fish Diseases, 32(4):335–350, April 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Athanassopoulou:2003:VER

[ABPK03] F. Athanassopoulou, C. Billinis, V. Psychas, and K. Karipoglou. Viral encephalopathy and retinopathy of Dicentrarchus labrax (L.) farmed in fresh water in Greece. Journal of Fish Diseases, 26(6):361–365, June 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Algoet:2009:SSF

[ABR^{+}09] M. Algöet, A. E. Bayley, E. G. Roberts, S. W. Feist, R. W. Wheeler, and D. W. Verner-Jeffreys. Susceptibility of selected freshwater fish species to a UK Lactococcus garvieae isolate. Journal of Fish Diseases, 32(10):825–834, October 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Alonso:2005:ILD

[ACGR^{+}05] M. C. Alonso, I. Cano, E. Garcia-Rosado, D. Castro, J. Lamas, J. L. Barja, and J. J. Borrego. Isolation of lymphocystis disease virus from sole, Solea senegalensis Kaup, and blackspot sea bream, Pagellus bogaraveo (Brünnich). Journal of Fish Diseases, 28(4):221–228, April 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Abollo:2001:HCB

[ACP01] E. Abollo, M. Calvo, and S. Pascual. Hepatic coccidiosis of the blue whiting, Micromesistius poutassou (Risso), and horse mackerel, Trachurus trachurus (L.), from Galician waters. Journal of Fish Diseases, 24(6):335–343, July 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anderson:2008:IIH

[ACS^{+}08] E. Anderson, S. Clouthier, W. Shewmaker, A. Weighall, and S. LaPatra. Inactivated infectious haematopoietic necrosis virus (IHNV) vaccines. Journal of Fish Diseases, 31(10):729– 745, October 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ahmed:2009:DCC

[ACS^{+}09] V. P. Ishaq Ahmed, V. Chandra, R. Sudhakaran, S. Rajesh Kumar, M. Sarathi, V. Sarath Babu, B. Ramesh, and A. S. Sahul Hameed. Development and characterization of cell lines derived from rohu, Labeo rohita (Hamilton), and catla, Catla catla (Hamilton). Journal of Fish Diseases, 32 (3):211–218, March 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Adams:2004:GPR

[AEN04] M. B. Adams, K. Ellard, and B. F. Nowak. Gross pathology and its relationship with histopathology of amoebic gill disease (AGD) in farmed Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 27(3):151–161, March 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Altinok:2001:ELS

[AG01] I. Altinok and J. M. Grizzle. Effects of low salinities on Flavobacterium columnare infection of euryhaline and freshwater stenohaline fish. Journal of Fish Diseases, 24(6):361– 367, July 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Aunsmo:2008:ASD

[AGM^{+}08] A. Aunsmo, A. Guttvik, P. J. Midtlyng, R. B. Larssen, Ø. Evensen, and E. Skjerve. Association of spinal deformity and vaccine-induced abdominal lesions in harvest-sized Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 31 (7):515–524, July 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Alcaide:2001:VHC

[AGSS^{+}01] E. Alcaide, C. Gil-Sanz, E. Sanjuán, D. Esteve, C. Amaro, and L. Silveira. Vibrio harveyi causes disease in seahorse, Hippocampus sp. Journal of Fish Diseases, 24(5):311–313, June 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Atkinson:2007:LCC

[AHB07] S. D. Atkinson, S. L. Hallett, and J. L. Bartholomew. The life cycle of Chloromyxum auratum (Myxozoa) from goldfish, Carassius auratus (L.), involves an antonactinomyxon actinospore. Journal of Fish Diseases, 30(3):149–156, March 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Avendano-Herrera:2006:CMT

[AHTM06] R. Avendaño-Herrera, A. E. Toranzo, and B. Magariños. A challenge model for Tenacibaculum maritimum infection in turbot, Scophthalmus maximus (L.). Journal of Fish Diseases, 29(6):371–374, June 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ariel:2009:CSE

[AJ09] E. Ariel and B. Bang Jensen. Challenge studies of European stocks of redfin perch, Perca fluviatilis L., and rainbow trout, Oncorhynchus mykiss (Walbaum), with epizootic haematopoietic necrosis virus. Journal of Fish Diseases, 33 (12):1017–1025, December 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Austin:2003:RUG

[AJA03] D. A. Austin, E. M. Jordan, and B. Austin. Recovery of an unusual Gram-negative bacterium from ulcerated rainbow trout, Oncorhynchus mykiss (Walbaum), in Scotland. Journal of Fish Diseases, 26(4):247–249, April 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kristmundsson:2006:GAM

[áKBH06] á Kristmundsson, S. H. Bambir, and S. Helgason. Gyrodactylus anarhichatis Mo & Lile (Monogenea: Gyrodactylidae) infection of farmed spotted wolffish, Anarhichas minor Olafsen, in Iceland. Journal of Fish Diseases, 29(6):365–370, June 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Andrade:2009:DMD

[AL09] T. P. D. Andrade and D. V. Lightner. Development of a method for the detection of infectious myonecrosis virus by reverse-transcription loop-mediated isothermal amplification and nucleic acid lateral flow hybrid assay. Journal of Fish Diseases, 32(11):911–924, November 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ascenso:2009:EPP

[ALAC09] R. M. T. Ascenso, R. B. Leite, R. Afonso, and M. L. Cancela. Expression pattern of Perkinsus olseni genes in response to bivalves with different susceptibility to perkinsosis. Journal of Fish Diseases, 32(7):633–636, July 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Amundrud:2009:MSL

[AM09] T. L. Amundrud and A. G. Murray. Modelling sea lice dispersion under varying environmental forcing in a Scottish sea loch. Journal of Fish Diseases, 32(1):27–44, January 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Arai:2007:MTP

[AMI^{+}07] H. Arai, Y. Morita, S. Izumi, T. Katagiri, and H. Kimura. Molecular typing by pulsed-field gel electrophoresis of Flavobacterium psychrophilum isolates derived from Japanese fish. Journal of Fish Diseases, 30(6):345–355, June 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Adams:2001:DSL

[AN01] M. B. Adams and B. F. Nowak. Distribution and structure of lesions in the gills of Atlantic salmon, Salmo salar L., affected with amoebic gill disease. Journal of Fish Diseases, 24(9): 535–542, December 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Adams:2003:AGD

[AN03] M. B. Adams and B. F. Nowak. Amoebic gill disease: sequential pathology in cultured Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 26(10):601–614, October 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Adams:2004:EAG

[AN04a] M. B. Adams and B. F. Nowak. Experimental amoebic gill disease of Atlantic salmon, Salmo salar L.: further evidence for the primary pathogenic role of Neoparamoeba sp. (Page, 1987). Journal of Fish Diseases, 27(2):105–113, February 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Adams:2004:SPA

[AN04b] M. B. Adams and B. F. Nowak. Sequential pathology after initial freshwater bath treatment for amoebic gill disease in cultured Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 27(3):163–173, March 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anderson:2006:RCF

[And06] C. D. Anderson. A review of causal factors and control measures for bloat in farmed salmonids with a suggested mechanism for the development of the condition. Journal of Fish Diseases, 29(8):445–453, August 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2001:A

[Ano01a] Anonymous. Announcement. Journal of Fish Diseases, 24(1): 61, January 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2001:AI

[Ano01b] Anonymous. Author index. Journal of Fish Diseases, 24(9): 561–564, December 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2001:Ea

[Ano01c] Anonymous. Editorial. Journal of Fish Diseases, 24(1):1, January 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2001:Eb

[Ano01d] Anonymous. Editorial. Journal of Fish Diseases, 24(9):495, December 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2001:SI

[Ano01e] Anonymous. Subject index. Journal of Fish Diseases, 24(9): 565–570, December 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2002:AI

[Ano02a] Anonymous. Author index. Journal of Fish Diseases, 25(12): 737–742, December 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2002:BRF

[Ano02b] Anonymous. Book review: Fish Pathology, Third Edition, Edited by Ronald J Roberts. Journal of Fish Diseases, 25 (2):125, February 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2002:SI

[Ano02c] Anonymous. Subject index. Journal of Fish Diseases, 25(12): 743–749, December 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2003:AI

[Ano03a] Anonymous. Author index. Journal of Fish Diseases, 26 (11–12):695–699, November 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2003:Ea

[Ano03b] Anonymous. Erratum. Journal of Fish Diseases, 26(1):63, January 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2003:Eb

[Ano03c] Anonymous. Erratum. Journal of Fish Diseases, 26(5):313, May 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2003:SI

[Ano03d] Anonymous. Subject Index. Journal of Fish Diseases, 26 (11–12):701–707, November 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2004:AI

[Ano04a] Anonymous. Author index. Journal of Fish Diseases, 27(12): 737–741, December 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2004:KI

[Ano04b] Anonymous. Keyword index. Journal of Fish Diseases, 27 (12):743–746, December 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2004:VI

[Ano04c] Anonymous. Volume index. Journal of Fish Diseases, 27 (12):i–ix, December 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2005:AI

[Ano05a] Anonymous. Author index. Journal of Fish Diseases, 28(12): 743–747, December 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2005:KI

[Ano05b] Anonymous. Keyword index. Journal of Fish Diseases, 28 (12):749–752, December 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2005:VI

[Ano05c] Anonymous. Volume index. Journal of Fish Diseases, 28 (12):i–vi, December 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2007:Eb

[Ano07a] Anonymous. Editorial. Journal of Fish Diseases, 30(5):255, May 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2007:Ea

[Ano07b] Anonymous. Erratum. Journal of Fish Diseases, 30(1):63, January 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2008:Ea

[Ano08a] Anonymous. Erratum. Journal of Fish Diseases, 31(2):159, February 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2008:Eb

[Ano08b] Anonymous. Erratum. Journal of Fish Diseases, 31(8):637, August 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:2008:Ec

[Ano08c] Anonymous. Erratum. Journal of Fish Diseases, 31(11):877, November 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Amagliani:2009:DMP

[AOA^{+}09] G. Amagliani, E. Omiccioli, F. Andreoni, R. Boiani, I. Bianconi, R. Zaccone, M. Mancuso, and M. Magnani. Development of a multiplex PCR assay for Photobacterium damselae subsp. piscicida identification in fish samples. Journal of Fish Diseases, 32(8):645–653, August 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Aoki:2008:PDS

[Aok08] Takashi Aoki. Professor Dr Syuzo Egusa (1920–2007). Journal of Fish Diseases, 31(6):480, June 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Austin:2003:ABV

[APR03] B. Austin, A. C. Pride, and G. A. Rhodie. Association of a bacteriophage with virulence in Vibrio harveyi. Journal of Fish Diseases, 26(1):55–58, January 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Aoki:2000:SSP

[APYH00] T. Aoki, C-I. Park, H. Yamashita, and I. Hirono. Speciesspecific polymerase chain reaction primers for Lactococcus garvieae. Journal of Fish Diseases, 23(1):1–6, January 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Agius:2003:MMC

[AR03] C. Agius and R. J. Roberts. Melano-macrophage centres and their role in fish pathology. Journal of Fish Diseases, 26(9): 499–509, September 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Arijo:2005:EDV

[ARC^{+}05] S. Arijo, R. Rico, M. Chabrillon, P. Diaz-Rosales, E. Martínez-Manzanares, Ma. C. Balebona, B. Magariños, A. E. Toranzo, and M. A. Moriñigo. Effectiveness of a divalent vaccine for sole, Solea senegalensis (Kaup), against Vibrio harveyi and Photobacterium damselae subsp. piscicida. Journal of Fish Diseases, 28(1):33–38, January 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anttila:2008:EGS

[ARKK08] P. Anttila, A. Romakkaniemi, J. Kuusela, and P. Koski. Epidemiology of Gyrodactylus salaris (Monogenea) in the River Tornionjoki, a Baltic wild salmon river. Journal of Fish Diseases, 31(5):373–382, May 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Arias:2003:CSE

[ASEK03] C. R. Arias, C. A. Shoemaker, J. J. Evans, and P. H. Klesius. A comparative study of Edwardsiella ictaluri parent (EILO) and E. ictaluri rifampicin-mutant (RE-33) isolates using lipopolysaccharides, outer membrane proteins, fatty acids, Biolog, API 20E and genomic analyses. Journal of Fish Diseases, 26(7):415–421, July 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Alonso:2005:IIH

[AST^{+}05] M. Alonso, D. A. Stein, E. Thomann, H. M. Moulton, J. C. Leong, P. Iversen, and D. V. Mourich. Inhibition of infectious haematopoietic necrosis virus in cell cultures with peptideconjugated morpholino oligomers. Journal of Fish Diseases, 28(7):399–410, July 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Aranguren:2002:ETE

[ATNF02a] R. Aranguren, C. Tafalla, B. Novoa, and A. Figueras. Experimental transmission of encephalopathy and retinopathy induced by nodavirus to sea bream, Sparus aurata L., using different infection models. Journal of Fish Diseases, 25 (6):317–324, June 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Aranguren:2002:NRT

[ATNF02b] R. Aranguren, C. Tafalla, B. Novoa, and A. Figueras. Nodavirus replication in a turbot cell line. Journal of Fish Diseases, 25(6):361–366, June 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Alne:2009:ISF

[ATT^{+}09] H. Alne, M. S. Thomassen, H. Takle, B. F. Terjesen, F. Grammes, M. Oehme, S. Refstie, T. Sigholt, R. K. Berge, and K-A. Rørvik. Increased survival by feeding tetradecylthioacetic acid during a natural outbreak of heart and skeletal muscle inflammation in S0 Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 32(11):953–961, November 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Adams:2008:ODA

[AVN08] M. B. Adams, M. Villavedra, and B. F. Nowak. An opportunistic detection of amoebic gill disease in blue warehou, Seriolella brama Günther, collected from an Atlantic salmon, Salmo salar L., production cage in south eastern Tasmania. Journal of Fish Diseases, 31(9):713–717, September 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Acosta:2009:ISP

[AVP^{+}09] F. Acosta, J. Vivas, D. Padilla, J. Vega, J. Bravo, V. Grasso, and F. Real. Invasion and survival of Photobacterium damselae subsp. piscicida in non-phagocytic cells of gilthead sea bream, Sparus aurata L. Journal of Fish Diseases, 32(6):535– 541, June 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ardelli:2001:CIC

[AW01] B. F. Ardelli and P. T. K. Woo. Conjugation of isometamidium chloride to antibodies and the use of the conjugate against the haemoflagellate, Cryptobia salmositica Katz, 1951: an immuno-chemotherapeutic strategy. Journal of Fish Diseases, 24(8):439–451, October 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ardelli:2002:ECS

[AW02] B. F. Ardelli and P. T. K. Woo. Experimental Cryptobia salmositica (Kinetoplastida) infections in Atlantic salmon, Salmo salar L.: cell-mediated and humoral immune responses against the pathogenic and vaccine strains of the parasite. Journal of Fish Diseases, 25(5):265–274, May 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Brunt:2005:UPC

[BA05] J. Brunt and B. Austin. Use of a probiotic to control lactococcosis and streptococcosis in rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 28(12):693– 701, December 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bowser:2005:FYP

[BAMG^{+}05] P. R. Bowser, N. Abou-Madi, M. M. Garner, S. L. Bartlett, S. G. Grimmett, G. A. Wooster, T. A. Paul, R. N. Casey, and J. W. Casey. Fibrosarcoma in yellow perch, Perca flavescens (Mitchill). Journal of Fish Diseases, 28(5):301–305, May 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Boiani:2009:ECP

[BAS^{+}09] R. Boiani, F. Andreoni, G. Serafini, I. Bianconi, R. Pierleoni, S. Dominici, F. Gorini, and M. Magnani. Expression and characterization of the periplasmic cobalamin-binding protein of Photobacterium damselae subsp. piscicida. Journal of Fish Diseases, 32(9):745–753, September 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bennett:2001:GPC

[BB01] S. M. Bennett and M. B. Bennett. Gill pathology caused by infestations of adult and preadult Dissonus manteri Kabata (Copepoda: Dissonidae) on coral trout, Plectropomus leopardus (Lacépède), (Serranidae). Journal of Fish Diseases, 24 (9):523–533, December 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bowater:2007:LMM

[BB07] R. O. Bowater and B. Burren. Lateral muscular myopathy associated with vitamin E deficiency in farmed barramundi, Lates calcarifer (Bloch). Journal of Fish Diseases, 30(2): 117–121, February 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bjork:2009:EWV

[BB09] S. J. Bjork and J. L. Bartholomew. The effects of water velocity on the Ceratomyxa shasta infectious cycle. Journal of Fish Diseases, 32(2):131–142, February 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Burger:2008:WRP

[BBA08] M. A. A. Burger, A. C. Barnes, and R. D. Adlard. Wildlife as reservoirs for parasites infecting commercial species: host specificity and a redescription of Kudoa amamiensis from teleost fish in Australia. Journal of Fish Diseases, 31(11): 835–844, November 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bjerkaas:2001:WTR

[BBBW01] E. Bjerkås, E. Bjørnestad, O. Breck, and R. Waagbø. Water temperature regimes affect cataract development in smolting Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 24 (5):281–291, June 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Breck:2005:HNG

[BBC^{+}05] O. Breck, E. Bjerkås, P. Campbell, J. D. Rhodes, J. Sanderson, and R. Waagbø. Histidine nutrition and genotype affect cataract development in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 28(6):357–371, June 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Birkbeck:2007:IFS

[BBFB07] T. H. Birkbeck, M. Bordevik, M. K. Frøystad, and Å. Baklien. Identification of Francisella sp. from Atlantic salmon, Salmo salar L., in Chile. Journal of Fish Diseases, 30(8): 505–507, August 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Baily:2005:PCE

[BBG^{+}05] J. E. Baily, M. J. Bretherton, F. M. Gavine, H. W. Ferguson, and J. F. Turnbull. The pathology of chronic erosive dermatopathy in Murray cod, Maccullochella peelii peelii (Mitchell). Journal of Fish Diseases, 28(1):3–12, January 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Buchmann:2004:EFT

[BBJ04] K. Buchmann, J. Bresciani, and C. Jappe. Effects of formalin treatment on epithelial structure and mucous cell densities in rainbow trout, Oncorhynchus mykiss (Walbaum), skin. Journal of Fish Diseases, 27(2):99–104, February 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Browman:2004:ELS

[BBK04] H. I. Browman, K. Boxaspen, and P. Kuhn. The effect of light on the settlement of the salmon louse, Lepeophtheirus salmonis, on Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 27(12):701–708, December 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Birkbeck:2000:RBA

[BBLC00] T. H. Birkbeck, B. Billcliffe, A. Laidler, and D. I. Cox. The relationship between Aeromonas sp. NCIMB 2263, a causative agent of skin lesions in Atlantic salmon, Vibrio marinus (Moritella marina) and Vibrio viscosus. Journal of Fish Diseases, 23(4):281–283, July 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bridle:2003:ICO

[BBN03] A. R. Bridle, R. Butler, and B. F. Nowak. Immunostimulatory CpG oligodeoxynucleotides increase resistance against amoebic gill disease in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 26(6):367–371, June 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Borucinska:2006:MAL

[BC06] J. D. Borucinska and J. N. Caira. Mode of attachment and lesions associated with trypanorhynch cestodes in the gastrointestinal tracts of two species of sharks collected from coastal waters of Borneo. Journal of Fish Diseases, 29(7):395–407, July 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Barker:2002:TMC

[BCB02] D. E. Barker, D. K. Cone, and M. D. B. Burt. Trichodina murmanica (Ciliophora) and Gyrodactylus pleuronecti (Monogenea) parasitizing hatchery-reared winter flounder, Pseudopleuronectes americanus (Walbaum): effects on host growth and assessment of parasite interaction. Journal of Fish Diseases, 25(2):81–89, February 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bigarre:2009:OBI

[BCB^{+}09] L. Bigarré, J. Cabon, M. Baud, M. Heimann, A. Body, F. Lieffrig, and J. Castric. Outbreak of betanodavirus infection in tilapia, Oreochromis niloticus (L.), in fresh water. Journal of Fish Diseases, 32(8):667–673, August 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bransden:2000:NTR

[BCM^{+}00] M. P. Bransden, J. Carson, B. L. Munday, J. H. Handlinger, C. G. Carter, and B. F. Nowak. Nocardiosis in tank-reared Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 23 (1):83–85, January 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bridle:2002:IAE

[BCM^{+}02] A. R. Bridle, P. B. B. Crosbie, R. N. Morrison, J. Kwang, and B. F. Nowak. The immuno-adjuvant effect of carp interleukin-1 beta on the humoral immune response of barramundi, Lates calcarifer (Bloch). Journal of Fish Diseases, 25(7):429–432, July 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bridle:2005:EGA

[BCMN05] A. R. Bridle, C. G. Carter, R. N. Morrison, and B. F. Nowak. The effect of β-glucan administration on macrophage respiratory burst activity and Atlantic salmon, Salmo salar L., challenged with amoebic gill disease — evidence of inherent resistance. Journal of Fish Diseases, 28(6):347–356, June 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Borucinska:2000:LAA

[BD00] J. Borucinska and A. Dunham. Lesions associated with attachment of the cestode Tentacularia sp. to the duodenospiral junction in the blue shark, Prionace glauca (L.), with a description of the intestinal morphology of the shark. Journal of Fish Diseases, 23(5):353–359, September 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Banerjee:2007:CFA

[BDSY07] S. Banerjee, T. N. Devaraja, M. Shariff, and F. M. Yusoff. Comparison of four antibiotics with indigenous marine Bacillus spp. in controlling pathogenic bacteria from shrimp and Artemia. Journal of Fish Diseases, 30(7):383–389, July 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bravo:2009:CSL

[BEL09] S. Bravo, F. Erranz, and C. Lagos. A comparison of sea lice, Caligus rogercresseyi, fecundity in four areas in southern Chile. Journal of Fish Diseases, 32(1):107–113, January 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Borucinska:2002:NOL

[BF02a] J. D. Borucinska and S. Frasca, Jr. Naturally occurring lesions and micro-organisms in two species of free-living sharks: the spiny dogfish, Squalus acanthias L., and the smooth dogfish, Mustelus canis (Mitchill), from the north-western Atlantic. Journal of Fish Diseases, 25(5):287–298, May 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Borucinska:2002:PFD

[BF02b] J. D. Borucinska and S. Frasca, Jr. Pancreatic fibrosis and ductal ectasia associated with the nematode Pancreatonema americanum sp. nov., in spiny dogfish, Squalus acanthias L., from the north-western Atlantic. Journal of Fish Diseases, 25(6):367–370, June 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Burr:2007:ATI

[BF07] S. E. Burr and J. Frey. Analysis of type III effector genes in typical and atypical Aeromonas salmonicida. Journal of Fish Diseases, 30(11):711–714, November 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Balboa:2007:PSG

[BFR07] S. Balboa, H. W. Ferguson, and J. L. Romalde. Phenotypic, serological and genetic characterization of Pseudomonas anguilliseptica strains isolated from cod, Gadus morhua L., in northern Europe. Journal of Fish Diseases, 30(11):657–664, November 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bjornsdottir:2004:EIT

[BGB^{+}04] B. Björnsdóttir, S. Gudmundsdóttir, S. H. Bambir, B. Magnadóttir, and B. K. Gudmundsdóttir. Experimental infection of turbot, Scophthalmus maximus (L.), by Moritella viscosa, vaccination effort and vaccine-induced side-effects. Journal of Fish Diseases, 27(11):645–655, November 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bjornsdottir:2005:EIT

[BGBG05] B. Björnsdóttir, S. Gudmundsdóttir, S. H. Bambir, and B. K. Gudmundsdóttir. Experimental infection of turbot, Scophthalmus maximus (L.), by Aeromonas salmonicida subsp. achromogenes and evaluation of cross protection induced by a furunculosis vaccine. Journal of Fish Diseases, 28(3):181– 188, March 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bain:2008:GAI

[BGR08] N. Bain, A. Gregory, and R. S. Raynard. Genetic analysis of infectious pancreatic necrosis virus from Scotland. Journal of Fish Diseases, 31(1):37–47, January 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Burkhardt-Holm:2002:PKD

[BH02] P. Burkhardt-Holm. Proliferative kidney disease: why is it of interest for the Swiss project ‘Fishnet’? Journal of Fish Diseases, 25(8):441–442, August 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Borucinska:2003:HCT

[BHB03] J. D. Borucinska, J. C. Harshbarger, and T. Bogicevic. Hepatic cholangiocarcinoma and testicular mesothelioma in a wild-caught blue shark, Prionace glauca (L.). Journal of Fish Diseases, 26(1):43–49, January 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Booy:2005:ISR

[BHO05] A. Booy, J. D. Haddow, and R. W. Olafson. Isolation of the salmonid rhamnose-binding lectin STL2 from spores of the microsporidian fish parasite Loma salmonae. Journal of Fish Diseases, 28(8):455–462, August 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bakopoulos:2004:EVG

[BHP^{+}04] V. Bakopoulos, A. Hanif, K. Poulos, M. Galeotti, A. Adams, and G. J. Dimitriadis. The effect of in vivo growth on the cellular and extracellular components of the marine bacterial pathogen Photobacterium damsela subsp. piscicida. Journal of Fish Diseases, 27(1):1–13, January 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Borucinska:2004:GNC

[BHRB04] J. D. Borucinska, J. C. Harshbarger, R. Reimschuessel, and T. Bogicevic. Gingival neoplasms in a captive sand tiger shark, Carcharias taurus (Rafinesque), and a wild-caught blue shark, Prionace glauca (L.). Journal of Fish Diseases, 27 (3):185–191, March 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Borucinska:2002:PAR

[BKNS02] J. Borucinska, N. Kohler, L. Natanson, and G. Skomal. Pathology associated with retained fishing hooks in blue sharks, Prionace glauca (L.), with implications for their conservation. Journal of Fish Diseases, 25(9):515–521, September 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Borucinska:2009:MTS

[BKSB09] J. D. Borucinska, K. Kotran, M. Shackett, and T. Barker. Melanomacrophages in three species of free-ranging sharks from the northwestern Atlantic, the blue shark Prionacae glauca (L.), the shortfin mako, Isurus oxyrhinchus Rafinesque, and the thresher, Alopias vulpinus (Bonnaterre). Journal of Fish Diseases, 32(10):883–891, October 2009. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Baillie:2009:UPM

[BLGR09] M. Baillie, F. Lees, G. Gettinby, and C. W. Revie. The use of prevalence as a measure of lice burden: a case study of Lepeophtheirus salmonis on Scottish Atlantic salmon, Salmo salar L., farms. Journal of Fish Diseases, 32(1):15–25, January 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Biacchesi:2007:FGS

[BLL^{+}07] S. Biacchesi, M. Le Berre, S. Le Guillou, A. Benmansour, M. Brémont, E. Quillet, and P. Boudinot. Fish genotype significantly influences susceptibility of juvenile rainbow trout, Oncorhynchus mykiss (Walbaum), to waterborne infection with infectious salmon anaemia virus. Journal of Fish Diseases, 30(10):631–636, October 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bowden:2003:ECP

[BLSE03] T. J. Bowden, K. Lockhart, D. A. Smail, and A. E. Ellis. Experimental challenge of post-smolts with IPNV: mortalities do not depend on population density. Journal of Fish Diseases, 26(5):309–312, May 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Buchmann:2001:IRR

[BLT01] K. Buchmann, J. L. Larsen, and B. Therkildsen. Improved recapture rate of vaccinated sea-ranched Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 24(4):245–248, May 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bermingham:2004:ERF

[BM04] M. L. Bermingham and M. F. Mulcahy. Environmental risk factors associated with amoebic gill disease in cultured salmon, Salmo salar L., smolts in Ireland. Journal of Fish Diseases, 27(10):555–571, October 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bermingham:2006:MAA

[BM06] M. L. Bermingham and M. F. Mulcahy. Microfauna associated with amoebic gill disease in sea-farmed Atlantic salmon, Salmo salar L., smolts. Journal of Fish Diseases, 29(8):455– 465, August 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bebak:2009:CEI

[BM09] J. Bebak and P. E. McAllister. Continuous exposure to infectious pancreatic necrosis virus during early life stages of rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 32(2):173–181, February 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Beran:2006:DMC

[BMD^{+}06] V. Beran, L. Matlova, L. Dvorska, P. Svastova, and I. Pavlik. Distribution of mycobacteria in clinically healthy ornamental fish and their aquarium environment. Journal of Fish Diseases, 29(7):383–393, July 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bakke-McKellep:2007:RSC

[BMFL^{+}07] A. M. Bakke-McKellep, M. K. Frøystad, E. Lilleeng, F. Dapra, S. Refstie, Å. Krogdahl, and T. Landsverk. Response to soy: T-cell-like reactivity in the intestine of Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 30(1):13– 25, January 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bakke-McKellep:2007:HDM

[BMKG^{+}07] A. M. Bakke-McKellep, E. O. Koppang, G. Gunnes, M. Sanden, G-I. Hemre, T. Landsverk, and Å. Krogdahl. Histological, digestive, metabolic, hormonal and some immune factor responses in Atlantic salmon, Salmo salar L., fed genetically modified soybeans. Journal of Fish Diseases, 30(2): 65–79, February 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Boscher:2006:ETS

[BML^{+}06] S. Kerbart Boscher, M. McLoughlin, A. Le Ven, J. Cabon, M. Baud, and J. Castric. Experimental transmission of sleeping disease in one-year-old rainbow trout, Oncorhynchus mykiss (Walbaum), induced by sleeping disease virus. Journal of Fish Diseases, 29(5):263–273, May 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bakke-McKellep:2000:CIE

[BMPB^{+}00] A. M. Bakke-McKellep, C. McL. Press, G. Baeverfjord, Å. Krogdahl, and T. Landsverk. Changes in immune and enzyme histochemical phenotypes of cells in the intestinal mucosa of Atlantic salmon, Salmo salar L., with soybean meal-induced enteritis. Journal of Fish Diseases, 23(2):115– 127, March 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bowers:2000:PRA

[BMS^{+}00] J. M. Bowers, A. Mustafa, D. J. Speare, G. A. Conboy, M. Brimacombe, D. E. Sims, and J. F. Burka. The physiological response of Atlantic salmon, Salmo salar L., to a single experimental challenge with sea lice, Lepeophtheirus salmonis. Journal of Fish Diseases, 23(3):165–172, May 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Barlic-Maganja:2002:CES

[BMŠH^{+}02] D. Barlič-Maganja, M. Štrancar, P. Hostnik, V. Jenčič, and J. Grom. Comparison of the efficiency and sensitivity of virus isolation and molecular methods for routine diagnosis of infectious haematopoietic necrosis virus and infectious pancreatic necrosis virus. Journal of Fish Diseases, 25(2):73–80, February 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bowman:2004:SGB

[BN04a] J. P. Bowman and B. Nowak. Salmonid gill bacteria and their relationship to amoebic gill disease. Journal of Fish Diseases, 27(8):483–492, August 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Butler:2004:VIB

[BN04b] R. Butler and B. F. Nowak. In vitro interactions between Neoparamoeba sp. and Atlantic salmon epithelial cells. Journal of Fish Diseases, 27(6):343–349, June 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Brunt:2007:DPC

[BNFA07] J. Brunt, A. Newaj-Fyzul, and B. Austin. The development of probiotics for the control of multiple bacterial diseases of rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 30(10):573–579, October 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Buchmann:2003:IRA

[BNN03] K. Buchmann, M. E. Nielsen, and C. V. Nielsen. Immune responses against Yersinia ruckeri have no effect on colonization of rainbow trout, Oncorhynchus mykiss (Walbaum), by Gyrodactylus derjavini (Mikailov, 1975). Journal of Fish Diseases, 26(3):183–186, March 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bader:2003:CCM

[BNS03] J. A. Bader, K. E. Nusbaum, and C. A. Shoemaker. Comparative challenge model of Flavobacterium columnare using abraded and unabraded channel catfish, Ictalurus punctatus (Rafinesque). Journal of Fish Diseases, 26(8):461–467, August 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bandin:2006:SFC

[BOB^{+}06] I. Bandín, J. G. Olveira, J. J. Borrego, C. P. Dopazo, and J. L. Barja. Susceptibility of the fish cell line SAF-1 to betanodavirus. Journal of Fish Diseases, 29(10):633–636, October 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bowden:2003:SSA

[Bow03] T. J. Bowden. A study of the susceptibility of Atlantic halibut, Hippoglossus hippoglossus (L.), to viral haemorrhagic septicaemia virus isolated from turbot, Scophthalmus maximus (L.). Journal of Fish Diseases, 26(4):207–212, April 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Breuil:2002:EVT

[BPBT02] G. Breuil, J. F. P. Pépin, S. Boscher, and R. Thiéry. Experimental vertical transmission of nodavirus from broodfish to eggs and larvae of the sea bass, Dicentrarchus labrax (L.). Journal of Fish Diseases, 25(12):697–702, December 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bakopoulos:2003:IMF

[BPV^{+}03] V. Bakopoulos, M. Pearson, D. Volpatti, L. Gousmani, A. Adams, M. Galeotti, and G. J. Dimitriadis. Investigation of media formulations promoting differential antigen expression by Photobacterium damsela spp. piscicida and recognition by sea bass, Dicentrarchus labrax (L.), immune sera. Journal of Fish Diseases, 26(1):1–13, January 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Brooks:2009:CDI

[Bro09] K. M. Brooks. Considerations in developing an integrated pest management programme for control of sea lice on farmed salmon in Pacific Canada. Journal of Fish Diseases, 32(1):59– 73, January 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bruno:2000:BR

[Bru00] David Bruno. BOOK REVIEWS. Journal of Fish Diseases, 23(1):91, January 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Becker:2004:ULC

[BS04] J. A. Becker and D. J. Speare. Ultraviolet light control of horizontal transmission of Loma salmonae. Journal of Fish Diseases, 27(3):177–180, March 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Becker:2005:IFR

[BSD05] J. A. Becker, D. J. Speare, and I. R. Dohoo. Influence of feeding ratio and size on susceptibility to microsporidial gill disease caused by Loma salmonae in rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 28 (3):173–180, March 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Becker:2006:IWT

[BSD06] J. A. Becker, D. J. Speare, and I. R. Dohoo. Interaction of water temperature and challenge model on xenoma development rates for Loma salmonae (Microspora) in rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 29 (3):139–145, March 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Becker:2002:EMD

[BSDD02] J. A. Becker, D. J. Speare, J. Daley, and P. Dick. Effects of monensin dose and treatment time on xenoma reduction in microsporidial gill disease in rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 25(11):673– 680, November 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bowden:2002:DRI

[BSE02] T. J. Bowden, D. A. Smail, and A. E. Ellis. Development of a reproducible infectious pancreatic necrosis virus challenge model for Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 25(9):555–563, September 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bosi:2005:CND

[BSG^{+}05] G. Bosi, A. P. Shinn, L. Giari, E. Simoni, F. Pironi, and B. S. Dezfuli. Changes in the neuromodulators of the diffuse endocrine system of the alimentary canal of farmed rainbow trout, Oncorhynchus mykiss (Walbaum), naturally infected with Eubothrium crassum (Cestoda). Journal of Fish Diseases, 28(12):703–711, December 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Balasubramanian:2006:SIW

[BSM^{+}06] G. Balasubramanian, R. Sudhakaran, S. Syed Musthaq, M. Sarathi, and A. S. Sahul Hameed. Studies on the inactivation of white spot syndrome virus of shrimp by physical and chemical treatments, and seaweed extracts tested in marine and freshwater animal models. Journal of Fish Diseases, 29 (9):569–572, September 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bonami:2005:WTD

[BSQW05] J-R. Bonami, Z. Shi, D. Qian, and J. Sri Widada. White tail disease of the giant freshwater prawn, Macrobrachium rosenbergii: separation of the associated virions and characterization of Mr NV as a new type of nodavirus. Journal of Fish Diseases, 28(1):23–31, January 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Borucinska:2008:MMC

[BSTW08] J. D. Borucinska, B. Schmidt, J. Tolisano, and D. Woodward. Molecular markers of cancer in cartilaginous fish: immunocytochemical study of PCNA, p-53, myc and ras expression in neoplastic and hyperplastic tissues from free ranging blue sharks, Prionace glauca (L.). Journal of Fish Diseases, 31(2): 107–115, February 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bilodeau:2003:PLC

[BSW03] A. L. Bilodeau, B. C. Small, and W. R. Wolters. Pathogen loads, clearance and plasma cortisol response in channel catfish, Ictalurus punctatus (Rafinesque), following challenge with Edwardsiella ictaluri. Journal of Fish Diseases, 26 (7):433–437, July 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Borucinska:2009:CHF

[BT09] J. D. Borucinska and M. Tafur. Comparison of histological features, and description of histopathological lesions in thyroid glands from three species of free-ranging sharks from the northwestern Atlantic, the blue shark, Prionace glauca (L.), the shortfin mako, Isurus oxyrhinchus Rafinesque, and the thresher, Alopias vulpinus (Bonnaterre). Journal of Fish Diseases, 32(9):785–793, September 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bowater:2003:DFI

[BTSH03] R. O. Bowater, A. Thomas, R. G. Shivas, and J. D. Humphrey. Deuteromycotic fungi infecting barramundi cod, Cromileptes altivelis (Valenciennes), from Australia. Journal of Fish Diseases, 26(11–12):681–686, November 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bakopoulos:2003:VTS

[BVG^{+}03] V. Bakopoulos, D. Volpatti, L. Gusmani, M. Galeotti, A. Adams, and G. J. Dimitriadis. Vaccination trials of sea bass, Dicentrarchus labrax (L.), against Photobacterium damsela subsp. piscicida, using novel vaccine mixtures. Journal of Fish Diseases, 26(2):77–90, February 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bowser:2004:PIH

[BWCM04] P. R. Bowser, G. A. Wooster, C-Y. Chen, and R. S. Mo. Polymicrobic infection of hybrid striped bass (Morone chrysops × Morone saxatilis) with three bacterial pathogens: a case report. Journal of Fish Diseases, 27(2):123–127, February 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Brzuzan:2007:BSD

[BWD^{+}07] P. Brzuzan, M. Woźny, S. Dobosz, H. Kuźmiński, M. K. Łuczyński, and M. Góra. Blue sac disease in larval whitefish, Coregonus lavaretus (L.): pathological changes in mRNA levels of CYP1A, ER α, and p53. Journal of Fish Diseases, 30 (3):169–173, March 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bebak-Williams:2002:EFD

[BWMSB02] J. Bebak-Williams, P. E. McAllister, G. Smith, and R. Boston. Effect of fish density and number of infectious fish on the survival of rainbow trout fry, Oncorhynchus mykiss (Walbaum), during epidemics of infectious pancreatic necrosis. Journal of Fish Diseases, 25(12):715–726, December 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bartlett:2008:NOB

[BWS^{+}08] S. L. Bartlett, G. A. Wooster, M. S. Sokolowski, A. D. M. Dove, and P. R. Bowser. Naturally occurring bacteraemia in American lobsters, Homarus americanus Milne-Edwards, in Long Island Sound. Journal of Fish Diseases, 31(1):19–25, January 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Crosbie:2007:DVA

[CAAN07] P. B. B. Crosbie, M. B. Adams, M. G. Attard, and B. F. Nowak. Development of an in vivo assay to assess attachment of Neoparamoeba sp. (an amphizoic gymnamoeba) to the gills of Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 30(3):187–190, March 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Costa:2007:IBC

[CAB^{+}07] J. Z. Costa, A. Adams, J. E. Bron, K. D. Thompson, W. G. Starkey, and R. H. Richards. Identification of B-cell epitopes on the betanodavirus capsid protein. Journal of Fish Diseases, 30(7):419–426, July 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Carrasco:2008:SHD

[CABF08] N. Carrasco, I. Arzul, F. C. J. Berthe, and M. D. Furones. In situ hybridization detection of initial infective stages of Marteilia refringens (Paramyxea) in its host Mytilus galloprovincialis. Journal of Fish Diseases, 31(2):153–157, February 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Carrasco:2008:CEI

[CAC^{+}08] N. Carrasco, I. Arzul, B. Chollet, M. Robert, J. P. Joly, M. D. Furones, and F. C. J. Berthe. Comparative experimental infection of the copepod Paracartia grani with Marteilia refringens and Marteilia maurini. Journal of Fish Diseases, 31 (7):497–504, July 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chase:2003:IDK

[CBDC^{+}03] J. C. Chase, M. H. Booy, J. A. Dawson-Coates, J. D. Haddow, L. R. Haines, D. J. Whitaker, R. W. Olafson, and T. W. Pearson. Immunological detection of Kudoa thyrsites spores in muscle tissues of farmed Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 26(7):427–431, July 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Corbel:2003:CSS

[CCB03] V. Corbel, F. Coste, and J-R. Bonami. CpSBV, a systemic virus of the edible crab, Cancer pagurus (L.). Journal of Fish Diseases, 26(2):121–126, February 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chiou:2009:CDI

[CCL09] P. P. Chiou, Y-C. Chen, and Y-S. Lai. Caspase-dependent induction of apoptosis in barramundi, Lates calcarifer (Bloch), muscle cells by grouper iridovirus. Journal of Fish Diseases, 33(12):997–1005, December 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cai:2006:IIS

[CCTL06] J. Cai, H. Chen, K. D. Thompson, and C. Li. Isolation and identification of Shewanella alga and its pathogenic effects on post-larvae of abalone Haliotis diversicolor supertexta. Journal of Fish Diseases, 29(8):505–508, August 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Crumlish:2007:DBF

[CDGF07] M. Crumlish, A. M. Diab, S. George, and H. W. Ferguson. Detection of the bacterium Flavobacterium psychrophilum from a natural infection in rainbow trout, Oncorhynchus mykiss (Walbaum), using formalin-fixed, wax-embedded fish tissues. Journal of Fish Diseases, 30(1):37–41, January 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Curtis:2001:NIJ

[CDI^{+}01] P. A. Curtis, M. Drawbridge, T. Iwamoto, T. Nakai, R. P. Hedrick, and A. P. Gendron. Nodavirus infection of juvenile white seabass, Atractoscion nobilis, cultured in southern California: first record of viral nervous necrosis (VNN) in North America. Journal of Fish Diseases, 24(5):263–271, June 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chen:2008:GDF

[CDL^{+}08] Y-C. Chen, M. A. Davis, S. E. LaPatra, K. D. Cain, K. R. Snekvik, and D. R. Call. Genetic diversity of Flavobacterium psychrophilum recovered from commercially raised rainbow trout, Oncorhynchus mykiss (Walbaum), and spawning Coho salmon, O. kisutch (Walbaum). Journal of Fish Diseases, 31 (10):765–773, October 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Crumlish:2000:DSM

[CDS^{+}00] M. Crumlish, T. T. Dung, T. Somsiri, V. Inglis, and K. D. Thompson. Development of a sampling method for the isolation of head kidney macrophages from fish at the pond-side. Journal of Fish Diseases, 23(4):289–293, July 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Crumlish:2002:IEI

[CDT^{+}02] M. Crumlish, T. T. Dung, J. F. Turnbull, N. T. N. Ngoc, and H. W. Ferguson. Identification of Edwardsiella ictaluri from diseased freshwater catfish, Pangasius hypophthalmus (Sauvage), cultured in the Mekong Delta, Vietnam. Journal of Fish Diseases, 25(12):733–736, December 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cutrin:2007:EPB

[CDT^{+}07] J. M. Cutrín, C. P. Dopazo, R. Thiéry, P. Leao, J. G. Olveira, J. L. Barja, and I. Bandín. Emergence of pathogenic betanodaviruses belonging to the SJNNV genogroup in farmed fish species from the Iberian Peninsula. Journal of Fish Diseases, 30(4):225–232, April 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cascon:2000:CCI

[CFA^{+}00] A. Cascón, J. Fregeneda, M. Aller, J. Yugueros, A. Temprano, C. Hernanz, M. Sánchez, L. Rodríguez-Aparicio, and G. Naharro. Cloning, characterization, and insertional inactivation of a major extracellular serine protease gene with elastolytic activity from Aeromonas hydrophila. Journal of Fish Diseases, 23(1):49–59, January 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cano:2009:ASD

[CFA^{+}09] I. Cano, P. Ferro, M. C. Alonso, C. Sarasquete, E. Garcia-Rosado, J. J. Borrego, and D. Castro. Application of in situ detection techniques to determine the systemic condition of lymphocystis disease virus infection in cultured gilt-head seabream, Sparus aurata L. Journal of Fish Diseases, 32(2): 143–150, February 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cepeda:2003:DFM

[CGMS03] C. Cepeda, S. García-Márquez, and Y. Santos. Detection of Flexibacter maritimus in fish tissue using nested PCR amplification. Journal of Fish Diseases, 26(2):65–70, February 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chambers:2008:IPR

[CGP08] E. Chambers, R. Gardiner, and E. J. Peeler. An investigation into the prevalence of Renibacterium salmoninarum in farmed rainbow trout, Oncorhynchus mykiss (Walbaum), and wild fish populations in selected river catchments in England and Wales between 1998 and 2000. Journal of Fish Diseases, 31 (2):89–96, February 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chin:2004:HRS

[CGW04] A. Chin, B. D. Glebe, and P. T. K. Woo. Humoral response and susceptibility of five full-sib families of Atlantic salmon, Salmo salar L., to the haemoflagellate, Cryptobia salmositica. Journal of Fish Diseases, 27(8):471–481, August 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Christianson-Heiska:2004:EWE

[CHWK^{+}04] I. Christianson-Heiska, P. Wahteristo, E-L. Kastilan, E. Bergelin, G. Bylund, and B. Isomaa. Effects of the wood extractive betulinol and 17 β-oestradiol on reproduction in zebrafish, Danio rerio (Hamilton) — complications due to a bacterial infection. Journal of Fish Diseases, 27(5):267–276, May 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Caballero:2004:HAL

[CIK^{+}04] M. J. Caballero, M. S. Izquierdo, E. Kjørsvik, A. J. Fernández, and G. Rosenlund. Histological alterations in the liver of sea bream, Sparus aurata L., caused by short- or longterm feeding with vegetable oils. Recovery of normal morphology after feeding fish oil as the sole lipid source. Journal of Fish Diseases, 27(9):531–541, September 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Crump:2008:CRI

[CK08] E. M. Crump and W. W. Kay. Congo red inhibition as a convenient diagnostic for Flavobacterium psychrophilum. Journal of Fish Diseases, 31(8):553–557, July 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Clewley:2002:IPS

[CKK02] A. Clewley, R. M. Kocan, and A. A. Kocan. An intraerythrocytic parasite from the spiny dogfish, Squalus acanthias L., from the Pacific Northwest. Journal of Fish Diseases, 25(11): 693–696, November 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chen:2001:MMA

[CKW^{+}01] S-C. Chen, R-J. Kou, C-T. Wu, P-C. Wang, and F-Z. Su. Mass mortality associated with a Sphaerospora-like myxosporidean infestation in juvenile cobia, Rachycentron canadum (L.), marine cage cultured in Taiwan. Journal of Fish Diseases, 24(4):189–195, May 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chang:2002:ETP

[CL02] C-I. Chang and W-Y. Liu. An evaluation of two probiotic bacterial strains, Enterococcus faecium SF68 and Bacillus toyoi, for reducing edwardsiellosis in cultured European eel, Anguilla anguilla L. Journal of Fish Diseases, 25(5):311–315, May 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chu:2005:MPA

[CL05] W-H. Chu and C-P. Lu. Multiplex PCR assay for the detection of pathogenic Aeromonas hydrophila. Journal of Fish Diseases, 28(7):437–441, July 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chen:2000:NSB

[CLL^{+}00] S-C. Chen, J-L. Lee, C-C. Lai, Y-W. Gu, C-T. Wang, H-Y. Chang, and K-H. Tsai. Nocardiosis in sea bass, Lateolabrax japonicus, in Taiwan. Journal of Fish Diseases, 23(5):299– 307, September 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chi:2001:CGN

[CLL01] S. C. Chi, B. J. Lo, and S. C. Lin. Characterization of grouper nervous necrosis virus (GNNV). Journal of Fish Diseases, 24 (1):3–13, January 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Callahan:2005:GRA

[CLN05] H. A. Callahan, R. W. Litaker, and E. J. Noga. Genetic relationships among members of the Ichthyobodo necator complex: implications for the management of aquaculture stocks. Journal of Fish Diseases, 28(2):111–118, February 2005. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Castro-Longoria:2008:DYH

[CLQAGCRP08] R. Castro-Longoria, N. Quintero-Arredondo, J. M. Grijalva-Chon, and J. Ramos-Paredes. Detection of the yellow-head virus (YHV) in wild blue shrimp, Penaeus stylirostris, from the Gulf of California and its experimental transmission to the Pacific white shrimp, Penaeus vannamei. Journal of Fish Diseases, 31(12):953–956, December 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chen:2002:LGC

[CLS^{+}02] S-C. Chen, L-L. Liaw, H-Y. Su, S-C. Ko, C-Y. Wu, H-C. Chaung, Y-H. Tsai, K-L. Yang, Y-C. Chen, T-H. Chen, G-R. Lin, S-Y. Cheng, Y-D. Lin, J-L. Lee, C-C. Lai, Y-J. Weng, and S-Y. Chu. Lactococcus garvieae, a cause of disease in grey mullet, Mugil cephalus L., in Taiwan. Journal of Fish Diseases, 25(12):727–732, December 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cutrin:2005:ICC

[CLVO^{+}05] J. M. Cutrín, C. López-Vázquez, J. G. Olveira, S. Castro, C. P. Dopazo, and I. Bandín. Isolation in cell culture and detection by PCR-based technology of IPNV-like virus from leucocytes of carrier turbot, Scophthalmus maximus (L.). Journal of Fish Diseases, 28(12):713–722, December 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chen:2003:OID

[CLW03] X. H. Chen, K. B. Lin, and X. W. Wang. Outbreaks of an iridovirus disease in maricultured large yellow croaker, Larimichthys crocea (Richardson), in China. Journal of Fish Diseases, 26(10):615–619, October 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cone:2004:RMB

[CM04] D. K. Cone and D. J. Marcogliese. Remodelling of mixed bone during infections of Myxobolus scleroperca (Myxozoa) in Percina caprodes (Rafinesque) (Percidae) from the St Lawrence River, Quebec. Journal of Fish Diseases, 27(9): 551–554, September 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Colquitt:2001:PFS

[CMD01] S. E. Colquitt, B. L. Munday, and M. Daintith. Pathological findings in southern bluefin tuna, Thunnus maccoyii (Castelnau), infected with Cardicola forsteri (Cribb, Daintith & Munday, 2000) (Digenea: Sanguinicolidae), a blood fluke. Journal of Fish Diseases, 24(4):225–229, May 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chilmonczyk:2002:PKD

[CMD02] S. Chilmonczyk, D. Monge, and P. De Kinkelin. Proliferative kidney disease: cellular aspects of the rainbow trout, Oncorhynchus mykiss (Walbaum), response to parasitic infection. Journal of Fish Diseases, 25(4):217–226, April 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Callahan:2002:TDR

[CN02] H. A. Callahan and E. J. Noga. Tricaine dramatically reduces the ability to diagnose protozoan ectoparasite (Ichthyobodo necator) infections. Journal of Fish Diseases, 25(7):433–437, July 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Crosbie:2004:IRB

[CN04] P. B. B. Crosbie and B. F. Nowak. Immune responses of barramundi, Lates calcarifer (Bloch), after administration of an experimental Vibrio harveyi bacterin by intraperitoneal injection, anal intubation and immersion. Journal of Fish Diseases, 27(11):623–632, November 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Costello:2009:GEC

[Cos09] Mark J. Costello. The global economic cost of sea lice to the salmonid farming industry. Journal of Fish Diseases, 32(1): 115–118, January 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Castro:2007:CHC

[CPB^{+}07] R. Castro, A. Paramá, J. L. Barja, J. Leiro, M. L. Sanmartin, and J. Lamas. Culture of the histophagous ciliate Philasterides dicentrarchi (Ciliophora: Scuticociliatia) in fish tissues. Journal of Fish Diseases, 30(4):239–242, April 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cote:2009:DCM

[CPRL09] I. Côté, B. T. Poulos, R. M. Redman, and D. V. Lightner. Development and characterization of a monoclonal antibody against Taura syndrome virus. Journal of Fish Diseases, 33 (12):989–996, December 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cook:2002:BFF

[CR02] M. A. Cook and M. B. Rust. Bioencapsulation of five forms of erythromycin by adult Artemia salina (L.). Journal of Fish Diseases, 25(3):165–170, March 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chabrillon:2005:IMI

[CRA^{+}05] M. Chabrillón, R. M. Rico, S. Arijo, P. Díaz-Rosales, M. C. Balebona, and M. A. Moriñigo. Interactions of microorganisms isolated from gilthead sea bream, Sparus aurata L., on Vibrio harveyi, a pathogen of farmed Senegalese sole, Solea senegalensis (Kaup). Journal of Fish Diseases, 28(9):531– 537, September 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chabrillon:2005:ASS

[CRBM05] M. Chabrillón, R. M. Rico, M. C. Balebona, and M. A. Moriñigo. Adhesion to sole, Solea senegalensis Kaup, mucus of microorganisms isolated from farmed fish, and their interaction with Photobacterium damselae subsp. piscicida. Journal of Fish Diseases, 28(4):229–237, April 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cook:2003:UEF

[CRM^{+}03] M. A. Cook, M. B. Rust, K. Massee, T. Majack, and M. E. Peterson. Uptake of erythromycin by first-feeding sockeye salmon, Oncorhynchus nerka (Walbaum), fed live or freezedried enriched adult Artemia or medicated pellets. Journal of Fish Diseases, 26(5):277–285, May 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chen:2001:PLM

[CTAR01] S-C. Chen, K. D. Thompson, A. Adams, and R. H. Richards. The production of a lymphokine (macrophage activating factor) by rainbow trout, Oncorhynchus mykiss (Walbaum), leucocytes stimulated with the extracellular products of Mycobacterium sp. Journal of Fish Diseases, 24(4):217–223, May 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Castro:2006:CET

[CTB^{+}06] N. Castro, A. E. Toranzo, J. L. Barja, S. Núñez, and B. Magariños. Characterization of Edwardsiella tarda strains isolated from turbot, Psetta maxima (L.). Journal of Fish Diseases, 29(9):541–547, September 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Colorni:2008:AAP

[CUHN08] A. Colorni, A. Ullal, G. Heinisch, and E. J. Noga. Activity of the antimicrobial polypeptide piscidin 2 against fish ectoparasites. Journal of Fish Diseases, 31(6):423–432, June 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Corrales:2009:LLD

[CUN09] J. Corrales, A. Ullal, and E. J. Noga. Lateral line depigmentation (LLD) in channel catfish, Ictalurus punctatus (Rafinesque). Journal of Fish Diseases, 32(8):705–712, August 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Contessi:2006:EIP

[CVGG06] B. Contessi, D. Volpatti, L. Gusmani, and M. Galeotti. Evaluation of immunological parameters in farmed gilthead sea bream, Sparus aurata L., before and during outbreaks of ‘winter syndrome’. Journal of Fish Diseases, 29(11):683–690, November 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cvitanich:2004:RSB

[Cvi04a] J. D. Cvitanich. Renibacterium salmoninarum bar forms: characterization, occurrence, and evidence of a host response to a R. salmoninarum infection. Journal of Fish Diseases, 27 (4):193–211, April 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cvitanich:2004:DDM

[Cvi04b] J. D. Cvitanich. Development of a direct microscopic counting method for kidney smears to estimate total numbers of microorganisms in fish kidney tissues. Journal of Fish Diseases, 27 (4):241–244, April 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cai:2007:CEG

[CWJL07] S. H. Cai, Z. H. Wu, J. C. Jian, and Y. S. Lu. Cloning and expression of the gene encoding an extracellular alkaline serine protease from Vibrio alginolyticus strain HY9901, the causative agent of vibriosis in Lutjanus erythopterus (Bloch). Journal of Fish Diseases, 30(8):493–500, August 2007. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chen:2002:AQW

[CWS02] Z. J. Chen, C. S. Wang, and H. H. Shih. An assay for quantification of white spot syndrome virus using a capture ELISA. Journal of Fish Diseases, 25(4):249–251, April 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chen:2000:PSL

[CWT^{+}00] S-C. Chen, P-C. Wang, M-C. Tung, K. D. Thompson, and A. Adams. A Piscirickettsia salmonis-like organism in grouper, Epinephelus melanostigma, in Taiwan. Journal of Fish Diseases, 23(6):415–418, November 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Criscione:2002:RDS

[CWW^{+}02] C. D. Criscione, V. Watral, C. M. Whipps, M. S. Blouin, S. R. M. Jones, and M. L. Kent. Ribosomal DNA sequences indicate isolated populations of Ichthyophonus hoferi in geographic sympatry in the north-eastern Pacific Ocean. Journal of Fish Diseases, 25(10):575–582, October 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chen:2006:EVN

[CYL^{+}06] S-P. Chen, H-L. Yang, H-Y. Lin, M-C. Chen, J-L. Wu, and J-R. Hong. Enhanced viability of a nervous necrosis virusinfected stable cell line over-expressing a fusion product of the zfBcl-x_{L} and green fluorescent protein genes. Journal of Fish Diseases, 29(6):347–354, June 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Crespo:2001:EHS

[CZP01] S. Crespo, C. Zarza, and F. Padrós. Epitheliocystis hyperinfection in sea bass, Dicentrarchus labrax (L.): light and electron microscope observations. Journal of Fish Diseases, 24 (9):557–560, December 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Corbel:2001:EIE

[CZS^{+}01] V. Corbel, Z. Zuprizal, C. Shi, Huang, Sumartono, J-M. Arcier, and J-R. Bonami. Experimental infection of European crustaceans with white spot syndrome virus (WSSV). Journal of Fish Diseases, 24(7):377–382, September 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dezfuli:2000:IDN

[DADB00] B. S. Dezfuli, S. Arrighi, C. Domeneghini, and G. Bosi. Immunohistochemical detection of neuromodulators in the intestine of Salmo trutta L. naturally infected with Cyathocephalus truncatus Pallas (Cestoda). Journal of Fish Diseases, 23(4):265–273, July 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Duremdez:2004:ISA

[DAMQ^{+}04] R. Duremdez, A. Al-Marzouk, J. A. Qasem, A. Al-Harbi, and H. Gharabally. Isolation of Streptococcus agalactiae from cultured silver pomfret, Pampus argenteus (Euphrasen), in Kuwait. Journal of Fish Diseases, 27(5):307–310, May 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Deveney:2005:PSF

[DBJN05] M. R. Deveney, T. J. Bayly, C. J. Johnston, and B. F. Nowak. A parasite survey of farmed Southern bluefin tuna, Thunnus maccoyii (Castelnau). Journal of Fish Diseases, 28(5):279– 284, May 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dupuy:2004:SAP

[DBR04] J-W. Dupuy, J-R. Bonami, and Ph Roch. A synthetic antibacterial peptide from Mytilus galloprovincialis reduces mortality due to white spot syndrome virus in palaemonid shrimp. Journal of Fish Diseases, 27(1):57–64, January 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

DiLeonardo:2005:CWI

[DBR^{+}05] V. A. Di Leonardo, V. Bonnichon, P. Roch, N. Parrinello, and J-R. Bonami. Comparative WSSV infection routes in the shrimp genera Marsupenaeus and Palaemon. Journal of Fish Diseases, 28(9):565–569, September 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dawson-Coates:2003:RBF

[DCCF^{+}03] J. A. Dawson-Coates, J. C. Chase, V. Funk, M. H. Booy, L. R. Haines, C. L. Falkenberg, D. J. Whitaker, R. W. Olafson, and T. W. Pearson. The relationship between flesh quality and numbers of Kudoa thyrsites plasmodia and spores in farmed Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 26 (8):451–459, August 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

DeLaRosa-Velez:2006:PDY

[DCTCB^{+}06] J. De La Rosa-Vélez, Y. Cedano-Thomas, J. Cid-Becerra, J. C. Méndez-Payán, C. Vega-Pérez, J. Zambrano-García, and J-R. Bonami. Presumptive detection of yellow head virus by reverse transcriptase-polymerase chain reaction and dot-blot hybridization in Litopenaeus vannamei and L. stylirostris cultured on the Northwest coast of Mexico. Journal of Fish Diseases, 29(12):717–726, December 2006. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Decostere:2001:VSP

[DDL^{+}01] A. Decostere, E. D’Haese, M. Lammens, H. Nelis, and F. Haesebrouck. In vivo study of phagocytosis, intracellular survival and multiplication of Flavobacterium psychrophilum in rainbow trout, Oncorhynchus mykiss (Walbaum), spleen phagocytes. Journal of Fish Diseases, 24(8):481–487, October 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dezfuli:2008:MCG

[DG08] B. S. Dezfuli and L. Giari. Mast cells in the gills and intestines of naturally infected fish: evidence of migration and degranulation. Journal of Fish Diseases, 31(11):845–852, November 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dezfuli:2003:IEH

[DGA^{+}03] B. S. Dezfuli, L. Giari, S. Arrighi, C. Domeneghini, and G. Bosi. Influence of enteric helminths on the distribution of intestinal endocrine cells belonging to the diffuse endocrine system in brown trout, Salmo trutta L. Journal of Fish Diseases, 26(3):155–166, March 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

DeKinkelin:2002:PIT

[DGF02] P. De Kinkelin, M. Gay, and S. Forman. The persistence of infectivity of Tetracapsula bryosalmonae-infected water for rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 25(8):477–482, August 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Deschamps:2009:CVS

[DGLS09] M-H. Deschamps, M. Girondot, L. Labbé, and J-Y. Sire. Changes in vertebral structure during growth of reared rainbow trout, Oncorhynchus mykiss (Walbaum): a new approach using modelling of vertebral bone profiles. Journal of Fish Diseases, 32(3):233–246, March 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dang:2009:VLP

[DGMD^{+}09] C. Dang, P. Gonzalez, N. Mesmer-Dudons, J-R. Bonami, N. Caill-Milly, and X. De Montaudouin. Virus-like particles associated with brown muscle disease in Manila clam, Ruditapes philippinarum, in Arcachon Bay (France). Journal of Fish Diseases, 32(7):577–584, July 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dezfuli:2002:HIU

[DGS^{+}02] B. S. Dezfuli, L. Giari, E. Simoni, G. Bosi, and M. Manera. Histopathology, immunohistochemistry and ultrastructure of the intestine of Leuciscus cephalus (L.) naturally infected with Pomphorhynchus laevis (Acanthocephala). Journal of Fish Diseases, 25(1):7–14, January 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Douglas-Helders:2001:DVN

[DHCHN01] M. Douglas-Helders, J. Carson, T. Howard, and B. Nowak. Development and validation of a new dot blot test for the detection of Paramoeba pemaquidensis (Page) in fish. Journal of Fish Diseases, 24(5):273–280, June 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Douglas-Helders:2002:WFS

[DHDCN02] G. M. Douglas-Helders, D. R. Dawson, J. Carson, and B. F. Nowak. Wild fish are not a significant reservoir for Neoparamoeba pemaquidensis (Page, 1987). Journal of Fish Diseases, 25(10):569–574, October 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Douglas-Helders:2005:EEF

[DHNB05] M. Douglas-Helders, B. Nowak, and R. Butler. The effect of environmental factors on the distribution of Neoparamoeba pemaquidensis in Tasmania. Journal of Fish Diseases, 28(10): 583–592, October 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Douglas-Helders:2003:TSD

[DHOM^{+}03] G. M. Douglas-Helders, D. P. O’Brien, B. E. McCorkell, D. Zilberg, A. Gross, J. Carson, and B. F. Nowak. Temporal and spatial distribution of paramoebae in the water column — a pilot study. Journal of Fish Diseases, 26(4):231– 240, April 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Diggles:2001:MJS

[Dig01] B. K. Diggles. A mycosis of juvenile spiny rock lobster, Jasus edwardsii (Hutton, 1875) caused by Haliphthoros sp., and possible methods of chemical control. Journal of Fish Diseases, 24(2):99–110, February 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dixon:2009:CRS

[DJW^{+}09] P. F. Dixon, C. L. Joiner, K. Way, R. A. Reese, G. Jeney, and Z. Jeney. Comparison of the resistance of selected families of common carp, Cyprinus carpio L., to koi herpesvirus: preliminary study. Journal of Fish Diseases, 33(12):1035–1039, December 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

DeKinkelin:2001:WTR

[DL01] P. De Kinkelin and B. Loriot. A water temperature regime which prevents the occurrence of proliferative kidney disease (PKD) in rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 24(8):489–493, October 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Durand:2002:QRT

[DL02] S. V. Durand and D. V. Lightner. Quantitative real time PCR for the measurement of white spot syndrome virus in shrimp. Journal of Fish Diseases, 25(7):381–389, July 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Drennan:2007:ELN

[DLS^{+}07] J. D. Drennan, S. E. LaPatra, C. A. Samson, S. Ireland, K. F. Eversman, and K. D. Cain. Evaluation of lethal and nonlethal sampling methods for the detection of white sturgeon iridovirus infection in white sturgeon, Acipenser transmontanus (Richardson). Journal of Fish Diseases, 30(6):367–379, June 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dalsgaard:2000:BPR

[DM00] I. Dalsgaard and L. Madsen. Bacterial pathogens in rainbow trout, Oncorhynchus mykiss (Walbaum), reared at Danish freshwater farms. Journal of Fish Diseases, 23(3):199–209, May 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Darwish:2009:EDA

[DM09] A. M. Darwish and A. J. Mitchell. Evaluation of diquat against an acute experimental infection of Flavobacterium columnare in channel catfish, Ictalurus punctatus (Rafinesque). Journal of Fish Diseases, 32(5):401–408, May 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dezfuli:2009:IRN

[DMG09] B. S. Dezfuli, M. Manera, and L. Giari. Immune response to nematode larvae in the liver and pancreas of minnow, Phoxinus phoxinus (L.). Journal of Fish Diseases, 32(4):383–390, April 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Darwish:2009:EPP

[DMS09] A. M. Darwish, A. J. Mitchell, and D. L. Straus. Evaluation of potassium permanganate against an experimental subacute infection of Flavobacterium columnare in channel catfish, Ictalurus punctatus (Rafinesque). Journal of Fish Diseases, 32 (2):193–199, February 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dykova:2005:NBS

[DNC^{+}05] I. Dyková, B. F. Nowak, P. B. B. Crosbie, I. Fiala, H. Pecková, M. B. Adams, B. Macháčková, and H. Dvořáková. Neoparamoeba branchiphila n. sp., and related species of the genus Neoparamoeba Page, 1987: morphological and molecular characterization of selected strains. Journal of Fish Diseases, 28(1):49–64, January 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dove:2005:UFE

[Dov05] A. D. M. Dove. Ultrastructural features of excretory calcinosis in the lobster, Homarus americanus Milne-Edwards. Journal of Fish Diseases, 28(5):313–316, May 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dezfuli:2007:SPI

[DPS^{+}07] B. S. Dezfuli, F. Pironi, E. Simoni, A. P. Shinn, and L. Giari. Selected pathological, immunohistochemical and ultrastructural changes associated with an infection by Diphyllobothrium dendriticum (Nitzsch, 1824) (Cestoda) plerocercoids in Coregonus lavaretus (L.) (Coregonidae). Journal of Fish Diseases, 30(8):471–482, August 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Daskalov:2000:IOL

[DRA00] H. Daskalov, P. A. W. Robertson, and B. Austin. Influence of oxidized lipids in diets on the development of rainbow trout fry syndrome. Journal of Fish Diseases, 23(1):7–14, January 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Diaz-Rosales:2006:SDC

[DRCA^{+}06] P. Díaz-Rosales, M. Chabrillón, S. Arijo, E. Martinez-Manzanares, M. A. Moriñigo, and M. C. Balebona. Superoxide dismutase and catalase activities in Photobacterium damselae spp. piscicida. Journal of Fish Diseases, 29(6):355– 364, June 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Diaz-Rosales:2008:EDA

[DRCA^{+}08] P. Díaz-Rosales, M. Chabrillón, R. T. Abdala, F. L. Figueroa, M. C. Balebona, and M. A. Moriñigo. Effect of dietary administration of Porphyridium cruentum on the respiratory burst activity of sole, Solea senegalensis (Kaup), phagocytes. Journal of Fish Diseases, 31(7):489–495, July 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Diaz-Rosales:2003:SAE

[DRCMB03] P. Díaz-Rosales, M. Chabrillón, M. A. Moriñigo, and M. C. Balebona. Survival against exogenous hydrogen peroxide of Photobacterium damselae subsp. piscicida under different culture conditions. Journal of Fish Diseases, 26(5):305–308, May 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dove:2005:STV

[DSBB05] A. D. M. Dove, M. S. Sokolowski, S. L. Bartlett, and P. R. Bowser. Spatio-temporal variation in serum chemistry of the lobster, Homarus americanus Milne-Edwards. Journal of Fish Diseases, 28(11):663–675, November 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

DeLasHeras:2008:SFV

[DSJPP08] A. I. De Las Heras, S. Rodríguez Saint-Jean, and S. I. Pérez-Prieto. Salmonid fish viruses and cell interactions at early steps of the infective cycle. Journal of Fish Diseases, 31 (8):535–546, July 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Deane:2005:EDK

[DW05a] E. E. Deane and N. Y. S. Woo. Evidence for disruption of Na^{+}-K^{+}-ATPase and hsp70 during vibriosis of sea bream, Sparus (= Rhabdosargus) sarba Forsskål. Journal of Fish Diseases, 28(4):239–251, April 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Deane:2005:MAI

[DW05b] E. E. Deane and N. Y. S. Woo. Modulation of β-actin, insulin-like growth factor 1 and glucose-6-phosphate dehydrogenase gene expression during vibriosis of sea bream, Sparus (= Rhabdosargus) sarba Forsskål. Journal of Fish Diseases, 28(10):593–601, October 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

DallaValle:2000:DSD

[DZP^{+}00] L. Dalla Valle, L. Zanella, P. Patarnello, L. Paolucci, P. Belvedere, and L. Colombo. Development of a sensitive diagnostic assay for fish nervous necrosis virus based on RT-PCR plus nested PCR. Journal of Fish Diseases, 23(5):321– 327, September 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Escobedo-Bonilla:2008:RMM

[EBASW^{+}08] C. M. Escobedo-Bonilla, V. Alday-Sanz, M. Wille, P. Sorgeloos, M. B. Pensaert, and H. J. Nauwynck. A review on the morphology, molecular characterization, morphogenesis and pathogenesis of white spot syndrome virus. Journal of Fish Diseases, 31(1):1–18, January 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Edgerton:2000:CIL

[Edg00] B. F. Edgerton. A compendium of idiopathic lesions observed in redclaw freshwater crayfish, Cherax quadricarinatus (von Martens). Journal of Fish Diseases, 23(2):103–113, March 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Elsayed:2006:IFP

[EEF06] E. E. Elsayed, A. E. Eissa, and M. Faisal. Isolation of Flavobacterium psychrophilum from sea lamprey, Petromyzon marinus L., with skin lesions in Lake Ontario. Journal of Fish Diseases, 29(10):629–632, October 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Elsayed:2006:IVH

[EFT^{+}06] E. Elsayed, M. Faisal, M. Thomas, G. Whelan, W. Batts, and J. Winton. Isolation of viral haemorrhagic septicaemia virus from muskellunge, Esox masquinongy (Mitchill), in Lake St Clair, Michigan, USA reveals a new sublineage of the North American genotype. Journal of Fish Diseases, 29(10):611– 619, October 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Evans:2002:CHG

[EKG^{+}02] J. J. Evans, P. H. Klesius, P. M. Gilbert, C. A. Shoemaker, M. A. Al Sarawi, J. Landsberg, R. Duremdez, A. Al Marzouk, and S. Al Zenki. Characterization of β-haemolytic Group B Streptococcus agalactiae in cultured seabream, Sparus auratus L., and wild mullet, Liza klunzingeri (Day), in Kuwait. Journal of Fish Diseases, 25(9):505–513, September 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Evans:2009:FIC

[EKS09] J. J. Evans, P. H. Klesius, and C. A. Shoemaker. First isolation and characterization of Lactococcus garvieae from Brazilian Nile tilapia, Oreochromis niloticus (L.), and pintado, Pseudoplathystoma corruscans (Spix & Agassiz). Journal of Fish Diseases, 32(11):943–951, November 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

El-Matbouli:2002:IWQ

[EMH02] M. El-Matbouli and R. W. Hoffmann. Influence of water quality on the outbreak of proliferative kidney disease — field studies and exposure experiments. Journal of Fish Diseases, 25(8):459–467, August 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

El-Matbouli:2005:DRA

[EMS05] M. El-Matbouli and H. Soliman. Development of a rapid assay for the diagnosis of Myxobolus cerebralis in fish and oligochaetes using loop-mediated isothermal amplification. Journal of Fish Diseases, 28(9):549–557, September 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ekman:2003:PIT

[EN03] E. Ekman and L. Norrgren. Pathology and immunohistochemistry in three species of salmonids after experimental infection with Flavobacterium psychrophilum. Journal of Fish Diseases, 26(9):529–538, September 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Eszterbauer:2000:DMB

[ESMB00] E. Eszterbauer, Cs Székely, K. Molnár, and F. Baska. Development of Myxobolus bramae (Myxosporea: Myxobolidae) in an oligochaete alternate host, Tubifex tubifex. Journal of Fish Diseases, 23(1):19–25, January 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Evelyn:2004:B

[Eve04] Trevor P. T. Evelyn. Biography. Journal of Fish Diseases, 27 (6):315–317, June 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Edgerton:2002:IBV

[EWBB02] B. F. Edgerton, H. Watt, J-M. Becheras, and J-R. Bonami. An intranuclear bacilliform virus associated with near extirpation of Austropotamobius pallipes Lereboullet from the Nant watershed in Ardéche, France. Journal of Fish Diseases, 25(9):523–531, September 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ernst:2005:ETS

[EWCT05] I. Ernst, I. D. Whittington, S. Corneillie, and C. Talbot. Effects of temperature, salinity, desiccation and chemical treatments on egg embryonation and hatching success of Benedenia seriolae (Monogenea: Capsalidae), a parasite of farmed Seriola spp. Journal of Fish Diseases, 28(3):157–164, March 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Farto:2006:PLP

[FAM^{+}06] R. Farto, Sp Armada, M. Montes, Mj Perez, and Tp Nieto. Presence of a lethal protease in the extracellular products of Vibrio splendidus–Vibrio lentus related strains. Journal of Fish Diseases, 29(12):701–707, December 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Florent:2009:FDB

[FBP09] R. L. Florent, J. Becker, and M. D. Powell. Further development of bithionol therapy as a treatment for amoebic gill disease in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 32(5):391–400, May 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Freeman:2003:HMI

[FBS03] M. A. Freeman, A. S. Bell, and C. Sommerville. A hyperparasitic microsporidian infecting the salmon louse, Lepeophtheirus salmonis: an rDNA-based molecular phylogenetic study. Journal of Fish Diseases, 26(11–12):667–676, November 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ferguson:2004:PAI

[FCM^{+}04] H. W. Ferguson, R. O. Collins, M. Moore, M. Coles, and D. D. MacPhee. Pseudomonas anguilliseptica infection in farmed cod, Gadus morhua L. Journal of Fish Diseases, 27 (4):249–253, April 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fenner:2006:DBJ

[FDG^{+}06] B. J. Fenner, Q. Du, W. Goh, R. Thiagarajan, H. K. Chua, and J. Kwang. Detection of betanodavirus in juvenile barramundi, Lates calcarifer (Bloch), by antigen capture ELISA. Journal of Fish Diseases, 29(7):423–432, July 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Forgan:2007:DPG

[FF07] L. G. Forgan and M. E. Forster. Development and physiology of gastric dilation air sacculitis in Chinook salmon, Oncorhynchus tshawytscha (Walbaum). Journal of Fish Diseases, 30(8):459–469, August 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fryer:2003:PSG

[FH03] J. L. Fryer and R. P. Hedrick. Piscirickettsia salmonis: a Gram-negative intracellular bacterial pathogen of fish. Journal of Fish Diseases, 26(5):251–262, May 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fang:2008:ETP

[FHY^{+}08] W. H. Fang, L. L. Hu, X. L. Yang, K. Hu, S. C. Liang, and S. Zhou. Effect of temperature on pharmacokinetics of enrofloxacin in mud crab, Scylla serrata (Forsskål), following oral administration. Journal of Fish Diseases, 31(3):171–176, March 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Figueiredo:2005:ICS

[FKA^{+}05] H. C. P. Figueiredo, P. H. Klesius, C. R. Arias, J. Evans, C. A. Shoemaker, D. J. Pereira, Jr., and M. T. D. Peixoto. Isolation and characterization of strains of Flavobacterium columnare from Brazil. Journal of Fish Diseases, 28(4):199–204, April 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ferguson:2005:ODR

[FKT^{+}05] H. W. Ferguson, R. T. Kongtorp, T. Taksdal, D. Graham, and K. Falk. An outbreak of disease resembling heart and skeletal muscle inflammation in Scottish farmed salmon, Salmo salar L., with observations on myocardial regeneration. Journal of Fish Diseases, 28(2):119–123, February 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fouz:2006:VVS

[FLA06] B. Fouz, J. L. Larsen, and C. Amaro. Vibrio vulnificus serovar A: an emerging pathogen in European anguilliculture. Journal of Fish Diseases, 29(5):285–291, May 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fang:2000:EPI

[FLGS00] H. M. Fang, K. C. Ling, R. Ge, and Y. M. Sin. Enhancement of protective immunity in blue gourami, Trichogaster trichopterus (Pallas), against Aeromonas hydrophila and Vibrio anguillarum by A. hydrophila major adhesin. Journal of Fish Diseases, 23(2):137–145, March 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Feist:2004:ODS

[FLHM04] S. W. Feist, M. Longshaw, R. H. Hurrell, and B. Mander. Observations of Dermocystidium sp. infections in bullheads, Cottus gobio L., from a river in southern England. Journal of Fish Diseases, 27(4):225–231, April 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Findlay:2000:IEL

[FM00] V. L. Findlay and B. L. Munday. The immunomodulatory effects of levamisole on the nonspecific immune system of Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 23(6): 369–378, November 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fujiwara-Nagata:2009:DEL

[FNE09] E. Fujiwara-Nagata and M. Eguchi. Development and evaluation of a loop-mediated isothermal amplification assay for rapid and simple detection of Flavobacterium psychrophilum. Journal of Fish Diseases, 32(10):873–881, October 2009. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Feist:2002:PKD

[FPG^{+}02] S. W. Feist, E. J. Peeler, R. Gardiner, E. Smith, and M. Longshaw. Proliferative kidney disease and renal myxosporidiosis in juvenile salmonids from rivers in England and Wales. Journal of Fish Diseases, 25(8):451–458, August 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Foltz:2009:DNS

[FPO^{+}09] J. R. Foltz, K. P. Plant, K. Overturf, K. Clemens, and M. S. Powell. Detection of Nucleospora salmonis in steelhead trout, Oncorhynchus mykiss (Walbaum), using quantitative polymerase chain reaction (qPCR). Journal of Fish Diseases, 32 (6):551–555, June 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fringuelli:2008:PAM

[FRW^{+}08] E. Fringuelli, H. M. Rowley, J. C. Wilson, R. Hunter, H. Rodger, and D. A. Graham. Phylogenetic analyses and molecular epidemiology of European salmonid alphaviruses (SAV) based on partial E2 and nsP3 gene nucleotide sequences. Journal of Fish Diseases, 31(11):811–823, November 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Foyle:2003:PAS

[FTE^{+}03] L. Foyle, T. Turnbull, A. Ellis, A. Barnes, A. Adams, and H. W. Ferguson. Pasteurellosis in Atlantic salmon, Salmo salar L.: immunohistochemistry of the naturally-occurring disease. Journal of Fish Diseases, 26(6):373–376, June 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ferguson:2001:BNF

[FTS^{+}01] H. W. Ferguson, J. F. Turnbull, A. Shinn, K. Thompson, T. T. Dung, and M. Crumlish. Bacillary necrosis in farmed Pangasius hypophthalmus (Sauvage) from the Mekong Delta, Vietnam. Journal of Fish Diseases, 24(9):509–513, December 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Feng:2001:CMG

[FW01] S. Feng and P. T. K. Woo. Cell membrane glycoconjugates on virulent and avirulent strains of the haemoflagellate Cryptobia salmositica (Kinetoplastida). Journal of Fish Diseases, 24 (1):23–32, January 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Freeman:2008:DPC

[FYO08] M. A. Freeman, H. Yokoyama, and K. Ogawa. Description and phylogeny of Ceratomyxa anko sp. n. and Zschokkella lophii sp. n. from the Japanese anglerfish, Lophius litulon (Jordan). Journal of Fish Diseases, 31(12):921–930, December 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fouz:2006:FDN

[FZA06] B. Fouz, C. Zarza, and C. Amaro. First description of nonmotile Yersinia ruckeri serovar I strains causing disease in rainbow trout, Oncorhynchus mykiss (Walbaum), cultured in Spain. Journal of Fish Diseases, 29(6):339–346, June 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Findlay:2000:ELT

[FZM00] V. L. Findlay, D. Zilberg, and B. L. Munday. Evaluation of levamisole as a treatment for amoebic gill disease of Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 23(3):193– 198, May 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gestal:2005:UGC

[GA05] C. Gestal and C. Azevedo. Ultrastructure of Goussia cruciata (Apicomplexa: Coccidia) infecting the liver of horse mackerel, Trachurus trachurus (L.), from Ibero-Atlantic waters. Journal of Fish Diseases, 28(3):125–132, March 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Garner:2008:RMW

[GAH^{+}08] M. M. Garner, S. D. Atkinson, S. L. Hallett, J. L. Bartholomew, R. W. Nordhausen, H. Reed, L. Adams, and B. Whitaker. Renal myxozoanosis in weedy sea dragons, Phyllopteryx taeniolatus (Lacépède), caused by Sinuolinea phyllopteryxa n. sp. Journal of Fish Diseases, 31(1):27–35, January 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Garcia-Abiado:2004:EDC

[GAMR^{+}04] M. A. Garcia-Abiado, G. Mbahinzireki, J. Rinchard, K. J. Lee, and K. Dabrowski. Effect of diets containing gossypol on blood parameters and spleen structure in tilapia, Oreochromis sp., reared in a recirculating system. Journal of Fish Diseases, 27(6):359–368, June 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Garcia:2005:E

[Gar05] Juan José Borrego Garcia. Erratum. Journal of Fish Diseases, 28(7):443, July 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Guy:2006:AII

[GBB^{+}06] D. R. Guy, S. C. Bishop, S. Brotherstone, A. Hamilton, R. J. Roberts, B. J. McAndrew, and J. A. Woolliams. Analysis of the incidence of infectious pancreatic necrosis mortality in pedigreed Atlantic salmon, Salmo salar L., populations. Journal of Fish Diseases, 29(11):637–647, November 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gudmundsdottir:2006:CSS

[GBGB06] B. K. Gudmundsdóttir, B. Björnsdóttir, S. Gudmundsdóttir, and S. H. Bambir. A comparative study of susceptibility and induced pathology of cod, Gadus morhua (L.), and halibut, Hippoglossus hippoglossus (L.), following experimental infection with Moritella viscosa. Journal of Fish Diseases, 29(8): 481–487, August 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Graham:2004:FOH

[GCG^{+}04] D. A. Graham, W. L. Curran, F. Geoghegan, F. McKiernan, and K. L. Foyle. First observation of herpes-like virus particles in northern pike, Esox lucius L., associated with bluespot-like disease in Ireland. Journal of Fish Diseases, 27 (9):543–549, September 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gross:2004:PAN

[GCN04] K. Gross, J. Carson, and B. Nowak. Presence of anti-Neoparamoeba sp. antibodies in Tasmanian cultured Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 27(2):81– 88, February 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Graham:2002:OVP

[GCR^{+}02] D. A. Graham, W. Curran, H. M. Rowley, D. I. Cox, D. Cockerill, S. Campbell, and D. Todd. Observation of virus particles in the spleen, kidney, gills and erythrocytes of Atlantic salmon, Salmo salar L., during a disease outbreak with high mortality. Journal of Fish Diseases, 25(4):227–234, April 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Goldschmidt-Clermont:2008:IBN

[GCWFB08] E. Goldschmidt-Clermont, T. Wahli, J. Frey, and S. E. Burr. Identification of bacteria from the normal flora of perch, Perca fluviatilis L., and evaluation of their inhibitory potential towards Aeromonas species. Journal of Fish Diseases, 31 (5):353–359, May 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Graham:2007:SSA

[GCWR07] D. A. Graham, K. Cherry, C. J. Wilson, and H. M. Rowley. Susceptibility of salmonid alphavirus to a range of chemical disinfectants. Journal of Fish Diseases, 30(5):269–277, May 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Grandes:2000:UAS

[GDG00] J. M. Fregeneda Grandes, M. Fernández Dıez, and J. M. Aller Gancedo. Ultrastructural analysis of Saprolegnia secondary zoospore cyst ornamentation from infected wild brown trout, Salmo trutta L., and river water indicates two distinct morphotypes amongst long-spined isolates. Journal of Fish Diseases, 23(2):147–160, March 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Grandes:2001:EPR

[GDG01] J. M. Fregeneda Grandes, M. Fernández Díez, and J. M. Aller Gancedo. Experimental pathogenicity in rainbow trout, Oncorhynchus mykiss (Walbaum), of two distinct morphotypes of long-spined Saprolegnia isolates obtained from wild brown trout, Salmo trutta L., and river water. Journal of Fish Diseases, 24(6):351–359, July 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Grossel:2003:PNS

[GDHM03] G. W. Grossel, I. Dykova, J. Handlinger, and B. L. Munday. Pentacapsula neurophila sp.n. (Multivalvulida) from the central nervous system of striped trumpeter, Latris lineata (Forster). Journal of Fish Diseases, 26(6):315–320, June 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gridelli:2003:GLS

[GDP^{+}03] S. Gridelli, A. Diana, A. Parmeggiani, M. Cipone, and R. Preziosi. Goitre in large and small spotted dogfish, Scyliorhinus stellaris (L.) and Scyliorhinus canicula (L.). Journal of Fish Diseases, 26(11–12):687–690, November 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Garver:2007:FDC

[GDR^{+}07] K. A. Garver, A. G. Dwilow, J. Richard, T. F. Booth, D. R. Beniac, and B. W. Souter. First detection and confirmation of spring viraemia of carp virus in common carp, Cyprinus carpio L., from Hamilton Harbour, Lake Ontario, Canada. Journal of Fish Diseases, 30(11):665–671, November 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gustafson:2008:EDT

[GEB^{+}08] L. Gustafson, S. Ellis, D. Bouchard, T. Robinson, F. Marenghi, J. Warg, and C. Giray. Estimating diagnostic test accuracy for infectious salmon anaemia virus in Maine, USA. Journal of Fish Diseases, 31(2):117–125, February 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gahlawat:2009:SLM

[GEC09] S. K. Gahlawat, A. E. Ellis, and B. Collet. A sensitive loopmediated isothermal amplification (LAMP) method for detection of Renibacterium salmoninarum, causative agent of bacterial kidney disease in salmonids. Journal of Fish Diseases, 32(6):491–497, June 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gustafson:2006:EEB

[GER^{+}06] L. Gustafson, S. Ellis, T. Robinson, F. Marenghi, and R. Endris. Efficacy of emamectin benzoate against sea lice infestations of Atlantic salmon, Salmo salar L.: evaluation in the absence of an untreated contemporary control. Journal of Fish Diseases, 29(10):621–627, October 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gustafson:2007:SNS

[GER^{+}07] L. Gustafson, S. Ellis, T. Robinson, F. Marenghi, P. Merrill, L. Hawkins, C. Giray, and B. Wagner. Spatial and non-spatial risk factors associated with cage-level distribution of infectious salmon anaemia at three Atlantic salmon, Salmo salar L., farms in Maine, USA. Journal of Fish Diseases, 30(2): 101–109, February 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gobeli:2009:PCS

[GGCFB09] S. Gobeli, E. Goldschmidt-Clermont, J. Frey, and S. E. Burr. Pseudomonas chlororaphis strain JF3835 reduces mortality of juvenile perch, Perca fluviatilis L., caused by Aeromonas sobria. Journal of Fish Diseases, 32(7):597–602, July 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gudmundsdottir:2003:AEP

[GHBW03] B. K. Gudmundsdóttir, Í. Hvanndal, B. Björnsdóttir, and U. Wagner. Analysis of exotoxins produced by atypical isolates of Aeromonas salmonicida, by enzymatic and serological methods. Journal of Fish Diseases, 26(1):15–29, January 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Graham:2003:RIB

[GJR^{+}03] D. A. Graham, V. A. Jewhurst, H. M. Rowley, M. F. McLoughlin, and D. Todd. A rapid immunoperoxidase-based virus neutralization assay for salmonid alphavirus used for a serological survey in Northern Ireland. Journal of Fish Diseases, 26(7):407–413, July 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Graham:2005:LSS

[GJR^{+}05] D. A. Graham, V. A. Jewhurst, H. M. Rowley, M. F. McLoughlin, H. Rodger, and D. Todd. Longitudinal serological surveys of Atlantic salmon, Salmo salar L., using a rapid immunoperoxidase-based neutralization assay for salmonid alphavirus. Journal of Fish Diseases, 28(6):373–379, June 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gollock:2005:EEA

[GKB05] M. J. Gollock, C. R. Kennedy, and J. A. Brown. European eels, Anguilla anguilla (L.), infected with Anguillicola crassus exhibit a more pronounced stress response to severe hypoxia than uninfected eels. Journal of Fish Diseases, 28(7):429– 436, July 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gibson-Kueh:2004:NSP

[GKCF04] S. Gibson-Kueh, M. Crumlish, and H. W. Ferguson. A novel ‘skinny pot-belly’ disease in Asian seabass fry, Lates calcarifer (Bloch). Journal of Fish Diseases, 27(12):731–735, December 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gibson-Kueh:2004:SID

[GKNLN^{+}04] S. Gibson-Kueh, G. H. Ngoh-Lim, P. Netto, J. Kurita, K. Nakajima, and M. L. Ng. A systemic iridoviral disease in mullet, Mugil cephalus L., and tiger grouper, Epinephelus fuscoguttatus Forsskål: a first report and study. Journal of Fish Diseases, 27(12):693–699, December 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gjessing:2009:NPS

[GKOF09] M. C. Gjessing, A. Kvellestad, K. Ottesen, and K. Falk. Nodavirus provokes subclinical encephalitis and retinochoroiditis in adult farmed Atlantic cod, Gadus morhua L. Journal of Fish Diseases, 32(5):421–431, May 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gunimaladevi:2004:DKH

[GKVS04] I. Gunimaladevi, T. Kono, M. N. Venugopal, and M. Sakai. Detection of koi herpesvirus in common carp, Cyprinus carpio L., by loop-mediated isothermal amplification. Journal of Fish Diseases, 27(10):583–589, October 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Grabowski:2004:SMA

[GLC04] L. D. Grabowski, S. E. LaPatra, and K. D. Cain. Systemic and mucosal antibody response in tilapia, Oreochromis niloticus (L.), following immunization with Flavobacterium columnare. Journal of Fish Diseases, 27(10):573–581, October 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gudmundsdottir:2003:PAA

[GLMG03] S. Gudmundsdóttir, S. Lange, B. Magnadóttir, and B. K. Gudmundsdóttir. Protection against atypical furunculosis in Atlantic halibut, Hippoglossus hippoglossus (L.); comparison of a commercial furunculosis vaccine and an autogenous vaccine. Journal of Fish Diseases, 26(6):331–338, June 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Griffin:2007:SCC

[GM07] B. R. Griffin and A. J. Mitchell. Susceptibility of channel catfish, Ictalurus punctatus (Rafinesque), to Edwardsiella ictaluri challenge following copper sulphate exposure. Journal of Fish Diseases, 30(10):581–585, October 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gagne:2007:IVH

[GMB^{+}07] N. Gagné, A-M. MacKinnon, L. Boston, B. Souter, M. Cook-Versloot, S. Griffiths, and G. Olivier. Isolation of viral haemorrhagic septicaemia virus from mummichog, stickleback, striped bass and brown trout in eastern Canada. Journal of Fish Diseases, 30(4):213–223, April 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gross:2004:ASS

[GMBN04] K. A. Gross, R. N. Morrison, R. Butler, and B. F. Nowak. Atlantic salmon, Salmo salar L., previously infected with Neoparamoeba sp. are not resistant to re-infection and have suppressed phagocyte function. Journal of Fish Diseases, 27 (1):47–56, January 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gahlawat:2004:NDT

[GME04] S. K. Gahlawat, E. S. Munro, and A. E. Ellis. A nondestructive test for detection of IPNV-carriers in Atlantic halibut, Hippoglossus hippoglossus (L.). Journal of Fish Diseases, 27(4):233–239, April 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gray:2002:DKH

[GML^{+}02] W. L. Gray, L. Mullis, S. E. LaPatra, J. M. Groff, and A. Goodwin. Detection of koi herpesvirus DNA in tissues of infected fish. Journal of Fish Diseases, 25(3):171–178, March 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Guselle:2006:IAG

[GMS06] N. J. Guselle, R. J. F. Markham, and D. J. Speare. Intraperitoneal administration of β-1,3/1,6-glucan to rainbow trout, Oncorhynchus mykiss (Walbaum), protects against Loma salmonae. Journal of Fish Diseases, 29(6):375–381, June 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Guselle:2007:TIA

[GMS07] N. J. Guselle, R. J. F. Markham, and D. J. Speare. Timing of intraperitoneal administration of β-1,3/1,6 glucan to rainbow trout, Oncorhynchus mykiss (Walbaum), affects protection against the microsporidian Loma salmonae. Journal of Fish Diseases, 30(2):111–116, February 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gregory:2009:EIA

[GMS^{+}09] A. Gregory, L. A. Munro, M. Snow, K. L. Urquhart, A. G. Murray, and R. S. Raynard. An experimental investigation on aspects of infectious salmon anaemia virus (ISAV) infection dynamics in seawater Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 32(6):481–489, June 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gregory:2007:DIP

[GMW^{+}07] A. Gregory, L. A. Munro, I. S. Wallace, N. Bain, and R. S. Raynard. Detection of infectious pancreatic necrosis virus (IPNV) from the environment in the vicinity of IPNVinfected Atlantic salmon farms in Scotland. Journal of Fish Diseases, 30(10):621–630, October 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Goodwin:2006:SFL

[Goo06] A. E. Goodwin. Steatitis, fin loss and skin ulcers of channel catfish, Ictalurus punctatus (Rafinesque), fingerlings fed salmonid diets. Journal of Fish Diseases, 29(1):61–64, January 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gonzalez:2004:EAV

[GOS04] S. F. Gonzalez, C. R. Osorio, and Y. Santos. Evaluation of the AQUARAPID-Va, AQUAEIA-Va and dot-blot assays for the detection of Vibrio anguillarum in fish tissues. Journal of Fish Diseases, 27(11):617–621, November 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gross:2005:CII

[GPB^{+}05] K. A. Gross, M. D. Powell, R. Butler, R. N. Morrison, and B. F. Nowak. Changes in the innate immune response of Atlantic salmon, Salmo salar L., exposed to experimental infection with Neoparamoeba sp. Journal of Fish Diseases, 28(5):293–299, May 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Goodwin:2005:STL

[GPN05] A. E. Goodwin, E. Park, and B. F. Nowak. Successful treatment of largemouth bass, Micropterus salmoides (L.), with epitheliocystis hyperinfection. Journal of Fish Diseases, 28 (10):623–625, October 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gottesfeld:2009:SLL

[GPRCH09] A. S. Gottesfeld, B. Proctor, L. D. Rolston, and C. Carr-Harris. Sea lice, Lepeophtheirus salmonis, transfer between wild sympatric adult and juvenile salmon on the north coast of British Columbia, Canada. Journal of Fish Diseases, 32 (1):45–57, January 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gupta:2008:MLD

[GPS^{+}08] S. K. Gupta, A. K. Pal, N. P. Sahu, R. Dalvi, V. Kumar, and S. C. Mukherjee. Microbial levan in the diet of Labeo rohita Hamilton juveniles: effect on non-specific immunity and histopathological changes after challenge with Aeromonas hydrophila. Journal of Fish Diseases, 31(9):649–657, September 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Graham:2007:FLC

[GRF^{+}07] D. A. Graham, H. M. Rowley, E. Fringuelli, G. Bovo, A. Manfrin, M. F. McLoughlin, C. Zarza, M. Khalili, and D. Todd. First laboratory confirmation of salmonid alphavirus infection in Italy and Spain. Journal of Fish Diseases, 30(9):569–572, September 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Govett:2004:GMG

[GRL04] P. D. Govett, D. S. Rotstein, and G. A. Lewbart. Gill metaplasia in a goldfish, Carassius auratus auratus (L.). Journal of Fish Diseases, 27(7):419–423, July 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Graham:2003:FIS

[GRW^{+}03] D. A. Graham, H. M. Rowley, I. W. Walker, J. H. Weston, E. J. Branson, and D. Todd. First isolation of sleeping disease virus from rainbow trout, Oncorhynchus mykiss (Walbaum), in the United Kingdom. Journal of Fish Diseases, 26(11–12): 691–694, November 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gomez:2004:PBD

[GSM^{+}04] D. K. Gomez, J. Sato, K. Mushiake, T. Isshiki, Y. Okinaka, and T. Nakai. PCR-based detection of betanodaviruses from cultured and wild marine fish with no clinical signs. Journal of Fish Diseases, 27(10):603–608, October 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Goodwin:2009:HHN

[GSMM09] A. E. Goodwin, J. Sadler, G. E. Merry, and E. N. Marecaux. Herpesviral haematopoietic necrosis virus (CyHV-2) infection: case studies from commercial goldfish farms. Journal of Fish Diseases, 32(3):271–278, March 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Graham:2007:BPS

[GSW^{+}07] D. A. Graham, C. Staples, C. J. Wilson, H. Jewhurst, K. Cherry, A. Gordon, and H. M. Rowley. Biophysical properties of salmonid alphaviruses: influence of temperature and pH on virus survival. Journal of Fish Diseases, 30(9):533–543, September 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gratacap:2009:LEP

[GTBA09] R. M. L. Gratacap, K. D. Thompson, I. R. Bricknell, and A. Adams. Lipopolysaccharide extraction: a phenol alternative. Journal of Fish Diseases, 32(9):811–814, September 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Garcia:2006:IPN

[GUE06] J. Garcia, K. Urquhart, and A. E. Ellis. Infectious pancreatic necrosis virus establishes an asymptomatic carrier state in kidney leucocytes of juvenile Atlantic cod, Gadus morhua L. Journal of Fish Diseases, 29(7):409–413, July 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gil:2000:FIS

[GVGR00] P. Gil, J. Vivas, C. S. Gallardo, and L. A. Rodrıguez. First isolation of Staphylococcus warneri, from diseased rainbow trout, Oncorhynchus mykiss (Walbaum), in Northwest Spain. Journal of Fish Diseases, 23(4):295–298, July 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Graff:2002:MSE

[GWF^{+}02] I. E. Graff, R. Waagbø, S. Fivelstad, C. Vermeer, Ø. Lie, and A. K. Lundebye. A multivariate study on the effects of dietary vitamin K, vitamin D_{3} and calcium, and dissolved carbon dioxide on growth, bone minerals, vitamin status and health performance in smolting Atlantic salmon Salmo salar L. Journal of Fish Diseases, 25(10):599–614, October 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Graham:2008:CCS

[GWJR08] D. A. Graham, C. Wilson, H. Jewhurst, and H. Rowley. Cultural characteristics of salmonid alphaviruses — influence of cell line and temperature. Journal of Fish Diseases, 31(11): 859–868, November 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hallett:2002:MCT

[HAEM02] S. L. Hallett, S. D. Atkinson, and M. El-Matbouli. Molecular characterisation of two aurantiactinomyxon (Myxozoa) phenotypes reveals one genotype. Journal of Fish Diseases, 25 (10):627–631, October 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hill:2009:INS

[HBB09] J. E. Hill, J. C. F. Baiano, and A. C. Barnes. Isolation of a novel strain of Bacillus pumilus from penaeid shrimp that is inhibitory against marine pathogens. Journal of Fish Diseases, 33(12):1007–1016, December 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Holzer:2006:KTS

[HBCS^{+}06] A. S. Holzer, I. Blasco-Costa, V. L. Sarabeev, M. O. Ovcharenko, and J. A. Balbuena. Kudoa trifolia sp. n. — molecular phylogeny suggests a new spore morphology and unusual tissue location for a well-known genus. Journal of Fish Diseases, 29(12):743–755, December 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Harriff:2007:EEZ

[HBK07] M. J. Harriff, L. E. Bermudez, and M. L. Kent. Experimental exposure of zebrafish, Danio rerio (Hamilton), to Mycobacterium marinum and Mycobacterium peregrinum reveals the gastrointestinal tract as the primary route of infection: a potential model for environmental mycobacterial infection. Journal of Fish Diseases, 30(10):587–600, October 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hayward:2009:SES

[HBN09] C. J. Hayward, N. J. Bott, and B. F. Nowak. Seasonal epizootics of sea lice, Caligus spp., on southern bluefin tuna, Thunnus maccoyii (Castelnau), in a long-term farming trial. Journal of Fish Diseases, 32(1):101–106, January 2009. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hu:2004:IAF

[HCFM04] G-B. Hu, R-S. Cong, T-J. Fan, and X-G. Mei. Induction of apoptosis in a flounder gill cell line by lymphocystis disease virus infection. Journal of Fish Diseases, 27(11):657–662, November 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hogge:2008:RMC

[HCJ08] C. I. Hogge, M. R. Campbell, and K. A. Johnson. Redescription and molecular characterization of Myxobolus kisutchi Yasutake & Wood, 1957. Journal of Fish Diseases, 31(9):707– 712, September 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hammell:2005:MPI

[HD05a] K. L. Hammell and I. R. Dohoo. Mortality patterns in infectious salmon anaemia virus outbreaks in New Brunswick, Canada. Journal of Fish Diseases, 28(11):639–650, November 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hammell:2005:RFA

[HD05b] K. L. Hammell and I. R. Dohoo. Risk factors associated with mortalities attributed to infectious salmon anaemia virus in New Brunswick, Canada. Journal of Fish Diseases, 28(11): 651–661, November 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hine:2002:ESD

[HDP^{+}02] P. M. Hine, B. K. Diggles, M. J. D. Parsons, A. Pringle, and B. Bull. The effects of stressors on the dynamics of Bonamia exitiosus Hine, Cochennec-Laureau & Berthe, infections in flat oysters Ostrea chilensis (Philippi). Journal of Fish Diseases, 25(9):545–554, September 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hernandez:2009:SRT

[HFI09] E. Hernández, J. Figueroa, and C. Iregui. Streptococcosis on a red tilapia, Oreochromis sp., farm: a case study. Journal of Fish Diseases, 32(3):247–252, March 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hontzeas:2001:IEM

[HFW01] N. Hontzeas, S. Feng, and P. T. K. Woo. Inhibitory effects of a monoclonal antibody (MAb-001) on in vitro oxygen consumption and multiplication of the pathogenic haemoflagellate, Cryptobia salmositica Katz. Journal of Fish Diseases, 24 (7):391–398, September 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hong:2008:ABI

[HGH^{+}08] J-R. Hong, B-J. Guan, G. M. Her, O. Evensen, N. Santi, and J-L. Wu. Aquatic birnavirus infection activates the transcription factor NF-κ B via tyrosine kinase signalling leading to cell death. Journal of Fish Diseases, 31(6):451–460, June 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hershberger:2007:LPH

[HGP^{+}07] P. K. Hershberger, J. Gregg, C. Pacheco, J. Winton, J. Richard, and G. Traxler. Larval Pacific herring, Clupea pallasii (Valenciennes), are highly susceptible to viral haemorrhagic septicaemia and survivors are partially protected after their metamorphosis to juveniles. Journal of Fish Diseases, 30(8):445–458, August 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hernandez-Herrera:2007:PSS

[HHCBR^{+}07] R. I. Hernandez-Herrera, V. Chappe-Bonnichon, P. Roch, J. Sri Widada, and J-R. Bonami. Partial susceptibility of the SSN-1 fish cell line to a crustacean virus: a defective replication study. Journal of Fish Diseases, 30(11):673–679, November 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hsu:2005:EPI

[HHL^{+}05] J-P. Hsu, C. Huang, C-M. Liao, S-L. Hsuan, H-H. Hung, and M-S. Chien. Engulfed pathogen-induced apoptosis in haemocytes of giant freshwater prawn, Macrobrachium rosenbergii. Journal of Fish Diseases, 28(12):729–735, December 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hong:2005:ABI

[HHW05] J-R. Hong, L-J. Huang, and J-L. Wu. Aquatic birnavirus induces apoptosis through activated caspase-8 and -3 in a zebrafish cell line. Journal of Fish Diseases, 28(3):133–140, March 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Haramoto:2009:DVC

[HKK^{+}09a] E. Haramoto, M. Kitajima, H. Katayama, T. Ito, and S. Ohgaki. Development of virus concentration methods for detection of koi herpesvirus in water. Journal of Fish Diseases, 32(3):297–300, March 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hossain:2009:LDV

[HKK^{+}09b] M. Hossain, S-R. Kim, S-I. Kitamura, D-W. Kim, S-J. Jung, T. Nishizawa, M. Yoshimizu, and M-J. Oh. Lymphocystis disease virus persists in the epidermal tissues of olive flounder, Paralichthys olivaceus (Temminch & Schlegel), at low temperatures. Journal of Fish Diseases, 32(8):699–703, August 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Haramoto:2007:DKH

[HKKO07] E. Haramoto, M. Kitajima, H. Katayama, and S. Ohgaki. Detection of koi herpesvirus DNA in river water in Japan. Journal of Fish Diseases, 30(1):59–61, January 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Han:2006:PET

[HKL^{+}06] H-J. Han, D-H. Kim, D-C. Lee, S-M. Kim, and S-I. Park. Pathogenicity of Edwardsiella tarda to olive flounder, Paralichthys olivaceus (Temminck & Schlegel). Journal of Fish Diseases, 29(10):601–609, October 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hirvela-Koski:2006:AGR

[HKPKS06] V. Hirvelä-Koski, T. Pohjanvirta, P. Koski, and A. Sukura. Atypical growth of Renibacterium salmoninarum in subclinical infections. Journal of Fish Diseases, 29(1):21–29, January 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Han:2006:CAD

[HKV^{+}06] H-J. Han, H. Kondo, G. Vivekanandhan, I. Hirono, and T. Aoki. Cloning of ATP-dependent protease ClpXP genes in Aeromonas veronii. Journal of Fish Diseases, 29(11):691– 695, November 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Huang:2000:IAT

[HLC00] S-L. Huang, I-C. Liao, and S-N. Chen. Induction of apoptosis in tilapia, Oreochromis aureus Steindachner, and in TO-2 cells by Staphylococcus epidermidis. Journal of Fish Diseases, 23(6):363–368, November 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Huang:2007:HEC

[HLWH07] J. N. Huang, L. Lin, S. P. Weng, and J. G. He. High expression of capsid protein of red-spotted grouper nervous necrosis virus in an avian cell line requires viral RNA2 non-coding regions. Journal of Fish Diseases, 30(7):439–444, July 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Heinecke:2007:MSG

[HMB07] R. D. Heinecke, T. Martinussen, and K. Buchmann. Microhabitat selection of Gyrodactylus salaris Malmberg on different salmonids. Journal of Fish Diseases, 30(12):733–743, December 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hashimoto:2008:PKC

[HMO^{+}08] E. Hashimoto, T. Miyadai, M. Ohtani, S-I. Kitamura, and M-J. Oh. Proliferation of kuchijirosho causative agent in a fugu-derived cell line. Journal of Fish Diseases, 31(6):443– 449, June 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Horton:2003:PHA

[HO03] T. Horton and B. Okamura. Post-haemorrhagic anaemia in sea bass, Dicentrarchus labrax (L.), caused by blood feeding of Ceratothoa oestroides (Isopoda: Cymothoidae). Journal of Fish Diseases, 26(7):401–406, July 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hirose:2009:MCT

[HOA09] E. Hirose, S-I. Ohtake, and K. Azumi. Morphological characterization of the tunic in the edible ascidian, Halocynthia roretzi (Drasche), with remarks on ‘soft tunic syndrome’ in aquaculture. Journal of Fish Diseases, 32(5):433–445, May 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hoffman:2000:B

[Hof00] Glenn Lyle Hoffman. BiographyBiography. Journal of Fish Diseases, 23(1):87–89, January 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Heuch:2009:TSV

[HOM^{+}09] P. A. Heuch, R. S. Olsen, R. Malkenes, C. W. Revie, G. Gettinby, M. Baillie, F. Lees, and B. Finstad. Temporal and spatial variations in lice numbers on salmon farms in the Hardanger fjord 2004–06. Journal of Fish Diseases, 32(1): 89–100, January 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hershberger:2008:IIS

[HPG08] P. K. Hershberger, C. A. Pacheco, and J. L. Gregg. Inactivation of Ichthyophonus spores using sodium hypochlorite and polyvinyl pyrrolidone iodine. Journal of Fish Diseases, 31 (11):853–858, November 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hogfors:2008:IRT

[HPMW08] E. Högfors, K-R. Pullinen, J. Madetoja, and T. Wiklund. Immunization of rainbow trout, Oncorhynchus mykiss (Walbaum), with a low molecular mass fraction isolated from Flavobacterium psychrophilum. Journal of Fish Diseases, 31 (12):899–911, December 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Heuch:2003:CEP

[HRG03] P. A. Heuch, C. W. Revie, and G. Gettinby. A comparison of epidemiological patterns of salmon lice, Lepeophtheirus salmonis, infections on farmed Atlantic salmon, Salmo salar L., in Norway and Scotland. Journal of Fish Diseases, 26(9): 539–551, September 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hayes:2007:PAP

[HSD07] P. M. Hayes, N. J. Smit, and A. J. Davies. Pathology associated with parasitic juvenile gnathiids feeding on the puffadder shyshark, Haploblepharus edwardsii (Voight). Journal of Fish Diseases, 30(1):55–58, January 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hossain:2008:PAL

[HSK^{+}08] M. Hossain, J-Y. Song, S-I. Kitamura, S-J. Jung, and M-J. Oh. Phylogenetic analysis of lymphocystis disease virus from tropical ornamental fish species based on a major capsid protein gene. Journal of Fish Diseases, 31(6):473–479, June 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Holzer:2003:TRS

[HSW03] A. S. Holzer, C. Sommerville, and R. Wootten. Tracing the route of Sphaerospora truttae from the entry locus to the target organ of the host, Salmo salar L., using an optimized and specific in situ hybridization technique. Journal of Fish Diseases, 26(11–12):647–655, November 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Huang:2001:DNB

[HTC^{+}01] Huang, Tan, Chang, Munday, Mathew, Ngoh, and Kwang. Detection of nodavirus in barramundi, Lates calcarifer (Bloch), using recombinant coat protein-based ELISA and RT–PCR. Journal of Fish Diseases, 24(3):135–141, March 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Humphrey:2007:SPF

[Hum07] J. D. Humphrey. Systemic pathology of fish: A text and atlas of normal tissues in teleosts and their response in disease. Journal of Fish Diseases, 30(6):381–382, June 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hsieh:2006:SHR

[HWT^{+}06] C-Y. Hsieh, Z-B. Wu, M-C. Tung, C. Tu, S-P. Lo, T-C. Chang, C-D. Chang, S-C. Chen, Y-C. Hsieh, and S-S. Tsai. In situ hybridization and RT-PCR detection of Macrobrachium rosenbergii nodavirus in giant freshwater prawn, Macrobrachium rosenbergii (de Man), in Taiwan. Journal of Fish Diseases, 29(11):665–671, November 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

He:2000:SDC

[HWZ^{+}00] J. G. He, S. P. Wang, K. Zeng, Z. J. Huang, and S-M. Chan. Systemic disease caused by an iridovirus-like agent in cultured mandarinfish, Siniperca chuatsi (Basilewsky), in China. Journal of Fish Diseases, 23(3):219–222, May 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Irianto:2002:PA

[IA02a] A. Irianto and B. Austin. Probiotics in aquaculture. Journal of Fish Diseases, 25(11):633–642, November 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Irianto:2002:UPC

[IA02b] A. Irianto and B. Austin. Use of probiotics to control furunculosis in rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 25(6):333–342, June 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Irianto:2003:UDP

[IA03] A. Irianto and B. Austin. Use of dead probiotic cells to control furunculosis in rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 26(1):59–62, January 2003.

CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Izumi:2005:DIF

[IFA05] S. Izumi, H. Fujii, and F. Aranishi. Detection and identification of Flavobacterium psychrophilum from gill washings and benthic diatoms by PCR-based sequencing analysis. Journal of Fish Diseases, 28(9):559–564, September 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Iwanowicz:2000:SRV

[IGH00] L. R. Iwanowicz, A. E. Goodwin, and N. Heil. A small RNA virus isolated from apparently healthy wild sandbar shiners, Notropis scepticus (Jordan & Gilbert). Journal of Fish Diseases, 23(5):349–352, September 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Iwanowicz:2001:ERG

[IGH01] Iwanowicz, Goodwin, and Harshbarger. Embryonal rhabdomyosarcoma of the giant gourami, Colisa fasciata (Bloch & Schneider). Journal of Fish Diseases, 24(3):177–179, March 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Itano:2006:EIN

[IKKS06] T. Itano, H. Kawakami, T. Kono, and M. Sakai. Experimental induction of nocardiosis in yellowtail, Seriola quinqueradiata Temminck & Schlegel by artificial challenge. Journal of Fish Diseases, 29(9):529–534, September 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Izumi:2003:NSF

[ILAW03] S. Izumi, H. Liu, F. Aranishi, and H. Wakabayashi. A novel serotype of Flavobacterium psychrophilum detected using antiserum against an isolate from amago, Oncorhynchus masou rhodurus Jordan & Gilbert, in Japan. Journal of Fish Diseases, 26(11–12):677–680, November 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Iwamoto:2001:CCC

[IMAN01] T. Iwamoto, K. Mori, M. Arimoto, and T. Nakai. A combined cell-culture and RT–PCR method for rapid detection of piscine nodaviruses. Journal of Fish Diseases, 24(4):231– 236, May 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Isshiki:2004:DMB

[INKS04] T. Isshiki, T. Nagano, K. Kanehira, and S. Suzuki. Distribution of marine birnavirus in cultured marine fish species from Kagawa Prefecture, Japan. Journal of Fish Diseases, 27(2): 89–98, February 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Iwamoto:2001:TSJ

[INM^{+}01] Iwamoto, Nakai, Mori, Arimoto, Mise, and Furusawa. Transfection of striped jack nervous necrosis virus (SJNNV) RNA into fish cells. Journal of Fish Diseases, 24(3):185–188, March 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Izumi:2007:PRG

[IOK^{+}07] S. Izumi, S. Ouchi, T. Kuge, H. Arai, T. Mito, H. Fujii, F. Aranishi, and A. Shimizu. PCR-RFLP genotypes associated with quinolone resistance in isolates of Flavobacterium psychrophilum. Journal of Fish Diseases, 30(3):141– 147, March 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Iglesias:2001:MPI

[IPA^{+}01] Iglesias, Paramá, Alvarez, Leiro, and Sanmartín. Myxobolid parasites infecting the gills of the Iberian nase, Chondrostoma polylepis (Steindachner), in Galicia (NW Spain), with a description of Myxobolus gallaicus sp. nov. Journal of Fish Diseases, 24(3):125–134, March 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Irianto:2003:OAF

[IRA03] A. Irianto, P. A. W. Robertson, and B. Austin. Oral administration of formalin-inactivated cells of Aeromonas hydrophila A3-51 controls infection by atypical A. salmonicida in goldfish, Carassius auratus (L.). Journal of Fish Diseases, 26(2): 117–120, February 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Itoh:2004:OII

[ITK^{+}04] N. Itoh, K. L. Tun, H. Komiyama, N. Ueki, and K. Ogawa. An ovarian infection in the Iwagaki oyster, Crassostrea nippona, with the protozoan parasite Marteilioides chungmuensis. Journal of Fish Diseases, 27(5):311–314, May 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Izumi:2007:IDP

[IYS^{+}07] S. Izumi, M. Yamamoto, K. Suzuki, A. Shimizu, and F. Aranishi. Identification and detection of Pseudomonas plecoglossicida isolates with PCR primers targeting the gyrB region. Journal of Fish Diseases, 30(7):391–397, July 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jamett:2001:CMA

[JAM^{+}01] A. Jamett, J. Aguayo, A. Miquel, I. Muller, R. Arriagada, M. I. Becker, P. Valenzuela, and L. O. Burzio. Characteristics of monoclonal antibodies against Piscirickettsia salmonis. Journal of Fish Diseases, 24(4):205–215, May 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jayasree:2009:IIC

[Jay09] S. Jayasree. Identification of immune cells interacting with Vibrio spp. and its in vitro post-phagocytic killing mechanism of haemocytes in the penaeid shrimp, Penaeus indicus H. Milne Edwards. Journal of Fish Diseases, 32(4):359–365, April 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jeffery:2007:ICH

[JBB^{+}07] K. R. Jeffery, K. Bateman, A. Bayley, S. W. Feist, J. Hulland, C. Longshaw, D. Stone, G. Woolford, and K. Way. Isolation of a cyprinid herpesvirus 2 from goldfish, Carassius auratus (L.), in the UK. Journal of Fish Diseases, 30(11):649–656, November 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jones:2002:IHP

[JD02] S. R. M. Jones and S. C. Dawe. Ichthyophonus hoferi Plehn & Mulsow in British Columbia stocks of Pacific herring, Clupea pallasi Valenciennes, and its infectivity to Chinook salmon, Oncorhynchus tshawytscha (Walbaum). Journal of Fish Diseases, 25(7):415–421, July 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Johnston:2008:GHC

[JDBN08] C. J. Johnston, M. R. Deveney, T. Bayly, and B. F. Nowak. Gross and histopathological characteristics of two lipomas and a neurofibrosarcoma detected in aquacultured southern bluefin tuna, Thunnus maccoyii (Castelnau), in South Australia. Journal of Fish Diseases, 31(4):241–247, April 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jonstrup:2009:FEV

[JGK^{+}09] S. P. Jonstrup, T. Gray, S. Kahns, H. F. Skall, M. Snow, and N. J. Olesen. FishPathogens.eu/ vhsv: a user-friendly viral haemorrhagic septicaemia virus isolate and sequence database. Journal of Fish Diseases, 32(11):925–929, November 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Johansen:2004:SSP

[JGS^{+}04] R. Johansen, S. Grove, A. K. Svendsen, I. Modahl, and B. Dannevig. A sequential study of pathological findings in Atlantic halibut, Hippoglossus hippoglossus (L.), throughout one year after an acute outbreak of viral encephalopathy and retinopathy. Journal of Fish Diseases, 27(6):327–341, June 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Johnson:2001:EMR

[JH01] K. A. Johnson and J. A. Heindel. Efficacy of manual removal and ivermectin gavage for control of Salmincola californiensis (Wilson) infestation of Chinook salmon, Oncorhynchus tshawytscha (Walbaum), captive broodstocks. Journal of Fish Diseases, 24(4):197–203, May 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jorgensen:2008:OGW

[JHKB08] L. Von Gersdorff Jørgensen, R. D. Heinecke, P. Kania, and K. Buchmann. Occurrence of gyrodactylids on wild Atlantic salmon, Salmo salar L., in Danish rivers. Journal of Fish Diseases, 31(2):127–134, February 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jones:2008:EDR

[JKB08] S. Jones, E. Kim, and W. Bennett. Early development of resistance to the salmon louse, Lepeophtheirus salmonis (Krøyer), in juvenile pink salmon, Oncorhynchus gorbuscha (Walbaum). Journal of Fish Diseases, 31(8):591–600, August 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jones:2006:EIL

[JKD06] S. Jones, E. Kim, and S. Dawe. Experimental infections with Lepeophtheirus salmonis (Krøyer) on threespine sticklebacks, Gasterosteus aculeatus L., and juvenile Pacific salmon, Oncorhynchus spp. Journal of Fish Diseases, 29(8):489–495, August 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jansson:2008:DBK

[JLSA08] E. Jansson, L. Lindberg, E. Säker, and A. Aspán. Diagnosis of bacterial kidney disease by detection of Renibacterium salmoninarum by real-time PCR. Journal of Fish Diseases, 31(10):755–763, October 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jakob:2009:MHA

[JNS^{+}09] E. Jakob, H. Neuhaus, D. Steinhagen, B. Luckhardt, and R. Hanel. Monitoring of Herpesvirus anguillae (HVA) infections in European eel, Anguilla anguilla (L.), in northern Germany. Journal of Fish Diseases, 32(6):557–561, June 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jung:2000:ILI

[JO00] S. J. Jung and M. J. Oh. Iridovirus-like infection associated with high mortalities of striped beakperch, Oplegnathus fasciatus (Temminck et Schlegel), in southern coastal areas of the Korean peninsula. Journal of Fish Diseases, 23(3):223– 226, May 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Johansson:2009:DIP

[JO09] T. Johansson and N. J. Olesen. Detection of infectious pancreatic necrosis virus from rainbow trout, Oncorhynchus mykiss (Walbaum), using the macrophage lysis method. Journal of Fish Diseases, 32(6):563–566, June 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jutfelt:2006:TVA

[JOG^{+}06] F. Jutfelt, R. E. Olsen, J. Glette, E. Ringø, and K. Sundell. Translocation of viable Aeromonas salmonicida across the intestine of rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 29(5):255–262, May 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jonsson:2004:DTU

[Jön04] K. I. Jönsson. On the disparate terminological use of the concept cryptobiosis. Journal of Fish Diseases, 27(3):175– 176, March 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jayaprakash:2006:IPS

[JPPS06] N. S. Jayaprakash, S. S. Pai, R. Philip, and I. S. B. Singh. Isolation of a pathogenic strain of Vibrio alginolyticus from necrotic larvae of Macrobrachium rosenbergii (de Man). Journal of Fish Diseases, 29(3):187–191, March 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jacobs:2009:RMM

[JSBK09] J. M. Jacobs, C. B. Stine, A. M. Baya, and M. L. Kent. A review of mycobacteriosis in marine fish. Journal of Fish Diseases, 32(2):119–130, February 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jutfelt:2008:IAS

[JSG^{+}08] F. Jutfelt, H. Sundh, J. Glette, L. Mellander, B. Thrandur Björnsson, and K. Sundell. The involvement of Aeromonas salmonicida virulence factors in bacterial translocation across the rainbow trout, Oncorhynchus mykiss (Walbaum), intestine. Journal of Fish Diseases, 31(2):141–151, February 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Johansen:2004:CNV

[JST^{+}04] R. Johansen, I. Sommerset, B. Tørud, K. Korsnes, M. J. Hjortaas, F. Nilsen, A. H. Nerland, and B. H. Dannevig. Characterization of nodavirus and viral encephalopathy and retinopathy in farmed turbot, Scophthalmus maximus (L.). Journal of Fish Diseases, 27(10):591–601, October 2004. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jung:2001:PCM

[JTM^{+}01] T. S. Jung, K. D. Thompson, D. J. Morris, A. Adams, and K. Sneddon. The production and characterization of monoclonal antibodies against Photobacterium damselae spp. piscicida and initial observations using immunohistochemistry. Journal of Fish Diseases, 24(2):67–77, February 2001. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jewhurst:2004:DAC

[JTR^{+}04] V. A. Jewhurst, D. Todd, H. M. Rowley, I. W. Walker, J. H. Weston, M. F. McLoughlin, and D. A. Graham. Detection and antigenic characterization of salmonid alphavirus isolates from sera obtained from farmed Atlantic salmon, Salmo salar L., and farmed rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 27(3):143–149, March 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Johnson:2004:IEA

[JZKS04] R. A. Johnson, J. Zabrecky, Y. Kiryu, and J. D. Shields. Infection experiments with Aphanomyces invadans in four species of estuarine fish. Journal of Fish Diseases, 27(5):287– 295, May 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Korkea-aho:2009:FAB

[KaVKT09] T. L. Korkea-aho, A. Vainikka, R. Kortet, and J. Taskinen. Factors affecting between-lake variation in the occurrence of epidermal papillomatosis in roach, Rutilus rutilus (L.). Journal of Fish Diseases, 32(3):263–270, March 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Korkea-aho:2006:FAI

[KaVT06] T. Korkea-aho, A. Vainikka, and J. Taskinen. Factors affecting the intensity of epidermal papillomatosis in populations of roach, Rutilus rutilus (L.), estimated as scale coverage. Journal of Fish Diseases, 29(2):115–122, February 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Knusel:2007:VIV

[KBEJ^{+}07] R. Knüsel, S. M. Bergmann, K. Einer-Jensen, J. Casey, H. Segner, and T. Wahli. Virus isolation vs RT-PCR: which method is more successful in detecting VHSV and IHNV in fish tissue sampled under field conditions? Journal of Fish Diseases, 30(9):559–568, September 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kim:2008:AGR

[KBH^{+}08] D-G. Kim, J-Y. Bae, G-E. Hong, M-K. Min, J-K. Kim, and I-S. Kong. Application of the rpoS gene for the detection of Vibrio anguillarum in flounder and prawn by polymerase chain reaction. Journal of Fish Diseases, 31(9):639–647, September 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Koubkova:2004:RAB

[KBPD04] B. Koubková, V. Baruš, M. Prokeš, and I. Dyková. Raphidascaris acus (Bloch, 1779) larvae infections of the stone loach, Barbatula barbatula (L.), from the River Haná, Czech Republic. Journal of Fish Diseases, 27(2):65–71, February 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kent:2003:TTD

[KBS03] M. L. Kent and J. K. Bishop-Stewart. Transmission and tissue distribution of Pseudoloma neurophilia (Microsporidia) of zebrafish, Danio rerio (Hamilton). Journal of Fish Diseases, 26(7):423–426, July 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kallert:2005:LCH

[KEEM^{+}05] D. M. Kallert, E. Eszterbauer, M. El-Matbouli, C. Erséus, and W. Haas. The life cycle of Henneguya nuesslini Schuberg & Schröder, 1905 (Myxozoa) involves a triactinomyxon-type actinospore. Journal of Fish Diseases, 28(2):71–79, February 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kennedy:2007:PHP

[Ken07] C. R. Kennedy. The pathogenic helminth parasites of eels. Journal of Fish Diseases, 30(6):319–334, June 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Koppang:2000:NPP

[KFM^{+}00] E. O. Koppang, M. Fjølstad, B. Melgård, M. Vigerust, and H. S. ørum. Non-pigment-producing isolates of Aeromonas salmonicida subspecies salmonicida: isolation, identification, transmission and pathogenicity in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 23(1):39–48, January 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kocan:2006:DIP

[KH06] R. Kocan and P. Hershberger. Differences in Ichthyophonus prevalence and infection severity between upper Yukon River and Tanana River Chinook salmon, Oncorhynchus tshawytscha (Walbaum), stocks. Journal of Fish Diseases, 29(8):497–503, August 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Khan:2004:DOM

[Kha04a] R. A. Khan. Disease outbreaks and mass mortality in cultured Atlantic cod, Gadus morhua L., associated with Trichodina murmanica (Ciliophora). Journal of Fish Diseases, 27(3): 181–184, March 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Khoa:2004:FII

[KHA04b] L. V. Khoa, K. Hatai, and T. Aoki. Fusarium incarnatum isolated from black tiger shrimp, Penaeus monodon Fabricius, with black gill disease cultured in Vietnam. Journal of Fish Diseases, 27(9):507–515, September 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kim:2005:DQR

[KHA05] M-J. Kim, I. Hirono, and T. Aoki. Detection of quinoloneresistance genes in Photobacterium damselae subsp. piscicida strains by targeting-induced local lesions in genomes. Journal of Fish Diseases, 28(8):463–471, August 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Koppang:2005:VAG

[KHH^{+}05] E. O. Koppang, E. Haugarvoll, I. Hordvik, L. Aune, and T. T. Poppe. Vaccine-associated granulomatous inflammation and melanin accumulation in Atlantic salmon, Salmo salar L., white muscle. Journal of Fish Diseases, 28(1):13–22, January 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kim:2004:BED

[KHK^{+}04] D-H. Kim, H-J. Han, S-M. Kim, D-C. Lee, and S-I. Park. Bacterial enteritis and the development of the larval digestive tract in olive flounder, Paralichthys olivaceus (Temminck & Schlegel). Journal of Fish Diseases, 27(9):497–505, September 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kocan:2009:ETD

[KHSW09] R. Kocan, P. Hershberger, G. Sanders, and J. Winton. Effects of temperature on disease progression and swimming stamina in Ichthyophonus-infected rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 32 (10):835–843, October 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kongtorp:2006:LSN

[KHTF06] R. T. Kongtorp, M. Halse, T. Taksdal, and K. Falk. Longitudinal study of a natural outbreak of heart and skeletal muscle inflammation in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 29(4):233–244, April 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kobayashi:2004:HSB

[KIIK04] T. Kobayashi, M. Imai, Y. Ishitaka, and Y. Kawaguchi. Histopathological studies of bacterial haemorrhagic ascites of ayu, Plecoglossus altivelis (Temminck & Schlegel). Journal of Fish Diseases, 27(8):451–457, August 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kania:2007:DBP

[KJB07] P. W. Kania, T. R. Jørgensen, and K. Buchmann. Differentiation between a pathogenic and a non-pathogenic form of Gyrodactylus salaris using PCR-RFLP. Journal of Fish Diseases, 30(2):123–126, February 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kim:2002:PAS

[KJC^{+}02] Y. J. Kim, S. J. Jung, T. J. Choi, H. R. Kim, K. V. Rajendran, and M. J. Oh. PCR amplification and sequence analysis of irido-like virus infecting fish in Korea. Journal of Fish Diseases, 25(2):121–124, February 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kim:2007:ESA

[KJL07] T. J. Kim, T. S. Jung, and J. I. Lee. Expression and serological application of a capsid protein of an iridovirus isolated from rock bream, Oplegnathus fasciatus (Temminck & Schlegel). Journal of Fish Diseases, 30(11):691–699, November 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kim:2008:VRB

[KJL08] T. J. Kim, E. J. Jang, and J. I. Lee. Vaccination of rock bream, Oplegnathus fasciatus (Temminck & Schlegel), using a recombinant major capsid protein of fish iridovirus. Journal of Fish Diseases, 31(8):547–551, July 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Korsnes:2009:TTN

[KKD^{+}09] K. Korsnes, E. Karlsbakk, M. Devold, A. H. Nerland, and A. Nylund. Tissue tropism of nervous necrosis virus (NNV) in Atlantic cod, Gadus morhua L., after intraperitoneal challenge with a virus isolate from diseased Atlantic halibut, Hippoglossus hippoglossus (L.). Journal of Fish Diseases, 32(8): 655–665, August 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kongtorp:2004:HSM

[KKT^{+}04] R. T. Kongtorp, A. Kjerstad, T. Taksdal, A. Guttvik, and K. Falk. Heart and skeletal muscle inflammation in Atlantic salmon, Salmo salar L.: a new infectious disease. Journal of Fish Diseases, 27(6):351–358, June 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kim:2007:SVG

[KL07] T. J. Kim and J. I. Lee. Sequence variation in the genes encoding the major capsid protein, the ATPase, and the RNA polymerase 2 (domain 6) of lymphocystis disease virus isolated from Schlegel’s black rockfish, Sebastes schlegelii Hilgendorf. Journal of Fish Diseases, 30(8):501–504, August 2007. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kocan:2006:IIC

[KLG^{+}06] R. Kocan, S. LaPatra, J. Gregg, J. Winton, and P. Hershberger. Ichthyophonus-induced cardiac damage: a mechanism for reduced swimming stamina in salmonids. Journal of Fish Diseases, 29(9):521–527, September 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kiryu:2001:ALK

[KM01] Y. Kiryu and C. M. Moffitt. Acute LD_{50} and kidney histopathology following injection of erythromycin (Erythro-200) and its carrier in spring Chinook salmon, Oncorhynchus tshawytscha (Walbaum). Journal of Fish Diseases, 24(7):409– 416, September 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kamilya:2006:AEM

[KMJM06] D. Kamilya, T. K. Maiti, S. N. Joardar, and B. C. Mal. Adjuvant effect of mushroom glucan and bovine lactoferrin upon Aeromonas hydrophila vaccination in catla, Catla catla (Hamilton). Journal of Fish Diseases, 29(6):331–337, June 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Karatas:2008:RTP

[KMS^{+}08] S. Karatas, J. Mikalsen, T. M. Steinum, T. Taksdal, M. Bordevik, and D. J. Colquhoun. Real time PCR detection of Piscirickettsia salmonis from formalin-fixed paraffin-embedded tissues. Journal of Fish Diseases, 31(10):747–753, October 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kang:2003:ECT

[KOK^{+}03] M. S. Kang, M. J. Oh, Y. J. Kim, K. Kawai, and S. J. Jung. Establishment and characterization of two new cell lines derived from flounder, Paralichthys olivaceus (Temminck & Schlegel). Journal of Fish Diseases, 26(11–12): 657–665, November 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kozinska:2007:DPS

[Koz07] A. Kozińska. Dominant pathogenic species of mesophilic aeromonads isolated from diseased and healthy fish cultured in Poland. Journal of Fish Diseases, 30(5):293–301, May 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kempter:2006:SPC

[KPWK06] J. Kempter, W. Piasecki, K. Wieski, and B. Krawczyk. Systematic position of copepods of the genus Achtheres (Crustacea: Copepoda: Siphonostomatoida) parasitizing perch, Perca fluviatilis L., and zander, Sander lucioperca (L.). Journal of Fish Diseases, 29(2):103–113, February 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kiran:2002:ESD

[KRJO02] R. B. Pramod Kiran, K. V. Rajendran, S. J. Jung, and M. J. Oh. Experimental susceptibility of different life-stages of the giant freshwater prawn, Macrobrachium rosenbergii (de Man), to white spot syndrome virus (WSSV). Journal of Fish Diseases, 25(4):201–207, April 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kaattari:2006:ESM

[KRKS06] I. M. Kaattari, M. W. Rhodes, S. L. Kaattari, and E. B. Shotts. The evolving story of Mycobacterium tuberculosis clade members detected in fish. Journal of Fish Diseases, 29 (9):509–520, September 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kumari:2006:DGP

[KS06] J. Kumari and P. K. Sahoo. Dietary β-1,3 glucan potentiates innate immunity and disease resistance of Asian catfish, Clarias batrachus (L.). Journal of Fish Diseases, 29(2):95–101, February 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kaltner:2007:CGD

[KSKEM07] H. Kaltner, M. Stippl, M. Knaus, and M. El-Matbouli. Characterization of glycans in the developmental stages of Myxobolus cerebralis (Myxozoa), the causative agent of whirling disease. Journal of Fish Diseases, 30(11):637–647, November 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Knuesel:2003:SVD

[KSW03] R. Knuesel, H. Segner, and T. Wahli. A survey of viral diseases in farmed and feral salmonids in Switzerland. Journal of Fish Diseases, 26(3):167–182, March 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kongtorp:2009:SET

[KT09] R. T. Kongtorp and T. Taksdal. Studies with experimental transmission of heart and skeletal muscle inflammation in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 32 (3):253–262, March 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kent:2001:IML

[KWD^{+}01] M. L. Kent, V. Watral, S. C. Dawe, P. Reno, J. R. Heidel, and S. R. M. Jones. Ichthyophonus and Mycobacterium-like bacterial infections in commercially-important rockfish, Sebastes spp., in the eastern North Pacific Ocean. Journal of Fish Diseases, 24(7):427–431, September 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kesarcodi-Watson:2009:TPG

[KWKLG09] A. Kesarcodi-Watson, H. Kaspar, M. J. Lategan, and L. Gibson. Two pathogens of Greenshell^{T M} mussel larvae, Perna canaliculus: Vibrio splendidus and a V. coralliilyticus/ neptunius-like isolate. Journal of Fish Diseases, 32(6):499– 507, June 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Leibowitz:2005:EPC

[LAZ05] M. Pimenta Leibowitz, R. Ariav, and D. Zilberg. Environmental and physiological conditions affecting Tetrahymena sp. infection in guppies, Poecilia reticulata Peters. Journal of Fish Diseases, 28(9):539–547, September 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lane:2000:SSS

[LB00] E. Lane and T. H. Birkbeck. Species specificity of some bacterial pathogens of bivalve molluscs is correlated with their interaction with bivalve haemocytes. Journal of Fish Diseases, 23(4):275–279, July 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mordue:2009:RHF

[LB09] A. J. Mordue (Luntz) and M. A. Birkett. A review of host finding behaviour in the parasitic sea louse, Lepeophtheirus salmonis (Caligidae: Copepoda). Journal of Fish Diseases, 32(1):3–13, January 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lees:2008:FAC

[LBGR08] F. Lees, M. Baillie, G. Gettinby, and C. W. Revie. Factors associated with changing efficacy of emamectin benzoate against infestations of Lepeophtheirus salmonis on Scottish salmon farms. Journal of Fish Diseases, 31(12):947–951, December 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lundstrom:2002:UPB

[LBN02] J. Lundström, H. Börjeson, and L. Norrgren. Ultrastructural pathology of Baltic salmon, Salmo salar L., yolk sac fry with the M74 syndrome. Journal of Fish Diseases, 25(3):143–154, March 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

LaPatra:2001:NRA

[LBO^{+}01] S. E. LaPatra, W. N. Batts, K. Overturf, G. N. Jones, W. D. Shewmaker, and J. R. Winton. Negligible risk associated with the movement of processed rainbow trout, Oncorhynchus mykiss (Walbaum), from an infectious haematopoietic necrosis virus (IHNV) endemic area. Journal of Fish Diseases, 24 (7):399–408, September 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lai:2008:CAI

[LCC^{+}08] Y-S. Lai, P. P. Chiou, W-J. Chen, Y-C. Chen, C-W. Chen, I-S. Chiu, S-D. Chen, Y-H. Cheng, and C-Y. Chang. Characterization of apoptosis induced by grouper iridovirus in two newly established cell lines from barramundi, Lates calcarifer (Bloch). Journal of Fish Diseases, 31(11):825–834, November 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lumsden:2002:GDA

[LCH^{+}02] J. S. Lumsden, P. Clark, S. Hawthorn, M. Minamikawa, S. G. Fenwick, M. Haycock, and B. Wybourne. Gastric dilation and air sacculitis in farmed Chinook salmon, Oncorhynchus tshawytscha (Walbaum). Journal of Fish Diseases, 25(3):155– 163, March 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lai:2001:VNM

[LCM^{+}01] Y-S. Lai, H-C. Chiu, S. Murali, I-C. Guo, S-C. Chen, K. Fang, and C-Y. Chang. In vitro neutralization by monoclonal antibodies against yellow grouper nervous necrosis virus (YGNNV) and immunolocalization of virus infection in yellow grouper, Epinephelus awoara (Temminck & Schlegel). Journal of Fish Diseases, 24(4):237–244, May 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Longshaw:2005:DDP

[LFF05] M. Longshaw, P. A. Frear, and S. W. Feist. Descriptions, development and pathogenicity of myxozoan (Myxozoa: Myxosporea) parasites of juvenile cyprinids (Pisces: Cyprinidae). Journal of Fish Diseases, 28(8):489–508, August 2005. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lategan:2003:AAA

[LG03] M. J. Lategan and L. F. Gibson. Antagonistic activity of Aeromonas media strain A199 against Saprolegnia sp., an opportunistic pathogen of the eel, Anguilla australis Richardson. Journal of Fish Diseases, 26(3):147–153, March 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Longshaw:2004:HPI

[LGF04] M. Longshaw, M. J. Green, and S. W. Feist. Histopathology of parasitic infections in greater pipefish, Syngnathus acus L., from an estuary in the UK. Journal of Fish Diseases, 27 (4):245–248, April 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lees:2008:CEP

[LGR08] F. Lees, G. Gettinby, and C. W. Revie. Changes in epidemiological patterns of sea lice infestation on farmed Atlantic salmon, Salmo salar L., in Scotland between 1996 and 2006. Journal of Fish Diseases, 31(4):259–268, April 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Leef:2007:ISV

[LHHP07] M. J. Leef, J. V. Hill, J. O. Harris, and M. D. Powell. Increased systemic vascular resistance in Atlantic salmon, Salmo salar L., affected with amoebic gill disease. Journal of Fish Diseases, 30(10):601–613, October 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lilley:2003:MCF

[LHP^{+}03] J. H. Lilley, D. Hart, V. Panyawachira, S. Kanchanakhan, S. Chinabut, K. Söderhäll, and L. Cerenius. Molecular characterization of the fish-pathogenic fungus Aphanomyces invadans. Journal of Fish Diseases, 26(5):263–275, May 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lonnstrom:2008:FPA

[LHW08] L-G. Lönnström, M. L. Hoffrén, and T. Wiklund. Flavobacterium psychrophilum associated with mortality of farmed perch, Perca fluviatilis L. Journal of Fish Diseases, 31(10): 793–797, October 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lai:2003:ECL

[LJL^{+}03] Y-S. Lai, J. A. C. John, C-H. Lin, I-C. Guo, S-C. Chen, K. Fang, C-H. Lin, and C-Y. Chang. Establishment of cell lines from a tropical grouper, Epinephelus awoara (Temminck & Schlegel), and their susceptibility to grouper irido- and nodaviruses. Journal of Fish Diseases, 26(1):31–42, January 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Levsen:2008:OPM

[LJM08] A. Levsen, A. Jørgensen, and T. A. Mo. Occurrence of postmortem myoliquefactive kudoosis in Atlantic mackerel, Scomber scombrus L., from the North Sea. Journal of Fish Diseases, 31(8):601–611, August 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

LaPatra:2008:PCC

[LKH08] S. LaPatra, R. Kocan, and P. Hershberger. Potential for cross-contamination of in vitro explant cultures initiated from Ichthyophonus-infected rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 31(4):317–320, April 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Longshaw:2002:DPK

[LLHF02] M. Longshaw, R-M. Le Deuff, A. F. Harris, and S. W. Feist. Development of proliferative kidney disease in rainbow trout, Oncorhynchus mykiss (Walbaum), following short-term exposure to Tetracapsula bryosalmonae infected bryozoans. Journal of Fish Diseases, 25(8):443–449, August 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

LaFrentz:2002:CSM

[LLJ^{+}02] B. R. LaFrentz, S. E. LaPatra, G. R. Jones, J. L. Congleton, B. Sun, and K. D. Cain. Characterization of serum and mucosal antibody responses and relative per cent survival in rainbow trout, Oncorhynchus mykiss (Walbaum), following immunization and challenge with Flavobacterium psychrophilum. Journal of Fish Diseases, 25(12):703–713, December 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

LaFrentz:2003:PIR

[LLJC03] B. R. LaFrentz, S. E. LaPatra, G. R. Jones, and K. D. Cain. Passive immunization of rainbow trout, Oncorhynchus mykiss (Walbaum), against Flavobacterium psychrophilum, the causative agent of bacterial coldwater disease and rainbow trout fry syndrome. Journal of Fish Diseases, 26(7):377– 384, July 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lu:2002:ASM

[LLWW02] Y. S. Lu, M. Li, Y. S. Wu, and J. G. Wang. Antigenic study of Myxobolus rotundus (Myxozoa: Myxosporea) using monoclonal antibodies. Journal of Fish Diseases, 25(5):307–310, May 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lu:2008:DCA

[LLZ08] A-J. Lü, Z-Q. Li, and Q-Y. Zhang. Detection of cutaneous antibodies in excised skin explants from grass carp, Ctenopharyngodon idella (Valenciennes), immune to Scophthalmus maximus rhabdovirus. Journal of Fish Diseases, 31 (8):559–565, August 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lumsden:2003:SNF

[LM03] J. S. Lumsden and S. Marshall. Sporadic neoplasms of farmed Chinook salmon, Oncorhynchus tshawytscha (Walbaum), from New Zealand. Journal of Fish Diseases, 26 (7):393–399, July 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lai:2001:PYG

[LMC^{+}01] Y-S. Lai, S. Murali, H-C. Chiu, H-Y. Ju, Y-S. Lin, S-C. Chen, I-C. Guo, K. Fang, and C-Y. Chang. Propagation of yellow grouper nervous necrosis virus (YGNNV) in a new nodavirussusceptible cell line from yellow grouper, Epinephelus awoara (Temminck & Schlegel), brain tissue. Journal of Fish Diseases, 24(5):299–309, June 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lumsden:2003:EPG

[LMG^{+}03] J. S. Lumsden, S. Marshall, M. Gillard, B. Wybourne, and M. Minamikawa. Experimental production of gastric dilation and its association with osmoregulatory stress and biogenic amines in Chinook salmon, Oncorhynchus tshawytscha (Walbaum). Journal of Fish Diseases, 26(8):469–476, August 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lai:2000:TIS

[LMJ^{+}00] Y-S. Lai, S. Murali, H-Y. Ju, M-F. Wu, I-C. Guo, S-C. Chen, K. Fang, and C-Y. Chang. Two iridovirus-susceptible cell lines established from kidney and liver of grouper, Epinephelus awoara (Temminck & Schlegel), and partial characterization of grouper iridovirus. Journal of Fish Diseases, 23(6):379–388, November 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Li:2005:PCC

[LMX^{+}05] H-X. Li, X-L. Meng, J-P. Xu, W. Lu, and J. Wang. Protection of crayfish, Cambarus clarkii, from white spot syndrome virus by polyclonal antibodies against a viral envelope fusion protein. Journal of Fish Diseases, 28(5):285–291, May 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lopez:2009:FIT

[LNM^{+}09] J. R. López, S. Núñez, B. Magariños, N. Castro, J. I. Navas, R. De La Herran, and A. E. Toranzo. First isolation of Tenacibaculum maritimum from wedge sole, Dicologoglossa cuneata (Moreau). Journal of Fish Diseases, 32(7):603–610, July 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Liu:2008:PAS

[LNZL08] G. Y. Liu, P. Nie, J. Zhang, and N. Li. Proteomic analysis of the sarcosine-insoluble outer membrane fraction of Flavobacterium columnare. Journal of Fish Diseases, 31(4):269–276, April 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lumsden:2008:ICG

[LRH^{+}08] J. S. Lumsden, S. Russell, P. Huber, B. A. Wybourne, V. E. Ostland, M. Minamikawa, and H. W. Ferguson. An immune-complex glomerulonephritis of Chinook salmon, Oncorhynchus tshawytscha (Walbaum). Journal of Fish Diseases, 31(12):889–898, December 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lindenstrom:2006:SEE

[LSDB06] T. Lindenstrøm, J. Sigh, M. B. Dalgaard, and K. Buchmann. Skin expression of IL-1 β in East Atlantic salmon, Salmo salar L., highly susceptible to Gyrodactylus salaris infection is enhanced compared to a low susceptibility Baltic stock. Journal of Fish Diseases, 29(2):123–128, February 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lima:2002:OCF

[LSM^{+}02] F. C. Lima, A. P. M. Souza, E. F. M. Mesquita, G. N. Souza, and V. C. J. Chinelli. Osteomas in cutlass fish, Trichiurus lepturus L., from Niteroi, Rio de Janeiro state, Brazil. Journal of Fish Diseases, 25(1):57–61, January 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Labella:2006:FIP

[LVA^{+}06] A. Labella, M. Vida, M. C. Alonso, C. Infante, S. Cardenas, S. Lopez-Romalde, M. Manchado, and J. J. Borrego. First isolation of Photobacterium damselae spp. damselae from cultured redbanded seabream, Pagrus auriga Valenciennes, in Spain. Journal of Fish Diseases, 29(3):175–179, March 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lopez-Vazquez:2007:EIT

[LVDBB07] C. López-Vázquez, C. P. Dopazo, J. L. Barja, and I. Bandín. Experimental infection of turbot, Psetta maxima (L.), with strains of viral haemorrhagic septicaemia virus isolated from wild and farmed marine fish. Journal of Fish Diseases, 30 (5):303–312, May 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lee:2006:HYR

[LVW^{+}06] L. E. J. Lee, S. J. Van Es, S. K. Walsh, D. J. Rainnie, N. Donay, R. Summerfield, and R. J. Cawthorn. High yield and rapid growth of Neoparamoeba pemaquidensis in co-culture with a rainbow trout gill-derived cell line RTgill-W1. Journal of Fish Diseases, 29(8):467–480, August 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lovy:2004:MCN

[LWWS04] J. Lovy, D. W. Wadowska, G. M. Wright, and D. J. Speare. Morphological characterization and notes on the life cycle of a newly discovered variant of Loma salmonae (Putz, Hoffman & Dunbar) from a natural infection of Chinook salmon, Oncorhynchus tshawytscha (Walbaum). Journal of Fish Diseases, 27(10):609–616, October 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lee:2004:MTG

[LYO04] S-J. Lee, H. Yokoyama, and K. Ogawa. Modes of transmission of Glugea plecoglossi (Microspora) via the skin and digestive tract in an experimental infection model using rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 27(8):435–444, August 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Liyanage:2003:DEP

[LYW03] Y. S. Liyanage, H. Yokoyama, and H. Wakabayashi. Dynamics of experimental production of Thelohanellus hovorkai (Myxozoa: Myxosporea) in fish and oligochaete alternate hosts. Journal of Fish Diseases, 26(10):575–582, October 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Leibowitz:2009:TSI

[LZ09] M. P. Leibowitz and D. Zilberg. Tetrahymena sp. infection in guppies, Poecilia reticulata Peters: parasite characterization and pathology of infected fish. Journal of Fish Diseases, 32 (10):845–855, October 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Morris:2004:LCT

[MA04] D. J. Morris and A. Adams. Localization of carbohydrate terminals on Tetracapsuloides bryosalmonae using lectin histochemistry and immunogold electron microscopy. Journal of Fish Diseases, 27(1):37–45, January 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Morris:2000:IPS

[MAF^{+}00] D. J. Morris, A. Adams, S. W. Feist, J. McGeorge, and R. H. Richards. Immunohistochemical and PCR studies of wild fish for Tetracapsula bryosalmonae (PKX), the causative organism of proliferative kidney disease. Journal of Fish Diseases, 23(2):129–135, March 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mercier:2000:CDF

[MALC00] C. Mercier, J. Aubin, C. Lefrançois, and G. Claireaux. Cardiac disorders in farmed adult brown trout, Salmo trutta L. Journal of Fish Diseases, 23(4):243–249, July 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Martin-Armas:2007:SUI

[MASS07] M. Martin-Armas, A-I. Sommer, and B. Smedsrød. Studies on uptake and intracellular processing of infectious pancreatic necrosis virus by Atlantic cod scavenger endothelial cells. Journal of Fish Diseases, 30(11):701–710, November 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Melena:2006:PEI

[MBB^{+}06] J. Melena, B. Bayot, I. Betancourt, Y. Amano, F. Panchana, V. Alday, J. Calderón, S. Stern, Ph Roch, and J-R. Bonami. Pre-exposure to infectious hypodermal and haematopoietic necrosis virus or to inactivated white spot syndrome virus (WSSV) confers protection against WSSV in Penaeus vannamei (Boone) post-larvae. Journal of Fish Diseases, 29(10): 589–600, October 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Michel:2002:CAA

[MBD^{+}02] C. Michel, J-F. Bernardet, P. Daniel, S. Chilmonczyk, M. Urdaci, and P. De Kinkelin. Clinical and aetiological aspects of a summer enteritic syndrome associated with the sporulating segmented filamentous bacterium ‘Candidatus Arthromitus’ in farmed rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 25(9):533–543, September 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Magnadottir:2002:AAS

[MBG^{+}02] B. Magnadóttir, S. H. Bambir, B. K. Gudmundsdóttir, L. Pilström, and S. Helgason. Atypical Aeromonas salmonicida infection in naturally and experimentally infected cod, Gadus morhua L. Journal of Fish Diseases, 25(10):583–597, October 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Millard:2006:DIH

[MBLK06] P. J. Millard, L. E. Bickerstaff, S. E. LaPatra, and C. H. Kim. Detection of infectious haematopoietic necrosis virus and infectious salmon anaemia virus by molecular padlock amplification. Journal of Fish Diseases, 29(4):201–213, April 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Morrison:2004:ILB

[MCN04] R. N. Morrison, P. B. B. Crosbie, and B. F. Nowak. The induction of laboratory-based amoebic gill disease revisited. Journal of Fish Diseases, 27(8):445–449, August 2004. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Molnar:2008:IHC

[MCS08] K. Molnár, G. Cech, and C. Székely. Infection of the heart of the common bream, Abramis brama (L.), with Myxobolus s.l. dogieli (Myxozoa, Myxobolidae). Journal of Fish Diseases, 31(8):613–620, August 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Madsen:2000:CSD

[MD00] L. Madsen and I. Dalsgaard. Comparative studies of Danish Flavobacterium psychrophilum isolates: ribotypes, plasmid profiles, serotypes and virulence. Journal of Fish Diseases, 23(3):211–218, May 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Madsen:2008:WRG

[MD08] L. Madsen and I. Dalsgaard. Water recirculation and good management: potential methods to avoid disease outbreaks with Flavobacterium psychrophilum. Journal of Fish Diseases, 31(11):799–810, November 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Merrifield:2009:SMA

[MDB^{+}09] D. L. Merrifield, A. Dimitroglou, G. Bradley, R. T. M. Baker, and S. J. Davies. Soybean meal alters autochthonous microbial populations, microvilli morphology and compromises intestinal enterocyte integrity of rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 32(9):755–766, September 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Munro:2008:CBN

[ME08] E. S. Munro and A. E. Ellis. A comparison between nondestructive and destructive testing of Atlantic salmon, Salmo salar L., broodfish for IPNV — destructive testing is still the best at time of maturation. Journal of Fish Diseases, 31 (3):187–195, March 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Moller:2005:IAM

[MEBD05] J. D. Møller, A. E. Ellis, A. C. Barnes, and I. Dalsgaard. Iron acquisition mechanisms of Flavobacterium psychrophilum. Journal of Fish Diseases, 28(7):391–398, July 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

McBeath:2006:RDP

[MEC^{+}06] S. J. McBeath, L. M. Ellis, P. F. Cook, L. Wilson, K. L. Urquhart, and I. R. Bricknell. Rapid development of polyclonal antisera against infectious salmon anaemia virus and its optimization and application as a diagnostic tool. Journal of Fish Diseases, 29(5):293–300, May 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Molnar:2009:MES

[MEM^{+}09] K. Molnár, E. Eszterbauer, Sz Marton, G. Cech, and Cs Székely. Myxobolus erythrophthalmi sp. n. and Myxobolus shaharomae sp. n. (Myxozoa: Myxobolidae) from the internal organs of rudd, Scardinius erythrophthalmus (L.), and bleak, Alburnus alburnus (L.). Journal of Fish Diseases, 32 (3):219–231, March 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Molnar:2002:MMB

[MES^{+}02] K. Molnár, E. Eszterbauer, C. Székely, Á. Dán, and B. Harrach. Morphological and molecular biological studies on intramuscular Myxobolus spp. of cyprinid fish. Journal of Fish Diseases, 25(11):643–652, November 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Marcogliese:2001:NSI

[MFLC01] D. J. Marcogliese, M. Fournier, A. Lacroix, and D. G. Cyr. Non-specific immune response associated with infections of lymphocystis disease virus in American plaice, Hippoglossoides platessoides (Fabricius). Journal of Fish Diseases, 24 (2):121–124, February 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

McLoughlin:2007:AIS

[MG07] M. F. McLoughlin and D. A. Graham. Alphavirus infections in salmonids — a review. Journal of Fish Diseases, 30(9): 511–531, September 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Munro:2009:ANA

[MG09] L. A. Munro and A. Gregory. Application of network analysis to farmed salmonid movement data from Scotland. Journal of Fish Diseases, 32(7):641–644, July 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Munro:2006:IPN

[MGAE06] E. S. Munro, S. K. Gahlawat, F. Acosta, and A. E. Ellis. In infectious pancreatic necrosis virus carrier Atlantic salmon, Salmo salar L., post-smolts, almost all kidney macrophages ex vivo contain a low level of non-replicating virus. Journal of Fish Diseases, 29(1):43–48, January 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Marino:2007:CEG

[MGB^{+}07] F. Marino, A. Germanà, S. Bambir, S. Helgason, G. De Vico, and B. Macrì. Calretinin and S-100 expression in goldfish, Carassius auratus (L.), schwannoma. Journal of Fish Diseases, 30(4):251–253, April 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Munro:2004:SND

[MGE04] E. S. Munro, S. K. Gahlawat, and A. E. Ellis. A sensitive non-destructive method for detecting IPNV carrier Atlantic salmon, Salmo salar L., by culture of virus from plastic adherent blood leucocytes. Journal of Fish Diseases, 27(3):129– 134, March 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Molina-Garza:2008:NSV

[MGGSREAG08] Z. J. Molina-Garza, L. Galaviz-Silva, J. L. Rosales-Encinas, and J. M. Alcocer-González. Nucleotide sequence variations of the major structural proteins (VP15, VP19, VP26 and VP28) of white spot syndrome virus (WSSV), a pathogen of cultured Litopenaeus vannamei in Mexico. Journal of Fish Diseases, 31(3):197–203, March 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Magnadottir:2006:ILJ

[MGL^{+}06] B. Magnadottir, B. K. Gudmundsdottir, S. Lange, A. Steinarsson, M. Oddgeirsson, T. Bowden, I. Bricknell, R. A. Dalmo, and S. Gudmundsdottir. Immunostimulation of larvae and juveniles of cod, Gadus morhua L. Journal of Fish Diseases, 29 (3):147–155, March 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

McClure:2004:AIS

[MHD^{+}04] C. A. McClure, K. L. Hammell, I. R. Dohoo, P. Nerette, and L. J. Hawkins. Assessment of infectious salmon anaemia virus prevalence for different groups of farmed Atlantic salmon, Salmo salar L., in New Brunswick. Journal of Fish Diseases, 27(7):375–383, July 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Madetoja:2001:PGC

[MHHK^{+}01] J. Madetoja, M-L. Hänninen, V. Hirvelä-Koski, I. Dalsgaard, and T. Wiklund. Phenotypic and genotypic characterization of Flavobacterium psychrophilum from Finnish fish farms. Journal of Fish Diseases, 24(8):469–479, October 2001. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Maki:2008:DRM

[MHKA08] T. Maki, I. Hirono, H. Kondo, and T. Aoki. Drug resistance mechanism of the fish-pathogenic bacterium Lactococcus garvieae. Journal of Fish Diseases, 31(6):461–468, June 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Magi:2009:CCG

[MIGR09] G. E. Magi, M. Iannaccone, C. Gili, and G. Rossi. Cardiac cholesterol granulomas in a piper gurnard, Trigla lyra (L.). Journal of Fish Diseases, 32(5):473–475, May 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Miwa:2007:MKH

[MIS07] S. Miwa, T. Ito, and M. Sano. Morphogenesis of koi herpesvirus observed by electron microscopy. Journal of Fish Diseases, 30(12):715–722, December 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Munday:2002:BIT

[MKM02] B. L. Munday, J. Kwang, and N. Moody. Betanodavirus infections of teleost fish: a review. Journal of Fish Diseases, 25 (3):127–142, March 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Munchan:2009:EXI

[MKW^{+}09] C. Munchan, O. Kurata, S. Wada, K. Hatai, A. Sano, K. Kamei, and N. Nakaoka. Exophiala xenobiotica infection in cultured striped jack, Pseudocaranx dentex (Bloch & Schneider), in Japan. Journal of Fish Diseases, 32(10):893–900, October 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Madetoja:2006:EIV

[MLB^{+}06] J. Madetoja, L-G. Lönnström, C. Björkblom, G. Uluköy, G. Bylund, C. Syvertsen, K. Gravningen, E-A. Norderhus, and T. Wiklund. Efficacy of injection vaccines against Flavobacterium psychrophilum in rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 29 (1):9–20, January 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Morris:2002:DIP

[MMA02a] D. C. Morris, D. J. Morris, and A. Adams. Development of improved PCR to prevent false positives and false negatives in the detection of Tetracapsula bryosalmonae, the causative agent of proliferative kidney disease. Journal of Fish Diseases, 25(8):483–490, August 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Morris:2002:MER

[MMA02b] D. C. Morris, D. J. Morris, and A. Adams. Molecular evidence of release of Tetracapsula bryosalmonae, the causative organism of proliferative kidney disease from infected salmonids into the environment. Journal of Fish Diseases, 25(8):501– 504, August 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Maniatis:2000:DPD

[MMAP00] K. Maniatis, D. J. Morris, A. Adams, and M. Pearson. Detection of Photobacterium damsela subspecies piscicida in fixed tissue sections using immunohistochemistry and antigen retrieval immunohistochemistry. Journal of Fish Diseases, 23 (5):343–347, September 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Michel:2002:MII

[MMB02] C. Michel, S. Messiaen, and J-F. Bernardet. Muscle infections in imported neon tetra, Paracheirodon innesi Myers: limited occurrence of microsporidia and predominance of severe forms of columnaris disease caused by an Asian genomovar of Flavobacterium columnare. Journal of Fish Diseases, 25 (5):253–263, May 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

McGurk:2005:MTB

[MMBA05] C. McGurk, D. J. Morris, J. E. Bron, and A. Adams. The morphology of Tetracapsuloides bryosalmonae (Myxozoa: Malacosporea) spores released from Fredericella sultana (Bryozoa: Phylactolaemata). Journal of Fish Diseases, 28 (5):307–312, May 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Madsen:2005:FPR

[MMD05] L. Madsen, J. D. Møller, and I. Dalsgaard. Flavobacterium psychrophilum in rainbow trout, Oncorhynchus mykiss (Walbaum), hatcheries: studies on broodstock, eggs, fry and environment. Journal of Fish Diseases, 28(1):39–47, January 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Marino:2006:LFN

[MMSM06] F. Marino, S. Monaco, A. Salvaggio, and B. Macrì. Lipoma in a farmed northern bluefin tuna, Thunnus thynnus (L.). Journal of Fish Diseases, 29(11):697–699, November 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Martinez-Manzanares:2008:SPD

[MMTPDR^{+}08] E. Martínez-Manzanares, S. T. Tapia-Paniagua, P. Díaz-Rosales, M. Chabrillón, and M. A. Moriñigo. Susceptibility of Photobacterium damselae subsp. piscicida strains isolated from Senegalese sole, Solea senegalensis Kaup, and gilthead seabream, Sparus aurata L., to several antibacterial agents. Journal of Fish Diseases, 31(1):73–76, January 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Morrison:2008:IDA

[MN08] R. N. Morrison and B. F. Nowak. Immunohistochemical detection of anterior gradient-2 in the gills of amoebic gill disease-affected Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 31(9):699–705, September 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

McLoughlin:2002:CHF

[MNM^{+}02] M. F. McLoughlin, R. N. Nelson, J. I. McCormick, H. M. Rowley, and D. B. Bryson. Clinical and histopathological features of naturally occurring pancreas disease in farmed Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 25 (1):33–43, January 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Munro:2006:SSC

[MO06] J. Munro and L. Owens. Sensitivity and specificity of current diagnostic tests for gill-associated virus in Penaeus monodon. Journal of Fish Diseases, 29(11):649–655, November 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mitchell:2009:EII

[MOG09] A. J. Mitchell, R. M. Overstreet, and A. E. Goodwin. Eustrongylides ignotus infecting commercial bass, Morone chrysops female × Morone saxatilis male, and other fish in the southeastern USA. Journal of Fish Diseases, 32(9):795– 799, September 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Madinabeitia:2009:HAL

[MOO^{+}09] I. Madinabeitia, S. Ohtsuka, J. Okuda, E. Iwamoto, T. Yoshida, M. Furukawa, N. Nakaoka, and T. Nakai. Homogeneity among Lactococcus garvieae isolates from striped jack, Pseudocaranx dentex (Bloch & Schneider), and its ectoparasites. Journal of Fish Diseases, 32(10):901–905, October 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mikalsen:2009:VPF

[MOR^{+}09] J. Mikalsen, A. B. Olsen, H. Rudra, T. Moldal, H. Lund, B. Djønne, Ø. Bergh, and D. J. Colquhoun. Virulence and pathogenicity of Francisella philomiragia subsp. noatunensis for Atlantic cod, Gadus morhua L., and laboratory mice. Journal of Fish Diseases, 32(4):377–381, April 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mansell:2005:EGM

[MPEN05] B. Mansell, M. D. Powell, I. Ernst, and B. F. Nowak. Effects of the gill monogenean Zeuxapta seriolae (Meserve, 1938) and treatment with hydrogen peroxide on pathophysiology of kingfish, Seriola lalandi Valenciennes, 1833. Journal of Fish Diseases, 28(5):253–262, May 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Madetoja:2003:PVF

[MPP^{+}03] J. Madetoja, P. Pylkkö, T. Pohjanvirta, L. Schildt, and S. Pelkonen. Putative virulence factors of atypical Aeromonas salmonicida isolated from Arctic charr, Salvelinus alpinus (L.), and European grayling, Thymallus thymallus (L.). Journal of Fish Diseases, 26(6):349–359, June 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Majack:2000:BEU

[MRM^{+}00] T. J. Majack, M. B. Rust, K. C. Massee, G. W. Kissil, R. W. Hardy, and M. E. Peterson. Bioencapsulation of erythromycin using adult brine shrimp, Artemia franciscana (Latreille). Journal of Fish Diseases, 23(1):71–76, January 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Munday:2003:DTT

[MSCH03] B. L. Munday, Y. Sawada, T. Cribb, and C. J. Hayward. Diseases of tunas, Thunnus spp. Journal of Fish Diseases, 26 (4):187–206, April 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mustafa:2000:ESS

[MSD^{+}00] A. Mustafa, D. J. Speare, J. Daley, G. A. Conboy, and J. F. Burka. Enhanced susceptibility of seawater cultured rainbow trout, Oncorhynchus mykiss (Walbaum), to the microsporidian Loma salmonae during a primary infection with the sea louse, Lepeophtheirus salmonis. Journal of Fish Diseases, 23 (5):337–341, September 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Montero:2004:GIA

[MST^{+}04] D. Montero, J. Socorro, L. Tort, M. J. Caballero, L. E. Robaina, J. M. Vergara, and M. S. Izquierdo. Glomerulonephritis and immunosuppression associated with dietary essential fatty acid deficiency in gilthead sea bream, Sparus aurata L., juveniles. Journal of Fish Diseases, 27(5):297–306, May 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Murali:2002:MCP

[MWG^{+}02] S. Murali, M-F. Wu, I-C. Guo, S-C. Chen, H-W. Yang, and C-Y. Chang. Molecular characterization and pathogenicity of a grouper iridovirus (GIV) isolated from yellow grouper, Epinephelus awoara (Temminck & Schlegel). Journal of Fish Diseases, 25(2):91–100, February 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Munday:2001:GDM

[MZF01] B. L. Munday, D. Zilberg, and V. Findlay. Gill disease of marine fish caused by infection with Neoparamoeba pemaquidensis. Journal of Fish Diseases, 24(9):497–507, December 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nya:2009:UDG

[NA09a] E. J. Nya and B. Austin. Use of dietary ginger, Zingiber officinale Roscoe, as an immunostimulant to control Aeromonas hydrophila infections in rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 32 (11):971–977, November 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nya:2009:UGA

[NA09b] E. J. Nya and B. Austin. Use of garlic, Allium sativum, to control Aeromonas hydrophila infection in rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 32(11):963–970, November 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nawawi:2008:GVA

[NBB08] R. A. Nawawi, J. Baiano, and A. C. Barnes. Genetic variability amongst Streptococcus iniae isolates from Australia. Journal of Fish Diseases, 31(4):305–309, April 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nowak:2004:GHW

[NDB^{+}04] B. F. Nowak, D. Dawson, L. Basson, M. Deveney, and M. D. Powell. Gill histopathology of wild marine fish in Tasmania: potential interactions with gill health of cultured Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 27 (12):709–717, December 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nerette:2005:ESS

[NDH05a] P. Nérette, I. Dohoo, and L. Hammell. Estimation of specificity and sensitivity of three diagnostic tests for infectious salmon anaemia virus in the absence of a gold standard. Journal of Fish Diseases, 28(2):89–99, February 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nerette:2005:ERR

[NDH^{+}05b] P. Nérette, I. Dohoo, L. Hammell, N. Gagné, P. Barbash, S. MacLean, and C. Yason. Estimation of the repeatability and reproducibility of three diagnostic tests for infectious salmon anaemia virus. Journal of Fish Diseases, 28(2):

101–110, February 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nematollahi:2003:FPI

[NDPH03] A. Nematollahi, A. Decostere, F. Pasmans, and F. Haesebrouck. Flavobacterium psychrophilum infections in salmonid fish. Journal of Fish Diseases, 26(10):563–574, October 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nielsen:2006:EIS

[NEG06] M. E. Nielsen and M. D. Esteve-Gassent. The eel immune system: present knowledge and the need for research. Journal of Fish Diseases, 29(2):65–78, February 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Norris:2008:HMR

[NFR08] A. Norris, L. Foyle, and J. Ratcliff. Heritability of mortality in response to a natural pancreas disease (SPDV) challenge in Atlantic salmon, Salmo salar L., post-smolts on a West of Ireland sea site. Journal of Fish Diseases, 31(12):913–920, December 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Naka:2005:MCF

[NHA05] H. Naka, I. Hirono, and T. Aoki. Molecular cloning and functional analysis of Photobacterium damselae subsp. piscicida haem receptor gene. Journal of Fish Diseases, 28(2):81–88, February 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Naka:2007:CCP

[NHA07] H. Naka, I. Hirono, and T. Aoki. Cloning and characterization of Photobacterium damselae spp. piscicida phospholipase: an enzyme that shows haemolytic activity. Journal of Fish Diseases, 30(11):681–690, November 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nowak:2006:EF

[NL06] B. F. Nowak and S. E. LaPatra. Epitheliocystis in fish. Journal of Fish Diseases, 29(10):573–588, October 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Najimi:2008:DSP

[NLO08] M. Najimi, M. L. Lemos, and C. R. Osorio. Distribution of small plasmids in Aeromonas salmonicida subsp. salmonicida strains isolated from NW Spain and Portugal: evidence of clonality in strains isolated from turbot, Psetta maxima (L.). Journal of Fish Diseases, 31(6):469–472, June 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nusbaum:2002:EIB

[NM02] K. E. Nusbaum and E. E. Morrison. Edwardsiella ictaluri bacteraemia elicits shedding of Aeromonas hydrophila complex in latently infected channel catfish, Ictalurus punctatus (Rafinesque). Journal of Fish Diseases, 25(6):343–350, June 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nomoto:2004:LGC

[NMJ^{+}04] R. Nomoto, L. I. Munasinghe, D-H. Jin, Y. Shimahara, H. Yasuda, A. Nakamura, N. Misawa, T. Itami, and T. Yoshida. Lancefield group C Streptococcus dysgalactiae infection responsible for fish mortalities in Japan. Journal of Fish Diseases, 27(12):679–686, December 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Novoa:2005:PSM

[NPF05] B. Novoa, D. Posada, and A. Figueras. Polymorphisms in the sequences of Marteilia internal transcribed spacer region of the ribosomal RNA genes (ITS-1) in Spain: genetic types are not related with bivalve hosts. Journal of Fish Diseases, 28 (6):331–338, June 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Noga:2000:PSD

[NSK00] E. J. Noga, R. Smolowitz, and L. H. Khoo. Pathology of shell disease in the blue crab, Callinectes sapidus Rathbun, (Decapoda: Portunidae). Journal of Fish Diseases, 23(6): 389–399, November 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nomoto:2006:CLG

[NUS^{+}06] R. Nomoto, N. Unose, Y. Shimahara, A. Nakamura, T. Hirae, K. Maebuchi, S. Harada, N. Misawa, T. Itami, H. Kagawa, and T. Yoshida. Characterization of Lancefield group C Streptococcus dysgalactiae isolated from farmed fish. Journal of Fish Diseases, 29(11):673–682, November 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ogut:2007:SMA

[OB07] H. Ogut and S. C. Bishop. A stochastic modelling approach to describing the dynamics of an experimental furunculosis epidemic in Chinook salmon, Oncorhynchus tshawytscha (Walbaum). Journal of Fish Diseases, 30(2):93–100, February 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Orozova:2009:IPR

[OBAA09] P. Orozova, M. Barker, D. A. Austin, and B. Austin. Identification and pathogenicity to rainbow trout, Oncorhynchus mykiss (Walbaum), of some aeromonads. Journal of Fish Diseases, 32(10):865–871, October 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ostland:2000:NMR

[OBHF00] V. E. Ostland, P. J. Byrne, G. Hoover, and H. W. Ferguson. Necrotic myositis of rainbow trout, Oncorhynchus mykiss (Walbaum): proteolytic characteristics of a crude extracellular preparation from Flavobacterium psychrophilum. Journal of Fish Diseases, 23(5):329–336, September 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Olivares-Fuster:2008:MTS

[OFKEA08] O. Olivares-Fuster, P. H. Klesius, J. Evans, and C. R. Arias. Molecular typing of Streptococcus agalactiae isolates from fish. Journal of Fish Diseases, 31(4):277–283, April 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ornsrud:2004:TEE

[ØGW04] R. Ørnsrud, L. Gil, and R. Waagbø. Teratogenicity of elevated egg incubation temperature and egg vitamin A status in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 27(4):213–223, April 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Oines:2007:CEN

[ØH07] Ø Øines and P. A. Heuch. Caligus elongatus Nordmann genotypes on wild and farmed fish. Journal of Fish Diseases, 30 (2):81–91, February 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Oh:2006:SMF

[OKK^{+}06] M-J. Oh, S-I. Kitamura, W-S. Kim, M-K. Park, S-J. Jung, T. Miyadai, and M. Ohtani. Susceptibility of marine fish species to a megalocytivirus, turbot iridovirus, isolated from turbot, Psetta maximus (L.). Journal of Fish Diseases, 29 (7):415–421, July 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Overturf:2006:QEW

[OL06] K. Overturf and S. LaPatra. Quantitative expression (Walbaum) of immunological factors in rainbow trout, Oncorhynchus mykiss (Walbaum), after infection with either Flavobacterium psychrophilum, Aeromonas salmonicida, or infectious haematopoietic necrosis virus. Journal of Fish Diseases, 29(4):215–224, April 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Olafsdottir:2003:LRA

[OLB03] S. H. Olafsdottir, H. P. Ø. Lassen, and K. Buchmann. Labile resistance of Atlantic salmon, Salmo salar L., to infections with Gyrodactylus derjavini Mikailov, 1975: implications for host specificity. Journal of Fish Diseases, 26(1):51–54, January 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Overturf:2001:RTP

[OLP01] K. Overturf, S. LaPatra, and M. Powell. Real-time PCR for the detection and quantitative analysis of IHNV in salmonids. Journal of Fish Diseases, 24(6):325–333, July 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Overturf:2003:EAV

[OLR03] K. Overturf, S. LaPatra, and P. N. Reynolds. The effectiveness of adenoviral vectors to deliver and express genes in rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 26(2):91–101, February 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

OShea:2006:DSA

[OMLF^{+}06] B. O’Shea, A. J. Mordue-Luntz, R. J. Fryer, C. C. Pert, and I. R. Bricknell. Determination of the surface area of a fish. Journal of Fish Diseases, 29(7):437–440, July 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Olsen:2006:NSG

[OMR^{+}06] A. B. Olsen, J. Mikalsen, M. Rode, A. Alfjorden, E. Hoel, K. Straum-Lie, R. Haldorsen, and D. J. Colquhoun. A novel systemic granulomatous inflammatory disease in farmed Atlantic cod, Gadus morhua L., associated with a bacterium belonging to the genus Francisella. Journal of Fish Diseases, 29(5):307–311, May 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ottem:2008:OFP

[ONI^{+}08] K. F. Ottem, A. Nylund, T. E. Isaksen, E. Karlsbakk, and Ø. Bergh. Occurrence of Francisella piscicida in farmed and wild Atlantic cod, Gadus morhua L., in Norway. Journal of Fish Diseases, 31(7):525–534, July 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ottesen:2007:ESP

[ONS07] O. H. Ottesen, E. J. Noga, and W. Sandaa. Effect of substrate on progression and healing of skin erosions and epidermal papillomas of Atlantic halibut, Hippoglossus hippoglossus (L.). Journal of Fish Diseases, 30(1):43–53, January 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ogut:2005:EEA

[OR05] H. Ogut and P. W. Reno. Evaluation of an experimental Aeromonas salmonicida epidemic in Chinook salmon, Oncorhynchus tshawytscha (Walbaum). Journal of Fish Diseases, 28(5):263–269, May 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Oines:2006:HPA

[ØSKH06] Ø Øines, J. H. Simonsen, J. A. Knutsen, and P. A. Heuch. Host preference of adult Caligus elongatus Nordmann in the laboratory and its implications for Atlantic cod aquaculture. Journal of Fish Diseases, 29(3):167–174, March 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Okamura:2002:BHT

[OW02] B. Okamura and T. S. Wood. Bryozoans as hosts for Tetracapsula bryosalmonae, the PKX organism. Journal of Fish Diseases, 25(8):469–475, August 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Plouffe:2004:ETT

[PB04] D. A. Plouffe and M. Belosevic. Enzyme treatment of Trypanosoma danilewskyi (Laveran & Mesnil) increases its susceptibility to lysis by the alternative complement pathway of goldfish, Carassius auratus (L.). Journal of Fish Diseases, 27 (5):277–285, May 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Poppe:2009:ICP

[PBIM09] T. T. Poppe, G. Bornø, L. Iversen, and E. Myklebust. Idiopathic cardiac pathology in seawater-farmed rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 32(9):807–810, September 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pillai:2006:RDM

[PBW06] D. Pillai, J-R. Bonami, and J. Sri Widada. Rapid detection of Macrobrachium rosenbergii nodavirus (Mr NV) and extra small virus (XSV), the pathogenic agents of white tail disease of Macrobrachium rosenbergii (De Man), by loopmediated isothermal amplification. Journal of Fish Diseases, 29(5):275–283, May 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Partridge:2004:SMJ

[PC04] G. J. Partridge and J. Creeper. Skeletal myopathy in juvenile barramundi, Lates calcarifer (Bloch), cultured in potassiumdeficient saline groundwater. Journal of Fish Diseases, 27(9): 523–530, September 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Paperna:2001:BAF

[PD01] I. Paperna and D. Di Cave. Branchiomycosis in an Amazonian fish, Baryancistrus sp. (Loricariidae). Journal of Fish Diseases, 24(7):417–420, September 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Penston:2009:ASF

[PD09] M. J. Penston and I. M. Davies. An assessment of salmon farms and wild salmonids as sources of Lepeophtheirus salmonis (Krøyer) copepodids in the water column in Loch Torridon, Scotland. Journal of Fish Diseases, 32(1):75–88, January 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Perlberg:2008:CSB

[PDOZ08] S. T. Perlberg, A. Diamant, R. Ofir, and D. Zilberg. Characterization of swim bladder non-inflation (SBN) in angelfish, Pterophyllum scalare (Schultz), and the effect of exposure to methylene blue. Journal of Fish Diseases, 31(3):215–228, March 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Penman:2004:AFB

[Pen04] D. J. Penman. Aquaculture and Fisheries Biotechnology: Genetic Approaches. Journal of Fish Diseases, 27(11):677, November 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pasnik:2005:ASA

[PEP^{+}05] D. J. Pasnik, J. J. Evans, V. S. Panangala, P. H. Klesius, R. A. Shelby, and C. A. Shoemaker. Antigenicity of Streptococcus agalactiae extracellular products and vaccine efficacy. Journal of Fish Diseases, 28(4):205–212, April 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Peeler:2008:AVP

[PFL^{+}08] E. J. Peeler, S. W. Feist, M. Longshaw, M. A. Thrush, and S. St-Hilaire. An assessment of the variation in the prevalence of renal myxosporidiosis and hepatitis in wild brown trout, Salmo trutta L., within and between rivers in South-West England. Journal of Fish Diseases, 31(10):719–728, October 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Preziosi:2006:SDS

[PGB^{+}06] R. Preziosi, S. Gridelli, P. Borghetti, A. Diana, A. Parmeggiani, M. L. Fioravanti, F. Marcer, I. Bianchi, M. Walsh, and I. Berzins. Spinal deformity in a sandtiger shark, Carcharias taurus Rafinesque: a clinical–pathological study. Journal of Fish Diseases, 29(1):49–60, January 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Purcell:2009:IHN

[PGC^{+}09] M. K. Purcell, K. A. Garver, C. Conway, D. G. Elliott, and G. Kurath. Infectious haematopoietic necrosis virus genogroup-specific virulence mechanisms in sockeye salmon, Oncorhynchus nerka (Walbaum), from Redfish Lake, Idaho. Journal of Fish Diseases, 32(7):619–631, July 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pallister:2007:DRT

[PGH^{+}07] J. Pallister, A. Gould, D. Harrison, A. Hyatt, J. Jancovich, and H. Heine. Development of real-time PCR assays for the detection and differentiation of Australian and European ranaviruses. Journal of Fish Diseases, 30(7):427–438, July 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pradeep:2009:FVD

[PKK09] B. Pradeep, I. Karunasagar, and I. Karunasagar. Fitness and virulence of different strains of white spot syndrome virus. Journal of Fish Diseases, 32(9):801–805, September 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Plant:2009:VRT

[PLC09] K. P. Plant, S. E. LaPatra, and K. D. Cain. Vaccination of rainbow trout, Oncorhynchus mykiss (Walbaum), with recombinant and DNA vaccines produced to Flavobacterium psychrophilum heat shock proteins 60 and 70. Journal of Fish Diseases, 32(6):521–534, June 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Penston:2008:STD

[PMZD08] M. J. Penston, C. P. Millar, A. Zuur, and I. M. Davies. Spatial and temporal distribution of Lepeophtheirus salmonis (Krøyer) larvae in a sea loch containing Atlantic salmon, Salmo salar L., farms on the north-west coast of Scotland. Journal of Fish Diseases, 31(5):361–371, May 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Powell:2002:CMR

[PNA02] M. D. Powell, B. F. Nowak, and M. B. Adams. Cardiac morphology in relation to amoebic gill disease history in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 25(4):209– 215, April 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pillai:2005:GSH

[PNS^{+}05] D. Pillai, C. M. Nair, K. R. Salin, A. Marques, J. S. Widada, and J-R. Bonami. Gross signs and histopathology of branchiostegal blister disease (balloon disease): an idiopathic disease of farmed Macrobrachium rosenbergii (De Man). Journal of Fish Diseases, 28(8):473–478, August 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Powell:2005:DRS

[POHJ05] M. Powell, K. Overturf, C. Hogge, and K. Johnson. Detection of Renibacterium salmoninarum in Chinook salmon, Oncorhynchus tshawytscha (Walbaum), using quantitative PCR. Journal of Fish Diseases, 28(10):615–622, October 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pereira:2002:OST

[PP02] N. M. Pereira and M. C. Peleteiro. Osteoma in the skin of a tawny nurse shark, Nebrius ferrugineus (Lesson). Journal of Fish Diseases, 25(9):565–567, September 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Padilla:2005:VFP

[PRG^{+}05] D. Padilla, F. Real, V. Gómez, E. Sierra, B. Acosta, S. Déniz, and F. Acosta. Virulence factors and pathogenicity of Hafnia alvei for gilthead seabream, Sparus aurata L. Journal of Fish Diseases, 28(7):411–417, July 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Poynton:2002:GSD

[PS02] S. L. Poynton and E. Sterud. Guidelines for species descriptions of diplomonad flagellates from fish. Journal of Fish Diseases, 25(1):15–31, January 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Powell:2006:WBN

[PSB06] M. D. Powell, D. J. Speare, and J. A. Becker. Whole body net ion fluxes, plasma electrolyte concentrations and haematology during a Loma salmonae infection in juvenile rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 29(12):727–735, December 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pedersen:2008:SHS

[PSLN^{+}08] K. Pedersen, H. F. Skall, A. M. Lassen-Nielsen, T. F. Nielsen, N. H. Henriksen, and N. J. Olesen. Surveillance of health status on eight marine rainbow trout, Oncorhynchus mykiss (Walbaum), farms in Denmark in 2006. Journal of Fish Diseases, 31(9):659–667, September 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pedersen:2009:PDS

[PSLN^{+}09] K. Pedersen, H. F. Skall, A. M. Lassen-Nielsen, L. Bjerrum, and N. J. Olesen. Photobacterium damselae subsp. damselae, an emerging pathogen in Danish rainbow trout, Oncorhynchus mykiss (Walbaum), mariculture. Journal of Fish Diseases, 32(5):465–472, May 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pakingking:2009:IRA

[PSP^{+}09] R. Pakingking, Jr., R. Seron, L. Dela Peña, K. Mori, H. Yamashita, and T. Nakai. Immune responses of Asian sea bass, Lates calcarifer Bloch, against an inactivated betanodavirus vaccine. Journal of Fish Diseases, 32(5):457–463, May 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Petterson:2009:SAA

[PSS09] E. Petterson, M. Sandberg, and N. Santi. Salmonid alphavirus associated with Lepeophtheirus salmonis (Copepoda: Caligidae) from Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 32(5):477–479, May 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pylkko:2006:EEB

[PSTV06] P. Pylkkö, L-R. Suomalainen, M. Tiirola, and E. Tellervo Valtonen. Evidence of enhanced bacterial invasion during Diplostomum spathaceum infection in European grayling, Thymallus thymallus (L.). Journal of Fish Diseases, 29(2): 79–86, February 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Poppe:2009:IDH

[PT09] T. Poppe and B. Tørud. Intramyocardial dissecting haemorrhage in farmed rainbow trout Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 33(12):1041–1043, December 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pourahmad:2009:DIA

[PTAR09] F. Pourahmad, K. D. Thompson, A. Adams, and R. H. Richards. Detection and identification of aquatic mycobacteria in formalin-fixed, paraffin-embedded fish tissues. Journal of Fish Diseases, 32(5):409–419, May 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Poppe:2007:SMN

[PTB07] T. T. Poppe, T. Taksdal, and P. H. Bergtun. Suspected myocardial necrosis in farmed Atlantic salmon, Salmo salar L.: a field case. Journal of Fish Diseases, 30(10):615–620, October 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Puttinaowarat:2002:IMS

[PTKA02] S. Puttinaowarat, K. D. Thompson, A. Kolk, and A. Adams. Identification of Mycobacterium spp. isolated from snakehead, Channa striata (Fowler), and Siamese fighting fish, Betta splendens (Regan), using polymerase chain reaction–reverse cross blot hybridization (PCR–RCBH). Journal of Fish Diseases, 25(4):235–243, April 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pourahmad:2008:EIL

[PTT^{+}08] F. Pourahmad, K. D. Thompson, J. B. Taggart, A. Adams, and R. H. Richards. Evaluation of the INNO-LiPA mycobacteria v2 assay for identification of aquatic mycobacteria. Journal of Fish Diseases, 31(12):931–940, December 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Padhi:2008:DPS

[PV08] A. Padhi and B. Verghese. Detecting positively selected codons in the glycoprotein of spring viraemia of carp virus (SVCV) isolates from the USA and China. Journal of Fish Diseases, 31(10):785–791, October 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pan:2008:IRE

[PWS^{+}08] X. Pan, T. Wu, Z. Song, H. Tang, and Z. Zhao. Immune responses and enhanced disease resistance in Chinese drum, Miichthys miiuy (Basilewsky), after oral administration of live or dead cells of Clostridium butyrium CB2. Journal of Fish Diseases, 31(9):679–686, September 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Poort:2006:MCM

[PWW^{+}06] M. J. Poort, C. M. Whipps, V. G. Watral, W. F. Font, and M. L. Kent. Molecular characterization of a Mycobacterium species in non-native poeciliids in Hawaii using DNA sequences. Journal of Fish Diseases, 29(3):181–185, March 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Padros:2006:PET

[PZD^{+}06] F. Padrós, C. Zarza, L. Dopazo, M. Cuadrado, and S. Crespo. Pathology of Edwardsiella tarda infection in turbot, Scophthalmus maximus (L.). Journal of Fish Diseases, 29(2):87–94, February 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Peddie:2002:BAD

[PZS02] S. Peddie, J. Zou, and C. J. Secombes. A biologically active IL-1β derived peptide stimulates phagocytosis and bactericidal activity in rainbow trout, Oncorhynchus mykiss (Walbaum), head kidney leucocytes in vitro. Journal of Fish Diseases, 25(6):351–360, June 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Qian:2003:ESV

[QSZ^{+}03] D. Qian, Z. Shi, S. Zhang, Z. Cao, W. Liu, L. Li, Y. Xie, I. Cambournac, and J-R. Bonami. Extra small virus-like particles (XSV) and nodavirus associated with whitish muscle disease in the giant freshwater prawn, Macrobrachium rosenbergii. Journal of Fish Diseases, 26(9):521–527, September 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Renault:2001:HLVa

[RA01] Renault and Arzul. Herpes-like virus infections in hatcheryreared bivalve larvae in Europe: specific viral DNA detection by PCR. Journal of Fish Diseases, 24(3):161–167, March 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roberts:2008:PSF

[RA08] R. J. Roberts and C. Agius. Pan-steatitis in farmed northern bluefin tuna, Thunnus thynnus (L.), in the eastern Adriatic. Journal of Fish Diseases, 31(2):83–88, February 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rorvik:2007:DCE

[RAG^{+}07] K-A. Rørvik, H. Alne, M. Gaarder, B. Ruyter, N. P. Måseide, J. V. Jakobsen, R. K. Berge, T. Sigholt, and M. S. Thomassen. Does the capacity for energy utilization affect the survival of post-smolt Atlantic salmon, Salmo salar L., during natural outbreaks of infectious pancreatic necrosis? Journal of Fish Diseases, 30(7):399–409, July 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rigos:2002:POA

[RAT^{+}02] G. Rigos, M. Alexis, A. E. Tyrpenou, I. Nengas, I. Piper, and G. Troisi. Pharmacokinetics of oxolinic acid in gilthead sea bream, Sparus aurata L. Journal of Fish Diseases, 25(7):401– 408, July 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Romestand:2003:SEL

[RB03] B. Romestand and J-R. Bonami. A sandwich enzyme linked immunosorbent assay (S-ELISA) for detection of Mr NV in the giant freshwater prawn, Macrobrachium rosenbergii (de Man). Journal of Fish Diseases, 26(2):71–75, February 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rodriguez:2003:WSS

[RBA^{+}03] J. Rodríguez, B. Bayot, Y. Amano, F. Panchana, I. De Blas, V. Alday, and J. Calderón. White spot syndrome virus infection in cultured Penaeus vannamei (Boone) in Ecuador with emphasis on histopathology and ultrastructure. Journal of Fish Diseases, 26(8):439–450, August 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Romero-Brey:2009:GAA

[RBBC^{+}09] I. Romero-Brey, I. Bandín, J. M. Cutrín, V. N. Vakharia, and C. P. Dopazo. Genetic analysis of aquabirnaviruses isolated from wild fish reveals occurrence of natural reassortment of infectious pancreatic necrosis virus. Journal of Fish Diseases, 32(7):585–595, July 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rorvik:2003:SED

[RDT^{+}03] K-A. Rørvik, A. Dehli, M. Thomassen, B. Ruyter, S. H. Steien, and R. Salte. Synergistic effects of dietary iron and omega-3 fatty acid levels on survival of farmed Atlantic salmon, Salmo salar L., during natural outbreaks of furunculosis and cold water vibriosis. Journal of Fish Diseases, 26 (8):477–485, August 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rodriguez:2001:AOP

[RF01] J. L. Tojo Rodriguez and M. T. Santamarina Fernandez. Attempts at oral pharmacological treatment of Ichthyophthirius multifiliis in rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 24(4):249–252, May 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ruehl-Fehlert:2005:PIF

[RFBD^{+}05] C. Ruehl-Fehlert, C. Bomke, M. Dorgerloh, X. Palazzi, and M. Rosenbruch. Pleistophora infestation in fathead minnows, Pimephales promelas (Rafinesque). Journal of Fish Diseases, 28(11):629–637, November 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Revie:2002:ESL

[RGTR02] C. W. Revie, G. Gettinby, J. W. Treasurer, and G. H. Rae. The epidemiology of the sea lice, Caligus elongatus Nordmann, in marine aquaculture of Atlantic salmon, Salmo salar L., in Scotland. Journal of Fish Diseases, 25(7):391–399, July 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roberts:2004:B

[RH04] Ronald J. Roberts and Ronald W. Hardy. Biography. Journal of Fish Diseases, 27(9):495–496, September 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roberts:2001:SDA

[RHS01] R. J. Roberts, R. W. Hardy, and S. H. Sugiura. Screamer disease in Atlantic salmon, Salmo salar L., in Chile. Journal of Fish Diseases, 24(9):543–549, December 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rodkhum:2006:PVR

[RHS^{+}06] C. Rodkhum, I. Hirono, M. Stork, M. Di Lorenzo, J. H. Crosa, and T. Aoki. Putative virulence-related genes in Vibrio anguillarum identified by random genome sequencing. Journal of Fish Diseases, 29(3):157–166, March 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Romero:2002:HSD

[RJ02] X. Romero and R. Jiménez. Histopathological survey of diseases and pathogens present in redclaw crayfish, Cherax quadricarinatus (Von Martens), cultured in Ecuador. Journal of Fish Diseases, 25(11):653–667, November 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ramasamy:2000:MDA

[RJB00] P. Ramasamy, R. Jayakumar, and G. P. Brennan. Muscle degeneration associated with cotton shrimp disease of Penaeus indicus. Journal of Fish Diseases, 23(1):77–81, January 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roberts:2004:CSC

[RJC04] R. J. Roberts, K. A. Johnson, and M. T. Casten. Control of Salmincola californiensis (Copepoda: Lernaeapodidae) in rainbow trout, Oncorhynchus mykiss (Walbaum): a clinical and histopathological study. Journal of Fish Diseases, 27(2): 73–79, February 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rigos:2000:TDF

[RKG^{+}00] G. Rigos, I. Kotzamanis, I. Gialamas, I. Nengas, and M. Alexis. Toxicity and digestibility of fumagillin DCH in gilthead sea bream, Sparus aurata L. Journal of Fish Diseases, 23(2):161–164, March 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Renault:2001:HLVb

[RLA01] T. Renault, C. Lipart, and I. Arzul. A herpes-like virus infecting Crassostrea gigas and Ruditapes philippinarum larvae in France. Journal of Fish Diseases, 24(6):369–376, July 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Revie:2009:ESI

[RLB09] C. Revie, F. Lees, and S. Bravo. Editorial for special issue. Journal of Fish Diseases, 32(1):1–2, January 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ramsrud:2007:DRA

[RLL^{+}07] A. L. Ramsrud, S. A. LaFrentz, B. R. LaFrentz, K. D. Cain, and D. R. Call. Differentiating 16S rRNA alleles of Flavobacterium psychrophilum using a simple PCR assay. Journal of Fish Diseases, 30(3):175–180, March 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Raida:2003:ERR

[RLNB03] M. K. Raida, J. L. Larsen, M. E. Nielsen, and K. Buchmann. Enhanced resistance of rainbow trout, Oncorhynchus mykiss (Walbaum), against Yersinia ruckeri challenge following oral administration of Bacillus subtilis and B. licheniformis (Bio-Plus2B). Journal of Fish Diseases, 26(8):495–498, August 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Romano:2004:ERC

[RM04] L. A. Romano and V. A. Marozzi. Epithelio-reticular cell thymoma in carp, Cyprinus carpio L.: an ultrastructural study. Journal of Fish Diseases, 27(6):369–373, June 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rodger:2007:EOP

[RM07] H. Rodger and S. Mitchell. Epidemiological observations of pancreas disease of farmed Atlantic salmon, Salmo salar L., in Ireland. Journal of Fish Diseases, 30(3):157–167, March 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ramstad:2008:SGI

[RM08] A. Ramstad and P. J. Midtlyng. Strong genetic influence on IPN vaccination-and-challenge trials in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 31(8):567–578, August 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ritchie:2009:CVI

[RMG^{+}09] R. J. Ritchie, J. T. McDonald, B. Glebe, W. Young-Lai, E. Johnsen, and N. Gagné. Comparative virulence of Infectious salmon anaemia virus isolates in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 32(2):157–171, February 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rodkhum:2008:GPA

[RMHA08] C. Rodkhum, T. Maki, I. Hirono, and T. Aoki. gyrA and parC associated with quinolone resistance in Vibrio anguillarum. Journal of Fish Diseases, 31(5):395–399, May 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ruane:2009:MDI

[RMSH09] N. M. Ruane, L. J. McCarthy, D. Swords, and K. Henshilwood. Molecular differentiation of infectious pancreatic necrosis virus isolates from farmed and wild salmonids in Ireland. Journal of Fish Diseases, 32(12):979–987, December 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic). See erratum [TCC10].

Rigos:2004:BOS

[RNAA04] G. Rigos, I. Nengas, M. Alexis, and F. Athanassopoulou. Bioavailability of oxytetracycline in sea bass, Dicentrarchus labrax (L.). Journal of Fish Diseases, 27(2):119–122, February 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roberts:2000:AFP

[Rob00] Rj Roberts. Applied Fish Pharmacology. Journal of Fish Diseases, 23(5):361, September 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roberts:2002:BRa

[Rob02a] Ronald J. Roberts. Book reviews. Journal of Fish Diseases, 25(7):439, July 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roberts:2002:BRb

[Rob02b] Ronald J. Roberts. Book reviews. Journal of Fish Diseases, 25(7):439–440, July 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roberts:2005:BFO

[Rob05a] Ronald J. Roberts. Bacteria from Fish and Other Aquatic Animals: A Practical Identification Manual. Journal of Fish Diseases, 28(10):627, October 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roberts:2005:BCC

[Rob05b] Ronald J. Roberts. Biology and culture of channel catfish. Journal of Fish Diseases, 28(9):571–572, September 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roberts:2006:LLL

[Rob06] R. J. Roberts. Leeches, lice and lampreys: A natural history of skin and gill parasites of fishes. Journal of Fish Diseases, 29 (5):317, May 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roberts:2007:E

[Rob07] Ronald J. Roberts. Editorial. Journal of Fish Diseases, 30 (9):509, September 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rodger:2007:EIB

[Rod07] H. D. Rodger. Erythrocytic inclusion body syndrome virus in wild Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 30(7):411–418, July 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roberts:2003:RTH

[RP03] S. D. Roberts and M. D. Powell. Reduced total hardness of fresh water enhances the efficacy of bathing as a treatment for amoebic gill disease in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 26(10):591–599, October 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rodrigues:2004:EDI

[RP04] P. N. S. Rodrigues and F. A. Pereira. Effect of dietary iron overload on Photobacterium damselae spp. piscicida pathogenicity in sea bass, Dicentrarchus labrax (L.). Journal of Fish Diseases, 27(11):673–676, November 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roberts:2005:IPN

[RP05] R. J. Roberts and M. D. Pearson. Infectious pancreatic necrosis in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 28(7):383–390, July 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Revie:2005:MMG

[RRG^{+}05] C. W. Revie, C. Robbins, G. Gettinby, L. Kelly, and J. W. Treasurer. A mathematical model of the growth of sea lice, Lepeophtheirus salmonis, populations on farmed Atlantic salmon, Salmo salar L., in Scotland and its use in the assessment of treatment strategies. Journal of Fish Diseases, 28 (10):603–613, October 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ramstad:2007:FVE

[RRKM07] A. Ramstad, A. B. Romstad, D. H. Knappskog, and P. J. Midtlyng. Field validation of experimental challenge models for IPN vaccines. Journal of Fish Diseases, 30(12):723–731, December 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ramsay:2004:TCG

[RSD04] J. M. Ramsay, D. J. Speare, and J. Daley. Timing of changes in growth rate, feed intake and feed conversion in rainbow trout, Oncorhynchus mykiss (Walbaum), experimentally infected with Loma salmonae (Microspora). Journal of Fish Diseases, 27(7):425–429, July 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rorvik:2000:DWF

[RSFS00] K-A. Rørvik, P. O. Skjervold, S. O. Fjæra, and S. H. Steien. Distended, water-filled stomach in seawater farmed rainbow trout, Oncorhynchus mykiss (Walbaum), provoked experimentally by osmoregulatory stress. Journal of Fish Diseases, 23(1):15–18, January 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ramsay:2001:TPL

[RSSD01] J. M. Ramsay, D. J. Speare, J. G. Sánchez, and J. Daley. The transmission potential of Loma salmonae (Microspora) in the rainbow trout, Oncorhynchus mykiss (Walbaum), is dependent upon the method and timing of exposure. Journal of Fish Diseases, 24(8):453–460, October 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ridgway:2007:IMN

[RST^{+}07] I. D. Ridgway, G. D. Stentiford, A. C. Taylor, R. J. A. Atkinson, T. H. Birkbeck, and D. M. Neil. Idiopathic muscle necrosis in the Norway lobster, Nephrops norvegicus (L.): aetiology, pathology and progression to bacteraemia. Journal of Fish Diseases, 30(5):279–292, May 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rigos:2002:PSF

[RTNA02] G. Rigos, A. Tyrpenou, I. Nengas, and M. Alexis. A pharmacokinetic study of flumequine in sea bass, Dicentrarchus labrax (L.), after a single intravascular injection. Journal of Fish Diseases, 25(2):101–105, February 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rolland:2003:RRP

[RW03] J. B. Rolland and J. R. Winton. Relative resistance of Pacific salmon to infectious salmon anaemia virus. Journal of Fish Diseases, 26(9):511–520, September 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roberts:2005:Ea

[RW05a] Ronald J. Roberts and Rodney Wootten. Editorial. Journal of Fish Diseases, 28(1):1, January 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roberts:2005:Eb

[RW05b] Ronald J. Roberts and Rodney Wootten. Editorial. Journal of Fish Diseases, 28(7):381, July 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roberts:2005:Ec

[RW05c] Ronald J. Roberts and Rodney Wootten. Editorial. Journal of Fish Diseases, 28(12):691, December 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roberts:2006:E

[RW06] Ronald J. Roberts and Rodney Wootten. Editorial. Journal of Fish Diseases, 29(3):129, March 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roberts:2008:E

[RW08] R. J. Roberts and R. Wootten. Editorial. Journal of Fish Diseases, 31(2):81, February 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ramsay:2009:HSE

[RWSK09] J. M. Ramsay, V. Watral, C. B. Schreck, and M. L. Kent. Husbandry stress exacerbates mycobacterial infections in adult zebrafish, Danio rerio (Hamilton). Journal of Fish Diseases, 32(11):931–941, November 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sung:2009:FAF

[SAC^{+}09] Y. Y. Sung, M. F. Ashame, S. Chen, T. H. MacRae, P. Sorgeloos, and P. Bossier. Feeding Artemia franciscana (Kellogg) larvae with bacterial heat shock protein, protects from Vibrio campbellii infection. Journal of Fish Diseases, 32 (8):675–685, August 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sudhakaran:2007:EVT

[SAH^{+}07] R. Sudhakaran, V. P. Ishaq Ahmed, P. Haribabu, S. C. Mukherjee, J. Sri Widada, J. R. Bonami, and A. S. Sahul Hameed. Experimental vertical transmission of Macrobrachium rosenbergii nodavirus (Mr NV) and extra small virus (XSV) from brooders to progeny in Macrobrachium rosenbergii and Artemia. Journal of Fish Diseases, 30(1):27–35, January 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sigh:2001:CIA

[SB01] J. Sigh and K. Buchmann. Comparison of immobilization assays and enzyme-linked immunosorbent assays for detection of rainbow trout antibody-titres against Ichthyophthirius multifiliis Fouquet, 1876. Journal of Fish Diseases, 24 (1):49–51, January 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Smail:2006:IPN

[SBB^{+}06] D. A. Smail, N. Bain, D. W. Bruno, J. A. King, F. Thompson, D. J. Pendrey, S. Morrice, and C. O. Cunningham. Infectious pancreatic necrosis virus in Atlantic salmon, Salmo salar L., post-smolts in the Shetland Isles, Scotland: virus identification, histopathology, immunohistochemistry and genetic comparison with Scottish mainland isolates. Journal of Fish Diseases, 29(1):31–41, January 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sitja-Bobadilla:2007:EHF

[SBDPAP07] A. Sitjà-Bobadilla, A. Diamant, O. Palenzuela, and P. Alvarez-Pellitero. Effect of host factors and experimental conditions on the horizontal transmission of Enteromyxum leei (Myxozoa) to gilthead sea bream, Sparus aurata L., and European sea bass, Dicentrarchus labrax (L.). Journal of Fish Diseases, 30(4):243–250, April 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Samuelsen:2003:SDP

[SBE03] O. B. Samuelsen, Ø. Bergh, and A. Ervik. A single-dose pharmacokinetic study of oxolinic acid and vetoquinol, an oxolinic acid ester, in cod, Gadus morhua L., held in sea water at 8 ^{o}C and in vitro antibacterial activity of oxolinic acid against Vibrio anguillarum strains isolated from diseased cod. Journal of Fish Diseases, 26(6):339–347, June 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Silverstein:2008:DSB

[SBG08] P. S. Silverstein, B. G. Bosworth, and P. S. Gaunt. Differential susceptibility of blue catfish, Ictalurus furcatus (Valenciennes), channel catfish, I. punctatus (Rafinesque), and blue × channel catfish hybrids to channel catfish virus. Journal of Fish Diseases, 31(1):77–79, January 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sanden:2005:EIT

[SBK^{+}05] M. Sanden, M. H. G. Berntssen, Å. Krogdahl, G-I. Hemre, and A-M. Bakke-McKellep. An examination of the intestinal tract of Atlantic salmon, Salmo salar L., parr fed different varieties of soy and maize. Journal of Fish Diseases, 28 (6):317–330, June 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Simko:2000:EIA

[SBM^{+}00] E. Simko, L. L. Brown, A. M. MacKinnon, P. J. Byrne, V. E. Ostland, and H. W. Ferguson. Experimental infection of Atlantic salmon, Salmo salar L., with infectious salmon anaemia virus: a histopathological study. Journal of Fish Diseases, 23 (1):27–32, January 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Seljestokken:2006:TEI

[SBM^{+}06] B. Seljestokken, Ø. Bergh, G. O. Melingen, H. Rudra, R. Hetlelid Olsen, and O. B. Samuelsen. Treating experimentally induced vibriosis (Listonella anguillarum) in cod, Gadus morhua L., with florfenicol. Journal of Fish Diseases, 29(12): 737–742, December 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sarathi:2008:AVM

[SBSH08] M. Sarathi, G. Balasubramanian, V. K. Sivakumar, and A. S. Sahul Hameed. Artemia is not a vector for monodon baculovirus (MBV) transmission to Penaeus monodon. Journal of Fish Diseases, 31(8):631–636, August 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stone:2002:EFB

[SBSR02] J. Stone, S. Boyd, C. Sommerville, and G. H. Rae. An evaluation of freshwater bath treatments for the control of sea lice, Lepeophtheirus salmonis (Krøyer), infections in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 25(6):371– 373, June 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Suomalainen:2009:IPC

[SBV09] L-R. Suomalainen, M. Bandilla, and E. T. Valtonen. Immunostimulants in prevention of columnaris disease of rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 32(8):723–726, August 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Schneck:2006:GDF

[SC06] J. L. Schneck and L. F. Caslake. Genetic diversity of Flavobacterium columnare isolated from fish collected from warm and cold water. Journal of Fish Diseases, 29(4):245– 248, April 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stephens:2001:PES

[SCJ^{+}01] F. J. Stephens, J. J. Cleary, G. Jenkins, B. Jones, S. R. Raidal, and J. B. Thomas. Pathogenesis and epidemiology of spontaneous exophthalmos in the West Australian dhufish, Glaucosoma hebraicum Richardson. Journal of Fish Diseases, 24(9):515–522, December 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stephens:2002:HOT

[SCJ^{+}02] F. J. Stephens, J. J. Cleary, G. Jenkins, B. Jones, S. R. Raidal, and J. B. Thomas. Haemoglobin and oxygen transport of the West Australian dhufish, Glaucosoma hebraicum Richardson, and other species. Journal of Fish Diseases, 25 (7):409–414, July 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Speare:2000:IMI

[SDD^{+}00] D. J. Speare, J. Daley, P. Dick, M. Novilla, and S. Poe. Ionophore-mediated inhibition of xenoma-expression in trout challenged with Loma salmonae (Microspora). Journal of Fish Diseases, 23(3):231–233, May 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Simko:2001:PPC

[SFPF01] E. Simko, K. Falk, T. T. Poppe, and H. W. Ferguson. Plasma protein changes in Atlantic salmon, Salmo salar L., with infectious salmon anaemia. Journal of Fish Diseases, 24(5): 293–298, June 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stone:2008:FRV

[SFT^{+}08] D. M. Stone, H. W. Ferguson, P. A. Tyson, J. Savage, G. Wood, M. J. Dodge, G. Woolford, P. F. Dixon, S. W. Feist, and K. Way. The first report of viral haemorrhagic septicaemia in farmed rainbow trout, Oncorhynchus mykiss (Walbaum), in the United Kingdom. Journal of Fish Diseases, 31(10):775–784, October 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Soliman:2003:SPW

[SGEM03] H. Soliman, K. Geissler, and M. El-Matbouli. SDS-PAGE and Western blot analysis of triactinomyxon spores of Myxobolus cerebralis, the cause of whirling disease in salmonid fish. Journal of Fish Diseases, 26(10):621–625, October 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Snyder:2002:ECC

[SGF02] G. S. Snyder, A. E. Goodwin, and D. W. Freeman. Evidence that channel catfish, Ictalurus punctatus (Rafinesque), mortality is not linked to ingestion of the hepatotoxin microcystin-LR. Journal of Fish Diseases, 25(5):275–285, May 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stingley:2003:CCV

[SGG03] R. L. Stingley, B. R. Griffin, and W. L. Gray. Channel catfish virus gene expression in experimentally infected channel catfish, Ictalurus punctatus (Rafinesque). Journal of Fish Diseases, 26(8):487–493, August 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Santos:2005:EFV

[SGMP^{+}05] Y. Santos, S. García-Marquez, P. G. Pereira, F. Pazos, A. Riaza, R. Silva, A. El Morabit, and F. M. Ubeira. Efficacy of furunculosis vaccines in turbot, Scophthalmus maximus (L.): evaluation of immersion, oral and injection delivery. Journal of Fish Diseases, 28(3):165–172, March 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sullivan:2007:ASDa

[SGRM07] M. Sullivan, D. R. Guy, R. J. Roberts, and N. J. Manchester. The aetiology of spinal deformity in Atlantic salmon, Salmo salar L.: influence of genetic factors on the frequency and severity in freshwater stages. Journal of Fish Diseases, 30 (12):753–758, December 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shinn:2001:MDG

[SGS01] A. P. Shinn, D. I. Gibson, and C. Sommerville. Morphometric discrimination of Gyrodactylus salaris Malmberg (Monogenea) from species of Gyrodactylus parasitising British salmonids using novel parameters. Journal of Fish Diseases, 24(2):83–97, February 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

St-Hilaire:2002:EPM

[SHBB^{+}02] S. St-Hilaire, M. Boichuk, D. Barnes, M. Higgins, R. Devlin, R. Withler, J. Khattra, S. Jones, and D. Kieser. Epizootiology of Parvicapsula minibicornis in Fraser River sockeye salmon, Oncorhynchus nerka (Walbaum). Journal of Fish Diseases, 25(2):107–120, February 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

St-Hilaire:2009:ART

[SHBJ^{+}09] S. St-Hilaire, N. Beevers, C. Joiner, R. P. Hedrick, and K. Way. Antibody response of two populations of common carp, Cyprinus carpio L., exposed to koi herpesvirus. Journal of Fish Diseases, 32(4):311–320, April 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shepherd:2009:SS

[She09] C. Jonathan Shepherd. Shmuel Sarig (1909–2009). Journal of Fish Diseases, 32(6):567, June 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Soto:2009:FSE

[SHFM09] E. Soto, J. P. Hawke, D. Fernandez, and J. A. Morales. Francisella sp., an emerging pathogen of tilapia, Oreochromis niloticus (L.), in Costa Rica. Journal of Fish Diseases, 32 (8):713–722, August 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shih:2002:DTW

[Shi02] H. H. Shih. Detection and titration of white spot syndrome virus using a Blue-Cell ELISA. Journal of Fish Diseases, 25 (3):185–189, March 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sterud:2000:SIU

[SHM00] E. Sterud, M. K. Hansen, and T. A. Mo. Systemic infection with Uronema-like ciliates in farmed turbot, Scophthalmus maximus (L.). Journal of Fish Diseases, 23(1):33–37, January 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Spanggaard:2000:PLV

[SHN^{+}00] B. Spanggaard, I. Huber, J. Nielsen, T. Nielsen, and L. Gram. Proliferation and location of Vibrio anguillarum during infection of rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 23(6):423–427, November 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sugiura:2004:PPD

[SHR04] S. H. Sugiura, R. W. Hardy, and R. J. Roberts. The pathology of phosphorus deficiency in fish — a review. Journal of Fish Diseases, 27(5):255–265, May 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sullivan:2007:SDC

[SHRM07] M. Sullivan, G. Hammond, R. J. Roberts, and N. J. Manchester. Spinal deformation in commercially cultured Atlantic salmon, Salmo salar L.: a clinical and radiological study. Journal of Fish Diseases, 30(12):745–752, December 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stueland:2005:MPC

[SHS05] S. Stueland, K. Hatai, and I. Skaar. Morphological and physiological characteristics of Saprolegnia spp. strains pathogenic to Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 28(8):445–453, August 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shi:2000:WSS

[SHZ^{+}00] Z. Shi, C. Huang, J. Zhang, D. Chen, and J. R. Bonami. White spot syndrome virus (WSSV) experimental infection of the freshwater crayfish, Cherax quadricarinatus. Journal of Fish Diseases, 23(4):285–288, July 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stephens:2006:TCF

[SI06] F. J. Stephens and M. Ingram. Two cases of fish mortality in low pH, aluminium rich water. Journal of Fish Diseases, 29 (12):765–770, December 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Starkey:2000:INS

[SIM^{+}00] W. G. Starkey, J. H. Ireland, K. F. Muir, A. P. Shinn, R. H. Richards, and H. W. Ferguson. Isolation of nodavirus from Scottish farmed halibut, Hippoglossus hippoglossus (L.). Journal of Fish Diseases, 23(6):418–422, November 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Saksida:2007:DLS

[SKCD07] S. Saksida, G. A. Karreman, J. Constantine, and A. Donald. Differences in Lepeophtheirus salmonis abundance levels on Atlantic salmon farms in the Broughton Archipelago, British Columbia, Canada. Journal of Fish Diseases, 30(6):357–366, June 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Snow:2005:ESA

[SKGR05] M. Snow, J. A. King, A. Garden, and R. S. Raynard. Experimental susceptibility of Atlantic cod, Gadus morhua (L.), and Atlantic halibut, Hippoglossus hippoglossus (L.), to different genotypes of viral haemorrhagic septicaemia virus. Journal of Fish Diseases, 28(12):737–742, December 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shima:2006:ALG

[SKI^{+}06] T. Shima, H. Kodama, T. Iwasaki, S. Watarai, and M. Asagi. Adherence of Lactococcus garvieae to the intestinal and brain gangliosides of the yellowtail, Seriola quinqueradiata Temminck and Schlegel. Journal of Fish Diseases, 29(4):249–253, April 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Savan:2005:LMI

[SKIS05] R. Savan, T. Kono, T. Itami, and M. Sakai. Loop-mediated isothermal amplification: an emerging technology for detection of fish and shellfish pathogens. Journal of Fish Diseases, 28(10):573–581, October 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shoemaker:2003:FDC

[SKLY03] C. A. Shoemaker, P. H. Klesius, C. Lim, and M. Yildirim. Feed deprivation of channel catfish, Ictalurus punctatus (Rafinesque), influences organosomatic indices, chemical composition and susceptibility to Flavobacterium columnare. Journal of Fish Diseases, 26(9):553–561, September 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Skall:2004:IWC

[SKO04] H. F. Skall, T. E. Kjær, and N. J. Olesen. Investigation of wild caught whitefish, Coregonus lavaretus (L.), for infection with viral haemorrhagic septicaemia virus (VHSV) and experimental challenge of whitefish with VHSV. Journal of Fish Diseases, 27(7):401–408, July 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Steinum:2008:FCA

[SKR^{+}08] T. Steinum, A. Kvellestad, L. B. Rønneberg, H. Nilsen, A. Asheim, K. Fjell, S. M. R. Nygård, A. B. Olsen, and O. B. Dale. First cases of amoebic gill disease (AGD) in Norwegian seawater farmed Atlantic salmon, Salmo salar L., and phylogeny of the causative amoeba using 18S cDNA sequences. Journal of Fish Diseases, 31(3):205–214, March 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shelby:2002:PIT

[SKSE02] R. A. Shelby, P. H. Klesius, C. A. Shoemaker, and J. J. Evans. Passive immunization of tilapia, Oreochromis niloticus (L.), with anti- Streptococcus iniae whole sera. Journal of Fish Diseases, 25(1):1–6, January 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sigh:2004:PCI

[SLB04] J. Sigh, T. Lindenstrøm, and K. Buchmann. The parasitic ciliate Ichthyophthirius multifiliis induces expression of immune relevant genes in rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 27(7):409–417, July 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Smith:2006:IHA

[SLC^{+}06] P. A. Smith, J. Larenas, J. Contreras, J. Cassigoli, C. Venegas, M. E. Rojas, A. Guajardo, S. Pérez, and S. Díaz. Infectious haemolytic anaemia causes jaundice outbreaks in seawater-cultured Coho salmon, Oncorhynchus kisutch (Walbaum), in Chile. Journal of Fish Diseases, 29(12):709–715, December 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Swan:2008:ILM

[SLC08] C. M. Swan, N. M. Lindstrom, and K. D. Cain. Identification of a localized mucosal immune response in rainbow trout, Oncorhynchus mykiss (Walbaum), following immunization with a protein-hapten antigen. Journal of Fish Diseases, 31 (5):383–393, May 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Smail:2008:IQI

[SM08] D. A. Smail and E. S. Munro. Isolation and quantification of infectious pancreatic necrosis virus from ovarian and seminal fluids of Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 31(1):49–58, January 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sousa:2001:MCP

[SME^{+}01] Sousa, Magariños, Eiras, Toranzo, and Romalde. Molecular characterization of Portuguese strains of Yersinia ruckeri isolated from fish culture systems. Journal of Fish Diseases, 24 (3):151–159, March 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sadler:2008:DKH

[SMG08] J. Sadler, E. Marecaux, and A. E. Goodwin. Detection of koi herpes virus (CyHV-3) in goldfish, Carassius auratus (L.), exposed to infected koi. Journal of Fish Diseases, 31(1):71– 72, January 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sudhakaran:2009:SNE

[SMK^{+}09] R. Sudhakaran, T. Mekata, T. Kono, K. Supamattaya, N. T. H. Linh, Y. Suzuki, M. Sakai, and T. Itami. A simple non-enzymatic method for the preparation of white spot syndrome virus (WSSV) DNA from the haemolymph of Marsupenaeus japonicus using FTA matrix cards. Journal of Fish Diseases, 32(7):611–617, July 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shao:2005:SER

[SMLZ05] M. Shao, Y. Ma, Q. Liu, and Y. Zhang. Secretory expression of recombinant proteins in an attenuated Vibrio anguillarum strain for potential use in vaccines. Journal of Fish Diseases, 28(12):723–728, December 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Snow:2009:DWA

[SMM09] M. Snow, P. McKay, and I. Matejusova. Development of a widely applicable positive control strategy to support detection of infectious salmon anaemia virus (ISAV) using Taqman real-time PCR. Journal of Fish Diseases, 32(2):151–156, February 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sugaya:2009:GHB

[SMN^{+}09] T. Sugaya, K. Mori, T. Nishioka, S. Masuma, M. Oka, K. Mushiake, Y. Okinaka, and T. Nakai. Genetic heterogeneity of betanodaviruses in juvenile production trials of Pacific bluefin tuna, Thunnus orientalis (Temminck & Schlegel). Journal of Fish Diseases, 32(10):815–824, October 2009. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Szekely:2001:CDC

[SMR01] Cs Székely, K. Molnár, and O. Rácz. Complete developmental cycle of Myxobolus pseudodispar (Gorbunova) (Myxosporea: Myxobolidae). Journal of Fish Diseases, 24(8):461–468, October 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sobhana:2002:EDV

[SMS02] K. S. Sobhana, C. V. Mohan, and K. M. Shankar. Effect of dietary vitamin C on the inflammatory response of mrigal, Cirrhinus mrigala (Hamilton), to Freund’s complete adjuvant. Journal of Fish Diseases, 25(3):179–184, March 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stentiford:2000:ROP

[SN00] G. D. Stentiford and D. M. Neil. A rapid onset, post-capture muscle necrosis in the Norway lobster, Nephrops norvegicus (L.) from the West coast of Scotland. Journal of Fish Diseases, 23(4):251–263, July 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Staroscik:2008:ISS

[SN08] A. M. Staroscik and D. R. Nelson. The influence of salmon surface mucus on the growth of Flavobacterium columnare. Journal of Fish Diseases, 31(1):59–69, January 2008. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shimahara:2008:GPC

[SNN^{+}08] Y. Shimahara, A. Nakamura, R. Nomoto, T. Itami, S-C. Chen, and T. Yoshida. Genetic and phenotypic comparison of Nocardia seriolae isolated from fish in Japan. Journal of Fish Diseases, 31(7):481–488, July 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Saksvik:2001:EME

[SNNB01] M. Saksvik, F. Nilsen, A. Nylund, and B. Berland. Effect of marine Eubothrium sp. (Cestoda: Pseudophyllidea) on the growth of Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 24(2):111–119, February 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Simonian:2009:PMM

[SNOR09] M. Simonian, S. V. Nair, W. A. O’Connor, and D. A. Raftos. Protein markers of Marteilia sydneyi infection in Sydney rock oysters, Saccostrea glomerata. Journal of Fish Diseases, 32 (4):367–375, April 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sundh:2009:EHR

[SOF^{+}09] H. Sundh, R-E. Olsen, F. Fridell, K. Gadan, Ø. Evensen, J. Glette, G-L. Taranger, R. Myklebust, and K. Sundell. The effect of hyperoxygenation and reduced flow in fresh water and subsequent infectious pancreatic necrosis virus challenge in sea water, on the intestinal barrier integrity in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 32(8):687– 698, August 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Skall:2005:VHS

[SOM05] H. F. Skall, N. J. Olesen, and S. Mellergaard. Viral haemorrhagic septicaemia virus in marine fish and its implications for fish farming — a review. Journal of Fish Diseases, 28(9): 509–529, September 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sanmartin:2008:VTP

[SPC^{+}08] M. L. Sanmartín, A. Paramá, R. Castro, S. Cabaleiro, J. Leiro, J. Lamas, and J. L. Barja. Vaccination of turbot, Psetta maxima (L.), against the protozoan parasite Philasterides dicentrarchi: effects on antibody production and protection. Journal of Fish Diseases, 31(2):135–140, February 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Seierstad:2005:IDL

[SPK^{+}05] S. L. Seierstad, T. T. Poppe, E. O. Koppang, A. Svindland, G. Rosenlund, L. Frøyland, and S. Larsen. Influence of dietary lipid composition on cardiac pathology in farmed Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 28 (11):677–690, November 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Seierstad:2005:ICT

[SPL05] S. L. Seierstad, T. Poppe, and S. Larsen. Introduction and comparison of two methods of assessment of coronary lesions in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 28(4):189–197, April 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Silva-Rubio:2008:FDS

[SRAHJ^{+}08] A. Silva-Rubio, R. Avendaño-Herrera, B. Jaureguiberry, A. E. Toranzo, and B. Magariños. First description of serotype O3 in Vibrio anguillarum strains isolated from salmonids in Chile. Journal of Fish Diseases, 31(3):235–239, March 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Snow:2003:ECD

[SRM^{+}03] M. Snow, R. S. Raynard, A. G. Murray, D. W. Bruno, J. A. King, R. Grant, I. R. Bricknell, N. Bain, and A. Gregory.

An evaluation of current diagnostic tests for the detection of infectious salmon anaemia virus (ISAV) following experimental water-borne infection of Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 26(3):135–145, March 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sullivan:2007:ASDb

[SRT^{+}07] M. Sullivan, S. W. J. Reid, H. Ternent, N. J. Manchester, R. J. Roberts, D. A. J. Stone, and R. W. Hardy. The aetiology of spinal deformity in Atlantic salmon, Salmo salar L.: influence of different commercial diets on the incidence and severity of the preclinical condition in salmon parr under two contrasting husbandry regimes. Journal of Fish Diseases, 30 (12):759–767, December 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stringer-Roth:2002:DCA

[SRYC02] K. M. Stringer-Roth, W. Yunghans, and L. F. Caslake. Differences in chondroitin AC lyase activity of Flavobacterium columnare isolates. Journal of Fish Diseases, 25(11):687–691, November 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sagstad:2007:ESI

[SSH^{+}07] A. Sagstad, M. Sanden, Ø. Haugland, A-C. Hansen, P. A. Olsvik, and G-I. Hemre. Evaluation of stress- and immuneresponse biomarkers in Atlantic salmon, Salmo salar L., fed different levels of genetically modified maize (Bt maize), compared with its near-isogenic parental line and a commercial suprex maize. Journal of Fish Diseases, 30(4):201–212, April 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shivappa:2008:DSV

[SSK^{+}08] R. B. Shivappa, R. Savan, T. Kono, M. Sakai, E. Emmenegger, G. Kurath, and J. F. Levine. Detection of spring viraemia of carp virus (SVCV) by loop-mediated isothermal amplification (LAMP) in koi carp, Cyprinus carpio L. Journal of Fish Diseases, 31(4):249–258, April 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Seierstad:2008:DIT

[SSL^{+}08a] S. L. Seierstad, A. Svindland, S. Larsen, G. Rosenlund, B. E. Torstensen, and Ø. Evensen. Development of intimal thickening of coronary arteries over the lifetime of Atlantic salmon, Salmo salar L., fed different lipid sources. Journal of Fish Diseases, 31(6):401–413, June 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Skinner:2008:GPA

[SSL^{+}08b] L. A. Skinner, P. M. Schulte, S. E. LaPatra, S. K. Balfry, and R. S. McKinley. Growth and performance of Atlantic salmon, Salmo salar L., following administration of a rhabdovirus DNA vaccine alone or concurrently with an oil-adjuvanted, polyvalent vaccine. Journal of Fish Diseases, 31(9):687–697, September 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sanchez:2000:NAT

[SSM00] J. G. Sanchez, D. J. Speare, and R. J. F. Markham. Normal and aberrant tissue distribution of Loma salmonae (Microspora) within rainbow trout, Oncorhynchus mykiss (Walbaum), following experimental infection at water temperatures within and outside of the xenoma-expression temperature boundaries. Journal of Fish Diseases, 23(4):235–242, July 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sanchez:2001:ATD

[SSM01] J. G. Sanchez, D. J. Speare, and R. J. F. Markham. Altered tissue distribution of Loma salmonae: effects of natural and acquired resistance. Journal of Fish Diseases, 24(1):33–40, January 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Song:2009:ADS

[SSO^{+}09] J-Y. Song, K. Sasaki, T. Okada, M. Sakashita, H. Kawakami, S. Matsuoka, H-S. Kang, K. Nakayama, S-J. Jung, M-J. Oh, and S-I. Kitamura. Antigenic differences of the scuticociliate Miamiensis avidus from Japan. Journal of Fish Diseases, 33(12):1027–1034, December 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stone:2000:DEF

[SSS^{+}00] J. Stone, I. H. Sutherland, C. Sommerville, R. H. Richards, and R. G. Endris. The duration of efficacy following oral treatment with emamectin benzoate against infestations of sea lice, Lepeophtheirus salmonis (Krøyer), in Atlantic salmon Salmo salar L. Journal of Fish Diseases, 23(3):185–192, May 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Starliper:2001:ISL

[Sta01] C. E. Starliper. Isolation of Serratia liquefaciens as a pathogen of Arctic char, Salvelinus alpinus (L.). Journal of Fish Diseases, 24(1):53–56, January 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sakai:2001:IEN

[STM^{+}01] M. Sakai, K. Taniguchi, K. Mamoto, H. Ogawa, and M. Tabata. Immunostimulant effects of nucleotide isolated from yeast RNA on carp, Cyprinus carpio L. Journal of Fish Diseases, 24(8):433–438, October 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Suomalainen:2005:IRC

[STV05] L-R. Suomalainen, M. A. Tiirola, and E. T. Valtonen. Influence of rearing conditions on Flavobacterium columnare infection of rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 28(5):271–277, May 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Suomalainen:2006:CAL

[STV06] L-R. Suomalainen, M. Tiirola, and E. T. Valtonen. Chondroitin AC lyase activity is related to virulence of fish pathogenic Flavobacterium columnare. Journal of Fish Diseases, 29(12):757–763, December 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Smith:2007:PES

[SWS07] J. L. Smith, R. Wootten, and C. Sommerville. The pathology of the early stages of the crustacean parasite, Lernaeocera branchialis (L.), on Atlantic cod, Gadus morhua L. Journal of Fish Diseases, 30(1):1–11, January 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shih:2001:CAM

[SWTC01] Shih, Wang, Tan, and Chen. Characterization and application of monoclonal antibodies against white spot syndrome virus. Journal of Fish Diseases, 24(3):143–150, March 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shi:2005:RCP

[SWZ^{+}05] Z. Shi, H. Wang, J. Zhang, Y. Xie, L. Li, X. Chen, B-F. Edgerton, and J-R. Bonami. Response of crayfish, Procambarus clarkii, haemocytes infected by white spot syndrome virus. Journal of Fish Diseases, 28(3):151–156, March 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sun:2007:GSA

[SXSN07] B. J. Sun, H. X. Xie, Y. Song, and P. Nie. Gene structure of an antimicrobial peptide from mandarin fish, Siniperca chuatsi (Basilewsky), suggests that moronecidins and pleurocidins belong in one family: the piscidins. Journal of Fish Diseases, 30(6):335–343, June 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shirakashi:2006:IDN

[SYYO06] S. Shirakashi, T. Yamada, T. Yamada, and K. Ogawa. Infection dynamics of Neoheterobothrium hirame (Monogenea) on juvenile olive flounder, Paralichthys olivaceus (Temminck & Schlegel), in coastal waters of Japan. Journal of Fish Diseases, 29(6):319–329, June 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tacon:2000:BR

[Tac00] Albert Tacon. BOOK REVIEWS. Journal of Fish Diseases, 23(1):92, January 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tuttle:2008:TCG

[TBB^{+}08] A. D. Tuttle, O. Burrus, M. A. Burkart, P. W. Scott, M. K. Stoskopf, and C. A. Harms. Three cases of gastric prolapse through the gill slit in sand tiger sharks, Carcharhinus taurus (Rafinesque). Journal of Fish Diseases, 31(4):311–315, April 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Treasurer:2007:AAS

[TBLC07] J. W. Treasurer, T. H. Birkbeck, L. A. Laidler, and D. I. Cox. Atypical Aeromonas salmonicida infection in naturallyand laboratory-challenged farmed haddock, Melanogrammus aeglefinus (L.). Journal of Fish Diseases, 30(5):313–318, May 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Taylor:2010:E

[TCC10] R. S. Taylor, P. B. Crosbie, and M. T. Cook. Erratum. Journal of Fish Diseases, 33(7):623, July 2010. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic). See [RMSH09].

Tobback:2007:YRI

[TDH^{+}07] E. Tobback, A. Decostere, K. Hermans, F. Haesebrouck, and K. Chiers. Yersinia ruckeri infections in salmonid fish. Journal of Fish Diseases, 30(5):257–268, May 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tierney:2004:RBF

[TF04] K. B. Tierney and A. P. Farrell. The relationships between fish health, metabolic rate, swimming performance and recovery in return-run sockeye salmon, Oncorhynchus nerka (Walbaum). Journal of Fish Diseases, 27(11):663–671, November 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tafalla:2001:CAA

[TFN01] Tafalla, Figueras, and Novoa. Cytotoxic activity against prelabelled RTG-2 cells in the turbot, Scophthalmus maximus (L.). Journal of Fish Diseases, 24(3):169–175, March 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tvete:2008:PCK

[TH08] T. Tvete and K. Haugan. Purification and characterization of a 630 kDa bacterial killing metalloprotease (KilC) isolated from plaice Pleuronectes platessa (L.), epidermal mucus. Journal of Fish Diseases, 31(5):343–352, May 2008.

CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Thomas-Jinu:2004:ACI

[TJG04a] S. Thomas-Jinu and A. E. Goodwin. Acute columnaris infection in channel catfish, Ictalurus punctatus (Rafinesque): efficacy of practical treatments for warmwater aquaculture ponds. Journal of Fish Diseases, 27(1):23–28, January 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Thomas-Jinu:2004:MGC

[TJG04b] S. Thomas-Jinu and A. E. Goodwin. Morphological and genetic characteristics of Flavobacterium columnare isolates: correlations with virulence in fish. Journal of Fish Diseases, 27(1):29–35, January 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tanaka:2003:SCJ

[TKNM03] S. Tanaka, I. Kuriyama, T. Nakai, and T. Miyazaki. Susceptibility of cultured juveniles of several marine fish to the sevenband grouper nervous necrosis virus. Journal of Fish Diseases, 26(2):109–115, February 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Temple:2009:DOF

[TL09] E. Temple and C. Langdon. Delivering oxytetracycline to first-feeding zebrafish, Danio rerio (Hamilton), and goby, Asterropteryx semipunctata Rüppell, larvae using lipid spray beads. Journal of Fish Diseases, 32(3):279–292, March 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Torres:2002:VDC

[TLC^{+}02] P. Torres, J. C. Lopez, V. Cubillos, C. Lobos, and R. Silva. Visceral diphyllobothriosis in a cultured rainbow trout, Oncorhynchus mykiss (Walbaum), in Chile. Journal of Fish Diseases, 25(6):375–379, June 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tanaka:2001:PIS

[TMA^{+}01] S. Tanaka, K. Mori, M. Arimoto, T. Iwamoto, and T. Nakai. Protective immunity of sevenband grouper, Epinephelus septemfasciatus Thunberg, against experimental viral nervous necrosis. Journal of Fish Diseases, 24(1):15–22, January 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Thiyagarajah:2001:NCS

[TMM01] A. Thiyagarajah, J. R. MacMillan, and A. D. Munson. Neoplasms in cultured sunshine bass. Journal of Fish Diseases, 24(9):551–556, December 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Thyssen:2002:VED

[TO02] A. Thyssen and F. Ollevier. In vitro evaluation of disc diffusion and agar dilution susceptibility testing of Photobacterium damselae spp. piscicida. Journal of Fish Diseases, 25 (4):245–248, April 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Taksdal:2007:PDF

[TOB^{+}07] T. Taksdal, A. B. Olsen, I. Bjerkås, M. J. Hjortaas, B. H. Dannevig, D. A. Graham, and M. F. McLoughlin. Pancreas disease in farmed Atlantic salmon, Salmo salar L., and rainbow trout, Oncorhynchus mykiss (Walbaum), in Norway. Journal of Fish Diseases, 30(9):545–558, September 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tun:2002:PCI

[TOW02] Tin Tun, K. Ogawa, and H. Wakabayashi. Pathological changes induced by three myxosporeans in the intestine of cultured tiger puffer, Takifugu rubripes (Temminck and Schlegel). Journal of Fish Diseases, 25(2):63–72, February 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

True:2009:DVQ

[TPF09] K. True, M. K. Purcell, and J. S. Foott. Development and validation of a quantitative PCR to detect Parvicapsula minibicornis and comparison to histologically ranked infection of juvenile Chinook salmon, Oncorhynchus tshawytscha (Walbaum), from the Klamath River, USA. Journal of Fish Diseases, 32(2):183–192, February 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Temprano:2005:PUY

[TRY^{+}05] A. Temprano, J. Riaño, J. Yugueros, P. González, L. de Castro, A. Villena, J. M. Luengo, and G. Naharro. Potential use of a Yersinia ruckeri O1 auxotrophic aroA mutant as a live attenuated vaccine. Journal of Fish Diseases, 28(7):419– 427, July 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tan:2001:QAE

[TSL^{+}01] L. T. Tan, S. Soon, K. L. Lee, M. Shariff, M. D. Hassan, and A. R. Omar. Quantitative analysis of an experimental white spot syndrome virus (WSSV) infection in Penaeus monodon Fabricius using competitive polymerase chain reaction. Journal of Fish Diseases, 24(6):315–323, July 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tucker:2000:ETS

[TSW00] C. S. Tucker, C. Sommerville, and R. Wootten. The effect of temperature and salinity on the settlement and survival of copepodids of Lepeophtheirus salmonis (Krøyer, 1837) on Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 23 (5):309–320, September 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Taylor:2006:EAS

[TSW06] N. G. H. Taylor, C. Sommerville, and R. Wootten. The epidemiology of Argulus spp. (Crustacea: Branchiura) infections in stillwater trout fisheries. Journal of Fish Diseases, 29 (4):193–200, April 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Torud:2006:MGS

[TTD^{+}06] B. Tørud, T. Taksdal, O. B. Dale, A. Kvellestad, and T. T. Poppe. Myocardial glycogen storage disease in farmed rainbow trout, Oncorhynchus mykiss (Walbaum). Journal of Fish Diseases, 29(9):535–540, September 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tanaka:2004:HSV

[TTM04] S. Tanaka, M. Takagi, and T. Miyazaki. Histopathological studies on viral nervous necrosis of sevenband grouper, Epinephelus septemfasciatus Thunberg, at the grow-out stage. Journal of Fish Diseases, 27(7):385–399, July 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tkachenko:2006:IIH

[TWOM06] A. Tkachenko, Y. Waguespack, J. Okoh, and E. May. Isolation of intact high-molecular weight glycoconjugates from the skin mucus of Morone saxatilis (Walbaum). Journal of Fish Diseases, 29(7):433–436, July 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Taylor:2009:IRF

[TWS09] N. G. H. Taylor, R. Wootten, and C. Sommerville. The influence of risk factors on the abundance, egg laying habits and impact of Argulus foliaceus in stillwater trout fisheries. Journal of Fish Diseases, 32(6):509–519, June 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tan:2009:MDT

[TXZ^{+}09] Y. Tan, Y. Xing, H. Zhang, Y. Feng, Y. Zhou, and Z-L. Shi. Molecular detection of three shrimp viruses and genetic variation of white spot syndrome virus in Hainan Province, China, in 2007. Journal of Fish Diseases, 32(9):777–784, September 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Temprano:2001:RIY

[TYH^{+}01] A. Temprano, J. Yugueros, C. Hernanz, M. Sánchez, B. Berzal, J. M. Luengo, and G. Naharro. Rapid identification of Yersinia ruckeri by PCR amplification of yruI–yruR quorum sensing. Journal of Fish Diseases, 24(5):253–261, June 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Urquhart:2009:ESS

[UBB^{+}09] K. Urquhart, T. J. Bowden, B-E. Buckett, J. Garcia, R. J. Fryer, and A. E. Ellis. Experimental study of the susceptibility of Atlantic cod, Gadus morhua (L.), to infection with an IPNV strain pathogenic for Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 32(5):447–456, May 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ucko:2004:NII

[UCD04] M. Ucko, A. Colorni, and A. Diamant. Nodavirus infections in Israeli mariculture. Journal of Fish Diseases, 27(8):459– 469, August 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Urquhart:2008:EID

[UMG^{+}08] K. Urquhart, A. G. Murray, A. Gregory, M. O’Dea, L. A. Munro, D. A. Smail, A. M. Shanks, and R. S. Raynard. Estimation of infectious dose and viral shedding rates for infectious pancreatic necrosis virus in Atlantic salmon, Salmo salar L., post-smolts. Journal of Fish Diseases, 31(12):879–887, December 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Uchii:2009:DIC

[UMIK09] K. Uchii, K. Matsui, T. Iida, and Z. Kawabata. Distribution of the introduced cyprinid herpesvirus 3 in a wild population of common carp, Cyprinus carpio L. Journal of Fish Diseases, 32(10):857–864, October 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

McCarthy:2008:ISH

[úMMUB08] ú M. McCarthy, K. L. Urquhart, and I. R. Bricknell. An improved in situ hybridization method for the detection of fish pathogens. Journal of Fish Diseases, 31(9):669–677, September 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Uran:2009:TRC

[USR^{+}09] P. A. Urán, J. W. Schrama, J. H. W. M. Rombout, J. J. Taverne-Thiele, A. Obach, W. Koppe, and J. A. J. Verreth. Time-related changes of the intestinal morphology of Atlantic salmon, Salmo salar L., at two different soybean meal inclusion levels. Journal of Fish Diseases, 32(9):733–744, September 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Valdebenito:2009:PSG

[VAH09] S. Valdebenito and R. Avendaño-Herrera. Phenotypic, serological and genetic characterization of Flavobacterium psychrophilum strains isolated from salmonids in Chile. Journal of Fish Diseases, 32(4):321–333, April 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Venegas:2004:DHA

[VCLS04] C. A. Venegas, J. R. Contreras, J. J. Larenas, and P. A. Smith. DNA hybridization assays for the detection of Piscirickettsia salmonis in salmonid fish. Journal of Fish Diseases, 27(7):431–433, July 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

VanGelderen:2009:EEP

[VCN09] R. Van Gelderen, J. Carson, and B. Nowak. Effect of extracellular products of Tenacibaculum maritimum in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 32(8):727– 731, August 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Villamil:2003:NSI

[VFAN03] L. Villamil, A. Figueras, R. Aranguren, and B. Novoa. Nonspecific immune response of turbot, Scophthalmus maximus (L.), experimentally infected with a pathogenic Vibrio pelagius. Journal of Fish Diseases, 26(6):321–329, June 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Villamil:2003:IHP

[VFT^{+}03] L. Villamil, A. Figueras, A. E. Toranzo, M. Planas, and B. Novoa. Isolation of a highly pathogenic Vibrio pelagius strain associated with mass mortalities of turbot, Scophthalmus maximus (L.), larvae. Journal of Fish Diseases, 26 (5):293–303, May 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Valheim:2000:VNB

[VHM^{+}00] M. Valheim, T. Håstein, E. Myhr, L. Speilberg, and H. W. Ferguson. Varracalbmi: a new bacterial panophthalmitis in farmed Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 23(1):61–70, January 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Valdes:2009:GCS

[VJR^{+}09] I. Valdés, B. Jaureguiberry, J. L. Romalde, A. E. Toranzo, B. Magariños, and R. Avendaño-Herrera. Genetic characterization of Streptococcus phocae strains isolated from Atlantic salmon, Salmo salar L., in Chile. Journal of Fish Diseases, 32(4):351–358, April 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Vine:2004:CAA

[VLK^{+}04] N. G. Vine, W. D. Leukes, H. Kaiser, S. Daya, J. Baxter, and T. Hecht. Competition for attachment of aquaculture candidate probiotic and pathogenic bacteria on fish intestinal mucus. Journal of Fish Diseases, 27(6):319–326, June 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Villavedra:2007:CEI

[VLT^{+}07] M. Villavedra, S. Lemke, J. To, K. Broady, M. Wallach, and R. L. Raison. Carbohydrate epitopes are immunodominant at the surface of infectious Neoparamoeba spp. Journal of Fish Diseases, 30(4):191–199, April 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Vincent:2006:AGD

[VMN06] B. N. Vincent, R. N. Morrison, and B. F. Nowak. Amoebic gill disease (AGD)-affected Atlantic salmon, Salmo salar L., are resistant to subsequent AGD challenge. Journal of Fish Diseases, 29(9):549–559, September 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

VanDerSalm:2000:EIE

[VNSB00] A. L. Van Der Salm, D. T. Nolan, F. A. T. Spanings, and S. E. Wendelaar Bonga. Effects of infection with the ectoparasite Argulus japonicus (Thiele) and administration of cortisol on cellular proliferation and apoptosis in the epidermis of common carp, Cyprinus carpio L., skin. Journal of Fish Diseases, 23(3):173–184, May 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

VanDerMarel:2008:CTA

[VSNS08] M. Van Der Marel, V. Schroers, H. Neuhaus, and D. Steinhagen. Chemotaxis towards, adhesion to, and growth in carp gut mucus of two Aeromonas hydrophila strains with different pathogenicity for common carp, Cyprinus carpio L. Journal of Fish Diseases, 31(5):321–330, May 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Voutilainen:2009:VEP

[VSST09] A. Voutilainen, M. Saarinen, A. Suonpää, and J. Taskinen. In vitro efficacy of praziquantel against the cercariae of Diplostomum sp., Rhipidocotyle fennica and R. campanula. Journal of Fish Diseases, 32(10):907–909, October 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Vatsos:2006:CRT

[VTA06] I. N. Vatsos, K. D. Thompson, and A. Adams. Colonization of rainbow trout, Oncorhynchus mykiss (Walbaum), eggs by Flavobacterium psychrophilum, the causative agent of rainbow trout fry syndrome. Journal of Fish Diseases, 29(7):441– 444, July 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:2002:SDD

[Wan02] G. T. Wang. Seasonal dynamics and distribution of Philometroides fulvidraconi (Philometridae) in the bullhead catfish, Pseudobagrus fulvidraco (Richardson). Journal of Fish Diseases, 25(10):621–625, October 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Woodland:2002:EOT

[WBNG02] J. E. Woodland, C. J. Brunner, A. D. Noyes, and J. M. Grizzle. Experimental oral transmission of largemouth bass virus. Journal of Fish Diseases, 25(11):669–672, November 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wahli:2005:ASC

[WBP^{+}05] T. Wahli, S. E. Burr, D. Pugovkin, O. Mueller, and J. Frey. Aeromonas sobria, a causative agent of disease in farmed perch, Perca fluviatilis L. Journal of Fish Diseases, 28(3): 141–150, March 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wolf:2008:RAP

[WCB08] D. P. Wolf, D. K. Cone, and M. D. B. Burt. Recent appearance and pathology of Actinobdella pediculata (Hirudinea) on freshwater drum, Aplodinotus grunniens Rafinesque, in Hay Bay, Lake Ontario. Journal of Fish Diseases, 31(3):229–233, March 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Whittington:2008:HCE

[WCJJ08] R. J. Whittington, M. Crockford, D. Jordan, and B. Jones. Herpesvirus that caused epizootic mortality in 1995 and 1998 in pilchard, Sardinops sagax neopilchardus (Steindachner), in Australia is now endemic. Journal of Fish Diseases, 31(2): 97–105, February 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:2009:PAS

[WCK^{+}09] C. S. Wang, S. Y. Chao, C. C. Ku, C. M. Wen, and H. H. Shih. PCR amplification and sequence analysis of the major capsid protein gene of megalocytiviruses isolated in Taiwan. Journal of Fish Diseases, 32(6):543–550, June 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:2007:RPA

[WCSC07] C. S. Wang, J. S. Chang, H. H. Shih, and S. N. Chen. RT-PCR amplification and sequence analysis of extra small virus associated with white tail disease of Macrobrachium rosenbergii (de Man) cultured in Taiwan. Journal of Fish Diseases, 30(3):127–132, March 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Whittington:2001:ISQ

[WCT^{+}01] I. D. Whittington, S. Corneillie, C. Talbot, J. A. T. Morgan, and R. D. Adlard. Infections of Seriola quinqueradiata Temminck & Schlegel and S. dumerili (Risso) in Japan by Benedenia seriolae (Monogenea) confirmed by morphology and 28S ribosomal DNA analysis. Journal of Fish Diseases, 24(7): 421–425, September 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:2009:NSI

[WCT^{+}09] P-C. Wang, S-D. Chen, M-A. Tsai, Y-J. Weng, S-Y. Chu, R-S. Chern, and S-C. Chen. Nocardia seriolae infection in the three striped tigerfish, Terapon jarbua (Forsskål). Journal of Fish Diseases, 32(4):301–310, April 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:2008:MRN

[WCW^{+}08] C. S. Wang, J. S. Chang, C. M. Wen, H. H. Shih, and S. N. Chen. Macrobrachium rosenbergii nodavirus infection in M. rosenbergii (de Man) with white tail disease cultured in Taiwan. Journal of Fish Diseases, 31(6):415–422, June 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Widada:2003:GBD

[WDC^{+}03] J. Sri Widada, S. Durand, I. Cambournac, D. Qian, Z. Shi, E. Dejonghe, V. Richard, and J. R. Bonami. Genome-based detection methods of Macrobrachium rosenbergii nodavirus, a pathogen of the giant freshwater prawn, Macrobrachium rosenbergii: dot-blot, in situ hybridization and RT-PCR. Journal of Fish Diseases, 26(10):583–590, October 2003. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wallace:2008:DIP

[WGM^{+}08] I. S. Wallace, A. Gregory, A. G. Murray, E. S. Munro, and R. S. Raynard. Distribution of infectious pancreatic necrosis virus (IPNV) in wild marine fish from Scottish waters with respect to clinically infected aquaculture sites producing Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 31 (3):177–186, March 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Weng:2007:RDC

[WGS^{+}07] S-P. Weng, Z-X. Guo, J-J. Sun, S-M. Chan, and J-G. He. A reovirus disease in cultured mud crab, Scylla serrata, in southern China. Journal of Fish Diseases, 30(3):133–139, March 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Waagbo:2003:CFA

[WHB^{+}03] R. Waagbø, K. Hamre, E. Bjerkås, R. Berge, E. Wathne, Ø. Lie, and B. Torstensen. Cataract formation in Atlantic salmon, Salmo salar L., smolt relative to dietary pro- and antioxidants and lipid level. Journal of Fish Diseases, 26 (4):213–229, April 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Watermann:2008:PMM

[WHD^{+}08] B. T. Watermann, M. Herlyn, B. Daehne, S. Bergmann, M. Meemken, and H. Kolodzey. Pathology and mass mortality of Pacific oysters, Crassostrea gigas (Thunberg), in 2005 at the East Frisian coast, Germany. Journal of Fish Diseases, 31(8):621–630, August 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Weng:2002:OID

[WHW^{+}02] S. P. Weng, J. G. He, X. H. Wang, L. Lü, M. Deng, and S-M. Chan. Outbreaks of an iridovirus disease in cultured tiger frog, Rana tigrina rugulosa, in southern China. Journal of Fish Diseases, 25(7):423–427, July 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Winton:2001:B

[Win01] James Winton. Biography. Journal of Fish Diseases, 24(2): 63–66, February 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wahli:2002:PKD

[WKB^{+}02] T. Wahli, R. Knuesel, D. Bernet, H. Segner, D. Pugovkin, P. Burkhardt-Holm, M. Escher, and H. Schmidt-Posthaus. Proliferative kidney disease in Switzerland: current state of knowledge. Journal of Fish Diseases, 25(8):491–500, August 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wiedenmayer:2008:MCA

[WKES08] A. A. Wiedenmayer, P. H. Klesius, J. J. Evans, and C. A. Shoemaker. The macrophage chemotactic activity of Edwardsiella tarda extracellular products. Journal of Fish Diseases, 31(5):331–342, May 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wu:2004:ADP

[WM04] J. L. Wu and K. Muroga. Apoptosis does not play an important role in the resistance of ‘immune’ Penaeus japonicus against white spot syndrome virus. Journal of Fish Diseases, 27(1):15–21, January 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wilson:2002:APB

[WMA^{+}02] L. Wilson, S. J. McBeath, K. L. Adamson, P. F. Cook, L. M. Ellis, and I. R. Bricknell. An alkaline phosphatase-based method for the detection of infectious salmon anaemia virus (ISAV) in tissue culture and tissue imprints. Journal of Fish Diseases, 25(10):615–619, October 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Watts:2002:IHI

[WMB02] M. Watts, B. L. Munday, and C. M. Burke. Investigation of humoral immune factors from selected groups of southern bluefin tuna, Thunnus maccoyii (Castelnau): implications for aquaculture. Journal of Fish Diseases, 25(4):191–200, April 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Woo:2003:CTS

[Woo03] P. T. K. Woo. Cryptobia (Trypanoplasma) salmositica and salmonid cryptobiosis. Journal of Fish Diseases, 26(11–12): 627–646, November 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Woo:2004:CIC

[Woo04] P. T. K. Woo. On cryptobiosis and infections by Cryptobia. Journal of Fish Diseases, 27(8):493–494, August 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wu:2000:IPM

[WP00] X. Wu and J. Pan. An intracellular prokaryotic microorganism associated with lesions in the oyster, Crassostrea ariakensis Gould. Journal of Fish Diseases, 23(6):409–414, November 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:2003:SEI

[WSKC03] C. S. Wang, H. H. Shih, C. C. Ku, and S. N. Chen. Studies on epizootic iridovirus infection among red sea bream, Pagrus major (Temminck & Schlegel), cultured in Taiwan. Journal of Fish Diseases, 26(3):127–133, March 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Weng:2002:OIR

[WWH^{+}02] S. P. Weng, Y. Q. Wang, J. G. He, M. Deng, L. Lu, H. J. Guan, Y. J. Liu, and S-M. Chan. Outbreaks of an iridovirus in red drum, Sciaenops ocellata (L.), cultured in southern China. Journal of Fish Diseases, 25(11):681–685, November 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Whipps:2003:CMS

[WWK03] C. M. Whipps, V. G. Watral, and M. L. Kent. Characterization of a Mycobacterium sp. in rockfish, Sebastes alutus (Gilbert) and Sebastes reedi (Westrheim & Tsuyuki), using rDNA sequences. Journal of Fish Diseases, 26(4):241–245, April 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:2005:NLY

[WYJ05] G-L. Wang, S-P. Yuan, and S. Jin. Nocardiosis in large yellow croaker, Larimichthys crocea (Richardson). Journal of Fish Diseases, 28(6):339–345, June 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:2001:OTC

[WYWL01] G. T. Wang, W. J. Yao, J. G. Wang, and Y. S. Lu. Occurrence of thelohanellosis caused by Thelohanellus wuhanensis (Myxosporea) in juvenile allogynogenetic silver crucian carp, Carassius auratus gibelio (Bloch), with an observation on the efficacy of fumagillin as a therapeutant. Journal of Fish Diseases, 24(1):57–60, January 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wang:2008:IIR

[WZJ^{+}08] C. Wang, X-H. Zhang, A. Jia, J. Chen, and B. Austin. Identification of immune-related genes from kidney and spleen of turbot, Psetta maxima (L.), by suppression subtractive hybridization following challenge with Vibrio harveyi. Journal of Fish Diseases, 31(7):505–514, July 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Xu:2009:SMC

[XCD09] J-Y. Xu, S-L. Chen, and H. Ding. Specific MHC class II B alleles associated with resistance to Edwardsiella tarda in turbot, Psetta maxima (L.). Journal of Fish Diseases, 32 (7):637–640, July 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Xu:2002:AMI

[XK02] D-H. Xu and P. H. Klesius. Antibody mediated immune response against Ichthyophthirius multifiliis using excised skin from channel catfish, Ictalurus punctatus (Rafinesque), immune to Ichthyophthirius. Journal of Fish Diseases, 25(5): 299–306, May 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Xu:2003:PEC

[XK03] D-H. Xu and P. H. Klesius. Protective effect of cutaneous antibody produced by channel catfish, Ictalurus punctatus (Rafinesque), immune to Ichthyophthirius multifiliis Fouquet on cohabited non-immune catfish. Journal of Fish Diseases, 26(5):287–291, May 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Xu:2006:ICP

[XKP06] D-H. Xu, P. H. Klesius, and V. S. Panangala. Induced cross-protection in channel catfish, Ictalurus punctatus (Rafinesque), against different immobilization serotypes of Ichthyophthirius multifiliis. Journal of Fish Diseases, 29 (3):131–138, March 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Xu:2002:CAE

[XKS02] D. H. Xu, P. H. Klesius, and R. A. Shelby. Cutaneous antibodies in excised skin from channel catfish, Ictalurus punctatus Rafinesque, immune to Ichthyophthirius multifiliis. Journal of Fish Diseases, 25(1):45–52, January 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Xu:2004:IRH

[XKS04] D-H. Xu, P. H. Klesius, and R. A. Shelby. Immune responses and host protection of channel catfish, Ictalurus punctatus (Rafinesque), against Ichthyophthirius multifiliis after immunization with live theronts and sonicated trophonts. Journal of Fish Diseases, 27(3):135–141, March 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Xu:2005:CAC

[XKS05] D-H. Xu, P. H. Klesius, and C. A. Shoemaker. Cutaneous antibodies from channel catfish, Ictalurus punctatus (Rafinesque), immune to Ichthyophthirius multifiliis (Ich) may induce apoptosis of Ich theronts. Journal of Fish Diseases, 28(4):213–220, April 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Xu:2006:AIM

[XKS06] D-H. Xu, P. H. Klesius, and C. A. Shoemaker. Apoptosis in Ichthyophthirius multifiliis is associated with expression of the Fas receptor of theronts. Journal of Fish Diseases, 29 (4):225–232, April 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Xu:2008:EIC

[XLZS08] H-J. Xu, J-G. Lin, C-X. Zhang, and S. Suzuki. Expression and immunogenic comparison of VP2 and VP3 from marine birnavirus. Journal of Fish Diseases, 31(4):297–304, April 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Xiao:2009:DVA

[XMM^{+}09] P. Xiao, Z. L. Mo, Y. X. Mao, C. L. Wang, Y. X. Zou, and J. Li. Detection of Vibrio anguillarum by PCR amplification of the empA gene. Journal of Fish Diseases, 32(3):293–296, March 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Xie:2004:CTM

[XNS04] H. X. Xie, P. Nie, and B. J. Sun. Characterization of two membrane-associated protease genes obtained from screening out-membrane protein genes of Flavobacterium columnare G_{4}. Journal of Fish Diseases, 27(12):719–729, December 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Xu:2007:ELB

[XSK07] D-H. Xu, C. A. Shoemaker, and P. H. Klesius. Evaluation of the link between gyrodactylosis and streptococcosis of Nile tilapia, Oreochromis niloticus (L.). Journal of Fish Diseases, 30(4):233–238, April 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Young:2009:PLC

[YCR^{+}09] K. M. Young, A. Czyrny, S. Russell, P. Huber, and J. S. Lumsden. Plasma ladderlectin concentration following sterile inflammation and Aeromonas salmonicida subsp . salmonicida infection. Journal of Fish Diseases, 32(7):569–576, July 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Young:2008:DDP

[YDNM08] N. D. Young, I. Dyková, B. F. Nowak, and R. N. Morrison. Development of a diagnostic PCR to detect Neoparamoeba perurans, agent of amoebic gill disease. Journal of Fish Diseases, 31(4):285–295, April 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yokoyama:2005:HPS

[YIT05] H. Yokoyama, N. Itoh, and S. Tanaka. Henneguya pagri n. sp. (Myxozoa: Myxosporea) causing cardiac henneguyosis in red sea bream, Pagrus major (Temminck & Schlegel). Journal of Fish Diseases, 28(8):479–487, August 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yuasa:2002:IRH

[YKR^{+}02] K. Yuasa, I. Koesharyani, D. Roza, K. Mori, M. Katata, and T. Nakai. Immune response of humpback grouper, Cromileptes altivelis (Valenciennes) injected with the recombinant coat protein of betanodavirus. Journal of Fish Diseases, 25(1):53–56, January 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yamashita:2009:NAL

[YMKN09] H. Yamashita, K. Mori, A. Kuroda, and T. Nakai. Neutralizing antibody levels for protection against betanodavirus infection in sevenband grouper, Epinephelus septemfasciatus (Thunberg), immunized with an inactivated virus vaccine. Journal of Fish Diseases, 32(9):767–775, September 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ye:2008:RVA

[YML^{+}08] J. Ye, Y. Ma, Q. Liu, D. L. Zhao, Q. Y. Wang, and Y. X. Zhang. Regulation of Vibrio alginolyticus virulence by the LuxS quorum-sensing system. Journal of Fish Diseases, 31 (3):161–169, March 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yamashita:2009:PCA

[YMN09] H. Yamashita, K. Mori, and T. Nakai. Protection conferred against viral nervous necrosis by simultaneous inoculation of aquabirnavirus and inactivated betanodavirus in the sevenband grouper, Epinephelus septemfasciatus (Thunberg). Journal of Fish Diseases, 32(2):201–210, February 2009. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yoganandhan:2004:PPA

[YMNH04] K. Yoganandhan, S. Syed Musthaq, R. B. Narayanan, and A. S. Sahul Hameed. Production of polyclonal antiserum against recombinant VP28 protein and its application for the detection of white spot syndrome virus in crustaceans. Journal of Fish Diseases, 27(9):517–522, September 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yoganandhan:2003:LEP

[YNH03] K. Yoganandhan, R. B. Narayanan, and A. S. Sahul Hameed. Larvae and early post-larvae of Penaeus monodon (Fabricius) experimentally infected with white spot syndrome virus (WSSV) show no significant mortality. Journal of Fish Diseases, 26(7):385–391, July 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yang:2006:SSM

[YSH^{+}06] B. Yang, X-L. Song, J. Huang, C-Y. Shi, Q-H. Liu, and L. Liu. A single-step multiplex PCR for simultaneous detection of white spot syndrome virus and infectious hypodermal and haematopoietic necrosis virus in penaeid shrimp. Journal of Fish Diseases, 29(5):301–305, May 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yanagida:2006:TED

[YSN^{+}06] T. Yanagida, M. Sameshima, H. Nasu, H. Yokoyama, and K. Ogawa. Temperature effects on the development of Enteromyxum spp. (Myxozoa) in experimentally infected tiger puffer, Takifugu rubripes (Temminck & Schlegel). Journal of Fish Diseases, 29(9):561–567, September 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yugueros:2001:MCY

[YTLN01] J. Yugueros, A. Temprano, J. M. Luengo, and G. Naharro. Molecular cloning of Yersinia ruckeri aroA gene: a useful taxonomic tool. Journal of Fish Diseases, 24(7):383–390, September 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yoganandhan:2005:SDM

[YWBH05] K. Yoganandhan, J. Sri Widada, J. R. Bonami, and A. S. Sahul Hameed. Simultaneous detection of Macrobrachium rosenbergii nodavirus and extra small virus by a single tube, one-step multiplex RT-PCR assay. Journal of Fish Diseases, 28(2):65–69, February 2005. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

You:2008:SGB

[YWS^{+}08] P. You, Y. Wang, X. Sun, X. Qiang, and D. Cone. Seasonality of Gyrodactylus brachymystacis Ergens on farmed rainbow trout, Oncorhynchus mykiss (Walbaum), in central China, with a report of an infection on wild Manchurian trout, Brachymystax lenok (Pallas). Journal of Fish Diseases, 31 (12):941–945, December 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

You:2006:PIG

[YYY^{+}06] P. You, B. Yuan, J. Yang, R. Easy, Z. Dong, and D. Cone. Pathogenic infections of Gyrodactylus brachymystacis (Monogenea) on Oncorhynchus mykiss (Walbaum) at a fish farm in the Qinling Mountain region of China. Journal of Fish Diseases, 29(5):313–316, May 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhang:2000:PVH

[ZA00] X-H. Zhang and B. Austin. Pathogenicity of Vibrio harveyi to salmonids. Journal of Fish Diseases, 23(2):93–102, March 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zorrilla:2003:VSI

[ZAC^{+}03] I. Zorrilla, S. Arijo, M. Chabrillon, P. Diaz, E. Martinez-Manzanares, M. C. Balebona, and M. A. Moriñigo. Vibrio species isolated from diseased farmed sole, Solea senegalensis (Kaup), and evaluation of the potential virulence role of their extracellular products. Journal of Fish Diseases, 26(2): 103–108, February 2003. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhang:2006:CLP

[ZASK06] Y. Zhang, C. R. Arias, C. A. Shoemaker, and P. H. Klesius. Comparison of lipopolysaccharide and protein profiles between Flavobacterium columnare strains from different genomovars. Journal of Fish Diseases, 29(11):657–663, November 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhang:2007:RLV

[ZB07] S. Zhang and J-R. Bonami. A roni-like virus associated with mortalities of the freshwater crab, Eriocheir sinensis Milne Edwards, cultured in China, exhibiting ‘sighs disease’ and black gill syndrome. Journal of Fish Diseases, 30(3):181– 186, March 2007. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zanoni:2008:OMS

[ZFF^{+}08] R. G. Zanoni, D. Florio, M. L. Fioravanti, M. Rossi, and M. Prearo. Occurrence of Mycobacterium spp. in ornamental fish in Italy. Journal of Fish Diseases, 31(6):433–441, June 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zilberg:2001:PSA

[ZGM01] D. Zilberg, A. Gross, and B. L. Munday. Production of salmonid amoebic gill disease by exposure to Paramoeba sp. harvested from the gills of infected fish. Journal of Fish Diseases, 24(2):79–82, February 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zimba:2001:CCI

[ZKG^{+}01] P. V. Zimba, L. Khoo, P. S. Gaunt, S. Brittain, and W. W. Carmichael. Confirmation of catfish, Ictalurus punctatus (Rafinesque), mortality from Microcystis toxins. Journal of Fish Diseases, 24(1):41–47, January 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zilberg:2000:PEA

[ZM00] D. Zilberg and B. L. Munday. Pathology of experimental amoebic gill disease in Atlantic salmon, Salmo salar L., and the effect of pre-maintenance of fish in sea water on the infection. Journal of Fish Diseases, 23(6):401–407, November 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zilberg:2001:RAS

[ZM01a] Zilberg and Munday. Responses of Atlantic salmon, Salmo salar L., to Paramoeba antigens administered by a variety of routes. Journal of Fish Diseases, 24(3):181–183, March 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zilberg:2001:EAP

[ZM01b] D. Zilberg and B. L. Munday. The effect of anti- Paramoeba antibodies on Paramoeba sp., the causative agent of amoebic gill disease. Journal of Fish Diseases, 24(6):345–350, July 2001. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zimba:2004:IEA

[ZRT04] P. V. Zimba, M. Rowan, and R. Triemer. Identification of euglenoid algae that produce ichthyotoxin(s). Journal of Fish Diseases, 27(2):115–117, February 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhang:2004:AIM

[ZS04] C. X. Zhang and S. Suzuki. Aquabirnaviruses isolated from marine organisms form a distinct genogroup from other aquabirnaviruses. Journal of Fish Diseases, 27(11):633–643, November 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhang:2004:PCN

[ZSZB04] S. Zhang, Z. Shi, J. Zhang, and J-R. Bonami. Purification and characterization of a new reovirus from the Chinese mitten crab, Eriocheir sinensis. Journal of Fish Diseases, 27(12): 687–692, December 2004. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zuo:2000:VHP

[ZW00] X. Zuo and P. T. K. Woo. In vitro haemolysis of piscine erythrocytes by purified metallo-protease from the pathogenic haemoflagellate, Cryptobia salmositica Katz. Journal of Fish Diseases, 23(3):227–230, May 2000. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhang:2006:CPD

[ZWW^{+}06] J. Y. Zhang, J. G. Wang, Y. S. Wu, M. Li, A. H. Li, and X. L. Gong. A combined phage display ScFv library against Myxobolus rotundus infecting crucian carp, Carassius auratus auratus (L.), in China. Journal of Fish Diseases, 29(1):1–7, January 2006. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhou:2002:UIS

[ZWZS02] Y-C. Zhou, J. Wang, B. Zhang, and Y-Q. Su. Ultrasonic immunization of sea bream, Pagrus major (Temminck & Schlegel), with a mixed vaccine against Vibrio alginolyticus and V. anguillarum. Journal of Fish Diseases, 25(6):325– 331, June 2002. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zhou:2008:IGC

[ZXZ^{+}08] S. M. Zhou, M. Q. Xie, X. Q. Zhu, Y. Ma, Z. L. Tan, and A. X. Li. Identification and genetic characterization of Streptococcus iniae strains isolated from diseased fish in China. Journal of Fish Diseases, 31(11):869–875, November 2008. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).